A Complete Bibliography of Publications in Dialectica

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

24 April 2024 Version 1.10

Title word cross-reference

$50.00 [2147]. ^{10^{1}0} [319]. ^{∘} [1515, 1527, 1533, 1539, 175, 401, 480]. ^{e} [198, 22]. th [1504]. ^{r} [128]. · [535, 382, 599]. F [1639].

-glycohydrolase [1064, 1064].

/Closed [1612]. /Functional [1612].

1 [1390]. 17 [1515]. 18 [1090, 1527]. 18th [1090]. 19 [1533]. 1945 [24]. 1949 [233]. 1953 [198]. 1954 [245, 253, 247, 255, 246, 254]. 1967 [626].

2 [1412]. 20 [1539]. 2002 [1629]. 2006 [1734]. 2008 [1802]. 2e [264]. 2nd [264].

4 [2147]. 43 [1262].

7 [2144, 2146]. 7th [1012].

9 [2145]. 978 [2144, 2145, 2147, 2146]. 978-0-19-871632-7 [2144]. 978-0-19-871662-4 [2147]. 978-0-7456-8906-7 [2146]. 978-1-137-54180-2 [2145]. 978-1-137-54181-9 [2145].

Êcole [176]. être [317, 1102, 921, 68].

A-Theory [2206, 1651]. A. [1071, 640]. Aarnio [1692]. Abgeschlossene [88]. Abhandlung [869]. abierta [757, 757]. aboutit [1284]. aboutit-elle [1284]. Abraham [1163]. Abschied [1219]. Absences [2046]. absorption [1059]. Abstract [546, 1442, 1361, 249]. Abstraction [1635, 859, 859]. Abstractions [1920]. Abstracts [481, 535, 607, 593, 606, 693, 724, 704, 705, 711, 751, 778, 378, 379, 304]. Abstrakte [1953]. absurdité [132]. absurdity [132]. Abusing [1414]. academic [988]. Acceptance [1479]. Access [2205, 2163, 1691]. Accommodating [1142]. According [1761, 963, 1102, 1355]. Account [2189, 1226, 1764, 1909, 1666, 1734, 1895, 1537, 2037, 1756, 1798, 2111]. Accounts [2082, 1215, 1387, 1223]. Accuracy [2211, 1963]. Accuracy-Dominance [2211]. achievement [313]. Achievements [1996]. Achille [1884]. Achinstein [1198]. Achtung [1035]. Acoustic [767, 150]. Acquaintance [2053, 1271]. acquired [618]. acquis [618]. Across [2238]. Acta [1112]. acte [1310]. Actes [320, 626]. Action [1116, 1076, 2054, 1251, 399, 1200]. Actions [1730, 925]. activité [370]. activités [374, 977]. activities [374, 146, 977, 146]. Activity [2147, 370, 1185]. Acts [1113]. Actualism [1643]. actualité [33, 21, 10]. Actually [1893, 1828]. Actually-Rigidified [1893]. actuel [635]. actuelle [366]. actuels [633]. Acyclic [1834]. Ad [961]. adage [575]. adäquate [281, 280, 250, 282]. addition [210]. address [19]. Adequate [281, 250, 282]. Adjointness [749]. Adjudication [2234]. adjustment [148]. administratif [572]. administrative [572]. Adopt [1342]. adresse [19]. Advances [2135]. adventure [292]. Advertisement [25, 477, 727]. Advices [1180, 1206, 1233, 1263, 1294]. Adzić [2168]. Aesthetic [1729, 1036]. Aesthetics [70, 73, 70, 762]. Aesthetik [70, 73]. Affairs [1739]. Affective [1730, 1725, 1724, 1727, 1729, 1726, 1728, 1731, 1724]. Affirmation [2071]. africaines [528]. After [1851, 76, 1875]. Again [1759]. Against [1808, 2193, 1765, 1217, 2167, 2098, 1534, 845, 1818, 1832, 2034, 1011, 1984, 1618, 2120, 2124]. age [186]. Agency [1716, 1746, 2127, 2140, 1732, 2085, 2001]. Agent [2037, 1853]. Agent-Relativity [1853]. Agential [2029]. Aggregates [1696]. Aging [938]. Agnosticism [1663]. agreement [1190]. Aim [1880]. Aims [1058]. AKRASIA [833]. Alan [2091, 1709]. Albert [233]. aléatoire [457]. Alethic [2035, 1902]. Alexander [281, 611, 281, 611]. Alexandra [2144]. Alexandre [280]. Alfred [302]. Algebra [1812]. Alien [1681]. Alienated [1870]. alla [1632, 1633, 1630, 1910, 1631]. allegans} [575]. allemand [499]. allemande [2050]. Allez [843]. Allgemeine [846]. allucinazioni [1884]. Alternative [1397, 818]. Alternatives [1386]. altri [2060]. am [769]. américain [904]. American [904]. Amie [2123]. Among [1004, 1177]. Amoral [2127]. Amore [2060]. amplifiante [105]. Amplifying [105]. analisi [150]. Analog [1957, 558, 557]. Analog/Digital [1957]. Analogical [555]. Analogie [969, 548, 558, 550, 551, 612, 556, 561]. Analogies [554, 555, 553, 554]. analogique [550, 555]. analogiques [558, 557]. analogue [550]. Analogy [1327, 552, 559, 969, 1004, 558, 550, 551, 553, 556, 561, 548, 612]. Analyse [1429, 368, 1064, 1165, 1119, 537]. Analyses [1061]. Analysis [699, 896, 1937, 1439, 1308, 971, 2215, 1754, 1543, 2231, 1109, 833, 891, 1056, 829, 1429, 517, 1120, 150, 1064, 368, 1165, 1119, 1223, 537]. Analytic [1515, 1527, 1533, 1539, 1892, 844, 375, 1629, 2011, 1478, 1431]. analytical [1368, 708, 1434]. Analytically [1009]. analyticité [1429]. Analyticity [1042, 1071, 1318, 1647, 1429]. analytique [1368, 844, 708]. Anaphoric [1894]. anatomical [189]. anatomiques [189]. Ancient [950, 974, 273, 900]. and‘linguistic [614]. Andrea [2169, 2060]. Andreas [1747, 2049, 244]. Andrej [1720]. Anfangsgründe [1486]. Anhang [1953]. Animal [1851]. animal^{≫} [144]. Animalism [1965]. Animals [1077, 1822]. Anjum [2003]. Anknüpfung [308, 315]. Annabelle [1967]. Annahme [1034]. Anne [1777]. année [330]. Announcement [939]. annual [33]. annuelle [1262, 1293, 1317, 33]. Anomalism [1871]. Anomalous [1444, 1466]. anomaly [89]. ansätze [475]. Answer [2182, 663, 2016, 282, 518, 194, 197, 640]. answers [671, 679, 642]. Anthony [2038]. anthropological [166]. Anthropologie [275, 184]. anthropologischen [166]. Anthropology [275, 184]. Anti [1643, 1789, 1255, 1387, 1678, 1244, 2166, 1742, 1237, 2194]. Anti-Actualism [1643]. Anti-Kant [1789]. Anti-Naturalism [1387]. Anti-Nihilist [2166]. Anti-Realism [1255, 1244, 1742, 1237, 2194]. Anti-Realist [1678]. Antidühring [76, 76]. Antike [1218]. antimécanisme [1176]. antimechanism [1176]. Antinomien [290]. Antinomies [573, 570, 574, 572, 580, 567, 577, 571, 566, 568, 581, 576, 569, 290, 578, 567, 571, 573, 581, 576, 570]. Antinomy [1831]. Antique [900]. Antirealism [1239]. Antiskeptical [1048]. Antonelli [2214]. Antwort [643, 657, 282]. Anweisungen [1182, 1234]. Anwendungen [1950]. any [372]. août [198, 264]. apercu [1139]. Apologies [1161]. apparatus [178]. Apparent [1796]. apparition [136]. Appeal [500, 503]. appearance [234, 136]. Appearing [1679]. Appendices [690]. appendix [235]. Application [848, 603, 558]. applications [359, 2103, 2134]. Applied [1065, 344]. appliqué [1065]. appliquée [803]. apport [435]. apprentissage [427]. Approach [2035, 1804, 1245, 1215, 1241, 1119, 165, 1358, 796, 151]. Approaches [1888, 475]. Approaching [2102]. approche [1119, 796]. appropriate [344]. Approximation [1012]. Apriori [765]. apriorism [512]. Apriority [1031]. Apt [1785]. Arapinis [2144]. Arbeit [207].

Arbeitshypothese [819]. Arcangeli [2204]. Arche [2215]. Arche-fossil [2215]. archetypal [285]. archetypischer [285]. Argue [2193]. Argument [1455, 1229, 1582, 1872, 1939, 1556, 983, 1918, 1832, 1555, 1030, 2174, 2116, 1954, 1550, 1671, 2034, 2099, 2231, 1958, 1043, 1655, 1987, 1669, 1818, 2185, 2016, 2219]. Arguments [1955, 1489, 1664, 1534, 2158, 1573, 1984]. Arianna [2120]. Aristoteles [211]. Aristotelian [1710, 1349, 1968]. aristotélicienne [1349]. Aristotle [1628, 211]. Arithmetic [1256, 1252, 1100]. Arithmetical [1610]. Armchair [2163]. Armour [2134]. Armour-Garb [2134]. Armstrong [1847, 2229]. Arnaud [294]. Arrow [1418, 1876]. Art [1461, 720, 773, 1969, 920, 467, 552, 169, 987, 720, 2070]. Artbegriff [27]. Article [1478, 440, 280]. articles [378, 379, 304]. Articulated [1932]. articulation [656, 659, 679]. Artificial [554]. artificielle [554]. Artiga [2190]. Artistic [1752]. arts [16]. Arvan [2145]. Ascent [1485, 1210]. Asia [902]. Asian [905]. asiatiques [905]. Asking [1940]. Aspect [367, 279, 903]. Aspects [560, 902, 982, 1367, 90, 291, 297, 780, 801, 1369, 203]. Aspekte [297]. aspirations [1462]. assay [1054]. Assent [1479, 1355]. Assertibility [1907, 1906]. Assertion [1800, 1914, 2216]. Assessment [1078, 2103]. Assignments [1817]. assimilation [64]. Associated [2162]. Association [939]. Assumption [1074, 1034, 1322]. Assumptions [864, 1076, 181, 1214, 1085]. ästhetischen [1036]. Asymmetry [957, 1694]. Athanasios [2012]. Atomenlehre [1789]. Atomic [366, 14]. atomique [14, 366]. Attachment [2071, 185]. Attack [2121]. Attention [1592]. Attitude [1036]. Attitudes [2025, 1113, 2100, 1565, 343]. attributes [31]. attributing [156]. Attributions [1817]. Attributive [1928]. attributs [31]. auditur [575]. Aufbau [141]. Auffassung [838]. Auflage [1953]. August [198, 264, 266, 265]. aujourd’hui [1404]. Auseinandersetzung [203]. Aussagen [473]. Aussagenlogik [1169]. Aussagenlogische [423]. Austauche [55]. Austin [983]. Austro [2050, 2050]. austro-allemande [2050]. Austro-German [2050]. Auteurs [1181, 305, 329, 361, 393, 1232, 1263, 1294]. Authenticity [1725, 1936, 2124]. authentiquement [877]. author [378, 379]. Authority [1445, 1108, 1168]. Authors [1180, 1182, 1206, 1233, 1234, 1263, 1294]. Autofondation [771]. Automatic [795]. automatique [795]. automne [245, 253]. Autonomie [195]. Autonomy [1992, 2095, 195]. Autopsie [1062]. Autopsy [1062]. Autoren [1182, 1234]. autour [1383]. autres [518]. Autumn [245, 253, 247, 255]. Aux [1181, 1280, 1232, 1263, 1294, 31, 4, 148, 573, 793, 803, 142]. Avant [909, 496, 565, 856, 872, 1344, 1126]. Avant-propos [909, 496, 565, 856, 872, 1344, 1126]. Avertissement [57, 114, 477, 727, 25]. avocat [502]. awareness [173]. Axel [2093]. axiological [1156]. Axiom [1484, 424]. axiomatic [372, 1121]. Axiomatics [271]. Axiomatik [271]. axiomatique [372, 1121]. axiomatiques [972]. Axiomatizable [774]. Axiomatizations [1666]. axiomatize [214]. axiomes [1186]. Axioms [930, 972, 1186, 1535, 424]. ayant [618].

B [2134, 236, 44, 651, 642, 632, 561, 1651]. B-Theory [1651]. Bachelard [1127]. Background [1641, 629]. Bad [2030, 1892, 1886, 2142]. Bader [2005]. Bailey [2091]. Banana [1735]. Band [1262, 1293, 1317]. Barcan [1482]. Barnes [1339]. Barry [1810, 1339]. Bartels [1747]. Based [2065, 1895, 1823]. Basic [1941, 1425, 1637, 1882, 545, 547, 2168, 738, 181, 250, 282, 161, 523, 2186]. Basingstoke [2145]. Basis [99, 1171, 147, 270]. Bastide [528]. Battle [1469]. battlefield [578]. Bayer [72]. Bayesian [1537]. Be [1988, 2141, 1068, 2117, 1252, 1639, 2142, 2124, 2018, 442, 363, 1535, 1102, 1529, 136, 1858, 1797]. Beables [1133]. Beantwortung [640]. Bearers [835, 883]. Becker [295]. Bedeutung [141, 220]. Beeinflussung [166]. Before [942, 503, 500, 501]. Begegnung [471]. Beginning [844, 843]. Beginnings [1486]. Begriff [17, 332, 88, 7]. Begriffes [841]. Begriffsschrift [1531]. Begründung [100, 413, 272]. Begrüssungsworte [58]. behavior [178, 933]. Behaviour [1089, 149]. Behaviourism [1170]. beiden [226]. Being [2013, 1919, 2147, 49, 921, 1775, 261, 68]. Beispiel [1520]. Beitrag [1747]. Bela [313]. belge [276]. Belgian [276]. Belgique [506, 505, 500, 21]. Belgium [506, 505, 500, 21]. Belief [1974, 1833, 2092, 2226, 1915, 1912, 1479, 1377, 1766, 1870, 1387, 1623, 1785, 1914, 945, 1092, 1889, 1043, 1913, 1074, 999, 1355]. Belief-Independent [1623]. Beliefs [2129, 2162]. Believe [1800, 1759, 2066, 1880, 1830]. Believing [1006]. Bell [1995, 1134]. Bellarmine [2234]. Bemerkungen [123, 1087, 73, 281, 1352, 432, 639, 207]. Ben [2022]. Benis [2083]. Benis-Sinaceur [2083]. Benjamin [1975]. Bennett [2157]. Benovsky [2118]. benützte [414]. Beobachtung [469]. Beobachtungssprache [411]. Beobachungsgrundlagen [338]. Berechtigung [116]. Berger [483]. Bergmann [1874]. Berkeley [208, 288]. Bernard [896, 1045]. Bernay [969]. Bernays [1751, 413, 940, 968, 413, 197, 969, 898]. Bernoulli [100]. Berto [1721]. Bertrand [1084, 810]. Best [1988, 2141, 1382, 1240, 1354]. Beth [591]. béton [1281]. Betrachtung [281, 1343, 1365, 1370, 1405, 1410]. Betrachtungen [624, 691, 725, 742, 777, 783, 788, 798, 811, 823, 827, 847, 855, 878, 889, 907, 960, 984, 1023, 1044, 1049, 1083, 1098, 1105, 1124, 1178, 1230, 1253, 1260, 1291, 1301, 1315, 1323, 1328, 1337, 1093, 1212, 989, 93, 1137, 1224, 818, 804, 933, 934, 935, 936, 937, 938, 1164, 1199, 333]. betreffend [640]. Better [1988, 1771]. Betti [2120]. Bettina [1747]. Between [2169, 1195, 1675, 409, 502, 957, 2162, 219, 1556, 210, 500, 577, 1174, 1824, 510, 504, 1084, 579, 530, 979, 874, 256]. Betweenness [930]. Bewusstsein [1200]. bewusstseinsfähiger [1747]. Beyer [1747]. Beyond [2023, 1812]. Beziehung [409]. Bibliographic [46, 158]. Bibliographical [283]. bibliographique [283]. bibliographiques [232, 46]. bibliographische [46, 158]. Bibliography [1714]. Biennial [939]. Big [2141]. Bilder [1747]. Bildung [611, 285]. Binding [2016]. Binyamin [1582]. biochemical [1054]. biochemistry [1053]. biochimie [1053]. biochimique [1054]. Biography [2146]. Biological [965, 781, 116, 1242, 780, 932, 333, 874]. Biologie [1053, 1334, 700, 28, 1058, 612, 876, 877, 1057]. biologique [965, 780, 932]. biologiques [874]. biologische [116]. biologischen [333]. biologists [142]. Biology [700, 1359, 2065, 782, 1089, 28, 1058, 612, 876, 877, 1334, 1057, 1053]. biomécanisme [353]. biomechanism [353]. biophysics [1053]. biophysique [1053]. birth [9]. bis [6]. Bisher [414]. Bivalence [1712]. Bizarre [1490]. Black [1267, 1269, 1265, 1270]. Blackburn [2116]. Blame [2180]. Blick [1747, 418]. Bloomsbury [2136]. Blumson [2022]. Bocheński [273]. Body [1677, 1202]. Boethius [235]. Boghossian [1808, 1857, 2234, 1905]. Bohr [2184, 808, 1248, 90, 808]. Bolzano [1789, 2125]. Bolzanos [1720]. Bonding [2114]. Bonino [1874, 1876]. Bonsack [534]. Book [379, 1667, 2144, 2004, 2001, 2120, 2022, 1746, 1911, 1748, 1877, 1974, 1966, 1941, 1774, 2091, 2051, 2081, 1808, 1790, 1884, 2170, 1628, 2092, 1940, 1783, 2013, 2003, 1810, 2090, 1995, 2012, 1993, 2039, 1951, 1747, 2184, 2049, 2082, 1632, 1633, 1849, 1630, 1782, 1976, 2169, 1786, 1649, 1847, 2204, 1776, 1910, 2145, 1699, 2122, 2093, 2040, 1789, 1788, 2072, 1760, 1763, 2103, 2061, 1721, 1874, 2083, 1975, 2186, 2134, 1994, 2048, 1629, 1631, 1848, 2019, 2015, 1785, 2147, 2118, 1737, 2011, 1787, 1773, 2202, 2101, 1777, 18, 304, 2146, 2157, 1885, 2158, 1876, 2021, 1952]. Book [1596, 2203, 2071, 2002, 1943, 2136, 1698, 2156, 2060, 1968, 2123, 1738, 1950, 2050, 2159, 2070, 2125, 1723, 2168, 1873, 1761, 1745, 2020, 2102, 1953, 1087, 2119, 1697, 2171, 1722, 1775, 1784, 1781, 2038, 2059, 2121, 1967, 1762, 1809, 1942, 2124, 2185, 1720, 2014, 1875, 611, 312, 39, 1777, 2050]. Books [692, 703, 726, 743, 908, 955, 991, 1024, 1050, 1080, 1099, 1125, 1154, 1179, 1205, 1231, 1261, 1292, 1316, 1338, 1360, 1380, 1400, 1423, 112, 138, 159, 267, 328, 360, 522, 592, 605, 38]. booktitle [2050]. Boole [41]. Bopp [39, 125]. botanique [30]. botany [30]. Both [1931, 1382, 1240]. Bottom [1660]. bouddhisme [902]. Bound [2016]. Bounds [1716]. Boydian [1313]. Bradley [1981, 1806, 1838, 1528]. Brain [1743, 815, 374, 146]. Brains [1419, 1502]. Breaking [1672]. Brentanian [2178]. Brentanians [1690]. Brentano [2020, 2203, 1690]. Brésil [528]. Brèves [241]. Brevissima [1910]. Brief [54, 241, 20]. Briefwechsel [350]. British [336, 693, 296, 314, 326, 342, 351, 378, 402, 458, 520, 546, 607, 724, 304]. Broad [1363]. Brogaard [2189]. Brouwer [975]. Brownian [104]. Brunner [1345, 1346]. brüsseler [256]. Brussels [198, 256]. Brutalist [2150]. Brute [2193, 2098]. Bruxelles [198]. Buch [611, 1087]. Buchbesprechung [327, 1595, 233]. Buchbesprechungen [1421, 1432, 1440, 1447, 1453, 1459, 1475, 1494, 1498, 1503, 1508, 1514, 1521, 1532, 1538, 1544, 1552, 1558, 1577, 1584, 1599, 1470]. bücher [138, 328, 692, 703, 726, 743, 799, 908, 955, 991, 1024, 1050, 1080, 1099, 1125, 1154, 1179, 1205, 1231, 1261, 1292, 1316, 1338, 1360, 1380, 1400, 1423]. Buddhism [902]. build [617]. Builders [1530]. Bunkum [1196]. Burge [1319]. Business [2078, 794]. Buts [1058]. bzw [1088].

C [210]. Caesar [1572, 1637, 1638, 1641]. cailloux [1281]. Calcul [1122, 270, 601, 99, 136]. Calculating [99, 601]. Calculation [271, 417, 270, 136, 272]. Calculus [687, 1122]. Calculuses [412]. Calibration [1919]. Call [2036]. called [809]. Cambridge [2146, 1474]. Cameron [1806]. Can [985, 2207, 1822, 1252, 1794, 1858, 1797, 1948, 363, 442]. Cannot [2167, 1645, 2021]. Canonical [1686]. Cantor [1336, 1455, 860]. Cantors [1455]. Capacities [1221, 1927]. Capital [1810]. Caputo [1722]. caractère [1379, 372, 1065]. Caractéristiques [366, 1139]. Caractérologie [110, 128, 175, 334, 401, 480]. caracterology [334, 401]. career [131]. Carlos [1774]. Carnap [1491, 1429]. Carpintero [1713, 1632, 1839, 1787]. Carrara [2144]. carré [1094]. carrière [131]. cas [620]. Casalegno [1911, 1904, 1909, 1910, 1791, 1903, 1905, 1907, 1906]. Case [2115, 2032, 2009, 506, 620]. case-law [506]. Cases [832]. Cassam [2048]. cassation [503, 500]. Categorial [2014]. Categorical [2088, 2232]. Categories [1025, 1165, 1165, 2013]. Category [2218, 1699, 734]. Catullus [767]. Causal [1930, 1376, 2046, 2198, 1626, 1259, 2099, 2128, 1818, 1057]. causale [1057]. Causalité [772, 1349]. Causality [85, 851, 841, 1493, 108, 1034, 1349, 772]. Causation [1795, 1743, 1622, 1555, 1018, 2210, 1222, 1428, 2210]. cause [1395, 287]. cause-effect [287]. Causes [2003, 1467]. Cavaillés [1367]. Ce [1285, 787]. cell [1055]. celui [357]. centenaire [9]. centenary [9, 41]. Centered [1387]. Centrone [1885]. Century [2184, 1786, 850, 1090]. cérébrales [374]. Certain [1474, 1321, 156]. certaines [156]. Certainty [2216, 96]. certitude [96]. cerveau [146]. César [1640]. chaines [557]. Chains [1806, 557]. chair [59]. Challenge [2029, 2215, 2075, 1841]. Challenging [2190]. champ [865]. chance [469]. Chances [1122]. Change [1718, 1779, 1655, 2014]. Changes [1474, 1114]. Changing [1657]. Character [2026, 1379]. Characteristics [1817, 1139, 462, 366]. Characterization [1957]. characterology [128, 175, 480, 110]. Characters [1562, 1567, 1570]. Charakterzüge [462]. Chardin [299]. Charge [1444]. Charity [1209, 1208, 1173]. Chemistry [1899]. chez [921, 1302, 1176, 1351, 1064, 715, 1122, 1258]. Choice [1672, 1484, 1354, 2118]. Choices [1128]. choix [1354]. choses [813, 1667]. Christian [1747]. Christine [1730, 1725, 1724, 1727, 1729, 1726, 1728, 1731]. Christopher [2093]. Chronicle [527, 128, 334, 401, 480]. Chronique [457, 128, 175, 334, 251, 263, 341, 527, 401, 480]. Chronology [175]. Chrysippean [1094]. Chrysippéen [1094]. Church [2122]. Circle [979, 981]. Ciuni [2015]. civilisation [170]. civilization [170]. clôture [494]. Claims [815]. Claims^{≫} [867]. Clandestine [1768]. Clarke [1872]. Class [2175, 964, 1612, 2191, 1142]. classement [569]. Classes [1883, 1095]. Classical [820, 2121, 498, 596, 1169]. classication [30]. Classicism [2121]. classics [1571]. classification [30, 538]. Classificational [802]. classique [498]. classiques [596]. Claude [453, 967, 1045]. Clavien [1730, 1725, 1724, 1727, 1729, 1726, 1728, 1731]. click [718]. Close [2189]. Closed [1612, 88, 2236]. closing [15, 494]. Closure [1842, 1901, 1742, 2089, 928]. Clusters [831, 1825]. CNC [1518]. Co [626]. Co-ordination [626]. Cocteau [134]. code [575, 575]. Cogito [410]. Cognate [1738]. Cognition [1018, 2188, 1777, 400, 1777, 1941]. Cognitive [696, 2232, 1971, 2177, 1553, 1016, 695, 406, 1200, 1969, 1970, 2089, 977]. cognitives [977]. Cognitivist [1989]. Coherence [2152, 1917, 238, 238]. Coherent [1917]. Coherentism [2090]. Coincide [1962]. Coincidence [1575]. Collected [1719]. Collective [2139, 2127, 1401, 1731]. colloque [626]. Color [1677, 1681, 1673, 1679, 1680]. Colors [1829, 1679, 1682]. Colour [1867, 1556]. Colours [1678, 1579]. Combination [1899]. combinatory [738]. Come [1427, 843]. comédie [1216]. comedy [1216]. Comité [863]. Comités [495]. Command [912]. Commandment [912]. comme [352]. commencement [844, 843]. Comment [617, 888, 877, 723, 1413, 448]. commentaires [1134]. Commentators [1763, 1784]. Comments [1683, 1654, 1656, 1684, 1866, 1685, 1658, 964, 1652, 723, 1575, 884, 886, 1134, 755]. Commitment [1438, 384, 388, 386, 387, 385]. Commitments [1501]. Committee [863]. Committees [495]. Common [1868, 1047, 1190, 999]. Common-Sense [1868]. communautaire [577]. communauté [903]. commune [999]. communicated [999]. Communication [824, 2153, 1827, 802, 848, 2161, 215, 72, 373]. Communications [34, 2225]. communiquée [999]. community [577, 903]. compagnie [134]. Companion [2091, 2136]. company [134]. comparative [353, 353]. comparée [353]. Comparison [1046, 1088]. Comparisons [2141]. Compatibility [1958]. Compatible [1929, 1665, 1871, 136]. Competence [1955]. Competition [1969]. complément [210]. complémentarite [1053, 91, 90, 92, 87, 91]. Complementarity [85, 91, 90, 127, 92, 108, 84, 1053, 87, 91]. Complete [1797, 803, 1322, 1322]. Completeness [2198, 421]. complètes [803]. Complex [1966, 1980]. Complexity [1033]. Component [1464, 2132, 1818, 613]. comportement [933]. compose [1583]. Composition [2182, 2112, 1765]. Compositionality [1794]. compréhension [397]. comprends [79]. Compte [862]. Computational [1403]. Computations [1612]. computer [793, 794]. computing [791, 801, 796]. Conceivability [2177, 1675, 2221]. Concept [1855, 1608, 1628, 98, 841, 1086, 2114, 1114, 487, 791, 17, 332, 1287, 368, 1325, 166, 27, 84, 385, 7, 622, 98]. Conception [838, 580, 1047, 1363, 2101, 2125, 1284]. Conceptions [1073, 1760, 1869, 1762, 927, 181]. Concepts [2108, 1695, 2053, 1609, 2153, 2042, 1841, 2173, 1962, 815, 1825, 1074, 1826, 15, 1541, 768]. Conceptual [1249, 1792, 1519, 1824, 1187, 1556]. conceptualism [2149]. concernant [671, 916, 280]. concerning [671, 916, 280, 566, 1201, 649, 640]. Concerns [2096]. Conclusions [906, 312]. conclusiveness [1906]. concordance [241, 238]. concrete [1281, 399, 399]. Condillac [785]. Conditional [976]. Conditionalization [1833]. Conditions [1939, 1295, 2186, 1074, 1321, 491, 186]. Conference [257, 257]. conférences [2203, 684]. Confidence [1976]. confines [1280]. confins [1280]. Confirm [1324]. Confirmation [1537, 849, 559]. Confusion [2208]. Congrés [245, 253, 264, 594, 198, 320, 22, 10]. Congress [198, 245, 253, 264, 247, 255, 266, 246, 254, 320, 265, 449, 22, 10, 594, 256]. Connaître [286, 311, 352]. connais [459]. Connaissance [446, 13, 106, 800, 442, 488, 486, 485, 492, 491, 548, 716, 792, 53, 79, 1350, 447, 550, 1186, 925, 468, 551, 515, 555, 43, 443, 445, 122]. Connaissances [807]. connection [308]. Connectionism [1359]. Connections [1668, 1675]. Connective [1856, 1222]. Connectives [1856]. conscience [489, 120, 934]. Consciousness [489, 1941, 1803, 2178, 986, 2208, 1200, 1690, 822, 120, 1995, 1952, 1952]. conseil [506, 505, 507, 504]. Conseils [1181, 1232, 1263, 1294]. Consequence [1582, 287, 274]. Consequences [1669, 1300]. Consequentialism [2004]. Conservation [1842, 1736, 1688]. conservative [634]. Consideration [281]. Considérations [82, 382, 376, 82, 683, 382, 683]. considered [352]. considérée [352]. consistance [388]. Consistency [2211, 388, 1615]. Conspiracy [2081]. constituants [239]. Constituents [1882, 239]. Constitution [1974, 1114]. Constitutive [2114, 2208]. Constitutivism [2226, 1766]. Constraint [2065]. Constraint-Based [2065]. Constraints [2042, 1408]. construct [1413]. Construction [1810, 982, 998]. Constructional [1171]. constructionnel [1171]. Constructive [1750, 971]. Constructivism [1808]. Constructivist [2049]. Constructivity [973, 1751]. Constructs [697]. construit [1413]. construit-on [1413]. contemporaine [170, 1287, 8, 1285, 1356]. contemporains [299]. Contemporary [1212, 873, 2136, 1968, 1022, 170, 299, 1287, 8, 1285, 1356]. Content [1974, 1376, 1986, 1666, 1377, 1922, 2200, 1373, 1999, 1556, 562]. Content-Relativism [1999]. Contents [56, 752, 779, 1262, 1293, 1317, 1659, 1700, 2176, 1372, 47, 190, 306, 1319]. Contenu [562]. contenus [1319]. Context [1952, 1641, 331]. contextes [1175]. Contexts [2068, 994, 1175]. Contextualism [1664, 1901, 1524, 2107]. Contextualist [1928]. contingence [1162]. Contingency [1562, 1188, 1162]. Contingentism [1780]. Continuants [2218, 2147]. continuation [534, 771]. continuité [374]. Continuity [1389, 287, 374]. Continuum [2091, 971, 1455, 735]. Contraddizione [1721]. Contradiction [37, 37, 323]. contradictions [579, 579]. Contrast [1872]. Contrastive [1625]. Contre [1618]. contribute [1334]. Contribution [67, 141, 435, 632, 67, 141]. Contributions [148, 146, 142, 148, 146, 142]. Control [1768, 982]. Controverse [1383]. Controversy [1457, 1383]. Convention [1758, 2088, 1017, 1189, 885, 1192, 1187, 1188, 1193, 959]. Convention-Relativism [2088]. Conventional [1188]. Conventionalism [1523]. Conventions [1186, 1186]. Conversation [1038]. conversations [203]. Converse [1935, 1482]. Convex [2175]. Coordinate [1982]. Coordinate-Free [1982]. Coordination [2188, 626]. Copenhagen [1358]. Corazza [1748]. Corbí [1976]. Core [1582, 541]. Corman [128]. Cornucopia [1601]. Correia [1975, 1707, 1738]. Correlationism [2215]. correlations [599]. correspondance [1084]. Correspondence [1644, 1197, 758, 1084, 1847]. Corrigendum [1014]. cosmology [564]. Could [2054, 2085, 1535]. Council [507, 506, 505, 504]. Count [1807]. Counter [1901, 220]. Counter-Closure [1901]. counter-terrorist [220]. Counterexample [2210]. Counterfactual [1665, 1694]. Counterfactuals [1927]. Counterpart [1627]. Counterpossibles [2189, 2237]. Counting [2112, 2106]. cour [503, 500]. Cours [785]. Course [2168, 785]. Court [503, 500]. Courte [793]. Couturat [1084]. Crane [2059]. crassa [1064]. creation [552, 1090, 1090]. Creative [986, 552]. creativity [553]. Credibility [1909, 892]. Crimson [1743]. Crisis [1446]. Crispin [2164]. critères [485, 492, 574]. Criteria [1594, 1342, 485, 492, 574]. Criterion [1469, 948]. Critical [1211, 1210, 1719, 1033, 1363, 1303, 2219, 1056, 308, 369, 517, 1054, 530, 1061]. Criticism [1033, 1325, 269]. Criticisms [1804, 1321, 208]. Critique [1540, 2189, 1029, 117, 1595, 1090]. critiques [624, 691, 725, 742, 777, 783, 788, 798, 811, 823, 827, 847, 855, 878, 889, 907, 960, 984, 1023, 1044, 1049, 1083, 1098, 1105, 1124, 1178, 1230, 1253, 1260, 1291, 1301, 1315, 1323, 1328, 1337, 1343, 1365, 1370, 1421, 1405, 1410, 1432, 1440, 1447, 1453, 1459, 1475, 1494, 1498, 1503, 1508, 1514, 1521, 1532, 1538, 1544, 1552, 1558, 1577, 1584, 1599, 1470, 1093, 1056, 1212, 989, 369, 93, 1137, 1224, 818, 269, 804, 933, 934, 935, 936, 937, 938, 1164, 1199, 1054, 1061]. Cross [848, 2014, 539]. Cross-Categorial [2014]. Cross-Cultural [848]. Cross-Disciplinary [848]. cross-induction [539]. Cross-Professional [848]. Crossed [2086]. Croyance [999, 1355]. Cuénod [548]. Cultural [848, 429]. Culturally [721]. culture [171]. Cultures [171]. Cum [1112]. Cumpa [1847]. Current [1753, 366, 1943, 635, 633]. Curry [236]. Curved [1816]. cybernetics [553, 142]. cybernétique [142]. Cycle [2195]. Cynical [174]. cynique [174].

D [214, 128]. D. [1102]. d’absorption [1059]. d’action [176]. d’Alembert [843]. Dame [2168]. Damming [1933]. Dan [2091]. d’analyse [1056]. dans [490, 506, 610, 505, 1095, 498, 618, 620, 1052, 920, 8, 330, 1186, 16, 1283, 1059, 228, 902, 509, 513, 504, 136, 385, 900, 686, 689, 719, 122, 561]. d’antinomies [569]. Darkness [1254]. Darwin [1851]. d’Asie [902]. data [140]. d’attribuer [156]. David [1810, 1847, 2229, 2147]. Davidson [1994, 1249, 1506, 1512, 1719, 1519, 1168, 1446, 2131, 1207, 1715, 1443, 1092, 1187, 884, 1251]. Davidsonian [1894]. d’axiomatiser [214]. day [137, 33, 21]. De-Fining [2197]. débat [632]. Debate [1457, 632, 1556]. Debates [1943]. Debatte [1747]. Deception [1944]. déchiffrement [561]. decidable [473]. deciphering [561]. Decision [1078, 1128]. Decisions [2141]. Deckert [2223]. déclic [718]. Dedekind [2083, 1325]. Dedekinds [1325]. Dedicated [1582, 1269]. déductifs [533]. Deduction [1025]. deductive [1002, 1413, 1413, 533]. déductive-nomologique [1413]. Deep [117]. Defeating [2074]. Defence [2000, 1047, 2116, 2209, 2132, 2185].

Defending [1500, 1780]. Defense [2067, 1872, 2090, 1909, 2002, 1574, 1251, 452, 452]. Defenses [1868]. Deference [1908]. Deferentialism [1965]. Definite [1273, 1741, 1483]. Definition [1135, 1383, 289, 94, 1383]. définitionnels [1171]. Definitive [1171]. Deflate [1794]. Deflating [1644]. deflation [1618, 1618]. Deflationism [1610, 1894, 1579, 1960]. Deflationist [2067, 2007]. Degrees [1912, 2210, 1914]. d’Einstein [1383, 599]. del [1911, 150, 1910, 1721]. Deliberation [1840]. Deliberations [1693]. della [1790]. dell’assurdo [1721]. Demands [2051]. Demarcation [805]. Demonstration [2046, 1134, 1134]. Demonstrative [1922]. Demonte [2014]. d’engagement [712]. Denis [2020]. Denken [115, 613, 261]. Denkens [333]. Denkweise [244]. dénonciation [709]. d’enseignement [667]. Dentith [2081]. denunciation [709]. Deonna [1993, 2002, 1730, 1725, 1724, 1727, 1729, 1726, 1728, 1731]. deontic [1008]. déontique [1008]. dépassement [719]. Dependence [1694, 1806, 1738]. Dependent [1468, 1548, 2154, 1372]. Depiction [1758]. depths [456]. depuis [21]. Derangement [1519]. Derive [1794]. Desargue [2032]. Descartes [1177, 1353, 1177, 1257, 513, 590, 1026]. Descent [1256, 1524]. Descripteurs [1375]. Description [1554]. Descriptions [1273, 1741, 1483]. Descriptivism [1893]. Descriptivist [1215, 1845]. descriptors [1375]. Designation [1648]. Desire [2226]. Desires [1793]. d’espèce [26, 28]. destination [54]. Destouches [215, 251]. d’état [506, 505, 507, 509, 504]. detection [587]. determinacy [882]. determinates [1593]. Determination [1939, 965, 834, 1062, 1322, 965]. Determine [1332]. Determinism [2207, 820]. Determinist [1693]. Detour [1330]. deutsche [537]. Deutung [1352, 471]. Deux [398, 964, 801, 38, 1153, 963, 768, 1306]. deuxième [320]. devant [170, 503, 500, 501]. Devaux [62]. Develop [985]. development [1520, 6, 374]. Developmental [896, 1773]. développement [374]. devenir [768]. Devitt [1211]. d’évolution [51]. d’existence [330, 1175]. d’experience [221]. Diagnosis [1497]. Diagrammatic [1687]. Dialectic [322, 967, 1295, 441, 922, 66, 6, 4, 766, 36, 122, 76, 16]. Dialectic-the [441]. Dialectica [616, 1734, 1802, 706, 1969, 1754, 20, 2032, 20, 54, 2205, 610, 642, 330, 863, 608]. Dialectical [240, 1296, 1295, 75, 154, 121, 95, 153, 187, 367, 371]. Dialectically [1860]. dialecticienne [240]. Dialectics [61, 60, 70, 384, 17, 64, 77, 45, 67, 8, 79, 78, 1127, 7, 60, 5, 1087]. dialectics^{≫} [74]. Dialectique [16, 71, 60, 70, 384, 61, 922, 580, 64, 66, 77, 367, 45, 67, 8, 79, 4, 154, 121, 78, 1127, 153, 766, 122, 187]. dialectique^{≫} [36]. dialectisation [120]. dialectization [120]. Dialektik [6, 7, 76, 60, 70, 17, 64, 77, 79, 5, 1087]. Dialektik^{≫} [74]. dialektische [765, 371]. dialektischen [75, 95]. Dialetheism [1717]. Dialetheist [1717]. Dialogic [1329]. Dialogical [797, 2037]. dialogische [1329]. Dialogue [1282, 394, 913, 200, 754, 3]. dialoguée [3]. dicti} [1042, 1009]. dicto [1042, 1009, 1042, 1009]. Didactic [807]. Didactique [807]. d’identité [241].

Dido [1566]. Diego [1776]. Dieu [1354]. Diez [2060]. Difference [1979]. Different [721, 973, 579]. differentiation [1064, 1064]. différents [579]. Digital [1957]. Dilemma [1717, 1227, 1298, 2131, 1981, 1625, 2155, 1463]. Dilemmas [1672, 2006]. Dimension [1296]. dimensional [1442, 1787]. Dimensionalism [1653]. Dimensions [2004]. d’important [1285]. d’individu [87]. d’information [791, 790]. Dingle [292]. Dingler [231]. d’inspiration [437, 452, 438]. d’intentionalité [1351]. d’interprétant [1081]. dire [1285]. Direct [1741, 1692, 1954]. directeur [863, 608]. Directions [1817]. director [863, 608]. Dis [2233]. Dis-Unity [2233]. Disagreement [2045, 2035, 1823, 2036, 291]. Disapproval [2180]. Disciplinary [848]. Discipline [985, 363, 617, 1143]. discordance [1053]. Discours [11, 925, 24, 1290, 1122, 284]. Discourse [2088, 1290, 1122, 284]. Discovery [1013]. Discussion [63, 69, 81, 206, 358, 356, 1019, 354, 534, 123, 308, 348, 67, 210, 396, 238, 216]. Discussions [2018, 1042, 1665, 2058, 1991, 959, 1627, 1249, 1992, 1594, 1342, 2069, 1228, 1359, 1902, 1648, 1217, 1666, 1209, 1162, 1163, 2056, 1168, 1161, 1736, 1452, 1086, 1625, 1901, 1102, 1092, 1581, 1949, 1626, 1259, 1574, 1123, 1223, 2057, 1009, 1593, 1369, 1013, 1364, 1948, 1300, 229]. disease [166, 165]. Disjunctive [2082, 2055]. Disjunctivism [1902, 1942, 2082]. Diskussion [123, 348, 63, 69, 81, 216]. Diskussionen [959]. Diskussionsbemerkung [229]. disorder [262]. Dispense [1342]. Displaced [2225]. Disposition [1820]. Disposition-manifestations [1820]. Dispositional [1740, 1479, 1407]. Dispositionalism [1920]. Dispositionalist [2080]. Dispositions [1426, 2079, 577]. Disquotationalism [1534]. Dissonance [2010]. Distal [1375]. distaux [1375]. Distinction [1957, 1604, 1824, 1924, 502, 500, 501, 510, 504]. Distinguishing [1972]. dite [809]. d’ivresse [509]. dixième [330]. dn [394]. Do [2045, 2109, 1427, 1890, 1727, 1726, 1770, 1340, 1645, 1583, 1880, 1688, 2063, 1962, 2028, 1879, 1970, 1413]. d’objectivité [14]. d’objet [1160]. Doctoral [1131]. Doctrine [846, 786, 1321, 513, 1193, 226, 225]. doctrines [226]. Dodging [2164]. Does [1235, 1332, 1939, 1688, 1772, 2044, 2089, 1334, 1284]. Doesn’t [1995, 2167, 1803]. Dogma [842, 1197]. Dogmas [1362, 1473]. Dogmatics [1469]. Dogmatism [2164]. Dogmatists [1970]. Domain [805, 357]. domaine [357]. domestic [577]. Dominance [2211, 1478]. Donald [1719, 1168, 884, 1994]. donné [134]. données [140]. Don’t [1759]. Doohm [2146]. Doomsday [2142]. d’organisation [374]. dosage [1054]. Dosen [2168]. d’Oswald [184]. d’ouverture [15, 200, 257]. Down [1660]. Doxastic [1805, 2079, 1772, 1798]. d’Oxford [1144]. Dr [696]. Dr. [128]. dream [1353]. dreams [1353]. Drei [613]. Drei-Komponenten-Modell [613]. Drieschner [1088, 1088]. Dritte [202]. dritten [203]. Driver [933]. Driving [2044, 509]. droit [503, 502, 572, 505, 497, 498, 499, 500, 567, 577, 501, 571, 566, 508, 568, 510, 581, 504, 576, 579, 570]. drunk [509]. drunkenness [509]. Dualism [1403, 2219, 370]. dualisme [370]. Dualistic [514]. Dualität [425]. dualité [210, 199, 399, 230, 214]. duality [614, 210, 199, 399, 230, 214, 425]. Dubarle [216, 215]. Dugas [232]. Duhem [1492]. Dummett [1299, 1434]. Dummettian [1244]. d’un [1056, 397, 10, 386, 39, 1160, 977]. Duncan [1942]. d’une [548, 1353, 637, 131, 192, 1064, 1062, 1174, 174, 1121, 806, 794]. d’unifier [464]. Duplication [1007]. Dupréel [53, 43]. durch [829, 1329, 475]. d’urgence [506]. Dürr [235]. Dyke [1809]. Dynamic [1481, 178]. dynamical [104]. dynamics [104].

e-doctrine [226]. e-Lehre [226]. E. [1603, 591, 207]. Eaditorial [191]. Eames [810]. Early [1885, 1339, 1434]. Easy [2158, 2123]. eBook [2145]. échanges [55]. Echoes [45, 20]. échos [45, 20]. Eckhart [1302]. Ecological [1507]. econometric [562]. économétriques [562]. écriture [561]. Ecumenical [2130]. Edgar [1950]. Edited [1974, 1966, 2091, 2013, 1810, 2012, 2082, 1847, 1874, 1975, 2015, 1787, 1952, 1943, 1968, 2020, 2014]. Editor [1714, 1706, 1701, 1703, 1705, 1707, 1709, 1711, 1712, 1702, 1710, 1704, 1708]. Editorial [1, 12, 2, 26, 48, 482, 625, 707, 728, 910, 2205, 1578, 1660, 1778, 380, 1553, 362, 609, 84, 177, 1051, 1278, 1571, 1516, 1545, 1621, 1850, 35, 26, 130, 139, 911, 276, 113, 35, 130, 139, 911, 276, 113]. Editors [1730, 1725, 1724, 1727, 1729, 1726, 1264, 1728, 1731]. eds [2144]. Education [669, 172, 174, 169, 633, 585, 630, 701, 681, 637, 632, 671, 168, 686, 583, 172, 174, 816, 611]. Educational [701]. effect [287]. Effectiveness [1100]. efficiency [116]. Effort [464]. Efird [1846]. Egidi [1874]. Egli [888]. Ehrlich [1810]. Eidgenössischen [58]. Eigenschaften [1055]. Eighty [426]. Einegange [1360]. einem [1087]. einfaches [613]. Einfachheitsprinzip [417]. Einfluss [285]. Eingegangene [743, 799, 908, 955, 991, 1024, 1050, 1080, 1099, 1125, 1154, 1179, 1205, 1231, 1261, 1292, 1316, 1338, 1380, 1400, 1423]. Eingelaufene [703, 726, 138, 692, 328]. Einheit [1177, 49, 808]. einige [422]. Einleitendes [258, 259, 260]. Einleitung [1477]. Einstein [233, 1021, 1383, 1248, 599]. Einstellung [1036]. Elaboration [1932]. electrical [374]. électriques [374]. electro [150, 144]. electro-acoustic [150]. electro-mechanical [144]. electroelastic [865]. électroélastique [865]. electromagnetic [865]. électromagnétique [865]. Element [1013]. élémentaire [787]. élémentaires [635]. Elementarteilchen [157]. elementary [635, 787, 157]. Elemente [155]. Elements [1150, 1306, 545, 547, 1306, 155]. elettroacustica [150]. elevated [363]. élevée [363]. Eliminating [773]. elimination [421]. Elizabeth [810]. elle [363, 1143, 1284, 551, 136]. elles [223, 602, 1095, 972]. éme [1090]. Emergence [1972, 978, 1339, 2156]. Emotion [1662, 1726, 1766, 2027, 2096, 1858, 2002]. Emotional [2026, 2087, 2200]. Emotions [1730, 2025, 2024, 1993, 2029, 2031, 1072, 2100, 2028, 1728, 2061]. Empathy [1746]. Empfindung [1649]. emphasis [553]. Empirical [1381, 1115, 338, 475, 1247]. Empiricism [733, 1204, 226, 300, 511, 339, 427, 512]. Empiricist [1451]. empirische [475]. empirischen [338]. empirisme [300, 427]. Empirismus [226]. emploi [382, 1059]. encounter [471]. End [1844, 534, 543, 740, 1304, 1396, 311, 542, 769]. End-Relational [1844]. Ende [1396, 769, 1304]. Endlichen [776]. Enemies [1500]. Energy [1842, 1736, 1688]. engagé [913]. Engaged [913]. Engagement [384, 1827, 388, 386, 387, 385, 712]. engineer [558]. Enigma [1733]. enough [928, 1698]. enquête [651, 642, 632, 641, 677, 630, 671, 679]. enquêtes [671, 700, 630]. enquiry [641, 642]. enseignement [641, 681, 651, 637, 642, 632, 671, 635, 677, 645, 647, 876, 686, 633, 689, 926, 585, 638, 650, 583, 630]. enseignements [626, 700, 659, 679]. enseigner [877]. Entail [1939, 2089]. enters [330]. entgegnungen [45, 20]. Entities [946, 1568, 1541]. Entitlement [1591]. entre [219, 210, 577, 330, 1174, 1084, 579, 874]. entreprise [794]. Entretiens [1144, 485, 4, 59, 257]. entscheidbarer [473]. Entwicklung [1520, 6]. enumeration [294, 294]. envers [526]. Environmental [1996]. enzyme [1064]. enzymes [1063, 1063]. épaisseur [599]. Epicure [1026]. Epikureismus [1218]. Epiphenomenal [2143]. Epiphenomenalism [1444, 929]. Epistemic [2088, 1041, 1069, 1888, 1996, 1792, 1221, 1860, 1510, 1926, 1040, 1524, 2079, 1525, 2017, 2234, 1390, 1412]. Epistemische [1412]. epistemischen [1747]. Epistemization [1457]. Epistemological [997, 1939, 982, 1156, 610, 1367, 812, 297, 876, 796, 1942]. Epistemologie [1225, 154, 153, 121, 268, 1224, 1281, 121, 428, 1283, 1314, 1285, 993, 1355, 806, 122, 209]. Epistemologies [896, 1161]. épistémologique [610, 1367, 796]. épistémologiques [876]. Epistemology [805, 985, 1904, 1238, 1307, 1228, 1359, 2082, 1245, 2072, 2027, 748, 1785, 2087, 1318, 1923, 1841, 2173, 1472, 1458, 2159, 1473, 1591, 2185, 1281, 154, 121, 428, 209, 1283, 1314, 1285, 825, 153, 268, 1409, 993, 1355, 806, 122, 1286, 1224, 1472, 2040]. Epistemology’^{≫} [1013, 121]. Equality [2004, 1213, 964]. Equilibrium [1646]. Equivalence [1213, 2118]. equivalent [149]. Equivalents [2192]. Equivocation [1792]. Erasmus [2001]. Erfahrung [211, 256, 227]. Erfassung [587]. ergo [410]. Eric [2072]. Erkenntnis [313, 49, 79]. Erkenntnislehre [846]. erkenntnistheoretische [297]. Erkenntnistheorie [308, 203, 315, 406]. Erkenntnistheorien [1200]. ‘Erklärung’ [994]. Ernest [274]. erörterung [123, 308]. Eros [1748]. Errata [954, 990]. Erratum [1118]. Error [2007, 1586, 1673, 1295, 2086]. ersten [912]. erweiterte [1953]. Erweiterung [414]. Erwin [1847]. ESAP [1406, 1411, 1417, 1422, 1433, 1437, 1441, 1448, 1454, 1460, 1465, 1471, 1476]. Esbozo [1873]. Escape [456]. Esfeld [1788, 2223]. espace [1356, 516]. espèce [31, 30]. Esprit [301, 1848, 1139]. esprits [919]. Esquisse [977]. Essai [15, 133, 1280, 120, 581, 569]. Essais [2203]. Essay [2022, 1976, 1734, 1483, 1802, 1280, 1969, 701, 120, 581, 133]. Essays [1974, 2013, 1719, 2147]. Essence [1706, 2173, 1845, 2057]. Essential [2209]. Essentialism [2080]. Essentials [1480]. est [1143, 551, 318, 214]. est-elle [1143, 551]. est-il [318, 214]. esthétique [70, 762]. états [1667]. Ethical [1463, 1943]. Ethics [2004, 1430, 875, 2072, 1968, 2070, 1035, 1034]. Ethik [1033, 1035, 1034]. Ethos [1150]. étinceile [988]. Etre [1310, 363]. étude [904, 566, 141, 1595]. études [624, 93, 691, 725, 742, 777, 811, 823, 827, 847, 855, 878, 889, 907, 960, 984, 1023, 1044, 1049, 1083, 1098, 1105, 1124, 1178, 1230, 1253, 1260, 1291, 1301, 1315, 1323, 1328, 1337, 1343, 1365, 1370, 1421, 1405, 1410, 1432, 1440, 1447, 1453, 1459, 1475, 1494, 1498, 1503, 1508, 1514, 1521, 1532, 1538, 1544, 1552, 1558, 1577, 1584, 1599, 1470, 1093, 1212, 989, 1137, 1224, 818, 804, 933, 934, 935, 936, 937, 938, 1164, 1199]. Euclid [930]. European [1515, 1527, 1533, 1539]. Evaluations [1729]. Evaluative [832, 2024, 2100]. Evans [534, 518]. Evanston [233]. Événement [284]. Event [2014, 284]. events [1467]. ever [1797]. Everett [2038]. Everything [1832, 1689]. Evidence [1887, 1963, 364, 364]. Evidentialism [1772]. Evidentialist [1918]. Evolution [966, 163, 933, 385, 1335, 2017, 51, 384, 41, 382, 509, 32, 187, 934, 41, 6, 163, 509, 32, 933, 187, 934]. Evolutionary [1888, 2202, 31, 1358]. evolutionism [299]. évolutionnismes [299]. evolutions [6]. évolutives [31]. Exact [338, 617, 587, 853]. exacte [617]. exakte [853]. exakten [587, 338]. Examination [1210, 1603, 203]. example [1520]. examples [569]. Exchange [1133]. Exchanges [55, 55]. Exclusion [2198, 1626, 24]. exemplarisches [611]. Exemplarist [2171]. Exemplary [611]. exemples [569]. Exist [1807, 2167]. Existence [1030, 2032, 1149, 488, 490, 330, 1085, 1175, 122, 1487]. existential [430, 1738]. Existentials [1563, 1569]. exist^{≫} [410]. Expecting [2018]. Experience [1922, 1679, 2026, 2087, 2200, 2183, 786, 221, 212, 218, 479, 219, 243, 221, 210, 201, 205, 307, 428, 415, 213, 224, 228, 918, 217, 230, 526, 227, 211, 256, 219, 210, 201, 205, 307, 428, 213, 224, 228, 918, 217, 230, 526, 415, 221, 212, 218]. Experiences [1829, 1556, 1319, 1319]. Experiencing [1764]. experiment [5, 469, 204, 157, 204, 338]. Experimental [875, 2135, 1056, 156, 1058, 1063, 876, 876]. experimentale [52, 1056, 1058, 1063]. expérimentales [156]. experimentalwissenschaft [109]. experimentierkunst [5]. Experiments [1589, 1887]. Explain [1803, 1934]. Explainability [1587]. Explaining [1588, 1951]. ’Explanation [994, 2115, 2090, 2077, 1076, 2065, 1215, 2032, 1010, 1000, 1198, 1326, 764, 1413, 1057]. Explanations [851, 1625]. Explanatory [2090, 1673, 1624, 1934, 1958, 1928]. explication [1326, 764, 1413, 1057, 1000]. explicative [167]. Explicit [1643]. exploration [748]. exposé [231]. exposés [62]. Express [2138]. Expressibility [2043]. Expressing [1730, 1542]. expression [1061]. expressions [1095, 514]. Expressivism [1921, 1860]. Expressivist [1992]. Extending [1458]. extension [414]. Extensionality [1973, 2068, 1391]. Extent [1937]. External [1195, 1005]. Externalism [1445, 1376, 1489, 1691, 1550, 1375, 1457]. externalisme [1375]. Externalist [1771]. externalized [1307]. extract [548]. extrait [548]. Extreme [1863].

F [242, 407, 2093, 2011, 432]. F. [341, 1368, 920, 1284, 1620, 935]. Fa [1488].

Fabio [1952]. Fabrice [1993, 1975, 2002, 1738, 1707]. Faces [2002]. Facets [1594]. Fact [1674, 502, 505, 500, 501, 508, 510, 766, 903, 504, 498]. facteurs [443]. factors [443]. Facts [1681, 1630, 1863, 2209, 1862, 497, 1186, 1467, 1312, 503, 499, 1617, 2120]. Faculty [1536]. Fails [1939]. Failure [2051, 1796, 1946, 1121]. Fairness [2145]. fait [503, 502, 505, 497, 498, 499, 500, 501, 1143, 508, 510, 766, 903, 504]. fait-elle [1143]. faith [843]. faits [1186]. Fall [246, 254]. Fallacy [1717, 1505, 961, 1809]. Fallibilism [1620]. false [981]. Falsehood [1565]. Falsification [957]. Falsifications [2049]. Famine [2128]. Farewell [1341]. fasc [771]. fascicles [771]. fat [127]. Fate [2040]. Father [299]. Fattori [1722]. Faultless [2035]. Favour [2231]. Faye [2184]. Fear [1808]. February [1504, 378, 304]. Federal [58]. Feedback [982]. Feedback-Control [982]. Feel [1995]. Feeling [2030, 1731, 1035]. Feels [815]. Fehldeutungen [1169]. Feibleman [440]. Feigl [815]. Feldman [2124]. Felka [2158]. Fellows [2106]. fenomenología [757]. Ferdinand [457, 608, 924, 771, 476, 879, 927, 1285, 923, 1127, 914, 608, 754, 915, 926, 1369, 924, 1279, 771]. fermeture [928, 15]. Fernand [1345, 1346]. few [518, 440, 280]. Feynman [1426]. Ficta [1561, 1783, 1782, 1784, 1781]. Fiction [1560, 2167, 1565, 1543, 1784, 1781, 1569, 460]. Fictional [1945, 1567, 1568, 1570]. Fictionalism [2143, 2167, 2134]. Fictions [1561, 1783, 697]. fidélité [196]. fidelity [196]. Field [1661, 865]. Fields [1815]. Fiengo [1940]. filosofia [1911, 1910, 1873, 1776]. filosofiche [1884]. filter [922]. filtre [922]. fin [534]. final [681, 82, 683]. finales [681, 82, 683]. finalité [133]. fine [127, 1714, 1706, 1701, 1703, 1705, 1707, 1709, 1711, 1712, 1702, 1710, 1704, 919, 1708]. Finean [1706]. finer [1556]. finer-grained [1556]. Fining [2197]. Finite [1750, 1256, 414, 776, 319]. finiten [414]. finitism [1751]. Fink [2231]. Finks [2195]. fins [919]. First [912, 1445, 1108, 1925, 2184, 1786, 1168, 1029, 2227, 1737, 911, 910, 123, 1924]. First- [1029]. First-Order [1924]. First-Order/Second-Order [1924]. Fisette [2020]. Fitch [325]. Fix [1817]. Fixing [1452]. Flam [1196]. Flim [1196]. Flim-Flam [1196]. Flow [986]. Focus [1884, 2015]. Fodor [1464, 1374]. foi [843]. Fold [1603]. Folk [1746]. Follow [1784, 1781]. Following [2133, 1159]. Folse [2184]. fonction [1064, 901]. fonctionnelles [557]. fonctions [185, 141, 140]. fondamentale [399]. fondateur [610, 608]. Fondation [252, 916, 806]. fondement [270, 712]. Fondements [168, 544, 99, 1171]. fonder [617]. Forced [1955]. Forced-March [1955]. Forces [1819, 1818]. forecasting [618]. foreword [1264, 1183, 828]. Form [1276, 2169, 1439, 618]. Formal [838, 1608, 866, 1602, 273, 423, 1601, 890, 1728, 2191, 1611, 412, 238, 300, 240, 239]. Formal-Ontological [1601]. formaler [838, 423]. formalisations [369]. Formalism [371]. formalismes [478]. formalisms [478]. Formalismus [371]. formalist [732, 236]. Formalizations [369]. Formation [1150, 807, 285]. forme [618]. formel [238]. formelle [300, 240]. formels [239]. formes [384]. Forming [1536, 2114]. Forms [1333, 384, 344, 400]. Formula [1482]. Formulation [1426, 1653]. Forrest [1773]. Forschers [19]. forschung [155]. Forschungsmethoden [523]. Fortschritt [536]. fossil [2215]. Fostering [1874]. Foundation [1426, 1606, 1699, 49, 712, 806, 252]. foundational [426]. Foundations [749, 18, 974, 1578, 734, 359, 739, 604, 277, 333, 168]. founder [610, 608]. founding [916]. Four [1653, 1699, 1603, 1442]. Four-Category [1699]. four-dimensional [1442]. Four-Dimensionalism [1653]. Four-Fold [1603]. Fourth [257]. Frage [75, 111, 308, 315, 270, 839]. Fragebogen [639]. Fragebogens [640]. Fragestellungen [942]. fragments [53, 43, 43]. frame [1442]. Frames [2151, 1820]. Framework [1560, 1142]. francais [504]. francaise [22]. France [813]. Francesco [1721]. Francois [1848]. Franz [1789, 2203]. Frascolla [1630, 1631]. Fraser [1705]. Frau [537]. Fréchette [2020]. Free [1518, 1771, 1859, 1982, 2140, 2054, 2185]. freedom [471, 1034, 490, 2185]. Frege [2083, 1229, 1548, 1572, 1638, 2068, 1854, 1402, 2075, 935, 1989, 1322, 1434, 1272, 1321]. Frege-Geach [1989]. Frege-Russell [1272]. Fregean [2192]. fregéen [1322]. frégéennes [813]. Freiheit [1747, 471, 1034]. Freiheits [166]. Freiheitsmodell [1747]. French [865, 785, 44, 35, 26, 9, 47, 55, 51, 60, 65, 130, 139, 170, 190, 206, 215, 198, 245, 253, 264, 252, 242, 237, 279, 299, 283, 306, 320, 341, 407, 442, 476, 495, 488, 487, 486, 493, 485, 494, 484, 489, 492, 490, 491, 534, 548, 594, 626, 641, 681, 700, 911, 223, 791, 503, 988, 403, 3, 651, 22, 652, 653, 502, 70, 384, 1056, 37, 756, 921, 62, 61, 410, 1144, 382, 737, 574, 506, 572, 446, 463, 540, 518, 602, 922, 1134, 1281, 124, 96, 1353, 222, 370]. French [464, 610, 479, 1302, 1347, 505, 1379, 637, 850, 64, 66, 300, 435, 716, 642, 632, 1342, 240, 465, 928, 1117, 219, 1287, 1383, 1367, 355, 1095, 1429, 558, 549, 792, 1374, 904, 1243, 444, 172, 185, 340, 497, 498, 53, 188, 671, 77, 14, 91, 90, 367, 618, 927, 859, 916, 353, 369, 620, 31, 1176, 1052, 965, 372, 507, 45, 388, 527, 529, 184, 221, 560, 598, 1368, 544, 731, 740, 762, 920, 499, 67, 163, 30, 397, 635, 105, 654, 15, 33, 152, 210, 596, 1351, 189, 500, 567, 577]. French [972, 127, 636, 655, 8, 28, 79, 133, 134, 132, 131, 201, 241, 301, 392, 459, 913, 168, 176, 747, 21, 110, 128, 175, 334, 401, 437, 452, 480, 1350, 98, 1371, 844, 13, 4, 38, 10, 82, 59, 92, 50, 119, 120, 154, 121, 145, 135, 192, 199, 194, 200, 205, 257, 231, 286, 270, 311, 307, 316, 312, 330, 337, 363, 396, 447, 428, 440, 550, 601, 617, 615, 683, 712, 714, 744, 784, 826, 863, 858, 415, 107, 1153, 1162, 1186, 1266, 325, 759, 925, 1224, 212, 280, 1064, 78, 106, 16, 213]. French [467, 501, 68, 1283, 1059, 1314, 148, 224, 468, 857, 571, 1008, 573, 708, 228, 1326, 193, 999, 818, 658, 659, 917, 1062, 1285, 722, 901, 677, 628, 715, 156, 660, 662, 1143, 310, 623, 478, 1147, 1348, 963, 1127, 1284, 1174, 1165, 871, 99, 551, 664, 720, 566, 1045, 645, 902, 238, 383, 470, 665, 575, 239, 269, 1139, 317, 390, 515, 1102, 141, 140, 374, 1549, 906, 899, 918, 1290, 730, 807, 508, 568, 509, 171, 294, 293, 647, 513, 590, 555, 1280, 702, 218, 386, 1122, 1777, 1090, 24, 1289, 1354, 32, 42, 43, 197]. French [364, 472, 510, 581, 1058, 153, 268, 427, 914, 862, 438, 766, 608, 612, 754, 753, 780, 804, 876, 932, 933, 934, 935, 936, 1053, 1288, 1404, 679, 668, 667, 174, 454, 19, 39, 125, 1375, 556, 718, 801, 72, 903, 554, 477, 36, 727, 387, 787, 404, 504, 1618, 136, 1054, 1356, 1119, 793, 803, 877, 993, 217, 366, 768, 385, 900, 399, 453, 434, 576, 443, 717, 800, 1084, 1306, 1355, 813, 816, 1322, 195, 276, 736, 2050, 377, 686, 633, 1319, 146, 569, 689, 926, 562, 584, 585, 196, 709, 1065, 638, 1171, 579, 1121]. French [533, 1081, 1160, 1258, 861, 1369, 795, 764, 806, 843, 1282, 719, 761, 919, 516, 1060, 905, 650, 670, 570, 1085, 1175, 357, 445, 122, 1026, 1094, 1193, 1216, 1279, 1487, 1061, 187, 113, 538, 1413, 794, 977, 1063, 142, 230, 583, 630, 561, 874, 1000, 1057, 1103, 557, 796, 809, 599, 771, 87, 101, 284, 352, 772, 790, 1349, 1300, 526, 588, 214]. French-Speaking [33]. Frequencies [1919]. Freud [1314, 1277]. friction [104]. Friederike [2144]. Friends [1500]. from‘Philosophy [704]. Frontiers [2238]. Function [1898, 2097, 1064, 901, 249, 248]. Functional [1750, 1753, 1612, 557]. Functionalism [1924]. functionality [738]. functions [185, 375, 141, 140, 420, 2083]. Fundamental [281, 1797, 178, 399, 80]. Fundamentalism [2193, 2098]. Fundamentality [1802]. Fundamentals [544, 789]. Fundamentis [1112]. Funktionen [420]. Funny [2113]. fuoco [1884]. Fuss [1653]. Future [1335]. Futuristics [812].

G [935, 1775]. G. [1008, 566, 2011]. Gabriel [2083]. Galileo [2234]. gambling [1520]. game [972, 1250]. Games [943, 1393, 1366, 1003, 763]. ganzen [461]. Gap [1934, 1958, 1928, 1409]. Garb [2134]. García [1713, 1839, 1787]. Gary [1943]. Gascoigne [2039]. Gaskin [1836, 1837, 1838]. Gasser [1966]. Gaston [483, 1127]. Gate [1895]. Gate-Based [1895]. Gavazzi [566]. Gaynesford [1737]. GC [1054]. GC/MS [1054]. Geach [1989]. Gedanken [611]. gedankliche [5]. Gefühls [1035]. Gegen [808]. Gegensätzlichkeit [49, 323]. Gegensätzlichkelt [220]. Gegenstände. [1953]. Geiste [1747]. geisteswissenschaftlichen [117]. Geleitwort [912, 2, 912]. gelöst [829]. Gender [1073]. Gene [1061]. genealogie [405]. genealogy [405]. génénetique [154]. General [603, 466, 629, 1391, 785, 507, 210, 573, 156, 293, 809, 507, 628, 363, 628]. générale [785, 210, 156, 293, 809]. généralisation [1326]. Generalités [363]. Generality [2083, 1623]. Generalization [1468, 1964, 1326]. Generalizations [1378]. généraux [573]. Genesis [975]. genetic [121, 268, 994, 154, 121, 153]. génétique [121, 153, 268, 1061]. genetique^{≫} [121]. genetischen [994]. Genevan [1865]. Gentzen [325]. Genuine [1271]. geologist [131]. géologue [131]. Geometric [971, 1812, 145]. géométrie [392, 312, 293, 516, 324, 335, 406]. géométrique [145]. geometrische [971]. Geometry [760, 1687, 392, 293, 324, 335, 406, 516, 312]. Georg [1966, 860]. Gerechtigkeit [1520]. German [2050, 409, 461, 100, 49, 75, 115, 216, 246, 254, 251, 271, 912, 1182, 1183, 1234, 1230, 1399, 164, 1390, 524, 1027, 462, 17, 73, 80, 111, 123, 202, 281, 308, 315, 348, 381, 611, 758, 789, 838, 869, 332, 1336, 1412, 1455, 1055, 829, 116, 1520, 411, 1511, 324, 335, 6, 839, 74, 86, 968, 499, 1189, 1304, 1396, 1486, 346, 413, 395, 313, 998, 414, 830, 1177, 893, 209, 371, 20, 1190, 970, 88, 417, 418, 1033, 537, 405, 94, 161, 265, 5, 368, 429, 523, 587, 613, 302, 285, 126, 808, 828]. German [841, 1035, 1185, 1264, 1352, 117, 469, 1169, 1329, 971, 1325, 1096, 298, 860, 1025, 323, 166, 1402, 244, 275, 295, 27, 536, 95, 262, 471, 763, 58, 541, 84, 204, 297, 278, 365, 420, 1036, 969, 338, 272, 1034, 1303, 1397, 853, 423, 406, 432, 769, 422, 309, 333, 1334, 290, 425, 473, 1200, 1114, 40, 994, 1087, 530, 639, 924, 475, 7, 76, 54, 1088, 776, 250, 282, 640, 162, 227, 207, 203, 229, 226, 220, 211, 225, 942, 261, 735, 157, 947, 949, 155, 151, 260, 256]. Geschichte [295]. Gesetzes [100]. Gesichtspunkte [203]. Gespräche [202]. Gesprächen [203]. Get [2117, 1948]. Getting [2003]. Gibbs [359]. Giberman [2221]. Giorgio [1790, 2170]. GIREP [642]. Giuliano [2015]. Given [2149, 134, 1312, 24]. gleichzeitig [473]. Global [171, 1416, 71]. Glücksspielen [1520]. Glüer [1994]. glycohydrolase [1064, 1064]. Go [843]. goal [376]. goal-structure [376]. Goals [2074]. God [1875, 1354]. Gödel [1751, 1429, 1176, 1751, 1753, 2047, 1551, 1754, 1573, 1352, 2168]. Gödels [1352]. Godless [2072]. Goes [2205]. Gonseth [432, 639, 461, 242, 407, 476, 457, 879, 462, 921, 549, 927, 1368, 920, 1517, 459, 755, 1283, 917, 1285, 923, 1127, 1284, 1290, 914, 608, 754, 915, 432, 926, 196, 1369, 639, 924, 1279, 1286, 771]. Gonsethian [1281, 1289]. gonséthienne [1281, 1289]. Gonseths [461, 924, 462]. Good [1988, 2181, 1903, 1593]. Goodman [1342, 1576, 234, 1171]. Gottfried [1941]. Gottschling [1747]. Govern [2078]. grained [1556]. Grammar [1203, 1300]. Graph [2069, 420]. Graphschemata [420]. Grasp [1855]. Grazer [1158]. Grèce [900]. Greece [900, 1431]. Greek [950, 974]. Green [849]. Grelling [853]. Grellingsche [853]. Grenzen [587]. Grice [1925, 1270]. grosse [127]. grossen [100]. Grössenbegriffs [368]. Grossmann [1847]. ground [455]. Grounding [1975, 2209, 2125]. Group [669, 1732]. groupes [31]. groups [31]. Grue [2055, 849, 1123]. Grue-Green [849]. Grue-Paradox [2055]. grundeigenschaften [423]. Grundfragen [161]. Grundlage [49]. Grundlagen [789, 116, 270, 333, 1950]. Grundlagen-Systeme-Anwendungen [1950]. Grundlagenforschung [282, 281, 280, 250]. Grundlagenprobleme [523]. Grundsätzliches [80]. Guest [1714, 1730, 1706, 1725, 1724, 1727, 1729, 1701, 1703, 1705, 1707, 1709, 1711, 1712, 1726, 1702, 1710, 1704, 1728, 1708, 1731]. Guide [2221]. Guido [1874, 1876]. Guillaume [2020]. Guilt [1731]. Gunderson [1250]. Gustav [1874]. Gustave [299]. Gymnasialfach [611]. Gymnasium [611].

H [242, 2013, 2184, 1008, 445]. H. [231, 18]. Haaparanta [2013]. Habermas [2146]. Haji [1927]. Hallucinations [1884]. Hamburgensis [1761]. Handlung [1200]. Hanson [867]. happening [987]. happy [186]. hardback [2144, 2145, 2147, 2146]. Harrington [2119]. Harris [814, 207]. Hartmann [122]. Haskell [236]. Haukioja [2135]. Haven’t [1082]. Haverkamp [2121]. having [618]. head [143]. heading [1131]. Heads [1612]. healing [166]. health [170]. Heather [1809]. Hegel [6]. Hegelian [51, 32, 42]. hégélien [51, 32, 42]. Heidegger [448, 921, 430]. Heil [1723]. Heilkunde [166].

Heinricii [327]. Heisenberg [808]. Heisenberg’schen [808]. Hell [587]. Hell^{≫} [587]. Help [1803, 2054]. Helped [1977]. Hempel [1010]. Henri [1399, 1546, 1399, 1547]. Herausgebers [1264]. Herausgegeben [233]. Herbert [292]. Herbst [246, 254]. Here [2117, 2071]. Hermann [275]. Hermeneutics [1002, 993]. herméneutique [993]. Hermits [1259]. Herr [1751]. herrn [123, 229]. Hetherington [2101]. heuristic [513]. heuristique [513]. Hidden [1305]. hierarchical [745]. Hierarchy [2068]. High [685]. Higher [1665, 1932, 2152, 180, 1438]. Higher-Level [1665, 1932]. Higher-Order [2152, 1438]. Highlights [797]. Hilbert [1751, 419, 1551]. hime [459]. Hintergrund [1225]. Hintikka [886]. Hints [894]. Hirnforschung [1747, 1747]. histoire [871, 513, 232]. Historical [490, 1082, 1018, 2202, 386, 1641, 491, 210, 108, 916, 94, 902]. Historicism [868, 869, 869]. Histories [1376]. historique [490, 491, 916, 386]. historiques [210, 902]. Historisches [94]. History [1021, 871, 295, 1364, 51, 513]. Hochberg [1977, 1984]. Hochschule [58]. Hoffmann [1949]. Hofmann [295, 295]. Hoist [2208]. Holism [1320, 1297, 1275]. Holmes [1543]. Holon [763]. Holonspiele [819, 763]. Homage [608, 771]. Hommage [863, 608, 19, 771, 610, 826]. homme [442, 444, 447, 754, 919, 357, 352]. homo [586]. Homogeneity [1626]. homologies [555]. Honest [1708, 919]. honnête [919]. Honneth [2093]. Honor [1483]. horizon [1287]. Horwich [1703]. Hossack [1863]. Hourya [2083]. Huemer [2102]. Hugh [2181]. Hughes [1654]. humaine [454, 1777]. Humaines [1224]. humains [1186]. Human [2001, 1746, 1976, 812, 986, 1777, 1089, 376, 1186, 613, 186, 373, 454]. humanisation [337]. humanism [770]. humanisme [451]. Humanist [386]. humaniste [386]. humanization [337]. Humanizing [797]. Humans [2048]. Hume [1397, 1397, 1355, 1214, 2150, 2091, 2040]. Humean [1815, 1796, 1555, 2233]. Humility [2122]. Hundred [41]. Hungarian [1087]. Husserl [1606, 1351, 1114, 1885]. Husserls [1114]. Husson [229]. Huygens [1122]. Hybridity [1854]. Hypothèse [804]. Hypothèses [871, 871]. Hypothesis [1381, 1455, 804, 735]. hypothetico [1002]. hypothetico-deductive [1002].

I. [273]. Iacona [2169, 2060]. Ian [2122]. Iconicity [1597]. Idea [1536, 8, 4, 92, 239, 597, 1034, 790]. Ideal [1837]. idéalisations [596]. idealisierten [758]. Idealism [1717, 1296, 1025, 2215, 1984]. idealizations [596]. idealized [758]. Ideas [1332, 1752, 887, 51, 41, 147, 6, 285, 297, 32, 87, 51]. Idee [84, 1034, 8, 4, 92, 239]. Ideen [297]. Ideengeschichte [51, 41, 6, 32]. idees [51, 41, 6, 32, 87]. Identification [792, 1945, 792]. Identifications [2009]. Identität [1747]. Identities [2075, 2034]. Identity [699, 1966, 1594, 1635, 1213, 2029, 2069, 1947, 1734, 851, 1799, 2095, 1446, 1935, 1979, 2147, 1435, 1958, 852, 1251, 241]. Ideological [475, 530]. ideologiekritische [530]. ideologienlehre [531]. ideologischen [475].

Idiodynamics [182]. Idolatry [1875]. idoneism [928, 79, 1404]. idonéisme [928, 79]. If [1970, 1851, 1794, 1797]. Ignorance [1254, 389, 1940]. Ignorantiam [961]. ihre [116, 313, 853]. ihrem [226]. II [486, 270, 1104, 628, 675, 547, 555, 42, 153]. III [51, 487, 22, 558, 154, 630]. Illative [1040]. Illinois [233]. Illusion [1946, 1747]. Illusions [1564]. Illusory [1678]. Im [942, 461, 115, 327, 1747, 275, 95, 309, 7, 261, 2117]. Image [2132]. Imagery [1747]. Imagery-Debatte [1747]. Images [77, 1115, 404, 77]. Imagination [2163, 2132]. Imaginative [2204]. Immanence [1309, 1462]. Immanuel [1649]. Immediate [400]. Immunity [2086]. Impact [1025, 237]. implication [409, 325, 1123]. Implications [532, 170, 366, 1063, 2092]. Implicit [870]. Implies [1235, 2207]. Implikation [409]. Imply [1940]. Importance [716, 30, 1035, 513, 716]. Impossibility [1755]. Impossible [2189, 2051, 1374]. In- [882]. inadéquat [382]. inadequate [382]. Inaugural [484]. inaugurale [484]. included [2215]. Incoherence [1152, 1295]. Incoherent [1734, 1860]. Incommensurability [1362]. Incompatibilism [1840, 2054]. Incompatibilist [1948]. incomplete [1095]. Incompleteness [2047, 1352]. incomplétes [1095]. Inconceivability [1675]. Inconditionnés [268]. inconsistencies [1340]. Inconstancy [1986]. Incorrigibility [983]. incorruptible [185]. incorruptibles [185]. inculture [748]. independence [735]. Independent [1623]. Indeterminacy [834, 2131, 1297, 1956, 1203, 287]. Indeterminism [2140, 2054]. Index [1659, 1700]. indexation [794]. Indexicality [1748, 1748]. Indexicals [1561, 951]. indexing [794]. Indian [2159]. Indirect [1757, 2068, 1926, 1537]. Indiscernibles [2069, 852]. Indispensable [1474]. individual [87]. Individuation [1251, 1258]. Induction [1646, 105, 1130, 1487, 539, 1487]. Inductive [1267]. inductivism [1011]. inébranlable [196]. inédite [1084]. inégalité [1134]. inequality [1134]. Infallibility [2053, 1949]. Infer [2076]. Inference [688, 1040, 1905, 105, 97]. Inferences [1609]. inferential [1583]. Inferentialism [1904]. Infinite [1256, 2178, 776]. Infinitist [1798]. Infinity [2102]. Influence [2210, 285, 166]. Informal [1000, 1142]. Informatics [800]. Information [2041, 2052, 2062, 2073, 2084, 2094, 2104, 2110, 2160, 2126, 2137, 2148, 2172, 2187, 2196, 564, 614, 2077, 1957, 802, 764, 713, 791, 540, 780, 803, 399, 790, 803, 764]. Information-Theoretic [2077]. Informational [802]. Informatique [794, 791, 801, 793, 800, 796]. Informative [2075]. Informed [845]. informelle [1000]. inganni [2060]. ingénieur [558]. Ingvar [1786]. Inhalt [752, 779, 1293, 1317, 1262]. Innen [1747]. Inner [2178, 1690]. Innocence [1597, 2008, 1598]. Inquiries [700]. Inquiry [1393, 1156, 322, 641, 677, 671, 679, 2204]. insbesondere [203]. Inscrutability [1166]. insertion [902]. inspiration [437, 452, 438]. Instances [1324]. instinct [514]. instinctifs [185]. instinctive [185]. institut [252]. Institute [252, 58]. Institutional [1998]. Institutions [1540]. Instructions [1182, 1234]. Instrumental [2226]. Instrumentalism [1389]. Integral [687, 686, 170]. intégrale [170, 686]. integrated [622].

Integration [902, 1841, 142, 142]. integrative [179, 375]. Intellectual [1020, 146, 2122]. intellectuelles [146]. intelligence [384, 932, 554]. Intelligent [1128]. intelligible [861]. intensional [775, 1175]. Intensionalization [1643]. intensionnels [1175]. Intensions [1962]. Intentality [1114]. Intention [1270]. Intentional [1670, 1109, 1111]. Intentionalität [1511, 1747, 1114]. Intentionalitätskonzepts [1114]. Intentionality [1116, 1730, 1389, 1511, 1725, 1724, 1727, 1729, 1726, 1115, 1407, 1107, 1889, 1114, 1728, 1731, 1351, 1117, 1775]. Intentions [1116, 1895]. Inter [2141]. Inter-World [2141]. Intérêt [794]. Interaction [1952]. Interactions [1716]. Interdefining [1795]. interdisciplinary [539]. intéréet [507]. interest [507, 794]. Interference [976]. intériorité [195]. interiority [195]. intermédiaire [447]. Internal [1419, 1209, 1195]. Internalism [1989, 2097, 1457]. Internalism/Externalism [1457]. International [266, 265, 198, 245, 253, 264, 252, 320, 449, 626, 1156, 10, 576, 570, 245, 253, 264, 252, 247, 255, 266, 246, 254, 916, 176, 265]. Internationale [320, 1155, 916, 176]. internationaler [265, 246, 254, 265]. interne [577]. Interpersonal [1716]. Interpretation [1592, 836, 1076, 1753, 1352, 1754, 884, 1358, 1250, 540, 825, 1306, 578, 540, 1306, 471]. Interpretationen [1088]. Interpretations [1750, 1140, 1088]. Interpretationsschemata [998]. Interpretative [1208]. interpreter [1081]. Interpretive [1215, 998]. Intersubjective [1945]. Intersubjectivity [1729]. Intersubjektivität [1747]. Intertheoretic [2034]. Intervention [215, 1795, 673, 672, 674, 675]. Interventionism [1929]. Interview [1384, 59]. interviews [485, 4, 257, 1144]. Intransitive [1834]. Intrinsic [1817, 2195]. Introducing [1629]. Introduction [1585, 1650, 2105, 1600, 1614, 2023, 1912, 1701, 134, 59, 2122, 1821, 1811, 1977, 1634, 1886, 1676, 1424, 1903, 686, 1132, 1749, 1559, 582, 548, 121, 793, 134, 59, 627, 686, 2170, 1993, 1786, 1994, 2119]. Introductory [260]. Introduzione [1630, 1910, 1632, 1633, 1631]. Introspective [2226]. Intuition [1589, 2176, 1644, 1590, 1591, 410, 859, 145]. Intuitionism [371, 975, 2121]. Intuitionismus [371]. intuitionist [418, 1026]. intuitionistischen [418]. intuitionnistes [1026]. Intuitions [2045, 1887, 1686, 1587]. intuitive [1134]. Invariance [948]. Invention [1017, 618]. Inventories [1612]. inversion [100]. investigation [641, 74, 530]. Investigations [431, 630]. Invisible [1960]. Invitation [83]. Irrationalism [1067]. irrationalité [1147]. Irrationality [1152, 2029, 1146, 1149, 1148, 1147]. irredeemably [1541]. Irreducibility [1428, 1549]. irréductibilité [1549]. Irreversibility [1418]. is/ought [1409]. Isaac [1810]. Isaye [194]. ISBN [2144, 2145, 2147, 2146]. Isolation [1055]. isomorphism [557]. isomorphisme [557]. Israel [611]. Issue [1714, 2041, 2052, 2062, 2073, 2084, 2094, 2104, 2110, 2160, 2126, 2137, 2148, 2172, 2187, 2196, 1730, 2105, 1706, 1725, 1724, 1727, 2023, 1729, 1701, 1703, 1705, 1707, 1709, 1711, 1712, 1726, 1702, 1710, 1691, 1704, 1728, 1708, 1731, 911, 910]. Issues [1073, 1156]. Italian [503, 150]. italienne [503]. Items [1612]. itinéaire [1008]. itinerary [1008]. IV [488, 671]. Ivo [1730, 1725, 1724, 1727, 1729, 1726, 1728, 1731]. ivresse [509]. IX [493].

J [251, 2013, 2184, 126, 2134, 1603, 453]. J. [215, 251, 121, 863, 126, 1103]. J.-Claude [453]. J.-L [251]. J.-L. [215, 251]. J.-M. [863]. Jaakko [886]. Jacksonism [188]. jacksonisme [188]. Jacob [1697]. Jaegwon [1698]. James [440, 2202, 2136, 2119]. January [626, 1504, 379]. janvier [626]. Jauch [863]. Javier [1847]. Jay [2071]. je [79, 459]. Jean [967, 896, 134, 122]. Jean-Paul [967, 122]. Jennifer [2070]. jeu [560, 972]. Jiri [2118]. Joerg [2159]. Johansson [1786]. John [1951, 302, 2103, 1723]. Johnston [1875]. Joint [1592]. Jonathan [1699]. Jos [295]. José [2060]. Josep [1976, 1787]. Journal [326, 336, 546, 607, 693, 724, 705, 711, 304, 296, 314, 342, 351, 378, 379, 402, 408, 433, 450, 481, 458, 520, 521, 1156, 276]. Joy [294]. Joyce [1918]. Judge [1822, 1879, 507, 501]. Judgement [1590]. Judgment [1040, 1989, 1617, 1322]. juge [507, 501]. jugement [1322]. Juhos [313, 313]. Jules [1504]. Julia [1968]. Julien [1993, 2002, 1730, 1725, 1724, 1727, 1729, 1726, 1728, 1731]. Julio [1640]. Julius [1572, 1637]. Jumble [1752]. juridiques [580, 228]. jurisprudence [506, 505, 504]. Jussi [2135]. Justice [2115, 964, 1520, 963, 472, 1193, 1193]. Justification [1888, 1891, 996, 1886, 1798, 2017, 100, 116, 1520, 413, 272].

K. [1176]. Kane [1859]. Kant [1649, 1033, 1789, 2070, 897, 1398, 1027, 1304, 1007, 1295, 1030, 1033, 1029, 1035, 1167, 1036, 1034, 1303, 1397, 331, 1032, 1031, 1026, 1091]. Kantian [308, 315, 309, 203]. kantische [308, 315]. kantischen [327, 309, 203]. Kantism [310]. kantisme [310]. Kants [1035, 1036, 1034, 1303]. Kapsner [2049]. Karen [2157]. Karl [1045, 235, 293]. Karsten [1746]. Katharina [2158]. Kathrin [1994, 1711]. Kausalität [1034, 841]. kein [1304, 347]. Kepler [285, 285]. Kernfragen [541]. Kevin [1706, 1995, 2040, 1714, 1701, 1703, 1705, 1707, 1709, 1711, 1712, 1702, 1710, 1857, 1704, 1708]. Kim [1698]. Kind [2080, 1726]. Kinds [1506, 2165, 1333]. Kistler [2156]. Kit [1714, 1706, 1701, 1703, 1705, 1707, 1709, 1711, 1712, 1702, 1710, 1704, 1708]. Kitcher [1624]. Kitcher-Style [1624]. klassischen [1169]. Klemmt [302]. Know [1048, 2066, 2101, 459]. Knowability [1742, 1733]. Knower [821]. Knower-Known [821]. Knowing [1939, 321, 1862, 352]. Knowledge [1116, 1665, 1445, 1071, 1808, 1890, 1865, 1636, 2217, 1271, 1744, 1951, 1936, 79, 1976, 135, 925, 1377, 2181, 468, 1861, 1523, 1195, 1691, 1246, 2186, 390, 807, 1946, 391, 1620, 1167, 389, 1588, 2085, 2057, 1070, 815, 867, 1525, 1074, 1254, 1339, 49, 442, 488, 486, 485, 492, 491, 548, 308, 315, 716, 792, 53, 313, 1350, 13, 447, 550, 1186, 551, 515, 43, 937, 445, 122, 203, 446, 286, 311, 106, 390, 443, 800, 2039, 810, 2186, 1785, 2101, 2048]. Known [821]. kognitiver [1357]. Komplementarität [84]. Komponenten [613]. Kongress [246, 254, 265]. Konsequenzen [1747]. Konstitution [1114]. Konstruktion [998]. Konstruktive [971]. Kontinuum [971]. Kontinuumhypothese [735]. Kontinuumshypothese [1455]. Konvention [1189]. Korrespondenz [758]. Koskinen [2013]. Koslicki [1711]. Kosta [2168]. Krankheitsbegriffs [166]. Krause [1720]. Kripke [1202, 1854, 1873]. Kristie [2015]. Kritik [117, 1325]. kritische [308, 1303]. kritischen [842]. Kuhn [2202, 1046, 867]. Kulturbedeutung [429]. Künne [1855, 1761, 1762, 1953]. Kurt [303].

L [251, 869, 1134, 2013, 2122, 935]. L. [215, 251]. Lab [1458]. ladenness [1201]. Lage [33, 21, 10]. L’Alternative [818]. Lando [2170]. Langage [761, 403, 1379, 716, 615, 415, 148, 720, 1848, 719, 1745, 1777]. Language [381, 1003, 1843, 1332, 950, 1519, 943, 1308, 1523, 1086, 1387, 1777, 817, 1869, 1200, 1366, 947, 403, 22, 1379, 411, 716, 74, 760, 150, 615, 830, 415, 759, 148, 1174, 720, 1189, 784, 719, 761, 942, 2169, 2135]. Language-Games [943, 1366, 1003]. Languages [838, 959, 721, 1392, 1513]. Langue [784, 22, 759, 1174]. large [100]. Lasonen [1692]. Lasonen-Aarnio [1692]. L’assemblée [33]. late [1491]. Later [1276]. Latin [904]. latino [904]. latino-américain [904]. Lauener [1546, 1399, 1547]. Lausanne [626, 626]. Law [1637, 100, 503, 502, 506, 572, 505, 497, 498, 499, 500, 567, 577, 200, 501, 571, 566, 568, 510, 581, 504, 576, 570]. Laws [1427, 2092, 1395, 1842, 1324, 946, 1449, 2078, 1436, 1591, 579]. lawyer [502]. lead [1284]. Learning [2232, 1971, 553, 427, 816, 514]. L’échec [1121]. lecteurs [824]. lecture [932, 24]. Lectures [684, 1649]. lee [489]. Left [2076]. Legacy [2070]. legal [228]. Lehre [226, 225]. Leibniz [9, 1657, 590]. Leistung [313]. Leistungsfähigkeit [116]. Leniewski [1139]. Léon [39]. Lesaniewski [1142]. Leser [824]. Leśniewski [1140]. L’Espace [813, 312]. l’esprit [1697, 2156]. lessons [679]. Letter [54, 548, 124, 125, 20, 196]. Lettre [124, 196, 548, 125]. lettura [1632, 1633, 1631, 1630]. L’étude [1288]. leur [478, 239]. Leurs [719]. Level [1665, 1932, 1797, 647, 877]. levels [221, 1153, 719, 786]. Lever [1967]. Levi [967]. Levi-Strauss [967]. l‘évolution [384, 382, 385]. Lewis [1796, 1846, 2210, 1493]. lews [719]. Lexical [2108, 1269, 1825]. Liberal [2067, 2199, 2190]. Libertarian [1518]. Libertarianism [1768]. Libertarians [1625]. Libertarismus [1747]. liberté [761, 490]. liberty [166, 761]. Library [233]. Lichte [95]. lichtsinns^{≫} [111]. L’Idonéisme [1404]. lie [709]. Lieber [1751, 1751]. lien [1174]. Life [600, 1276, 1415, 1022, 713, 1333, 737, 376, 1052, 148, 186, 1288, 136, 436]. lifelong [816]. Light [1450, 1321, 1254, 140, 95]. light^{≫} [111]. Like [1995]. Lill [2003]. limitations [1056, 478]. limites [65, 491, 1056, 464, 435, 1058, 794, 485, 492]. Limits [2021, 2071, 65, 491, 464, 435, 587, 1058, 794, 485, 492]. Linda [2171]. linéaire [561]. Linear [1754, 561]. linguaggio [150, 1911, 1910]. Linguistic [1955, 1331, 1330, 1804, 1200, 831, 890, 708, 722, 456]. Linguistics [1843, 1000, 785, 1334, 2144]. linguistique [785, 708, 722, 1000]. link [315, 1174]. L’institution [1745]. L’intentionnalité [1117]. L’Intentionnalité. [1697]. Lisa [2051]. list [495]. Liste [305, 329, 361, 393, 840, 495]. Literatur [624, 691, 725, 742, 777, 783, 788, 798, 811, 823, 827, 847, 855, 878, 889, 907, 960, 984, 1023, 1044, 1049, 1083, 1098, 1105, 1124, 1178, 1230, 1253, 1260, 1291, 1301, 1315, 1323, 1328, 1337, 1343, 1365, 1370, 1405, 1410, 1093, 1212, 989, 93, 1137, 1224, 818, 804, 933, 934, 935, 936, 937, 938, 1164, 1199]. Literaturblatt [251, 263, 341, 457, 480]. literature [719]. littératures [719]. Living [966, 233, 1394, 178, 397]. Livres [692, 703, 726, 743, 799, 908, 955, 991, 1024, 1050, 1080, 1099, 1125, 1154, 1179, 1205, 1231, 1261, 1292, 1316, 1338, 1360, 1380, 1400, 1423, 138, 328, 840]. Local [1133]. Location [1628, 2179, 1605, 2063]. Locical [339]. Locke [302, 302]. Lockie [2185]. locuteur [1165]. loffener [461]. Logic [60, 805, 2144, 61, 1904, 814, 1992, 864, 976, 2169, 1481, 1015, 1393, 2174, 1771, 1079, 2083, 1615, 1257, 1631, 2037, 1754, 1505, 2021, 2113, 1663, 2168, 890, 1775, 1669, 449, 103, 41, 300, 738, 91, 369, 1350, 126, 1143, 1169, 775, 235, 274, 441, 97, 422, 473, 162, 225, 759, 2013, 1849, 273, 2121, 1885]. Logical [1464, 1609, 1630, 1439, 1308, 1484, 1863, 1692, 1591, 1000, 300, 968, 439, 1350, 440, 1139, 1085, 1429, 560, 598, 2169]. Logically [1734, 1483]. logician [510]. Logicians [845]. logicien [510]. Logicism [1198]. Logicisms [2083]. Logico [1632, 1633, 1630, 1631]. Logico-Philosophicus [1632, 1633, 1630, 1631]. Logics [2049]. Logik [126, 473, 225, 369, 60, 162]. Logique [300, 759, 60, 61, 1429, 91, 369, 1350, 1143, 343, 1085, 1000]. logiques [560, 598, 1139]. logischen [968]. Logos [54, 51, 32, 42, 40]. loi [200]. Lonergan [896]. Look [2026]. Looking [1878]. Loose [1490]. Loss [1976]. Lösung [853]. Lottery [1915]. Louis [1084]. Louise [2014]. Love [1773]. Lowe [1699, 1603]. Luck [1996, 1886, 1525]. Ludwig [1088, 1088]. lueur [140]. Luminosity [1916]. L’univers [457]. Lupasco [38]. Lying [1925]. Lynøe [1786].

M [2004, 1810, 1847, 2122, 235, 273, 432]. M. [242, 407, 518, 124, 549, 53, 241, 38, 154, 440, 863, 280, 43, 197, 39, 125, 72, 453, 432, 196]. maître [1302]. MacBride [2224, 1705, 1604]. MacFarlane [2103]. Machame [2012]. machine [357, 352]. Machines [1101]. Maciá [1787]. Mackie [1136]. Macmillan [2145]. Made [2123]. Magic [931, 1708, 389]. Magnetism [685]. Magnitudes [1257]. main [378, 379, 304]. Making [1730, 1128, 768, 2157]. Malcolm [983]. Malebranche [803]. Man [1851, 1177, 442, 444, 1396, 173, 447, 754, 919, 357, 352, 710]. Manifest [1333]. Manifestations [91, 1820, 91]. Manifeste [176]. Manifesto [176]. Manipulation [1518]. Mannoury [19]. Manuel [1787, 1873, 1713]. MANUSCRITO [1157]. Many [1594, 2138]. March [1955]. Marco [2083]. Marconi [1633, 1776]. Marcum [2202, 2136]. Marcus [1747, 2082, 2145, 1483, 1478]. marge [131]. Margherita [2204]. marginality [1286]. Maria [2011]. Mark [1810, 2186, 1407, 1875]. Martin [235]. Martine [1747]. Marx [1151]. Marxism [7]. Marxismus [7]. Mass [2230]. Massimiliano [2144]. master [1302]. Material [2033, 2197, 2070, 1123]. Materialism [2034, 75]. Materialismus [75]. matériel [2156]. Mathematica [1095]. Mathematical [1020, 281, 1814, 1244, 244, 2158, 1641, 2121, 701, 681, 671, 603, 859, 598, 395, 550, 416, 1283, 250, 282, 370, 689]. Mathematics [626, 669, 700, 611, 2229, 346, 1753, 854, 2032, 295, 1885, 969, 2021, 276, 377, 633, 585, 974, 1031, 630, 223, 651, 103, 348, 789, 222, 637, 632, 289, 671, 369, 635, 656, 316, 733, 1266, 634, 224, 659, 1143, 623, 702, 679, 926, 638, 639, 236, 631, 640, 858, 860, 861, 583]. Mathematik [611, 1219, 369, 348, 789, 346, 860, 295, 969]. Mathematikunterricht [639, 640]. Mathématique [702, 626, 681, 651, 370, 671, 859, 1283, 659, 679, 633, 377]. mathématiques [700, 223, 222, 637, 632, 369, 598, 635, 316, 550, 858, 1266, 224, 1143, 623, 689, 926, 585, 638, 861, 583, 630, 276]. Mathématisation [857]. mathematische [395, 244]. mathematischen [281, 280, 250, 282]. Mathematization [857]. matières [47, 190, 306]. Matter [2077, 1674, 2044, 600]. Matthew [2081]. Matthias [1966]. Maturation [183]. maturity [186]. Max [2156, 1267, 1274, 1269, 1265]. Maxim [1925]. Maximilian [1737]. maxims [142]. May [512]. Mayr [2001]. McBride [2186]. McDowell [2176, 1556]. McKenna [1927]. McKinsey [1771]. McMahon [2070]. McNally [2014]. Me [2141, 1653]. Mead [821]. Meaning [2045, 1320, 2105, 1117, 1269, 2138, 2056, 1086, 1270, 1129, 1737, 1203, 1647, 1828, 1321, 1907, 1906, 1300, 55, 91, 429, 1165, 239, 220]. Meaning-Scepticism [1647]. Meaningful [1940, 1311]. meaningless [981]. Means [1394, 444, 352]. Measure [1778, 98]. Measurement [1491, 1522, 429, 1358]. Mécanique [1122, 353, 232]. mécanisme [551, 146]. mecanismes [717]. Mechanical [866, 144]. Mechanics [1426, 1450, 1486, 1122, 24, 614, 353, 86, 604, 89, 1088, 18]. Mechanik [1486, 86]. Mechanism [873, 551, 146]. Mechanisms [717, 148]. médecine [163]. mediante [150]. Medicine [163, 164, 162, 2136, 1786]. medieval [274]. meditation [256]. Medizin [164, 162]. Meeker [2040]. Meeting [939, 484, 33]. Meetings [1106]. Meillassoux [2215]. meilleur [1354]. Meinong [1511]. Meinongs [1511]. Member [863]. Membre [863]. mémoire [146]. Memorandum [23]. Memories [914]. Memory [1764, 1550, 1582, 1269, 146]. Mengenlehre [413, 290]. Menger [293]. Menschen [1177]. menschliche [613]. mensonge [709]. Mental [2143, 1899, 1743, 1622, 2227, 2064, 178, 5, 1549, 404, 1941]. mentale [1747]. mentales [404]. Mention [1938]. Mercier [299]. Merely [1490]. Mereological [1765, 1501]. Mereology [1710, 2221, 2170]. Merricks [1627]. Message [2117]. Messens [429]. messo [1884]. mesure [98]. Meta [2183, 701, 539, 2118]. Meta-Metaphysics [2183, 2118]. meta-science [539]. Meta-theory [701]. Metaethics [1512, 1663]. Metalinguistic [2225]. Metaphor [1269, 1104, 1097, 953]. Metaphors [1039]. Metaphysical [1419, 1847, 1699, 2118, 1845, 492, 1486, 1975]. Metaphysics [64, 1865, 2013, 2229, 1308, 2033, 1861, 2072, 1863, 1819, 1780, 2063, 1980, 1303, 2183, 1775, 455, 345, 1396, 67, 193, 1096, 936, 195, 122, 256, 463, 64, 1348, 2159, 1809, 2118]. Metaphysik [1396, 109, 1788, 256, 64, 1096, 1303, 1720]. Metaphysique [52, 1348, 64, 492, 463, 67, 936, 195, 122, 193]. Métaphysische [1486]. Method [301, 1033, 1079, 698, 1000, 548, 439, 152, 1002, 313, 119, 784, 149, 1279, 284, 440]. Méthode [784, 548, 152, 301, 119, 744, 1279, 1000, 284, 313]. Methodenkomplexität [1033]. Méthodes [1060, 557, 558, 325, 667, 555]. Methodical [536]. methodischen [123, 536, 151]. methodological [123, 780, 530, 874, 151, 312, 906]. Methodologie [251, 353, 369, 369, 927, 154, 121, 363, 601, 153, 1288]. méthodologiques [312, 363, 906, 780, 874]. methodologisch [530]. methodologisch-ideologiekritische [530]. Methodology [805, 251, 449, 927, 353, 757, 154, 121, 363, 601, 744, 755, 153, 1288, 532, 369]. Methods [524, 523, 558, 325, 555, 667, 182, 557, 1060]. Metodología [757]. Michael [1849, 1788, 2102, 1211]. Michel [1745]. Micro [851]. Micro-Explanations [851]. micromycologie [1060]. micromycology [1060]. mictrophysics [597]. middle [639]. Miéville [242, 241]. Might [1643]. Mikko [2061]. Mill [1574]. Miller [2015]. Millian [1543]. Milos [2168]. Mind [1748, 1677, 1743, 1202, 1255, 978, 1388, 786, 1403, 143, 1549]. Mind-Body [1677, 1202]. Minds [1799, 1574, 1101, 919]. mineralogical [151]. minéralogie [152, 27]. mineralogischen [151]. mineralogy [152, 27]. Minimal [1921, 1295, 2174]. Minimalism [1575]. Minimalist [1807]. Minor [1324]. miopie [1884]. mirabilis} [590]. Miracle [1694]. miracles [2115]. Misanthrope [1216]. Miscellaneous [476, 824, 476]. Misconceptions [1169]. Mises [1121]. Misidentification [2086]. Mistakes [849]. Misunderstanding [1573]. misunderstandings [545, 547]. Mitteilung [824]. Mittelschulen [639]. Mixtures [2230]. Mm. [548, 62]. Modal [1586, 1481, 1864, 1505, 2034, 1841, 1987, 2111]. Modalitätenlogik [422]. modalités [1094]. Modalities [1681, 1094]. Modality [1706, 422]. Modals [1926]. mode [715]. Model [1116, 1832, 951, 2031, 1097, 1392, 2140, 1056, 722, 613, 421, 584, 977]. model-completeness [421]. Model-Theoretic [1832, 1392]. modèle [1056, 722, 584, 977]. Modèles [533, 550, 555, 562, 689, 905]. Modell [613]. Modelled [1858]. models [550, 555, 689, 562, 533, 905]. Modern [974, 786, 164, 384, 88, 2236, 708, 702]. moderne [708, 702]. modernen [164, 88]. modernes [384]. Modes [1071, 2024, 1603, 185, 715]. Modest [1768]. Modified [738]. Modularity [1826]. Modulation [2108]. Mögliche [1747]. Molecular [782]. Molino [1103]. Moltmann [2144]. Momentum [1842]. monde [1667, 77, 1280, 487]. mondiale [171]. mondo [1884]. Money [1834]. Money-Pump [1834]. Monism [1444, 1802, 1528, 1466]. Monsieur [19]. Monstrosity [2225]. Montaigne [709, 709]. Moody [274]. Moore [1047, 1005, 1274]. Moorean [2010]. Moral [2018, 1672, 2207, 2127, 1033, 2145, 2027, 2116, 174, 2180, 2006, 2171, 1774, 2051]. morale [174, 230]. Morality [1726, 1032, 2051, 1976]. Moralphilosophie [1033]. morals [230]. Morgan [41]. morpho [128]. morpho-psychologie [128]. morpho-psychology [128]. Morscher [1950]. Mortality [1150]. Most [1752]. moteur [933]. moteurs [145]. motion [104]. Motivation [1727, 192, 183, 192]. motor [145, 142]. motrice [142]. mots [560, 1314]. mouvement [55, 21, 1289]. movement [55, 21, 1289]. Moya [1774]. moyen [352]. moyens [444]. Mr. [123, 229]. MS [1054]. Mugnai [1658]. Müller [2146]. Mulligan [1706, 1714, 1701, 1703, 1705, 1707, 1709, 1711, 1712, 1702, 1710, 1857, 1704, 2050, 1708]. Multi [2037]. Multi-Agent [2037]. Multilocation [1973]. Multiple [1757]. Mumford [2003]. Murdered [1643]. Murderer [1643]. Music [544]. musical [762, 920]. musicale [762]. musique [544]. Muslim [563]. my [1582, 2106]. Mycenaean [561]. mycénienne [561]. Myopias [1884]. Mystery [1810]. mystical [1348]. Mysticism [1302, 903]. Mysticisme [1302]. Mystique [903, 1348]. Myth [2149, 470, 1216]. mythe [470, 1216]. Mythic [1032].

n [175, 401, 480, 1515, 1527, 1533, 1539]. nach [216, 76]. Nachwort [980]. NAD [1064]. naissance [9]. Naive [1239, 1119]. name [1258]. Names [1271, 1564, 1495, 1543, 1945]. Naming [1580]. napoléon [575, 575]. Narrative [2115]. Narrow [1373]. National [171]. nationales [171]. Natur [1027, 1788]. Natural [959, 1506, 2080, 1972, 1666, 2169, 1699, 2032, 986, 978, 1952, 1303, 1333, 947, 1461, 313, 530, 76, 297]. Natural-Kind [2080]. Naturalism [2199, 1387, 1394, 1397, 1943, 1386, 1374, 1185, 736, 2150]. naturalisme [1374, 736]. Naturalismus [1185, 1397]. naturalist [397]. naturaliste [397]. Naturalization [1228, 1512]. Naturalized [1389, 1359, 2040, 1949, 1409, 1472, 1473]. Naturalizing [1398, 1358]. Naturally [741]. Naturalness [2175]. Nature [1867, 1427, 834, 1607, 1324, 1208, 1449, 1797, 1107, 1335, 822, 491, 1027, 737, 289, 372, 133, 416, 27]. Naturphilosophie [1788]. Naturwissenschaft [1303, 88]. Naturwissenschaften [530, 76, 947]. naturwissenschaftliche [313, 297]. naturwissenschaftlicher [285]. Near [1698]. Necessary [1668]. nécessité [65, 490, 1162]. Necessity [490, 1017, 1555, 1588, 1188, 1364, 1579, 1162]. Need [1770, 65]. Needn’t [2142]. needs [180]. negation [325, 325]. Negative [1563, 1569, 122, 122]. Neglected [1756]. Neil [2039]. Nelson [1342, 234, 1171]. nemo [575]. Neo [283, 1710, 1025, 1690, 283]. Neo-Aristotelian [1710]. Neo-Brentanians [1690]. Neo-Positivism [1025]. Neo-Scholastic [283]. néo-scolastique [283]. nervous [148, 375]. Neuchâtel [637, 638]. neuen [1087]. Neuer [1789]. neuere [422, 541]. Neural [1971]. neuronal [147]. neurophysiological [140]. neurophysiologiques [140]. Neuroscience [1553]. Neurospora [1064]. Neutral [1900, 1704]. News [1406, 1411, 1417, 1422, 1433, 1437, 1441, 1448, 1454, 1460, 1465, 1471, 1476, 21]. Newsletter [1533, 1539, 1515, 1527]. Newtonian [766]. newtonienne [766]. nicht [414, 473]. Nick [2121]. Nida [1747]. Niels [1786, 2184]. Nigarjuna [1336]. Nightmares [1313]. Nihilist [2166]. Nikolaï [122, 122]. Nikolaj [1974]. niveau [647, 877]. niveaux [221, 1153, 719]. No [1779, 1948, 1304, 128, 334, 317, 1797]. Nobel [24, 24]. noch [414]. Noema [1109]. Nolfi [826]. Nom [1258]. Nominalism [2175, 1984, 1642, 1669]. nominalistic [1495]. nomological [1413]. nomologique [1413]. Non [1929, 1668, 2035, 2149, 604, 802, 1775, 2150, 37, 1556, 1383, 149, 317, 1119, 1721, 1906]. Non-Alethic [2035]. Non-being [1775]. Non-Classificational [802]. non-conceptual [1556]. Non-conceptualism [2149]. Non-conclusiveness [1906]. Non-Contraddizione [1721]. non-contradiction [37]. non-equivalent [149]. Non-Informational [802]. Non-naturalism [2150]. Non-quantal [604]. Non-Reductive [1929]. non-séparabilité [1383]. non-separability [1383]. non-standard [1119]. Non-Transferability [1668]. Noncognitivist [1430]. Nondoxasticism [1944]. Nonexistent [2038]. Nonrational [1150]. Nonsense [1629]. nonstandard [1120, 1119]. Normal [176, 332, 29]. Normale [176]. Normalen [332]. Normalisation [1366]. Norman [235]. Normative [1898, 2072, 2231, 2097, 2130, 1520, 1185]. normativer [1520]. Normativité [1549]. Normativity [2045, 1908, 2037, 2133, 1549]. norme [241]. Normenlogik. [1950]. normes [1186]. Norms [2088, 1184, 1186]. nos [824]. Noson [2021]. Notation [1576]. Note [283, 1807, 1071, 916, 167, 438, 1306, 1057, 453, 122, 92, 438, 1057]. Notes [2018, 783, 788, 798, 1042, 1665, 2058, 1991, 959, 1627, 1249, 687, 1992, 1594, 1342, 2069, 1228, 1359, 1902, 1648, 1217, 1849, 232, 1666, 1209, 1162, 1163, 2056, 1168, 1161, 722, 1736, 1452, 1086, 1625, 1901, 1102, 1092, 1581, 1949, 932, 1626, 1259, 1574, 1123, 1223, 2057, 1009, 1593, 1369, 1013, 1364, 1948, 1300, 739, 46, 158, 1342, 722, 1102, 1369, 1300]. Notice [449, 1719, 2219]. Notion [1674, 1022, 26, 506, 558, 91, 90, 367, 1351, 28, 747, 29, 88, 2236, 1142, 108, 508, 509, 1081, 1160, 790]. Notions [85, 973, 14, 15, 1738]. Notizen [959, 46, 158]. Notre [2168]. Nottelmann [1974]. Noumena [958]. nouveau [444, 284]. nouvelle [1054]. Nouvelles [1165]. Nuccetelli [1943]. nuclear [466]. number [378, 379, 970, 1325, 304, 319]. Numbers [1636, 2158, 100, 589, 405]. numéro [911]. Nummer [912].

o [226, 1983]. o-doctrine [226]. o-Lehre [226]. O-Roles [1983]. Obituary [1791]. Object [1468, 958, 1372, 1608, 1548, 1029, 1167, 773, 1111, 298, 1160]. Object-dependent [1372]. objectif [16]. Objection [1530, 1989, 1948]. Objective [1110, 13, 16, 195, 13]. objective. [195]. Objectivism [1680]. objectivist [540]. objectiviste [540]. Objectivité [620, 623, 465, 618, 768]. Objectivity [896, 866, 1715, 1823, 1136, 465, 14, 618, 1462, 620, 623, 768]. Objects [1636, 1607, 1670, 1113, 2033, 2151, 1361, 1863, 1247, 1728, 2059, 1998, 552]. objektrelation [298]. objets [1667]. Obligation [2207]. Obligations [2211]. Obliqua [1596]. O’Brien [2091]. Observation [618, 1166, 1305, 867, 411, 618, 469]. Observations [338, 1201]. Occasion [1997, 41]. Occasion-Sensitivity [1997]. occidentale [385]. occidentaux [905]. Occurrent [2218]. Occurring [741]. œuvre [610, 915, 935]. œuvres [803]. Off [1997, 1771]. Off-Target [1997]. offener [1185]. of‘Mental [1115].

Old [1646, 186]. Older [1457]. Omissions [1991, 2128]. omniscience [1350]. One [1394]. Only [1851, 1386, 424]. Onofri [2153]. ontogenése [454]. ontogenesis [454]. Ontological [1972, 1408, 1030, 1874, 1602, 1977, 1802, 1438, 1601, 1723, 1442, 239]. Ontologie [1953, 949, 341, 1143, 1085]. ontologies [741]. ontologique [239]. Ontology [2080, 1804, 973, 1425, 2151, 1699, 946, 1868, 1142, 1603, 1435, 2201, 1140, 1661, 1141, 341, 1143, 1085, 949, 2123]. Opacity [1388, 1341, 1490]. Open [283, 2205, 192, 1954, 1612, 377, 80, 757, 762, 396, 615, 1185, 1284, 539, 1288, 806, 124, 757, 1368, 468, 843]. Opening [257, 928, 15, 200, 918, 932]. openness [479, 415, 307, 428]. Operational [805, 864]. Operationism [226]. Operationismus [226]. Opposition [1087, 49]. Oppy [2213]. Optics [1122]. Optimum [556, 556]. Optique [1122]. Oratio [1596]. Order [965, 2152, 1770, 1438, 1924]. Order/Second [1924]. Ordered [887]. Ordering [1983]. ordinaire [403]. Ordinary [1523, 403]. ordination [626]. ordnung [262]. ordre [965]. O’Regan [1995]. organelles [1055]. Organicism [873]. organisation [145, 140]. organism [622]. Organisms [966, 178]. Organization [1329, 145, 140, 374, 344]. organon [756]. Origin [1364, 454]. origine [454]. origines [45, 813]. Origins [1586, 1377, 2011, 45, 813]. Oskar [295]. Österreichischen [2020]. Oswald [184]. Otero [1873]. Other [1799, 1574]. Ought [2207, 1844, 1299, 1191, 1192, 1409]. Oui [317]. Our [2043, 1855, 1332, 1636, 49, 824, 986]. Outer [2021]. Outlawed [1449]. outline [745, 977]. outlines [236]. Outright [1912, 1913]. ouverte [283, 124, 1368, 762, 192, 615, 750, 468, 1284, 1288, 377, 806, 843, 80]. ouverture [479, 928, 307, 428, 415, 918, 932]. ouvrage [121, 312, 39]. ouvrages [38]. ouvrir [396]. Overcoming [1308, 475, 719, 1308]. Overdetermination [1526]. Overview [157, 1139]. Own [2208]. Oxford [2144, 2147, 1144]. Oxford/New [2147].

P [215, 194, 1064, 233, 1620]. Package [1752]. Paderewski [1852]. Padrot [826]. Paglieri [1952]. Pains [1924]. Palgrave [2145]. Panpsychism [1803, 2114, 1581, 929]. Panpsychists [1899, 2208]. Panza [2083]. Paolo [1911, 1910, 1791, 1903, 1906]. pape [11]. Paper [1151, 884, 696, 886]. Papers [1163, 2015]. Papineau [1695]. parable [553]. Paradies [1219]. Paradigm [832, 1439]. Paradigmatology [848]. paradigmes [1122]. paradigms [1122]. Paradox [1915, 2055, 395, 2215, 1075, 979, 1383, 853, 599]. paradoxal [279]. Paradoxe [1390, 1383, 599]. Paradoxes [1390, 699, 1412, 1878]. paradoxical [279]. Paradoxie [853]. Paradoxien [1412]. Paraphrase [1039]. Paraphrasing [1039]. parler [36]. Parsimony [2228]. Part [1390, 1412, 165, 1143]. Partial [1734]. Partially [887]. participant [681]. participants [495, 681]. particle [597, 787]. Particles [157, 603]. Particular [2151, 1604, 572, 203]. Particularized [1981]. Particulars [1882, 1456, 1603, 1442]. particule [787]. particulier [572]. partie [1143].

Parties [1306]. Parts [1306, 1655, 1987]. Pasadena [1964]. Passive [1793]. Past [1764, 1948]. Pathology [2134]. Patterns [2065, 938]. Pauen [1747]. Paul [1808, 968, 967, 940, 968, 1703, 413, 1857, 122]. Pauli [285, 285]. pebbles [1281]. péché [66]. pedagogical [170]. Pédagogie [377, 1282, 988, 730]. pédagogiques [170]. Pedagogy [1282, 988, 730]. Peels [1735]. Peer [2036]. Peg [2076]. Peirce [1283, 881, 1258]. Peircean [1081, 1160]. peircéenne [1081, 1160]. Penetration [1971, 1969, 1970]. Pensée [240, 222, 1379, 920, 68, 238, 1289, 385, 1319, 719, 1349]. pentecôtisme [904]. Pentecostalism [904]. Perceived [1799]. Perception [1930, 2232, 1662, 1727, 1726, 2235, 2027, 2026, 427, 982, 1858, 1969, 1756, 804, 427, 2082, 2012]. perceptions [141]. Perceptual [1757, 2232, 1971, 1681, 983, 2031]. Perceptualism [2023]. Père [299]. Perelman [123]. Perelmans [123]. perennis} [461, 461]. Pérez [1873]. Perfect [1681]. Performative [1495]. Performative-nominalistic [1495]. Performatives [891]. perhaps [317]. Perils [2164]. periodicals [794]. périodiques [794]. permanente [816]. Permeation [2232]. Permissible [2128]. Permission [1985]. Perry [1530]. Persist [2063]. Persistence [1973, 2063, 1998]. Person [1445, 1108, 1168, 1195, 1737, 1747]. Personal [1547, 2095, 1150, 2016, 1546, 900, 1966, 2019]. Personalität [1747]. personality [148, 149, 180, 182, 187]. Personalized [2206]. personnalité [148, 187]. personnelle [900]. Persons [1480]. Perspectival [1468, 1378]. Perspective [2206, 1662, 966, 2049, 1638, 1283, 399, 1968]. Perspectives [1167, 2161, 2184]. Perspectivity [2235]. pertinence [1287]. pessimism [132]. pessimisme [132]. Petard [2208]. Peter [2012, 2122, 1773]. Peters [1968]. Peterson [2004]. Peut [36, 442, 363, 317]. peut-être [317]. peut-elle [363]. peut-il [442]. Peut-on [36]. Ph [62]. Phædo [400]. Phänomen [365]. phase [543, 542]. phenomena [1408, 365]. Phenomenal [1695, 2053, 1829, 1878, 2026, 2114, 1962]. Phenomenological [1241, 1847]. phenomenologicus [586]. Phénoménologie [479]. Phenomenologigal [694]. Phenomenology [870, 2177, 1115, 757, 479]. phenomenon [108, 1286]. Philippe [457]. philosophe [610]. philosophenkongresses [256]. Philosopher [256, 610, 1509, 233]. philosopher-scientist [233]. Philosophers [2018, 233, 1127, 1026]. philosophes [1127, 1026]. philosophia [461]. Philosophic [708, 762, 615, 1284, 793, 806, 18]. Philosophica [1156]. Philosophical [9, 115, 985, 2162, 1887, 1884, 2170, 2092, 53, 1993, 511, 2204, 841, 2134, 1196, 1149, 43, 19, 1148, 1277, 1812, 833, 2119, 461, 487, 1156, 322, 1353, 829, 850, 1578, 21, 291, 478, 1122, 1369, 795, 1193, 794, 164, 53, 392, 284]. Philosophicus [1632, 1633, 1630, 1631]. Philosophie [246, 254, 320, 403, 80, 381, 916, 1368, 1747, 468, 265, 1329, 275, 1218, 969, 769, 1334, 1114, 198, 245, 253, 264, 252, 283, 3, 22, 462, 123, 756, 435, 340, 163, 33, 8, 110, 437, 452, 844, 10, 50, 192, 316, 750, 107, 383, 438, 2050, 843, 1697, 874, 1000, 588, 1185, 95, 924, 302, 2020, 1848]. philosophie^{≫} [461]. Philosophies [721, 107, 123]. philosophies^{≫} [123]. philosophique [9, 487, 1353, 53, 33, 392, 21, 10, 478, 1122, 43, 19, 1193, 284]. philosophiques [850, 1369, 795, 794]. Philosophische [1158, 829, 33, 21, 10, 164]. philosophischen [115, 841]. Philosophisches [53, 43, 9]. Philosophy [137, 198, 245, 253, 264, 247, 255, 266, 246, 254, 283, 320, 326, 336, 521, 546, 607, 693, 724, 705, 711, 751, 939, 1515, 1527, 1533, 1539, 1714, 2144, 3, 1491, 462, 2022, 381, 756, 1212, 2081, 1210, 710, 1706, 340, 916, 1368, 1892, 163, 2184, 694, 33, 1701, 1703, 1705, 1707, 1709, 1711, 2082, 1553, 1712, 745, 21, 1786, 844, 10, 192, 107, 950, 468, 1033, 1702, 2122, 2135, 265, 1710, 854, 1079, 1811, 1329, 1629, 1704, 539, 383, 1363, 2202, 186, 304, 1885, 969, 1397, 817, 2136, 1478, 1114, 786]. Philosophy [1431, 924, 475, 1708, 1669, 1000, 588, 461, 296, 314, 342, 351, 378, 402, 449, 458, 520, 593, 22, 80, 770, 435, 1461, 762, 8, 110, 437, 452, 50, 316, 708, 302, 1284, 441, 431, 95, 438, 1571, 769, 430, 2050, 1334, 987, 456, 436, 806, 843, 1434, 236, 512, 874, 252, 778, 403, 33, 21, 1185, 275, 2011, 379, 408, 433, 450, 481, 535, 606]. Philosphical [1590]. Pholosophy [345]. Physical [367, 1425, 1814, 946, 1241, 2198, 1581, 746, 1383, 1280, 365, 532, 1487]. Physicalism [1929, 1896, 1972, 1736, 1688, 1581, 1698]. Physicalists [2099]. Physics [626, 700, 595, 621, 1243, 2184, 820, 684, 2063, 1797, 947, 251, 641, 651, 850, 642, 1287, 14, 603, 620, 1052, 598, 596, 656, 659, 677, 644, 645, 647, 218, 278, 679, 667, 1356, 769, 689, 648, 649, 650, 214, 24]. Physik [251, 769, 947, 278]. Physikalische [365]. physiological [1063]. physiologiques [1063]. physique [626, 641, 700, 651, 850, 642, 1287, 1383, 1243, 14, 367, 620, 1052, 598, 596, 659, 677, 645, 647, 1280, 218, 24, 679, 667, 1356, 689, 650, 1487, 214]. phytobiologie [1056]. phytobiology [1056]. Piaget [896, 154, 121, 126]. Pickering [1557]. Pictorial [1019]. Pictures [1630, 2022]. piece [460]. Pierre [1697]. Piety [1972]. Piguet [453]. Pills [1924]. Pitoëff [134]. Place [1215, 2033, 1107, 1628]. Plan [374, 374]. plastic [16]. Plasticity [1971]. plastiques [16]. platinische [346]. Plato [6, 211]. Platon [6, 211]. platonic [1455, 348, 346]. platonische [348]. Platonism [1244]. platonismus [349]. platonistisches [1455]. platonisums [347]. plausibility [1120]. plausible [97]. play [560]. Plourde [1877]. Pluralism [1961, 1680]. Plurality [2144, 1149, 1148]. Plurals [1932]. plus [1300]. Podolsky [1383]. Poetic [2115, 1353, 560, 13]. poétique [1353, 560, 13]. Point [1876, 1723, 502, 1242, 510, 732]. Points [1903]. Polemics [845]. Political [252, 2145, 1156, 74, 78]. Politics [1146, 1147]. politique [252, 78, 1147]. politischen [74]. politischer [1225]. Polity [2146]. Polysemy [2107]. Popper [1012, 1046, 1045, 1096, 1011, 1458, 1013]. Poppers [1096]. population [1174]. portée [562]. portrait [9, 53, 43, 9, 53, 43]. portraits [9, 53, 43]. pose [497]. posent [497]. position [461, 348, 414, 386]. Positions [1983]. Positive [2114, 965, 1324]. Positivism [1025, 808]. Positivismus [808]. positivistische [531]. possibilité [156]. Possibility [1740, 1937, 2095, 455, 156]. Possible [1985, 1493, 1312, 1122, 36, 214]. possibles [1122]. posteriori [1555]. Postmoderne [1219]. Poston [2090]. Potter [1849]. Pour [540, 396, 1314, 730, 932, 9, 574, 558, 444, 53, 1165, 871, 43, 796]. pouvoir [464]. Power [1673, 2046, 1960, 464]. Powerful [1882]. Powers [1931, 1645, 2003]. pp [2144, 2145, 2147, 2146]. Practical [1730, 1725, 1724, 1727, 1729, 1744, 1726, 1728, 1731, 962]. Practicalist [2101]. practice [512, 515]. Prädikatenvariablen [970]. Pragmatic [1560, 1198, 1557]. Pragmatics [1915, 951, 775, 1174, 1230]. Pragmatik [1357, 1230]. pragmatique [1174]. Pragmatism [1619, 1283]. pragmatisme [1283]. Pragmatist [2070]. Pratique [515]. preach [512]. préalable [630]. Précis [1835, 1861, 1760, 1781]. Precise [1541, 1689]. Predicate [1856, 970]. Predicates [1947, 1391, 2016]. Predication [1030, 1980, 1452]. Preface [457, 753, 753]. Preference [1268]. Preferences [1834]. Preissmann [124]. prejudices [905]. préjugés [905]. préliminaire [206]. préliminaires [270]. Preliminary [97, 630, 270, 206, 641, 630]. premier [911]. premiere [256]. premieres [107, 123]. préparatoires [630]. Prepositional [1565]. Prescriptivity [1805]. Present [2015, 2183]. Présentation [1371, 231, 1371]. presentations [62]. Presentism [2183, 2015]. président [59]. Press [2144, 2147, 2146]. présupposés [1085]. Pretense [2134]. preuve [364]. prévision [618]. Priest [1775]. Príhonský [1789]. Primary [835, 1962, 107]. primitif [382]. Primitive [1959, 2133, 382]. Primitiveness [2118]. Primordial [1070, 1071]. principe [65, 210, 241, 199, 238, 1354, 214]. Principes [964, 174, 573, 963, 550]. Principia [1095]. Principio [1721]. Principle [1846, 1767, 2077, 417, 1484, 1329, 1794, 1173, 979, 2150, 65, 210, 981, 241, 199, 238, 825, 1354, 89, 214, 24]. Principles [1041, 1842, 373, 178, 755, 573, 963, 550, 174]. Prinzip [1329]. Prior [2036]. Priorean [1706]. Priori [1028, 1489, 846, 1691, 307, 1523, 307]. Priorities [701]. Priority [1530, 1519, 1802]. Pritchard [1942]. Privacy [1967]. private [570]. privé [570]. privilege [193, 193]. Privileged [1691]. prix [24]. Prize [24, 1734, 2032, 1802]. Probabilism [1918]. probabiliste [620]. probabilistic [620]. probabilitaire [102]. Probabilité [96, 98]. probabilités [279, 602, 270, 601, 99, 390, 136]. Probabilities [1686, 1770, 99, 2036, 602, 270, 601, 390, 541, 136]. Probability [271, 103, 688, 976, 417, 892, 279, 98, 94, 390, 391, 541, 389, 272, 1088, 96, 95, 278, 261]. probable [468, 101]. Problem [1028, 1646, 1933, 2141, 73, 1677, 1110, 1117, 2012, 2080, 844, 1007, 1012, 1572, 1718, 956, 1202, 2192, 841, 1196, 1640, 1115, 2178, 1623, 275, 2209, 1881, 1524, 2079, 1945, 833, 831, 883, 2191, 1970, 1906, 170, 1522, 1592, 731, 740, 734, 201, 1350, 312, 615, 415, 937, 454, 146, 843, 516, 475, 250, 282, 1641]. Problematik [841]. Probleme [829, 170, 201, 454, 1117, 731, 740, 1350, 844, 312, 615, 415, 146, 843, 516]. Problèmes [633, 497, 1059, 148, 876, 142, 1697]. Problems [281, 1991, 1489, 1130, 523, 633, 815, 829, 497, 1059, 148, 1201, 876, 142]. Problemstellung [281, 280, 250, 282]. procedure [498, 498]. Proceedings [320, 626]. Process [797, 1898, 698, 812, 248, 2140, 2231, 179, 249, 430, 772]. Process-Requirements [1898, 2231]. Processes [1623, 552]. processing [795]. processus [772]. professeur [608, 1346, 771]. professeurs [276]. Professional [848]. Professor [1151, 608, 771]. Program [386, 1367, 1751]. Programme [1551, 1339, 1367, 419, 386]. Programming [1226]. progrès [1054]. Progress [536, 1054]. Projection [298]. projective [182]. Prolegomena [1238]. Promising [2227]. prononcé [24]. Proof [1226, 1252, 1005, 1075, 364]. Proof-Theoretic [1226]. proofs [1120]. Proper [1589, 1033, 894, 1495, 1483, 1258]. Properties [2007, 1817, 1622, 1931, 2032, 1922, 1657, 1581, 2209, 1980, 1924, 1055, 600, 1443]. Property [1372, 1734, 1495]. Property-theoretical [1495]. Proportionality [1991]. Propos [544, 681, 988, 62, 572, 637, 909, 731, 740, 38, 10, 121, 566, 1102, 496, 565, 856, 872, 39, 72, 1344, 1126, 585, 445, 144, 809]. proposals [681]. proposed [1088]. Proposition [1835, 1839, 1195, 1838, 1836, 1321]. Propositional [2200, 1881, 1169, 235]. Propositions [2010, 1548, 1324, 2138, 1863, 1838, 883, 1617, 681, 2111, 681]. propre [1258]. propriam [575]. propriétés [1667]. Prospective [744, 222, 222, 744]. prospectives [816]. Prospects [885, 1335]. protein [1062]. protéine [1062]. protestant [917, 917]. Protophysics [893]. Protophysik [893]. Providential [1886]. provisional [372]. provisions [577]. provisoire [372]. Prozesse [1357]. pseudo [572]. pseudo-antinomies [572]. Pseudoscience [1196]. psychanalyse [871]. psychanalytique [1314]. psyche [189]. Psychiatrie [332]. psychiatry [332]. psychic [140, 1280]. psychique [1280]. psychiques [140]. psychisme [189]. psycho [1374]. psycho-semantics [1374]. Psychoanalysis [871]. psychoanalytic [1314]. psychobiologie [434]. Psychobiology [434]. Psychodynamics [118]. Psychological [1468, 1378, 2011, 2089, 175, 181]. Psychologie [116, 1355, 128, 119, 932, 298]. psychologique [175]. psychologists [142]. Psychology [1398, 1746, 1843, 117, 1173, 116, 345, 128, 119, 589, 298, 932, 1355]. Psychopaths [2009]. Psychophysical [1217]. psychophysiological [716]. psychophysiologie [404]. psychophysiologique [716]. psychophysiology [404]. Psychosemantics [1681]. psychosémantique [1374]. psychosomatic [165]. Psychotherapie [161]. psychotherapy [161]. public [571, 576]. Publications [799, 476, 1399]. Publikationen [1399]. puissance [1310]. Pump [1834]. Pumps [1589]. Pun [548]. Punny [2113]. Pure [1773, 344, 223]. Purely [175]. purement [175]. purpose [444, 133]. Pursuing [1972]. Putnam [1249, 1513, 1247]. Puzzle [2058, 2228, 1938, 1682, 2075, 1964, 1779, 1490]. Puzzles [890].

Quackery [1196]. Qualia [2143]. Qualität [587]. qualitative [524]. Qualities [1931]. Quality [1037, 1925, 587]. Quantal [820, 604]. Quanten [86]. Quanten-mechanik [86]. Quantenmechanik [1088]. Quantenwelt [808]. Quantification [837, 1979, 2201, 1438, 886, 1175, 1569, 949]. Quantified [1481, 1563]. Quantifiers [837, 887, 421, 886]. Quantifikation [949]. Quantifying [1832]. quantique [1383, 598, 596, 599]. quantitative [524]. Quantum [1426, 1450, 2077, 866, 976, 287, 894, 18, 24, 1522, 614, 1383, 598, 1408, 413, 604, 808, 825, 89, 290, 1358, 1088, 599, 86, 596, 1088]. quantum-measurement [1358]. Quasi [2115, 1748]. Quasi-Indexicality [1748]. Quasi-miracles [2115]. Quassim [2048]. Quatrièmes [257]. Que [1499, 548, 79, 459, 1326, 1285]. quel [602, 551]. Quelques [572, 1367, 355, 90, 372, 120, 440, 280, 780, 569, 230, 874, 1056, 518, 850, 660, 902, 269, 1122, 876, 1054]. Qu’est [787, 710, 1220, 1702]. Qu’est-ce [787]. Question [2182, 997, 888, 1954, 942, 75, 111, 308, 315, 921, 839, 119, 1175, 809]. Questionnaire [647, 648, 583, 651, 653, 671, 659, 677, 661, 644, 645, 679, 667, 639, 650, 640, 647, 631]. Questions [962, 1940, 161, 660]. qu’ils [497]. Quine [1229, 1430, 1037, 1210, 1461, 1517, 1512, 1509, 1674, 1227, 1298, 881, 2131, 1452, 1492, 1166, 1416, 1463, 868, 949, 869]. Quinean [1529]. Quine^{≫} [869]. Quotation [1869]. qu’une [787].

R [1746, 72, 2071]. R. [215, 194]. rêves [1353]. rôle [561]. Rabbits [1172]. Radical [1592, 836, 2040, 1250, 1339, 884]. radioisotopes [1059]. Raffaele [2002]. Raftopoulos [2012]. Raison [928, 384, 388, 386, 387, 385]. raisonnement [102, 410]. Ralf [2005]. Ramsden [810]. Ramsey [1616, 1620, 1619, 1617]. rang [363]. Rani [2003]. Ranking [2092, 569]. rapport [219, 671, 677, 679, 628]. rapports [210, 719, 874, 671]. Rasmussen [2019]. Rational [1068, 1078, 1077, 1827, 1727, 1909, 996, 1092, 2085, 2079]. Rationalism [1864, 386, 339]. rationalisme [386]. Rationalismus [842]. rationalist [1090]. rationaliste [1090]. rationalité [1302, 1153, 1147]. Rationality [1730, 1725, 1724, 1727, 1729, 1076, 2152, 1726, 1073, 956, 2235, 1092, 1732, 1070, 1728, 1074, 1731, 1302, 1153, 1724, 1147, 1071]. rationnel [318]. rattachements [185]. Raum [225]. Ravens [1123]. Raw [815]. Rawls [964, 963, 1193]. Raymond [869, 869]. Re [1502, 169, 1112, 1801]. re-statement [169]. reach [1461, 1938]. reactions [466]. read [1571]. readers [824]. Reading [1430, 1854, 932]. Real [1696, 1420, 589, 1111, 101]. Realism [1239, 897, 1211, 1235, 1757, 1320, 1506, 1248, 2012, 1381, 1419, 1209, 1847, 1242, 1255, 1244, 1557, 1236, 1382, 1742, 1237, 1171, 2194, 1170, 1240, 1176, 2072, 2158]. Réalisme [1176, 1171]. Realismus [1747]. Realist [1238, 1678]. réaliste [1284]. Realistic [1245, 1284]. Realität [860, 365]. réalité [1287, 1383, 618, 620, 858, 857, 623, 470, 768, 1063, 1243]. Reality [944, 1220, 950, 857, 1702, 1241, 860, 323, 945, 982, 821, 942, 947, 948, 619, 595, 621, 1287, 1383, 618, 324, 335, 620, 86, 858, 623, 470, 365, 768, 1063, 1243, 1810, 1975]. Really [1990]. Realm [2204]. Reason [1145, 384, 388, 386, 2021, 387, 385, 928, 2090]. Reasonable [1006]. Reasoning [962, 1069, 1744, 2065, 1732, 410]. Reasons [1800, 1844, 2097, 2130, 1853]. Reboul [1777]. Recalcitrance [2031]. Recalcitrant [2029]. Recalled [1978]. Recanati [1848]. recasting [894]. Receieved [159]. Received [112, 138, 267, 328, 360, 522, 592, 605, 692, 703, 726, 908, 955, 991, 1024, 1050, 1080, 1099, 1125, 1154, 1179, 1205, 1231, 1261, 1292, 1316, 1338, 1360, 1380, 1400, 1423, 743]. receiving [24]. réception [24]. recherche [366, 1065, 919]. Recherches [312]. Rechtfertigung [1520]. reciprocal [166]. recognition [1322]. Recollective [1764]. Recombination [1846, 1767]. Reconciliation [873]. reconnaissance [578]. Reconstruction [1389]. Reconstructions [2077]. recours [573]. Recovered [1598]. Recta [1596]. Rector [58]. Recursion [774]. Recursion-theoretic [774]. recursive [420]. recus [138, 328, 692, 726, 743, 799, 840, 908, 955, 991, 1024, 1050, 1080, 1099, 1125, 1154, 1179, 1205, 1231, 1261, 1292, 1316, 1338, 1360, 1380, 1400, 1423, 703]. Red [1743, 1995]. redefinition [1556]. Redeployed [1321]. Rediscovering [1746]. Redistribution [2155]. Reduction [1226, 782, 424, 2156]. Reductionism [1228, 1998]. Reductive [1929, 2099]. Reevaluating [2031]. referat [216, 229, 258, 259, 260]. Reference [1932, 1757, 1110, 944, 1854, 1908, 1741, 1692, 1954, 1946, 1554, 1845, 1101, 2191, 1820, 216, 182, 229, 2012]. Reference-frames [1820]. Referentia [1369]. Referential [1930, 1464, 1280, 1928, 1369]. Referential/Attributive [1928]. réferentiel [1280]. Referring [1570]. Reflecting [1748]. Reflections [1855, 1904, 710, 966, 1854, 2083, 1150, 1906, 397, 241, 390, 1054, 444, 30, 816, 333, 924, 796]. Reflective [1646, 1785]. Réflexions [444, 30, 796, 355, 397, 241, 390, 1054, 816]. Reform [626, 700, 701, 641, 637, 634, 638]. réforme [626, 641, 700, 637, 585, 638]. Reformed [2199]. Reforming [585]. Refutation [897, 1576, 1187]. Refutations [1489]. regard [72]. Regarding [815, 891, 1251]. Regelbrauch [1190]. Regelmässigkeit [1027]. règle [472]. règles [156]. Regress [2178, 1806, 2079, 1838, 1528, 2219]. regressive [123, 107]. régressive^{≫} [123]. Regressiveness [851]. Regret [2071]. regularity [1027]. Reichenbach [18, 1522, 2191]. Reidemeister [303]. Reine [1303]. Reinhardt [1847]. Reject [2099]. Rejoinder [1632, 1633]. Rektors [58]. rekursive [420]. Relata [1769]. Relata-Specific [1769]. Related [1941, 890, 1568]. Relation [2162, 1796, 1542, 1195, 1739, 2114, 1222, 1823, 298]. Relation-Based [1823]. Relational [1764, 1844, 1043, 1474]. Relations [1384, 1900, 1015, 1982, 1981, 1602, 1704, 1983, 2063, 1984, 1601, 1528, 1769, 61, 793, 1306]. Relationship [61, 979, 409, 219, 210, 530, 226]. relationships [719]. Relative [2112, 1990, 2103, 421, 534, 1194]. Relativism [1808, 1999, 64, 275, 2234, 2088, 1342, 1187, 2016]. relativisme [64]. relativismo [1776]. Relativismus [275]. relativist [1284]. relativiste [1284]. Relativitätstheorie [893]. Relativité [534]. Relativity [893, 685, 746, 1853, 517, 466, 534, 423]. Relevance [1792, 1624, 2128, 1287]. Relevant [1736]. Reliabilism [1665, 1623]. Reliability [1675]. Reliable [2018, 2217]. Reliably [1933]. Relief [2128]. religieux [903]. Religion [904, 900, 1875, 899, 904]. religions [528]. religious [903]. Remarks [543, 696, 2237, 289, 1047, 929, 1573, 1237, 1087, 979, 681, 1042, 1056, 123, 572, 671, 372, 210, 120, 231, 270, 440, 280, 97, 230, 73, 281, 1352, 432, 813, 639, 207].

Remarque [92, 680]. Remarques [676, 678, 813, 682, 681, 1056, 671, 372, 398, 210, 120, 231, 270, 440, 280, 230]. remembrance [1546]. Reminiscence [1547]. Rendering [1929]. rendu [862]. René [232]. Reouverture [394]. Replace [2085]. Replacement [782]. Replies [1839, 1866, 1763, 1857, 1784, 651, 642, 632]. Reply [669, 2218, 1877, 1627, 2153, 1846, 2217, 1909, 661, 1771, 1927, 1604, 1692, 646, 666, 2221, 1806, 1949, 1151, 2005, 2190, 651, 407, 549, 269, 125, 453]. Réponse [242, 407, 652, 653, 518, 549, 654, 636, 655, 194, 658, 659, 660, 662, 664, 665, 269, 197, 668, 125, 453, 432, 670, 1103, 154, 584]. Réponses [645, 667, 650, 671, 677, 679, 630, 651, 642, 632]. report [671, 677, 629, 628]. Reports [983, 867, 874, 671]. Repräsentation [839, 1747]. Represent [2028]. représentatifs [1347]. Representation [2022, 1616, 1941, 1824, 1687, 1019, 839]. Representational [1809]. Representationalism [2067, 2190]. Representations [1020]. Representative [1347]. Representing [1957, 1578]. repression [509]. Required [2057]. Requirements [1898, 2231]. Research [875, 694, 523, 748, 536, 366, 1065, 475, 312, 250, 282, 155, 281]. researchers [748]. Resemblance [1734, 1755, 2022]. Residual [1403]. Resistance [1277]. resistances [850, 850]. Résolutions [681, 681]. Resolving [2058, 574]. résoudre [574]. respect [1035]. Response [2164, 1713, 1703, 1705, 1707, 1709, 1711, 241, 154, 2154, 644, 898, 1769, 432, 241, 215, 242, 652, 653, 654, 636, 655, 658, 659, 660, 662, 664, 665, 668, 670, 1103, 584]. Response-Dependent [2154]. Responses [1997, 630, 677, 651, 632, 645, 667, 650]. Responsibilities [701]. Responsibility [1774, 2139, 2154, 1510]. Restricted [1569]. Result [1755]. résultats [671]. results [671]. Resumed [1321]. retreat [1096]. Retrieved [1321]. retrospect [41]. Revelation [1867, 1896, 2208]. Reverse [2195, 526]. Reverse-Cycle [2195]. Review [379, 1595, 1667, 2144, 2004, 2001, 2120, 2022, 1746, 1911, 1748, 1974, 1966, 1941, 1774, 2091, 2051, 2081, 1808, 1790, 1884, 2170, 1628, 2092, 1940, 2013, 2003, 2224, 1810, 2090, 1995, 2012, 1993, 2039, 1951, 1747, 2184, 2049, 2082, 1849, 1976, 2169, 1786, 1649, 1847, 2204, 1776, 1910, 2145, 1699, 2122, 2093, 2040, 1789, 1788, 2072, 2103, 2061, 1721, 1874, 2083, 1975, 2186, 2134, 1994, 2048, 1629, 2212, 1848, 2019, 2015, 1785, 2147, 2118, 1737, 2011, 1787, 1773, 2202, 2101, 1777, 18, 304, 2146, 2157, 1885, 2158, 2021, 1952, 2203, 2071, 2002, 1943, 2136, 1698, 2156, 2060]. Review [2213, 1968, 2123, 1738, 1950, 2050, 2159, 2070, 2125, 1723, 2222, 2168, 1873, 1745, 2214, 2020, 2102, 1953, 2119, 1697, 2171, 1722, 1775, 2038, 2059, 2121, 1967, 2223, 1809, 1942, 2124, 2185, 1720, 2014, 1875]. Reviewing [129, 624, 691, 725, 742, 777, 783, 788, 798, 811, 823, 827, 847, 855, 878, 889, 907, 960, 984, 1023, 1044, 1049, 1083, 1098, 1105, 1124, 1178, 1158, 1155, 1157, 1156, 1230, 1253, 1260, 1291, 1301, 1315, 1323, 1328, 1337, 1343, 1365, 1370, 1405, 1410, 1093, 1212, 989, 93, 1137, 1224, 818, 1218, 804, 933, 934, 935, 936, 937, 938, 1164, 1199]. Reviews [251, 263, 341, 457, 1421, 1432, 1440, 1447, 1453, 1459, 1475, 1494, 1498, 1503, 1508, 1514, 1521, 1532, 1538, 1544, 1552, 1558, 1577, 1584, 1599, 1470, 527, 528, 529, 401, 480, 519, 273, 56]. révisibilité [65]. Revision [2036, 65]. Revisited [1930, 1681, 1718, 1893, 1191, 1704, 1474, 1731]. Revolutions [2202, 781]. revues [56]. Ribot [188]. Richard [1121]. Richtungen [117]. Riddle [2055]. right [508]. Rightness [1342, 2145]. Rigid [1648]. Rigidified [1893]. rigor [558]. rigueur [558]. Risk [1078, 2004]. rivali [1721]. Road [1331]. Robert [1940, 1737, 2185]. Roberto [2015]. Robinson [1163, 2219]. Robinson-Selected [1163]. robot [442, 447]. robots [145]. Robust [2072]. Rodogno [2002]. Roger [528]. Role [115, 1226, 1086, 969, 1952, 770, 1556, 745, 561]. Roles [1983, 1583]. Rolle [115, 969]. romaine [498]. Roman [498]. romande [33]. Ronde [862]. Rønnow [2019]. Rønnow-Rasmussen [2019]. Room [2201]. Roots [1335, 1032]. Rosaria [1874]. Rosen [1383]. Rosenthal [696]. Roski [2125]. Round [2076, 862]. Rückzug [1096]. Rule [2133, 472]. Rule-Following [2133]. Rules [1226, 1267, 1159, 156]. Rümelin [1747]. Ruminations [1865]. Russell [810, 1876, 981, 1439, 1273, 1981, 1084, 288, 979, 1272]. Russellian [1717, 1478]. Ruth [1483].

S [2021, 1718, 1671]. S5 [1481]. Sa [915, 330, 1062]. sachgerechten [1033]. Safety [1996]. Said [1712, 2138]. saint [901]. Salerno [2189]. Salmela [2061]. Salva [1932]. Same [1878]. Samuelson [2122]. Sandra [696]. Sandu [2083]. santé [170]. sapiential [770]. Sartre [967, 122]. Saussurian [722]. saussurien [722]. savant [610]. Saving [1875]. savoir [135]. say [1285]. Saying [1759]. says [1488]. Sceptical [1415]. Scepticism [1497, 2040, 1647, 1524, 2194, 1590, 1774]. Schaar [2011]. Schaffer [2016]. schema [444, 444]. schemas [603]. schemata [998, 420]. Schematic [758]. schematische [758]. Schematism [1028, 1091, 331]. Schemes [1249, 424]. Scherrer [407, 432]. Schilpp [233]. Schlick [846]. Schnieder [1839, 1975]. Scholastic [283]. Scholastics [1177]. Scholastikern [1177]. Scholz [327]. School [669, 688, 685, 634, 649, 639, 669, 176]. Schrödinger [599]. Schwarz [184]. Science [245, 253, 264, 247, 255, 266, 246, 254, 237, 704, 705, 711, 751, 778, 939, 1491, 381, 869, 455, 985, 814, 1021, 710, 770, 797, 1461, 1536, 1951, 467, 1500, 265, 117, 748, 383, 2202, 936, 2021, 1303, 1200, 806, 868, 588, 947, 449, 521, 593, 103, 756, 463, 525, 345, 313, 8, 110, 286, 311, 363, 755, 88, 2236, 310, 539, 513, 590, 766, 344, 36, 793, 434, 719, 563, 794, 512, 226, 151, 237, 320, 916, 50, 467, 383, 936, 806, 588, 693, 408, 1699, 2122, 326, 336, 379, 433]. Science [450, 535, 546, 607, 606, 724, 304, 296, 314, 342, 351, 378, 402, 481, 458, 520]. Sciences [524, 1224, 1425, 297, 338, 1022, 245, 253, 264, 618, 8, 50, 363, 552, 228, 118, 156, 532, 530, 719, 76, 874, 320, 916, 1746]. Scientia [590]. Scientific [1843, 1332, 2199, 1381, 1951, 2077, 956, 1847, 698, 1392, 1013, 1339, 170, 116, 435, 927, 439, 511, 437, 452, 363, 440, 78, 1198, 292, 285, 1127, 1284, 730, 438, 538, 486, 397, 475]. scientifique [486, 435, 927, 397, 437, 452, 363, 78, 1127, 730, 438, 451, 538]. scientifiques [1284]. scientist [610, 233]. Scientists [1324]. scienttfiques [170]. scolastique [283]. Scope [1145, 763, 562]. scotiste [1349]. scotistic [1349]. script [561]. Scritti [1911]. Se [1833, 1801]. Séance [494, 484, 200]. Search [2065, 919]. Searle [1540]. Seay [1943]. Second [266, 320, 1770, 265]. Second-Order [1770, 1924]. secondaire [647, 877, 686]. Secondary [688, 1962, 647, 877, 686, 649]. Seeing [2030, 1727, 1726, 1159, 1172]. seen [325]. Sein [261]. Selbstorganisation [1329]. Selected [1163]. Selection [978]. Selective [1529]. Self [1930, 1944, 1941, 1636, 1936, 1976, 2074, 2181, 2027, 1329, 1946, 2085, 1101, 2161, 2048]. Self-Consciousness [1941]. Self-Deception [1944]. Self-Defeating [2074]. Self-Knowledge [1936, 1976, 2181, 1946, 2085, 2048]. Self-Organization [1329]. Self-Reference [1101]. Self-Referential [1930]. Self-Related [1941]. Self-Representation [1941]. Self-Subsistent [1636]. seligen [1789]. Sellars [1010, 736]. selon [963, 1102, 964, 1355, 1369]. Semantic [1489, 1320, 1507, 1526, 1755, 1908, 1824, 1920, 2188, 1598, 717, 1642, 839]. Semantical [891]. Semantics [1464, 1451, 1481, 1894, 1392, 1141, 1272, 1611, 1374, 1174, 2049, 1787]. Semantik [1953]. sémantique [1174]. sémantiques [717]. Semantische [839]. Semipalatinsk [321]. sens [602, 91, 1165]. Sensation [1496]. Sense [1730, 1827, 1868, 1854, 1629, 1047, 1321, 111, 602, 309, 1223]. sensible [1310]. Sensitivity [1997, 2103]. sensori [145]. sensori-moteurs [145]. sensorielles [141]. sensorio [142]. sensorio-motor [142]. sensory [145, 141]. sensory-motor [145]. Sentence [1939, 1530, 1252, 1166]. Sentences [2043]. Sententialism [2000]. Sentimentalism [1729, 2006]. séparabilité [1383]. separability [1383]. sequence [1062, 1062]. Sera [1499]. serait [136]. serait-elle [136]. Seriously [1856, 1651]. Seriousness [2096]. service [148]. servir [442]. ses [65, 170, 813, 1369]. session [494, 200]. Set [1750, 1484, 739]. Set-Theoretical [1484]. set-theory [739]. Seymour [1745]. Shadow [1829]. Shadow-Experiences [1829]. Shame [2002]. Shaping [1952]. Sharing [2161]. Sheer [1708]. Ship [2220]. Shoemaker [2154]. Short [1994, 793]. Shortcut [1330]. Should [1068, 1342, 1890, 2099, 1830, 1571]. Shouldn’t [2124]. show [134]. Shown [1082]. sidelines [131]. Sider [1653]. siècle [850, 1090]. Sigmund [1314]. sign [1160]. signal [399]. Signals [1311]. signe [1160]. Significance [1672, 1021, 1324, 1001, 1812, 156, 478, 1193]. significant [74]. signification [1117, 156, 478, 239, 1193]. significato [1911]. signifique [55]. signifische [74]. Simon [2124]. Simple [1966, 1920, 613, 424]. Simplicity [417]. Simulation [717, 717]. simultaneity [517]. simultaneously [473]. sin [66]. Sinaceur [2083]. since [21]. Sincerely [1759]. Singular [1947, 1555, 945, 1554, 1223]. Sinn [1100]. Sinne [309, 327]. Sinnesfunktionen [141]. situated [1578]. Situation [465, 1065, 465]. Six [1167]. Sixth [939]. Size [746, 368]. Skepsis [1218]. Skeptical [1489, 1664, 1237]. Skepticism [1194, 1462]. Skeptics [1469]. Skeptizismus [1747]. Smith [1810, 1643]. so-called [809]. Soames [2212]. Social [1540, 524, 1810, 748, 1952, 475, 1436, 1156, 1379, 1224, 148, 118, 901, 108, 568, 532, 579, 530, 900]. social-political [1156]. sociale [148, 901, 900]. sociaux [579]. société [237, 904, 33, 276, 899]. sociétés [22, 902]. Societies [902, 22]. Society [1515, 1527, 1533, 1539, 33, 276, 237, 904, 899]. Sociological [523, 904]. Sociologie [172, 340, 77, 527, 529]. sociologique [904]. Sociology [77, 1339, 527, 529, 172, 340]. socioloigie [538]. socioloigy [538]. sociopsychological [165]. Solipsist [1420]. Solution [1592, 2191, 1490, 853]. Solutions [2080]. solved [1522, 829]. solving [250, 282]. Somatosensation [2086]. Some [1042, 962, 1804, 755, 208, 1978, 1252, 1221, 291, 1602, 1387, 902, 181, 1201, 1565, 929, 1019, 1984, 1237, 649, 1159, 849, 890, 1690, 815, 979, 1906, 1056, 850, 1500, 660, 269, 1122, 186, 876, 1054, 845, 723, 142, 572, 710, 1367, 355, 90, 372, 120, 780, 422, 569, 230, 874]. Something [2167, 1698, 2057, 2132]. Sommaire [752, 779]. sommaires [56]. son [1582, 370, 915, 903, 146]. Songe [474, 1353]. sont [223, 602, 1095, 972]. sont-elles [223, 602, 1095, 972]. Sorites [1955, 1878]. Sosa [1785]. Sound [1995]. source [922, 551]. sous [903]. Souvenirs [914]. Sozial [530]. Sozial- [530]. sozialforschung [475]. sozialphilosophie [475]. Sozialwissenschaften [524]. Soziologie [77]. soziologischen [523]. Space [1037, 1816, 813, 2233, 1356, 516, 288, 225, 312]. Spanish [757]. spannung [256]. spark [988]. speak [36]. speaker [1165]. Speaking [1172, 1567, 33]. Special [1714, 1730, 2182, 2105, 1706, 1725, 1724, 1727, 2023, 1729, 1701, 1703, 1705, 1707, 1709, 1711, 1712, 1726, 1702, 1710, 1704, 1728, 1708, 1731, 182]. species [26, 31, 30, 28]. Specific [1769]. Specifying [1295]. Specious [2183]. spectacle [134]. Speculative [1363, 1461]. speech [829, 925]. Speiser [244, 244]. Spekulation [387]. Sphere [1815]. Spiele [729]. Spieltheorie [391]. Spinning [1815]. Spinoza [715, 1122]. spirit [1139, 301]. Spiritual [117, 918]. spiritualité [904, 899]. spirituality [904, 899]. spirituelle [918]. Spohn [2092]. sporadic [1064]. sporale [1064]. Sprach [1200]. Sprach- [1200]. Sprachanalyse [829, 1334]. Sprache [381, 830, 1200, 947, 411, 1189, 942]. Sprachen [838]. Sprachgebrauch [74]. sprechen [830]. Spurious [1947]. square [1094]. St [2181]. stades [374]. Stage [698]. Stages [1580, 374]. Stance [2162]. standard [241, 1119]. Standing [2180]. standpoint [251, 511]. Standpunkt [251, 348, 346]. Standpunktes [414]. Stanisław [1139]. State [506, 505, 504, 2014, 507]. statement [169, 260]. Statements [1958, 473]. States [2227, 1739]. Static [31]. statiques [31]. Statistical [688, 105]. Statistics [1613, 395, 277, 1058]. Statistik [395]. statistique [105, 1058]. Status [696, 1076, 1505, 695, 1436, 1065]. Steal [2167]. Steenbergen [341]. Stefan [2125, 1966, 2146]. Stefania [1885]. Stefano [1722]. stellar [104]. Stellenwert [1035]. stellung [461]. Stéphane [38]. Stephen [2003, 2101]. stereotype [537]. Stereotyps [537]. stimuli [149]. Stoa [1218]. Stochastic [2140]. stochastics [106]. stochastique [106]. Stoneham [1846]. Storie [1884]. Stories [1884]. Stout [2011, 2218]. strategic [712, 712, 712, 712].

Strategies [2065]. Strategische [729]. Strauss [967]. strenger [409]. Strevens [2210]. Strict [1917, 409]. Strike [1361]. strikter [409]. Stroll [1071]. Strömungen [164]. Strong [1339]. Structural [896, 1382, 1240]. Structuralism [1416]. Structuralist [1116, 1204]. Structure [967, 1829, 1664, 1975, 1825, 1667, 376, 439, 392, 747, 192, 440, 234, 763, 1349, 1974, 2014]. Structures [1020, 1940, 598, 758, 189]. structuring [150]. Struktur [763]. Strukturale [1747]. Strukturen [758]. strutturazione [150]. Students [685]. studiata [150]. studied [150]. Studien [2020, 1158]. Studies [691, 725, 742, 777, 783, 788, 798, 811, 823, 827, 847, 855, 878, 889, 907, 960, 984, 1023, 1044, 1049, 1083, 1098, 1105, 1124, 1178, 1158, 1155, 1157, 1156, 1230, 1253, 1260, 1291, 1301, 1315, 1323, 1328, 1337, 1343, 1365, 1370, 1405, 1410, 1093, 1212, 989, 1137, 1224, 818, 1218, 804, 933, 934, 935, 936, 937, 938, 1164, 1199, 129, 624, 93, 422, 426]. Study [669, 1211, 1843, 2032, 904, 149, 566, 141, 1288]. Stueber [1746]. Style [1624, 1989]. Subject [611, 833, 348, 298]. subject-object-relation [298]. subjectif [16]. Subjective [1013, 602, 16]. subjectives [602]. Subjectivism [2064]. subjectivité [14]. Subjectivity [448, 14]. Subjects [2033]. subjekt [298]. subjekt-objektrelation [298]. Subjektivität [1747]. subordination [508]. Subsectivity [2076]. Subsequent [1364]. Subsistent [1636]. Substance [2080, 1657]. Substantive [1612]. substitutional [949, 1175]. substitutionelle [949]. substitutionnelle [1175]. substratum [146]. substructure [392]. Subtraction [1671]. Success [1951]. Successful [2009]. succinct [862]. Suffering [1976]. sufficient [455]. suffisante [928]. Suggestion [891]. suggestions [355]. Suggestives [1926]. suite [534, 740, 311, 771]. sum^{≫} [410]. Summa [1761]. Summaries [1131]. Summary [862]. Supermaxim [1038]. Superordinate [262]. Supervaluationism [1526]. Supervenience [1815, 1796, 1217, 2116, 1871, 1466]. Support [2045, 1267, 1772]. Supposed [957, 1694]. supposition [1322, 2204]. Supreme [503, 500]. Surfaces [952]. Surprises [1092]. Survey [1212, 1073, 536, 630, 651, 642, 632, 671]. Susana [1943]. Swamping [1933]. Swiggers [1042]. Swiss [33, 58]. Symbol [619]. symbolic [41]. Symmetrical [1672]. symmetry [767]. Symposium [626, 1155, 789, 1877, 1783, 1632, 1633, 1630, 1782, 1760, 1763, 1631, 1876, 1596, 1761, 1784, 1781, 1762, 629]. Synonymy [1016]. Syntax [381, 714]. Syntaxe [714]. Synthese [371, 55, 581, 493]. Synthesis [371, 55, 581, 493]. Synthetic [1028]. System [367, 982, 87]. systematic [1035, 994]. systematics [382]. systématique [382]. systematische [1035]. systematischen [994]. Systeme [838, 1950, 367, 87, 423]. Systémes [1306, 1139, 579, 533]. systemic [1289]. systémique [1289]. Systems [1988, 838, 866, 1148, 423, 1306, 1004, 412, 745, 416, 1139, 424, 579, 533]. Szientismus [1200].

T [2171, 959, 885]. Table [47, 190, 306, 1262, 1293, 1317, 47, 190, 306, 862]. Tacit [2039]. Take [1977]. Taken [1694]. Taking [1856, 1651]. Tale [1813].

Talk [952, 1490, 830]. Talking [1670, 1335, 2158]. Talks [202]. tant [548, 1326]. Target [1997]. Tarski [1207]. Tasks [1920]. Taste [2016]. Tätigkeit [1185]. taxinomiques [31]. taxonomic [31]. Taxonomy [1012]. Tea [460]. teach [877]. Teachers [276]. Teaching [687, 641, 651, 642, 632, 671, 635, 656, 677, 644, 645, 647, 876, 667, 689, 926, 648, 649, 638, 639, 650, 631, 640, 377]. technical [399, 209]. technique [337, 1054, 399]. Technische [209]. Technischen [58]. technology [337, 58]. Ted [2090]. Tegtmeier [1847]. Teil [1390, 1412]. Teilhard [299]. tel [79, 459]. Teleological [2182]. Téléologie [965]. Teleology [1022, 965]. Teleosemantics [1956]. Telicity [2014]. Tell [2021]. Telos [54, 40]. Temporal [1694, 2183, 1655, 1306, 599]. temporelle [1306, 599]. Temps [714, 715, 731, 740, 920, 1356, 761]. tendances [31, 366]. tenebrae [578]. Tense [1651]. Tension [1497, 256]. Tensions [578]. tenth [330]. Teorie [1790, 1721]. term [74, 1737]. terme [382]. Termination [1478]. Terminus [74]. Terms [1947, 1391, 945, 2230, 1496, 695, 1223, 696]. Teroni [1993, 2002]. terrorist [220]. Tessman [2051]. Test [2220, 569, 15]. textes [795]. Textiles [931]. texts [795]. th [198]. the‘British [705, 711]. Their [2147, 719, 116, 973, 478, 239, 344, 2208, 226]. Theism [1773]. Themas [348]. theme [1280, 1280, 485, 485]. Themes [1267, 2020]. Themselves [958]. theological [1290, 903]. théologique [1290, 903]. Theorem [1616, 100, 2032, 1352, 2047]. théorèmes [369]. Theorems [100, 1551, 369]. Theoretic [1226, 1832, 2077, 1392, 774]. Theoretical [696, 1484, 695, 815, 475, 895, 376, 411, 1287, 90, 618, 747, 1326, 302, 1495, 464, 905]. Theoreticity [1135]. théorétique [747]. theoretische [411, 302]. Theorie [1511, 224, 1036, 201, 205, 213, 228, 807, 217, 279, 865, 223, 540, 219, 53, 221, 210, 212, 218, 43, 399, 230, 204, 227, 211, 157]. Theorie“ [88]. Theorien [1520, 285]. Theories [1507, 1207, 1204, 1149, 774, 1019, 1403, 1998, 116, 1520, 285, 1284, 1284, 737, 737, 2081]. théorique [464, 1287, 618, 1326]. théoriques [90, 905]. Theorists [1082, 1556]. Theory [1930, 2007, 1048, 1194, 1627, 1740, 2092, 1511, 1890, 2206, 1078, 243, 1750, 1576, 2069, 1844, 2077, 834, 951, 893, 1327, 2154, 837, 1029, 1046, 810, 1273, 1859, 894, 2227, 2026, 1678, 2118, 807, 2042, 1495, 1554, 1167, 1036, 1166, 406, 1200, 2130, 849, 821, 1617, 1928, 1651, 157, 865, 279, 223, 308, 315, 1128, 540, 614, 219, 53, 603, 221, 734, 210, 413, 745, 212, 970, 88, 2236, 359, 559, 1198, 739, 1452, 1201, 95, 541, 218, 43, 182, 399, 290, 886, 1358, 230, 701, 203, 201, 205, 213, 224, 228, 2145, 204]. Theory [217, 227, 211, 2171]. theory-ladenness [1201]. Theory-relative [1194]. There [2218, 1215, 1252, 1923, 1779, 2132, 1793, 1425, 1149, 1881, 1797]. Therefore [2138, 410]. Thesen [203]. Theses [1131, 203]. Theseus [2220]. Thesis [1166, 1251]. thick [1593]. thickness [599]. thin [1593]. Thing [1726]. Things [958, 1978, 2197, 2063, 2157, 813]. Think [2100, 1669, 410]. Thinking [115, 1327, 1982, 244, 2057, 1335, 2106, 222, 68, 613, 261]. Third [22, 1197, 9, 203, 202, 2047]. Thomas [2202]. Thomasson [2123]. Thornton [2039]. Thought [1116, 1589, 1887, 240, 950, 1946, 1823, 385, 2161, 1826, 1379, 920, 238, 1289, 1319, 333, 719, 1349, 2059]. Thoughts [1468, 2043, 869, 1378, 1312, 2106, 868, 355, 649, 611]. Three [178, 1653, 1314, 613, 1026]. Three- [1653]. three-component [613]. Tiefenpsychologie [117]. Ties [1672]. Tim [2039, 2059]. Time [2206, 1735, 893, 1268, 1651, 713, 564, 731, 740, 920, 1356, 309, 761, 714, 715, 1418, 2119]. TOC [2041]. Today [1067, 10, 1404, 987]. Toil [1708]. Token [1576]. tolerate [1340]. Too [1988, 1970]. Top [1660]. Topologie [327, 309]. topology [309]. Torrengo [2015]. toute [372]. Tractarian [1630]. Tractatus [1632, 1633, 1631, 1607, 1630, 1631, 1579]. tradition [1286]. Traditional [1013]. Traditions [956, 905]. trägt [1334]. Tragweite [763]. traité [529]. traitement [795, 1085]. Trans [812]. Trans-epistemological [812]. Transaction [821]. transcendantaux [268]. Transcendental [958, 1028, 1398, 1296, 1295, 1029, 1025, 1392, 1397, 1091, 2185, 1034, 1033, 268]. Transcendentalism [1386, 1185]. transfer [298]. Transferability [1668]. transfinie [294]. transfinite [294]. Transition [275, 599]. Transitions [1384]. Translation [834, 882, 2131]. Transparency [1805, 1772]. transtemporalen [1747]. Transzendentale [1033]. transzendentalen [1034]. Transzendentalismus [1185]. Transzendentalphilosophie [1397]. Travaux [1347]. Travel [1735]. travers [325]. treatise [869, 529]. treatment [1085]. treaty [578]. treibt [1396]. Trends [366, 164, 31]. Triangulation [1592, 1506]. tribute [826, 610, 863, 19]. Trois [1026, 1314]. troisième [9]. Trolley [2044]. Trope [1984]. Tropes [1668, 1882, 2165]. True [1988, 1252, 1579, 1009, 2061]. Truly [1451, 877]. Truth [1959, 835, 1485, 1211, 1855, 1560, 446, 1210, 1342, 2008, 1939, 1725, 1209, 1990, 2056, 1610, 1760, 2103, 1207, 1246, 1615, 1692, 1402, 1856, 1236, 1401, 1187, 883, 1960, 1761, 1733, 1830, 1321, 1762, 44, 300, 397, 239, 274, 1618]. Truth-Bearers [883]. Truth-Conditions [1939, 1321]. Truthmakers [1482]. Truths [1565, 1312]. Trying [1859]. Tune [1919]. Turn [1330, 1874, 1977]. turpitudinem [575]. Turri [2058, 2217]. Tuske [2159]. Twelve [2147]. twentieth [850]. Twenty [2184, 1786]. Twenty-First [2184, 1786]. Two [1991, 2080, 1648, 1130, 739, 1868, 1813, 2042, 1534, 781, 1473, 2132, 38, 1153, 963, 460, 768, 1306, 1223, 1214, 226, 801, 1787]. Two-Component [2132]. Two-Dimensional [1787]. Type [1576, 1958]. Type-Identity [1958]. Type-Token [1576]. Types [1750, 1648]. Typology [1018].

U.I.P.S. [264, 266, 265]. U.S.A [408]. U.S.A. [379, 450, 481]. überblick [157]. Übereinkunft [1190]. Übergang [275]. Überlegungen [924]. übertragung [298]. überwindung [475]. UK [2145]. Umfrageforschung [536]. Umgekehrt [1334]. Umkehrung [100]. Unabhängigkeit [735]. Unbewussten [115, 297]. Unbounded [1773]. unconditioned [268]. Unconscious [115, 297]. Unddoch [349]. Undefinability [1402]. Undefinierbarkeit [1402]. Under-determinacy [882]. Under-Determination [834]. Underdeterminacy [2043].

Underdetermination [1227, 1298]. Underdetermines [2138]. understand [79]. Understanding [2115, 710, 1939, 525, 1995, 1951, 1972, 2065, 1215, 996, 1086, 1975, 995, 1072, 1001, 1166, 2188, 1794, 1905, 397, 1408, 2001]. Unending [710]. Unendlichen [776]. Unendorsed [2129]. Unentscheidbarkeitstheorem [1352]. ungarischen [1087]. Ungrounded [1806]. unification [649]. Unificationism [1624]. Unifier [1739]. unify [464]. Union [320, 916, 320, 916]. unis [610]. unité [135]. United [1838, 610]. Unity [1992, 2080, 1837, 1835, 1839, 135, 1177, 2209, 1881, 1838, 1836, 2233, 49, 2144, 808]. univers [1052, 136, 318]. Universal [1765, 1324, 1604]. Universalism [1883]. universalité [26]. universality [26]. Universals [1456, 1603, 1863, 831, 883]. Universe [1052, 136]. Universes [1686]. universitaire [988]. University [2144, 2147, 1897]. Unknown [1681]. unmögliche [1747]. unordnung [262]. Unprovable [1252]. unpublished [1084]. unsere [824]. unserer [49]. unseres [49]. Untersuchung [74, 530]. Untersuchungen [422]. unused [414]. unwavering [196]. upon [553]. urgency [506]. Urs [888]. USA [669, 433]. usage [74]. Use [1589, 1021, 1414, 1753, 1129, 382, 1190, 1059, 573, 553]. used [442]. useful [1128]. Using [1940, 1115]. Utility [1999]. Utterances [2138].

V [489, 489, 1637]. vague [1266, 1541]. Vagueifiers [1526]. Vagueness [1695, 1655, 1987, 1916, 2219]. Valeur [435, 392]. valeurs [489, 388]. validité [65]. Validity [1309, 65]. Vallicella [1839, 1769]. valuation [220]. Value [1715, 392, 435, 2019]. Values [489, 388, 2028, 1853, 1136]. Varia [231, 226, 220]. Variable [1073, 1831]. variables [970]. Variation [1681, 2065]. Variations [1852]. Varieties [1269, 1525]. Varzi [1884, 1656, 1575]. Vat [1419, 1502]. Vector [1817]. Vectorial [1815]. Vectors [1816, 1814, 1813, 1811, 1812]. velocity [534]. Verena [1747]. Verfahren [524]. Vergil [1566]. Vergleich [1088]. Verhältnis [530, 226]. Verification [957]. verità [1776, 1722, 1911, 1790]. Verità- [1790]. Veritate [1761, 1932]. Vérité [397, 44, 446, 300, 239, 1618]. Vernunft [387]. vernunftkritische [1033]. Vers [756, 1282, 538]. versa [1334]. Version [1229, 1281]. Verstehen [997]. Versuch [469, 338]. Versus [1847, 1270, 1010, 805, 1932, 1185, 1469, 1013]. Very [1536, 1546]. vestal [988]. vestale [988]. VI [490, 683]. via [1754]. Viable [1557]. vice [1334]. Vicious [979, 981]. Vicious-Circle [979]. Victor [2181]. vie [737, 1052, 148, 915, 136]. viendra [843]. View [1807, 1507, 1204, 774, 2071, 1043, 2132, 2014, 502, 634, 1242, 510, 732, 563, 418, 1723]. Viewpoint [346]. Views [2106]. Vii [491, 684, 491]. VIII [492]. VIII. [690]. Vindicated [1627]. Vindication [1528]. Vindications [2211]. vingtième [850]. Violation [1925]. Violeta [2014]. Virtue [2152, 1510, 1785]. Virtues [1150]. vitesse [534]. vivant [397, 1288]. Vocation [797]. Vol [1262, 1293, 1317]. volant [509]. Volitions [2009]. Volpe [1790]. Voltolini [1782]. Volume [1659, 1700]. Voluntariness [1889]. vom [297, 1219, 225, 776]. vor [1225]. Vorfeld [942]. Vorgehen [151]. vorgeschlagenen [1088]. vormittage [256].

Vorstellungen [285]. Vorwort [941, 1066, 1183, 1385, 880, 828, 1138, 1264, 160, 992]. Vosgerau [1941]. vous [843]. vrai [101]. vs [837, 887, 881, 1856, 886, 2121]. vue [502, 510]. vues [325]. Vuillemin [1504].

W [407, 432, 591]. W. [285]. Wager [2166]. Wahrheit [1402]. Wahrscheinlichkeit [1088, 278, 261]. Wahrscheinlichkeitsdefinition [94]. Wahrscheinlichkeitsrechnung [270, 271, 417, 272]. Wahrscheinlichkeitstheorie [95, 541]. Walde [1747]. Wallace [2071]. Wallimann [1730, 1725, 1724, 1727, 1729, 1726, 1728, 1731]. Wandlungen [1114]. Wanting [1859]. Warte [418]. Waryn [347]. Was [1396, 1334]. Watkins [1666]. Wave [1961, 600, 1266]. Way [1982]. Ways [1774, 1679]. Weak [2179, 1217, 1926]. wechselseitige [166]. Wehrle [457]. Weight [1844]. Wein [275, 275]. Weir [1709]. Weiss [869, 869]. Weizsäcker [1088]. welchem [1396]. Welcome [58]. Welt [1747]. werden [829]. Werk [968]. Wertung [220]. Wertungen [1225]. Wesen [1747]. Wesens [49]. Western [385, 905]. Weyl [1325]. Weyls [1325]. Wh [1445]. Wh-Knowledge [1445]. Where [1427, 1897, 1456, 1129]. Whitehead [448, 430, 448]. Who [1116, 2044]. whole [461, 165, 622]. Widerspruch [1087]. Wielenberg [2072]. Wieweit [829]. Wiggins [2147]. Will [2141, 1518, 1859, 1773, 2185, 333, 843]. Willaschek [1747, 2082, 2222]. Willens [333]. Willensfreiheit [1747, 391]. Williamson [1904]. Willing [1859]. Winding [1331]. Winner [1734, 1802, 1969, 2032]. Wires [2086]. Wirklichkeit [942, 947, 324, 335, 86, 323]. Wissenschaft [151, 1649]. Wissenschaften [246, 254, 265, 381, 338]. Wissenschaftslehren [226]. Wissenschaftstheoretische [475]. wissenschaftstheoretischen [116]. Withholding [2058]. Within [1394, 1556, 1142]. Without [944, 1891, 1771, 1092, 1173, 2108, 1783, 1362, 2165, 2085]. Wittenberg [281, 639, 640, 281, 611, 280, 639, 640]. Wittgenstein [2050, 1155, 1003, 1497, 1530, 1276, 1215, 1579, 1275, 1573, 813, 1849, 2050]. Wittgenstein’s [1876]. Wolfgang [2092, 1761, 1953]. woman [537]. Wonderful [590]. Won’t [2117]. Woodbridge [2134]. Word [54, 2105]. Words [832, 2109, 58, 560, 1314]. Work [1696, 1276, 610, 968, 121, 207]. Works [1568, 1347, 935, 803, 1951]. World [2141, 1884, 1220, 2139, 1691, 986, 1420, 2070, 1005, 77, 808, 1312, 1280, 563, 1976]. Worldmate [1796]. Worlds [2189, 1985, 1382, 1513, 1962, 1493, 1240, 1312]. Worms [1220]. Worry [1918]. Wort [54]. Would [136]. Wouldn’t [1694]. Wright [2164, 1951, 1008, 1907]. Writings [1339]. Wrong [2029]. Würdigung [19].

xi [2145, 198]. XI-th [198]. xv [2144].

Yablo [1743]. Yanofsky [2021]. year [330]. years [41, 426]. Yes [1237, 317]. York [2147]. Yourself [2124].

Zagzebski [2171]. zahl [405]. Zahlbegriff [1325]. Zahlen [100]. Zahlentheorie [970]. Zaslawsky [1102]. Zeit [327, 893, 309]. Zeitschriftenschau [56]. Zellorganellen [1055]. Ziel [54]. Zimmerman [1652]. zipperdynamics [895]. Zombie [2177]. Zufall [469]. Zugang [1747]. Zum [1200, 251, 73, 1397, 639, 475, 776, 229, 17, 789, 968, 417, 151]. Zur [1033, 1357, 624, 691, 725, 777, 783, 788, 798, 811, 823, 827, 847, 855, 878, 889, 907, 912, 960, 984, 1023, 1044, 1049, 1083, 1098, 1105, 1124, 1178, 1230, 1253, 1260, 1291, 1301, 1315, 1323, 1328, 1337, 1343, 1365, 1370, 1405, 1410, 1093, 80, 381, 1212, 989, 1747, 93, 1137, 1224, 818, 94, 1352, 275, 804, 933, 934, 935, 936, 937, 938, 1164, 1199, 432, 422, 1334, 2020, 203, 75, 251, 271, 111, 123, 308, 315, 348, 1520, 6, 270, 405, 368, 126, 841, 117, 32, 272, 225, 51, 41]. Zurich [245, 253, 264, 247, 255, 266, 246, 254, 265, 246, 254, 265, 245, 253, 264, 202, 4, 257, 203, 202, 203]. Zurn [2093]. Zusammenhängen [994]. zweiter [265]. zwischen [409, 350, 530, 256].

References

Anonymous:1947:E

[1] Anonymous. Editorial. Dialectica: International Review of Philosophy of Knowledge, 1(1):4–8, February 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1947:EG

[2] Anonymous. Editorial: Geleitwort. Dialectica: International Review of Philosophy of Knowledge, 1(1):9–10, February 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bachelard:1947:PDF

[3] Gaston Bachelard. La philosophie dialoguée. (French) [Dialogue philosophy]. Dialectica: International Review of Philosophy of Knowledge, 1(1): 11–20, February 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1947:LDE

[4] F. Gonseth. L’idée de dialectique aux entretiens de Zurich. (French) [The idea of dialectic at the interviews in Zurich]. Dialectica: International Review of Philosophy of Knowledge, 1(1):21–37, February 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Konig:1947:DGE

[5] H. König. Dialektik als gedankliche experimentierkunst. (German) [Dialectics as a mental experiment]. Dialectica: International Review of Philosophy of Knowledge, 1(1):38–44, February 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Durr:1947:EIE

[6] Karl Dürr. The evolutions of ideas — l’évolution des idées — Zur Ideengeschichte: Die Entwicklung der Dialektik von Platon bis Hegel. (German) [The development of the dialectic from Plato to Hegel]. Dialectica: International Review of Philosophy of Knowledge, 1(1):45–62, February 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Walter:1947:BDI

[7] Emil J. Walter. Der Begriff der Dialektik im Marxismus. (German) [The concept of dialectics in Marxism]. Dialectica: International Review of Philosophy of Knowledge, 1(1):63–71, February 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gagnebin:1947:IDP

[8] H.-S. Gagnebin. Sur l’idée de dialectique dans la philosophie des sciences contemporaine. (French) [On the idea of dialectics in the philosophy of contemporary science]. Dialectica: International Review of Philosophy of Knowledge, 1(1):72–75, February 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1947:PPP

[9] Anonymous. Philosophical portraits — portrait philosophique — philosophisches Portrait: pour le troisième centenaire de la naissance de Leibniz (1646–1716). (French) [On the third centenary of the birth of Leibniz (1646–1716)]. Dialectica: International Review of Philosophy of Knowledge, 1(1):77–96, February 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1947:PTA

[10] F. Gonseth. Philosophy of today — l’actualité philosophique — Die philosophische Lage: à propos d’un congrès international de philosophie. (French) [On an international congress of philosophy]. Dialectica: International Review of Philosophy of Knowledge, 1(1):98–108, February 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1947:DDP

[11] Anonymous. Discours du pape. Dialectica: International Review of Philosophy of Knowledge, 1(1):108–111, February 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1947:EE

[12] Anonymous. Editorial: editorial. Dialectica: International Review of Philosophy of Knowledge, 1(2):115–116, May 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1947:COC

[13] F. Gonseth and J.-P. Gonseth. Connaissance objective et connaissance poétique. (French) [Objective knowledge and poetic knowledge]. Dialectica: International Review of Philosophy of Knowledge, 1(2):117–126, May 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Destouches-Fevrier:1947:NOS

[14] Paulette Destouches-Février. Les notions d’objectivité et de subjectivité en physique atomique. (French) [The notions of objectivity and subjectivity in atomic physics]. Dialectica: International Review of Philosophy of Knowledge, 1(2):127–146, May 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fiala:1947:ENO

[15] F. Fiala. Essai sur les notions d’ouverture et de fermeture. (French) [Test on the concepts of opening and closing]. Dialectica: International Review of Philosophy of Knowledge, 1(2):147–158, May 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hainard:1947:DOD

[16] Robert Hainard. Dialectique de l’objectif et du subjectif dans les arts plastiques. (French) [Dialectic of the objective and the subjective in the plastic arts]. Dialectica: International Review of Philosophy of Knowledge, 1(2):159–171, May 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1947:BDG

[17] Paul Bernays. Zum Begriff der Dialektik. (German) [The concept of dialectics]. Dialectica: International Review of Philosophy of Knowledge, 1(2):172–175, May 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pauli:1947:BRH

[18] Wolfgang Pauli. Book review: H. Reichenbach, Philosophic Foundations of Quantum Mechanics. Dialectica: International Review of Philosophy of Knowledge, 1(2):176–178, May 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Preissmann:1947:PAA

[19] A. Preissmann. Philosophical address — adresse philosophique — Würdigung eines Forschers: Hommage à Monsieur Mannoury. (French) [Tribute to Monsieur Mannoury]. Dialectica: International Review of Philosophy of Knowledge, 1(2):189–190, May 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gunther:1947:EEE

[20] Werner Günther. Echoes — échos — entgegnungen: Brief an Dialectica. (German) [Letter to Dialectica]. Dialectica: International Review of Philosophy of Knowledge, 1(2):191–194, May 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gerard:1947:PDA

[21] Jacques Gérard. Philosophy of the day — l’actualité philosophique — Die philosophische Lage: Le mouvement philosophique en Belgique depuis 1939. (French) [Philosophy of the day — philosophical news — Die philosophische Lage: The philosophical movement in Belgium since 1939]. Dialectica: International Review of Philosophy of Knowledge, 1 (2):195–199, May 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Barzin:1947:ICS

[22] M. Barzin and L. Noël. III^{e} congrès des sociétés de philosophie de langue française. (French) [third congress of philosophy societies in the French language]. Dialectica: International Review of Philosophy of Knowledge, 1(2):200–203, May 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1947:MM

[23] Anonymous. Memorandum: memorandum. Dialectica: International Review of Philosophy of Knowledge, 1(2):203, May 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pauli:1947:EPQ

[24] Wolfgang Pauli. Exclusion Principle and quantum mechanics: Discours prononcé à la réception du prix Nobel de physique 1945. (French) [Lecture given on receiving the 1945 Nobel Prize in Physics]. Dialectica: International Review of Philosophy of Knowledge, 1(2):204, May 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Redaction:1947:AA

[25] La Rédaction. Advertisement: avertissement. Dialectica: International Review of Philosophy of Knowledge, 1(3):206, August 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1947:EUN

[26] Anonymous. Editorial: universalité de la notion d’espèce. (French) [Editorial: universality of the notion of species]. Dialectica: International Review of Philosophy of Knowledge, 1(3):207–213, August 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Niggli:1947:AMG

[27] Paul Niggli. Der Artbegriff in der Mineralogie. (German) [The concept of nature in mineralogy]. Dialectica: International Review of Philosophy of Knowledge, 1(3):214–228, August 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gagnebin:1947:NEB

[28] Elie Gagnebin. La notion d’espèce en biologie. (French) [The notion of species in biology]. Dialectica: International Review of Philosophy of Knowledge, 1(3):229–242, August 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1947:NDN

[29] F. Gonseth. La notion du normal. Dialectica: International Review of Philosophy of Knowledge, 1(3):243–252, August 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Favarger:1947:RIE

[30] Claude Favarger. Réflexions sur l’importance de l’espèce et de la classication en botanique. (French) [Reflections on the importance of species and classification in botany]. Dialectica: International Review of Philosophy of Knowledge, 1(3):253–263, August 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dubois:1947:ASE

[31] Georges Dubois. Des attributs statiques de l’espèce aux tendances évolutives des groupes taxinomiques. (French) [Static attributes of species to the evolutionary trends of taxonomic groups]. Dialectica: International Review of Philosophy of Knowledge, 1(3):264–276, August 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pensa:1947:EIE

[32] M. Pensa. The evolution of ideas — l’évolution des idées — Zur Ideengeschichte: Le logos hégélien. (French) [The evolution of ideas — the evolution of ideas — Zur Ideengeschichte: The Hegelian logos]. Dialectica: International Review of Philosophy of Knowledge, 1(3):277–287, August 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fiala:1947:PDA

[33] F. Fiala. Philosophy of the day — l’actualité philosophique — Die philosophische Lage: L’assemblée annuelle de la société romande de philosophie. (French) [The annual meeting of the French-Speaking Swiss Society of Philosophy]. Dialectica: International Review of Philosophy of Knowledge, 1(3):288, August 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1947:C

[34] Anonymous. Communications. Dialectica: International Review of Philosophy of Knowledge, 1(4):290, November 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1947:EFE

[35] Anonymous. Éditorial. (French) [Editorial]. Dialectica: International Review of Philosophy of Knowledge, 1(4):291–292, November 1947. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Redaction:1947:PPF

[36] La Rédaction and Ferdinand Gonseth. Peut-on parler de ^{≪}science dialectique^{≫}?. (French) [Is it possible to speak of “dialectic science”?]. Dialectica: International Review of Philosophy of Knowledge, 1(4):293– 304, November 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1947:CNC

[37] Paul Bernays. Contradiction et non-contradiction. (French) [Contradiction and non-contradiction]. Dialectica: International Review of Philosophy of Knowledge, 1(4):305–309, November 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1947:PDO

[38] F. Gonseth. À propos de deux ouvrages de M. Stéphane Lupasco. (French) [On two books by M. Stéphane Lupasco]. Dialectica: International Review of Philosophy of Knowledge, 1(4):309–316, November 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Preissmann:1947:POM

[39] A. Preissmann. À propos d’un ouvrage de M. Léon Bopp. (French) [On a book by M. Léon Bopp]. Dialectica: International Review of Philosophy of Knowledge, 1(4):316–318, November 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Taubes:1947:LTG

[40] Jacob Taubes. Logos und Telos. (German) [Logos and Telos]. Dialectica: International Review of Philosophy of Knowledge, 1(4):319–330, November 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Beth:1947:EIE

[41] Evert W. Beth. The evolution of ideas — l’évolution des idées — zur Ideengeschichte: Hundred years of symbolic logic — a retrospect on the occasion of the Boole–de Morgan centenary. Dialectica: International Review of Philosophy of Knowledge, 1(4):331–346, November 1947. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pensa:1947:LHI

[42] M. Pensa. Le logos hégélien II. (French) [The Hegelian logos II]. Dialectica: International Review of Philosophy of Knowledge, 1(4):347–353, November 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Perelman:1947:PPP

[43] Ch. Perelman. Philosophical portraits — portrait philosophique — Philosophisches Portrait: fragments pour la théorie de la connaissance de M. E. Dupréel. (French) [Fragments for the theory of knowledge of M. E. Dupréel]. Dialectica: International Review of Philosophy of Knowledge, 1(4):354–361, November 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1947:BVF

[44] Anonymous. B. De la vérité. (French) [B. On truth]. Dialectica: International Review of Philosophy of Knowledge, 1(4):362–366, November 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dupreel:1947:EEE

[45] E. Dupréel. Echoes — échos — entgegnungen: sur les origines de la dialectique. (French) [On the origins of dialectics]. Dialectica: International Review of Philosophy of Knowledge, 1(4):367–370, November 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nolfi:1947:BNN

[46] P. Nolfi. Bibliographic notes notes bibliographiques bibliographische notizen. Dialectica: International Review of Philosophy of Knowledge, 1 (4):371–374, November 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1947:TMF

[47] Anonymous. Table des matières. (French) [Table of contents]. Dialectica: International Review of Philosophy of Knowledge, 1(4):375–376, November 1947. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1948:EE

[48] Anonymous. Editorial: editorial. Dialectica: International Review of Philosophy of Knowledge, 2(1):6–10, February 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1948:EGG

[49] Anonymous. Einheit in der Gegensätzlichkeit als eine Grundlage unseres Wesens und unserer Erkenntnis. (German) [Unity in opposition as a foundation of our being and our knowledge]. Dialectica: International Review of Philosophy of Knowledge, 2(1):11–24, February 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1948:SPF

[50] F Gonseth. Les sciences et la philosophie. (French) [Science and philosophy]. Dialectica: International Review of Philosophy of Knowledge, 2(1): 25–44, February 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1948:EIE

[51] Anonymous. The evolution of ideas — d’évolution des idees — zur Ideengeschichte III le logos hégélien. (French) [The evolution of ideas — of evolution of ideas — on history of ideas. III. The Hegelian logos]. Dialectica: International Review of Philosophy of Knowledge, 2(1):45– 62, February 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1948:ME

[52] Anonymous. Metaphysique experimentale. Dialectica: International Review of Philosophy of Knowledge, 2(1):47–62, February 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

DelaNecessiteo:1948:PPP

[53] C. De la Nécessitéo. Philosophical portraits — portrait philosophique — Philosophisches Portrait: fragments pour la théorie de la connaissance de M. E. Dupréel. (French) [Philosophical portraits — philosophical portrait — Philosophisches Portrait: Fragments for the theory of knowledge of M. E. Dupréel]. Dialectica: International Review of Philosophy of Knowledge, 2(1):63–77, February 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Walter:1948:WLZ

[54] Emil J. Walter and H. Guggenheimer. Wort (Logos) und Ziel (Telos)? : Brief an Dialectica. (German) [Word (logos) and destination (telos)? : Letter to Dialectica]. Dialectica: International Review of Philosophy of Knowledge, 2(1):78–80, February 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1948:EEA

[55] Anonymous. Exchanges — échanges — Austauche: le mouvement signifique et synthèse. (French) [Exchanges — exchanges — Austauche: the movement of meaning and synthesis]. Dialectica: International Review of Philosophy of Knowledge, 2(1):80–82, February 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1948:CRS

[56] Anonymous. Contents of reviews — sommaires de revues – Zeitschriftenschau. Dialectica: International Review of Philosophy of Knowledge, 2 (1):83–84, February 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1948:A

[57] Anonymous. Avertissement. Dialectica: International Review of Philosophy of Knowledge, 2(2):89–90, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pallmann:1948:BRE

[58] H. Pallmann. Begrüssungsworte des Rektors der Eidgenössischen Technischen Hochschule. (German) [Welcome words of the Rector of the Swiss Federal Institute of Technology]. Dialectica: International Review of Philosophy of Knowledge, 2(2):91–92, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1948:IDP

[59] F Gonseth. Introduction du président des entretiens. (French) [Introduction of interview chair]. Dialectica: International Review of Philosophy of Knowledge, 2(2):93–94, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1948:LDL

[60] Anonymous. Logic and dialectics — logique et dialectique — Logik und Dialektik: Dialectique et logique. (French) [Dialectics and logic]. Dialectica: International Review of Philosophy of Knowledge, 2(2):95–108, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Beth:1948:RDL

[61] E. W. Beth. Les relations de la dialectique à la logique. (French) [The relationship of dialectics to logic]. Dialectica: International Review of Philosophy of Knowledge, 2(2):109–119, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Beth:1948:PEM

[62] E. W. Beth and F. Gonseth. À propos des exposés de Mm. Ph. Devaux. (French) [About the presentations of Mm. Ph. Devaux]. Dialectica: International Review of Philosophy of Knowledge, 2(2):120–125, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1948:DDDa

[63] Anonymous. Discussion discussion diskussion. Dialectica: International Review of Philosophy of Knowledge, 2(2):126–142, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Carcano:1948:MDM

[64] P. Filiasi Carcano. Metaphysics and dialectics — metaphysique et dialectique — Metaphysik und Dialektik: L’assimilation du relativisme. (French) [Metaphysics and dialectics — metaphysics and dialectics — Metaphysik und Dialektik: The assimilation of relativism]. Dialectica: International Review of Philosophy of Knowledge, 2(2):143–156, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1948:NDP

[65] Anonymous. La nécessité du principe de révisibilité et ses limites de validité. (French) [The need for the principle of revision and its limits of validity]. Dialectica: International Review of Philosophy of Knowledge, 2 (2):157–160, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Castelli:1948:DDP

[66] E. Castelli. La dialectique du péché. (French) [The dialectic of sin]. Dialectica: International Review of Philosophy of Knowledge, 2(2):161–163, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ewing:1948:CDM

[67] A. C. Ewing. Contribution to the discussion on métaphysique et dialectique. (French) [Contribution to the discussion on metaphysics and dialectics]. Dialectica: International Review of Philosophy of Knowledge, 2(2):164–168, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hegel:1948:EPF

[68] Hegel. Être et pensée. (French) [Being and thinking]. Dialectica: International Review of Philosophy of Knowledge, 2(2):169–184, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1948:DDDb

[69] Anonymous. Discussion discussion diskussion. Dialectica: International Review of Philosophy of Knowledge, 2(2):185–198, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bayer:1948:ADE

[70] R. Bayer. Aesthetics and dialectics — esthétique et dialectique — Aesthetik und Dialektik: Esthétique et dialectique. (French) [Aesthetics and dialectics]. Dialectica: International Review of Philosophy of Knowledge, 2(2):199–204, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hainard:1948:DDG

[71] Robert Hainard. Dialectique du global. Dialectica: International Review of Philosophy of Knowledge, 2(2):205–208, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Puyvallee:1948:PCM

[72] R. de Bengy Puyvalleé. À propos de la communication de M. R. Bayer. (French) [With regard to M. R. Bayer]. Dialectica: International Review of Philosophy of Knowledge, 2(2):209–210, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1948:BPA

[73] P. Bernays. Bemerkungen zum Problem der Aesthetik. (German) [Remarks on the problem of aesthetics]. Dialectica: International Review of Philosophy of Knowledge, 2(2):211, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Einleitung:1948:TFD

[74] I. Einleitung. Der Terminus ^{≪}Dialektik^{≫}in politischen Sprachgebrauch, eine signifische Untersuchung. (German) [The term ^{≪}dialectics^{≫}in political language usage, a significant investigation]. Dialectica: International Review of Philosophy of Knowledge, 2(2):213–221, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1948:FDM

[75] Anonymous. Zur Frage des dialektischen Materialismus. (German) [On the question of dialectical materialism]. Dialectica: International Review of Philosophy of Knowledge, 2(2):222–228, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Walter:1948:DNN

[76] Emil J. Walter. Die Dialektik in den Naturwissenschaften nach dem Antidühring. (German) [The dialectic in the natural sciences after Antidühring]. Dialectica: International Review of Philosophy of Knowledge, 2(2):229–247, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Delhorbe:1948:SDS

[77] Florian Delhorbe. Sociology and dialectics — sociologie et dialectique — Soziologie und Dialektik: Images du monde. (French) [Images of the world]. Dialectica: International Review of Philosophy of Knowledge, 2 (2):249–261, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Guillaume:1948:DSD

[78] G. Guillaume. De la dialectique scientifique à la dialectique politique. (French) [From scientific dialectics to political dialectics]. Dialectica: International Review of Philosophy of Knowledge, 2(2):262–266, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gagnebin:1948:KDC

[79] S. Gagnebin. Knowledge and dialectics — connaissance et dialectique — Erkenntnis und Dialektik: l’idonéisme tel que je le comprends. (French) [Idoneism as I understand it]. Dialectica: International Review of Philosophy of Knowledge, 2(2):267–272, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1948:GPO

[80] Paul Bernays. Grundsätzliches zur Philosophie ouverte. (German) [Fundamental to open philosophy]. Dialectica: International Review of Philosophy of Knowledge, 2(2):273–279, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1948:DDDc

[81] Anonymous. Discussion discussion diskussion. Dialectica: International Review of Philosophy of Knowledge, 2(2):280–294, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1948:CFF

[82] F Gonseth. Considérations finales. (French) [Final considerations]. Dialectica: International Review of Philosophy of Knowledge, 2(2):295–303, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bayer:1948:I

[83] R. Bayer. Invitation. Dialectica: International Review of Philosophy of Knowledge, 2(2):304, May 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pauli:1948:EIK

[84] W. Pauli. Editorial: Die Idee der Komplementarität. (German) [The concept of complementarity]. Dialectica: International Review of Philosophy of Knowledge, 2(3–4):307–311, November 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bohr:1948:NCC

[85] Niels Bohr. On the notions of causality and complementarity. Dialectica: International Review of Philosophy of Knowledge, 2(3–4):312–319, November 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Einstein:1948:QMW

[86] A. Einstein. Quanten-mechanik und Wirklichkeit. (German) [Quantum mechanics and reality]. Dialectica: International Review of Philosophy of Knowledge, 2(3–4):320–324, November 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deBroglie:1948:CII

[87] Louis de Broglie. Sur la complémentarité des idées d’individu et de système. (French) [On the complementarity of ideas of the individual and the system]. Dialectica: International Review of Philosophy of Knowledge, 2(3–4):325–330, November 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL http://www.jstor.org/stable/

42963891.

Heisenberg:1948:BGA

[88] Werner Heisenberg. Der Begriff ,,Abgeschlossene Theorie“in der modernen Naturwissenschaft. (German) [The notion of a “closed theory” in modern science]. Dialectica: International Review of Philosophy of Knowledge, 2(3–4):331–336, November 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). English translation in [2236].

Reichenbach:1948:PAQ

[89] Hans Reichenbach. The principle of anomaly in quantum mechanics. Dialectica: International Review of Philosophy of Knowledge, 2(3–4): 337–350, November 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Destouches:1948:QAT

[90] Jean-Louis Destouches. Quelques aspects théoriques de la notion de complémentarité I. la complémentarité de Bohr. (French) [Some theoretical aspects of the notion of complementarity. I. The complementarity of Bohr]. Dialectica: International Review of Philosophy of Knowledge, 2(3–4):351–382, November 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Destouches-Fevrier:1948:MSN

[91] P. Destouches-Fevrier. Manifestations et sens de la notion de complémentarité I. Complémentarité et logique. (French) [Manifestations and meaning of the notion of complementarity I. Complementarity and logic]. Dialectica: International Review of Philosophy of Knowledge, 2(3–4):383–412, November 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1948:RIC

[92] F. Gonseth. Remarque sur l’idée de complémentarité. (French) [Note on the idea of complementarity]. Dialectica: International Review of Philosophy of Knowledge, 2(3–4):413–420, November 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fierz:1948:RSE

[93] M. Fierz. Reviewing studies — études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 2(3–4):421–422, November 1948. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jecklin:1949:HWG

[94] H. Jecklin. Historisches zur Wahrscheinlichkeitsdefinition. (German) [Historical probability definition]. Dialectica: International Review of Philosophy of Knowledge, 3(1–2):5–15, June 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nolfi:1949:WIL

[95] P. Nolfi. Die Wahrscheinlichkeitstheorie im Lichte der dialektischen Philosophie. (German) [Probability theory in the light of dialectical philosophy]. Dialectica: International Review of Philosophy of Knowledge, 3(1–2):16–23, June 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Borel:1949:PCF

[96] Emile Borel. Probabilité et certitude. (French) [Probability and certainty]. Dialectica: International Review of Philosophy of Knowledge, 3 (1–2):24–27, June 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Polya:1949:PRL

[97] G. Pólya. Preliminary remarks on a logic of plausible inference. Dialectica: International Review of Philosophy of Knowledge, 3(1–2):28–35, June 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gini:1949:CMP

[98] Corrado Gini. Concept et mesure de la probabilité. (French) [Concept and measure of probability]. Dialectica: International Review of Philosophy of Knowledge, 3(1–2):36–54, June 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Levy:1949:FDC

[99] Paul Lévy. Les fondements du calcul des probabilités. (French) [The basis of calculating probabilities]. Dialectica: International Review of Philosophy of Knowledge, 3(1–2):55–64, June 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anderson:1949:BGG

[100] O. Anderson. Die Begründung des Gesetzes der grossen Zahlen und die Umkehrung des Theorems von Bernoulli. (German) [The justification of the law of large numbers and the inversion of Bernoulli’s theorem]. Dialectica: International Review of Philosophy of Knowledge, 3(1–2): 65–77, June 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deFinetti:1949:VPF

[101] B. de Finetti. Le vrai et le probable. (French) [The real and the probable]. Dialectica: International Review of Philosophy of Knowledge, 3(1–2): 78–92, June 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Baptist:1949:RP

[102] J. H. Baptist. Le raisonnement probabilitaire. Dialectica: International Review of Philosophy of Knowledge, 3(1–2):93–103, June 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bartlett:1949:PLM

[103] M. S. Bartlett. Probability in logic, mathematics and science. Dialectica: International Review of Philosophy of Knowledge, 3(1–2):104–113, June 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chandrasekhar:1949:BMD

[104] S. Chandrasekhar. Brownian motion, dynamical friction and stellar dynamics. Dialectica: International Review of Philosophy of Knowledge, 3 (1–2):114–126, June 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Feraud:1949:IAI

[105] L. Féraud. Induction amplifiante et inference statistique. (French) [Amplifying induction and statistical inference]. Dialectica: International Review of Philosophy of Knowledge, 3(1–2):127–152, June 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hagstroem:1949:CSF

[106] K.-G. Hagstrœm. Connaissance et stochastique. (French) [Knowledge and stochastics]. Dialectica: International Review of Philosophy of Knowledge, 3(1–2):153–172, June 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonsrth:1949:PPP

[107] F. Gonsrth and Ch. Perelman. Philosophies premieres et philosophie regressive. (French) [Primary philosophy and regressive philosophy]. Dialectica: International Review of Philosophy of Knowledge, 3(3):175–191, September 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lion:1949:SPN

[108] Aline Lion. The social phenomenon: the notion of complementarity and historical causality. Dialectica: International Review of Philosophy of Knowledge, 3(3):192–206, September 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Guggenheimer:1949:ME

[109] H. Guggenheimer. Metaphysik als experimentalwissenschaft. Dialectica: International Review of Philosophy of Knowledge, 3(3):207–218, September 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gex:1949:CSP

[110] Maurice Gex. Caractérologie, science et philosophie. (French) [Characterology, science and philosophy]. Dialectica: International Review of Philosophy of Knowledge, 3(3):219–235, September 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1949:FFL

[111] Adolphe Bernays. Zur Frage des ^{≪}lichtsinns^{≫}. (German) [On the question of ^{≪}sense of light^{≫}]. Dialectica: International Review of Philosophy of Knowledge, 3(3):236–242, September 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1949:RB

[112] Anonymous. Received books. Dialectica: International Review of Philosophy of Knowledge, 3(3):242, September 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Walter:1949:EFE

[113] E. J. Walter. Éditorial. (French) [Editorial]. Dialectica: International Review of Philosophy of Knowledge, 3(4):247–253, December 1949. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1949:A

[114] Anonymous. Avertissement. Dialectica: International Review of Philosophy of Knowledge, 3(4):247, December 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1949:RUI

[115] Anonymous. Die Rolle des Unbewussten im philosophischen Denken. (German) [The role of the unconscious in philosophical thinking]. Dialectica: International Review of Philosophy of Knowledge, 3(4):254–271, December 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Brun:1949:BPI

[116] R. Brun. Über biologische Psychologie ihre wissenschaftstheoretischen Grundlagen, ihre Berechtigung und Leistungsfähigkeit. (German) [On biological psychology, their scientific theories, their justification and efficiency]. Dialectica: International Review of Philosophy of Knowledge, 3(4):272–290, December 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lincke:1949:KGR

[117] Harold Lincke. Zur Kritik der geisteswissenschaftlichen Richtungen in der Tiefenpsychologie. (German) [On the critique of spiritual science in deep psychology]. Dialectica: International Review of Philosophy of Knowledge, 3(4):291–313, December 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kardiner:1949:PSS

[118] Abram Kardiner. Psychodynamics and the social sciences. Dialectica: International Review of Philosophy of Knowledge, 3(4):314–323, December 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1949:QMP

[119] F. Gonseth. La question de la méthode en psychologie. (French) [The question of method in psychology]. Dialectica: International Review of Philosophy of Knowledge, 3(4):324–337, December 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1949:QRE

[120] J.-P. Gonseth. Quelques remarques sur un essai de dialectisation de la conscience. (French) [Some remarks on an essay on the dialectization of consciousness]. Dialectica: International Review of Philosophy of Knowledge, 3(4):337–338, December 1949. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1950:LGM

[121] F. Gonseth. L’épistémologie génétique et la méthodologie dialectique: à propos de l’ouvrage récent de J. Piaget: ^{≪}introduction à l’epistemologie genetique^{≫}. (French) [Genetic epistemology and dialectical methodology: on the recent work of J. Piaget: introduction to genetic epistemology^{≫}]. Dialectica: International Review of Philosophy of Knowledge, 4(1):5–20, March 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vuillemin:1950:DNC

[122] J. Vuillemin. La dialectique négative dans la connaissance et l’existence: (note sur l’épistémologie et la métaphysique de Nikolaï Hartmann et de Jean-Paul Sartre). (French) [The negative dialectic in knowledge and existence: (note on the epistemology and metaphysics of Nikolaı Hartmann and Jean-Paul Sartre)]. Dialectica: International Review of Philosophy of Knowledge, 4(1):21–42, March 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1950:MDB

[123] Paul Bernays. Zur methodischen Diskussion (Bemerkungen zu herrn Perelmans ^{≪}erörterung philosophies premières et philosophie régressive^{≫}). (German) [On methodological discussion (remarks on Mr. Perelman’s ^{≪}discussion on first and regressive philosophies^{≫})]. Dialectica: International Review of Philosophy of Knowledge, 4(1):43–45, March 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bopp:1950:LOM

[124] Léon Bopp. Lettre ouverte à M. Preissmann. (French) [Open letter to M. Preissmann]. Dialectica: International Review of Philosophy of Knowledge, 4(1):46–47, March 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See response [125].

Preissmann:1950:RLM

[125] A. Preissmann. Réponse à la lettre de M. Bopp. (French) [Reply to M. Bopp’s letter]. Dialectica: International Review of Philosophy of Knowledge, 4(1):48, March 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See [124].

Krouner:1950:LJP

[126] F. Kröuner. Zur Logik von J. Piaget. (German) [On the logic of J. Piaget]. Dialectica: International Review of Philosophy of Knowledge, 4 (1):49–64, March 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Frere:1950:FGC

[127] Hubert Frere. De la fine à la grosse complementarity. (French) [From fine to fat complementarity]. Dialectica: International Review of Philosophy of Knowledge, 4(1):65–67, March 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gex:1950:CCN

[128] Maurice Gex. Chronique de caractérologie no 1: la morpho-psychologie du d^{r} Corman. (French) [Chronicle of characterology no. 1: the morphopsychology of Dr. Corman]. Dialectica: International Review of Philosophy of Knowledge, 4(1):68–76, March 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1950:RS

[129] Anonymous. Reviewing studies. Dialectica: International Review of Philosophy of Knowledge, 4(1):77–78, March 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1950:EFEa

[130] Anonymous. Éditorial. (French) [Editorial]. Dialectica: International Review of Philosophy of Knowledge, 4(2):83–84, June 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gagnebin:1950:MCG

[131] Samuel Gagnebin. En marge d’une carrière de géologue. (French) [On the sidelines of a career as a geologist]. Dialectica: International Review of Philosophy of Knowledge, 4(2):85–120, June 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gagnebin:1950:LDP

[132] Elie Gagnebin. L’absurdité du pessimisme. (French) [The absurdity of pessimism]. Dialectica: International Review of Philosophy of Knowledge, 4(2):121–132, June 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gagnebin:1950:EFN

[133] Elie Gagnebin. Essai sur la finalité de la nature. (French) [Essay on the purpose of nature]. Dialectica: International Review of Philosophy of Knowledge, 4(2):133–143, June 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gagnebin:1950:ISJ

[134] Elie Gagnebin. Introduction à un spectacle de Jean Cocteau: donné par la compagnie Pitoëff. (French) [Introduction to a show by Jean Cocteau: given by the company Pitoëff]. Dialectica: International Review of Philosophy of Knowledge, 4(2):144–147, June 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1950:UDS

[135] Ferdinand Gonseth. De l’unité du savoir. (French) [From unity of knowledge]. Dialectica: International Review of Philosophy of Knowledge, 4(2): 148–157, June 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rivier:1950:LVU

[136] W. Rivier. L’apparition de la vie dans l’univers serait-elle compatible avec le calcul des probabilités?. (French) [Would the appearance of life in the Universe be compatible with the calculation of probabilities?]. Dialectica: International Review of Philosophy of Knowledge, 4(2):158–162, June 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1950:PD

[137] Anonymous. Philosophy of the day. Dialectica: International Review of Philosophy of Knowledge, 4(2):163, June 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1950:RBL

[138] Anonymous. Received books livres reçus eingelaufene bücher. Dialectica: International Review of Philosophy of Knowledge, 4(2):164, June 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1950:EFEb

[139] Anonymous. Éditorial. (French) [Editorial]. Dialectica: International Review of Philosophy of Knowledge, 4(3):167, September 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Monnier:1950:LFP

[140] Marcel Monnier. L’organisation des fonctions psychiques à la lueur des données neurophysiologiques. (French) [The organization of psychic functions in the light of neurophysiological data]. Dialectica: International Review of Philosophy of Knowledge, 4(3):169–179, September 1950. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Monnieh:1950:CEF

[141] Marcel Monnieh. Contribution à l’étude des fonctions sensorielles et des perceptions: Aufbau und Bedeutung der Sinnesfunktionen. (French) [Contribution to the study of sensory functions and perceptions: Aufbau und Bedeutung der Sinnesfunktionen]. Dialectica: International Review of Philosophy of Knowledge, 4(3):180–185, September 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wiener:1950:CPI

[142] Norber Wiener. Contributions aux problèmes de l’intégration sensorio motrice et de la cybernétique: some maxims for biologists and psychologists. (French) [Contributions to the problems of sensorio-motor integration and cybernetics: some maxims for biologists and psychologists]. Dialectica: International Review of Philosophy of Knowledge, 4(3):186– 191, September 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

McCulloch:1950:WMH

[143] Warren S. McCulloch. Why the mind is in the head? Dialectica: International Review of Philosophy of Knowledge, 4(3):192–205, September 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Walter:1950:EMP

[144] W. Grey Walter. An electro-mechanical a propos ^{≪}animal^{≫}. Dialectica: International Review of Philosophy of Knowledge, 4(3):206–213, September 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1950:LRS

[145] Ferdinan Gonseth. L’organisation des robots sensori-moteurs et l’intuition géométrique. (French) [The organization of sensory-motor robots and geometric intuition]. Dialectica: International Review of Philosophy of Knowledge, 4(3):214–223, September 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sherrington:1950:CAP

[146] Charles Sherrington. Contributions au problème du substratum des activitiés intellectuelles et de la mémoire: le cerveau et son mécanisme. (French) [Contributions to the problem of the substratum of intellectual activities and memory: the brain and its mechanism]. Dialectica: International Review of Philosophy of Knowledge, 4(3):224–236, September 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Brazier:1950:NBI

[147] Mary A. B. Brazier. A neuronal basis for ideas. Dialectica: International Review of Philosophy of Knowledge, 4(3):237–242, September 1950. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Herrick:1950:CPP

[148] C. Judson Herrick. Contributions aux problèmes de la personnalité et du langage au service de la vie sociale: mechanisms of nervous adjustment. (French) [Contributions to the problems of personality and language in the service of social life: mechanisms of nervous adjustment]. Dialectica: International Review of Philosophy of Knowledge, 4(3):243–247, September 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kluyer:1950:MEN

[149] Heinrich Klüyer. The method of equivalent and non-equivalent stimuli in the study of behaviour and personality. Dialectica: International Review of Philosophy of Knowledge, 4(3):248–254, September 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gemelli:1950:SDL

[150] Agostino Gemelli. La strutturazione del linguaggio studiata mediante l’analisi elettroacustica. (Italian) [The structuring of language studied by electro-acoustic analysis]. Dialectica: International Review of Philosophy of Knowledge, 4(3):255–268, September 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vonNiggli:1950:MVM

[151] Paul von Niggli. Zum methodischen Vorgehen in der mineralogischen Wissenschaft. (German) [The methodological approach in mineralogical science]. Dialectica: International Review of Philosophy of Knowledge, 4(4):271–283, December 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fiala:1950:MMF

[152] F. Fiala. De la méthode en minéralogie. (French) [On the mineralogy method]. Dialectica: International Review of Philosophy of Knowledge, 4(4):284–286, December 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Piaget:1950:EGM

[153] Jean Piaget. Epistémologie génétique et méthodologie dialectique II. (French) [Genetic epistemology and dialectical methodology II]. Dialectica: International Review of Philosophy of Knowledge, 4(4):287–295, December 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1950:EGM

[154] Ferdinand Gonseth. Epistémologie génénetique et méthodologie dialectique. III: réponse à M. Piaget. (French) [Genetic epistemology and dialectical methodology. III: Response to M. Piaget]. Dialectica: International Review of Philosophy of Knowledge, 4(4):296–304, December 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vonMiescher:1950:EFG

[155] Karl von Miescher. Elemente der forschung. (German) [Elements of research]. Dialectica: International Review of Philosophy of Knowledge, 4(4):305–321, December 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Konig:1950:PAS

[156] Hans König. Sur la possibilité d’attribuer une signification générale à certaines règles des sciences expérimentales. (French) [On the possibility of attributing a general significance to certain rules of the experimental sciences]. Dialectica: International Review of Philosophy of Knowledge, 4(4):322–330, December 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vonGrutter:1950:EUE

[157] Felix von Grütter. Elementarteilchen, ein überblick über experiment und theorie. (German) [Elementary particles, an overview of experiment and theory]. Dialectica: International Review of Philosophy of Knowledge, 4(4):331–339, December 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schafroth:1950:BNB

[158] M. R. Schafroth. Bibliographic notes: bibliographische notizen. Dialectica: International Review of Philosophy of Knowledge, 4(4):340–341, December 1950. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1950:RBR

[159] Anonymous. Receieved books: receieved books. Dialectica: International Review of Philosophy of Knowledge, 4(4):341–342, December 1950. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mitscherlich:1951:VV

[160] A. Mitscherlich and P. Rossier. Vorwort: Vorwort. Dialectica: International Review of Philosophy of Knowledge, 5(1):5–7, March 1951. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jung:1951:GPG

[161] C. G. Jung. Grundfragen der Psychotherapie. (German) [Basic questions of psychotherapy]. Dialectica: International Review of Philosophy of Knowledge, 5(1):8–24, March 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vWeizsaucker:1951:MLG

[162] Viktor v. Weizsaucker. Medizin und Logik. (German) [Medicine and logic]. Dialectica: International Review of Philosophy of Knowledge, 5 (1):25–58, March 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ey:1951:LMP

[163] Henri Ey. L’évolution de la médecine et la philosophie. (French) [The evolution of medicine and philosophy]. Dialectica: International Review of Philosophy of Knowledge, 5(1):59–83, March 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bally:1951:PSM

[164] G. Bally, G. Bally, H. Ey, C. G. Jung, A. Mitscherlich, P.-H. Rossier, J. Ruesch, and W. Voneizsaucker. Philosophische Strömungen in der modernen Medizin. (German) [Philosophical trends in modern medicine]. Dialectica: International Review of Philosophy of Knowledge, 5(1):84–96, March 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ruesch:1951:PWS

[165] Jurgen Ruesch. Part and whole: the sociopsychological and psychosomatic approach to disease. Dialectica: International Review of Philosophy of Knowledge, 5(2):99–125, June 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mitscherlich:1951:WBF

[166] Alexander Mitscherlich. Die wechselseitige Beeinflussung des Freiheits und Krankheitsbegriffs in einer anthropologischen Heilkunde. (German) [The reciprocal influence of liberty and the concept of disease in an anthropological healing]. Dialectica: International Review of Philosophy of Knowledge, 5(2):126–145, June 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1951:NE

[167] F. Gonseth. Note explicative. Dialectica: International Review of Philosophy of Knowledge, 5(2):147, June 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gattegno:1951:FEF

[168] C. Gattegno. Fondements de l’education. (French) [Foundations of education]. Dialectica: International Review of Philosophy of Knowledge, 5(2):148–152, June 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Read:1951:ETA

[169] Herbert Read. Education through art a re-statement. Dialectica: International Review of Philosophy of Knowledge, 5(2):153–165, June 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1951:SIS

[170] Anonymous. La santé intégrale ses implications scienttfiques et pédagogiques devant le probléme de la civilisation contemporaine. (French) [Integral health: its scientific and pedagogical implications for the problem of contemporary civilization]. Dialectica: International Review of Philosophy of Knowledge, 5(2):166–180, June 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1951:CNC

[171] Philippe Müller. Cultures nationales et culture mondiale. (French) [National cultures and global culture]. Dialectica: International Review of Philosophy of Knowledge, 5(2):181–187, June 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

DeCoster:1951:SEF

[172] S. De Coster. Sociologie et éducation. (French) [Sociology and education]. Dialectica: International Review of Philosophy of Knowledge, 5 (2):188–193, June 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gattegno:1951:AM

[173] C. Gattegno. The awareness of man. Dialectica: International Review of Philosophy of Knowledge, 5(2):194–205, June 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Polin:1951:PEM

[174] Raymond Polin. Principes d’une éducation morale cynique. (French) [Principles of a cynical moral education]. Dialectica: International Review of Philosophy of Knowledge, 5(2):206–212, June 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gex:1951:CCC

[175] Maurice Gex. Chronique de caractérologie n^{∘} 2: Caractérologie purement psychologique. (French) [Chronology of characterology n^{∘} 2: Purely psychological characterology]. Dialectica: International Review of Philosophy of Knowledge, 5(2):213–222, June 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gattegno:1951:MAE

[176] C. Gattegno. Manifeste d’action de l’Êcole Normale Internationale. (French) [Manifesto of the International Normal School]. Dialectica: International Review of Philosophy of Knowledge, 5(2):223–226, June 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Paulus:1951:E

[177] Jean Paulus. Editorial. Dialectica: International Review of Philosophy of Knowledge, 5(3–4):231–238, December 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Alexander:1951:TFD

[178] Franz Alexander. Three fundamental dynamic principles of the mental apparatus and of the behavior of living organisms. Dialectica: International Review of Philosophy of Knowledge, 5(3–4):239–245, December 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

French:1951:IP

[179] Thomas M. French. The integrative process. Dialectica: International Review of Philosophy of Knowledge, 5(3–4):246–256, December 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Maslow:1951:HNP

[180] A. H. Maslow. Higher needs and personality. Dialectica: International Review of Philosophy of Knowledge, 5(3–4):257–265, December 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Murray:1951:SBP

[181] Henry A. Murray. Some basic psychological assumptions and conceptions. Dialectica: International Review of Philosophy of Knowledge, 5 (3–4):266–292, December 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rosenzweig:1951:IPT

[182] Saul Rosenzweig. Idiodynamics in personality theory with special reference to projective methods. Dialectica: International Review of Philosophy of Knowledge, 5(3–4):293–311, December 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Buhler:1951:MM

[183] Charlotte Bühler. Maturation and motivation. Dialectica: International Review of Philosophy of Knowledge, 5(3–4):312–361, December 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Duyckaerts:1951:LOS

[184] Fr. Duyckaerts. L’anthropologie d’Oswald Schwarz. (French) [The anthropology of Oswald Schwarz]. Dialectica: International Review of Philosophy of Knowledge, 5(3–4):362–375, December 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

DeGreeff:1951:MRI

[185] E. De Greeff. Les modes de rattachements instinctifs, fonctions incorruptibles. (French) [The instinctive modes of attachment, incorruptible functions]. Dialectica: International Review of Philosophy of Knowledge, 5(3–4):376–392, December 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Paulus:1951:PHL

[186] Jean Paulus. Philosophy of human life, some of the conditions of a happy maturity and old age. Dialectica: International Review of Philosophy of Knowledge, 5(3–4):393–401, December 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wallon:1951:LDP

[187] Henri Wallon. L’évolution dialectique de la personnalité. (French) [The dialectical evolution of personality]. Dialectica: International Review of Philosophy of Knowledge, 5(3–4):402–412, December 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Delay:1951:RJF

[188] Jean Delay. Ribot et le jacksonisme. (French) [Ribot and Jacksonism]. Dialectica: International Review of Philosophy of Knowledge, 5(3–4): 413–444, December 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fontes:1951:PSA

[189] Vítor Fontes and René Zazzo. Le psychisme et les structures anatomiques. (French) [The psyche and anatomical structures]. Dialectica: International Review of Philosophy of Knowledge, 5(3–4):445–470, December 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1951:TMF

[190] Anonymous. Table des matières. (French) [Table of contents]. Dialectica: International Review of Philosophy of Knowledge, 5(3–4):471, December 1951. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1952:E

[191] F. Gonseth. Eaditorial. Dialectica: International Review of Philosophy of Knowledge, 6(1):5–8, March 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1952:MSP

[192] F. Gonseth. Motivation et structure d’une philosophie ouverte. (French) [Motivation and structure of an open philosophy]. Dialectica: International Review of Philosophy of Knowledge, 6(1):9–29, March 1952. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Isaye:1952:PMF

[193] Gaston Isaye. Le privilège de la mètaphysique. (French) [The privilege of metaphysics]. Dialectica: International Review of Philosophy of Knowledge, 6(1):30–52, March 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1952:RAR

[194] F. Gonseth. Réponse au R. P. Isaye. (French) [Answer to R. P. Isaye]. Dialectica: International Review of Philosophy of Knowledge, 6(1):53–66, March 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sciacca:1952:AMI

[195] M. F. Sciacca. Autonomie de la métaphysique et intériorité objective.. (French) [Autonomy of metaphysics and objective interiority]. Dialectica: International Review of Philosophy of Knowledge, 6(1):67–80, March 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sottiaux:1952:LMG

[196] Ie R. P. Sottiaux. Lettre à M. Gonseth sur la fidélité inébranlable. (French) [Letter to M. Gonseth on unwavering fidelity]. Dialectica: International Review of Philosophy of Knowledge, 6(1):81–91, March 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Perelman:1952:RMB

[197] Ch. Perelman. Réponse à M. Bernays. (French) [Answer to M. Bernays]. Dialectica: International Review of Philosophy of Knowledge, 6(1):92– 95, March 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1952:XCI

[198] Anonymous. XI^{e} congrès international de philosophie Bruxelles, 20–26 août 1953. (French) [XI-th International Congress of Philosophy Brussels, 20–26 August 1953]. Dialectica: International Review of Philosophy of Knowledge, 6(1):96–100, March 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1952:PDF

[199] F. Gonseth. Le principe de dualité. (French) [The principle of duality]. Dialectica: International Review of Philosophy of Knowledge, 6(2): 103–106, June 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1952:SOL

[200] Ferdinand Gonseth. A. Séance d’ouverture la loi du dialogue. (French) [AT. Session opening the law of dialogue]. Dialectica: International Review of Philosophy of Knowledge, 6(2):107–119, June 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gagnebin:1952:TEP

[201] S. Gagnebin. Théorie et expérience le probléme. (French) [Theory and experience the problem]. Dialectica: International Review of Philosophy of Knowledge, 6(2):120–129, June 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1952:DGZ

[202] Paul Bernays. Dritte Gespräche von Zürich. (German) [Third talks of Zurich]. Dialectica: International Review of Philosophy of Knowledge, 6 (2):130–137, June 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vanBernays:1952:GEI

[203] P. van Bernays. Gesichtspunkte der Erkenntnistheorie (insbesondere zur Auseinandersetzung mit der kantischen Erkenntnistheorie): Thesen zu den dritten Gesprächen von Zürich. (German) [Aspects of the theory of knowledge (in particular, the examination of the Kantian theory of knowledge): Theses on the third conversations of Zurich]. Dialectica: International Review of Philosophy of Knowledge, 6(2):137–140, June 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pauli:1952:TEG

[204] Wolfgang Pauli. Theorie und Experiment. (German) [Theory and experiment]. Dialectica: International Review of Philosophy of Knowledge, 6 (2):141–142, June 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1952:TEF

[205] F. Gonseth. Théorie et expérience. (French) [Theory and experience]. Dialectica: International Review of Philosophy of Knowledge, 6(2):143–144, June 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1952:DPF

[206] Anonymous. Discussion préliminaire. (French) [Preliminary discussion]. Dialectica: International Review of Philosophy of Knowledge, 6(2):145– 165, June 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vanBernays:1952:BAH

[207] P. van Bernays. Bemerkungen zu der Arbeit von E. Harris. (German) [Remarks on the work of E. Harris]. Dialectica: International Review of Philosophy of Knowledge, 6(2):166, June 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Harris:1952:SRC

[208] Errol E. Harris. Some recent criticisms of Berkeley. Dialectica: International Review of Philosophy of Knowledge, 6(2):167–185, June 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Guggenheimer:1952:TEG

[209] E. Guggenheimer and H. Guggenheimer. Technische epistemologie. (German) [Technical epistemology]. Dialectica: International Review of Philosophy of Knowledge, 6(2):186–190, June 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Filiasi-Carcano:1952:CPD

[210] Paolo Filiasi-Carcano. C. Le principe de dualité en complément à la discussion générale: remarques historiques sur les rapports entre la théorie et l’expérience. (French) [C. the principle of duality in addition to the general discussion: historical remarks on the relationship between theory and experience]. Dialectica: International Review of Philosophy of Knowledge, 6(3):195–209, September 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vanMoser:1952:TEP

[211] Simon van Moser. Theorie und Erfahrung bei Platon und Aristoteles. (German) [Theory and experience in Plato and Aristotle]. Dialectica: International Review of Philosophy of Knowledge, 6(3):210–221, September 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Guggenheimer:1952:ETF

[212] H. Guggenheimer. Expérience et théorie. (French) [Experience and theory]. Dialectica: International Review of Philosophy of Knowledge, 6 (3):222–223, September 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hainard:1952:TEF

[213] Robert Hainard. Théorie et expérience. (French) [Theory and experience]. Dialectica: International Review of Philosophy of Knowledge, 6 (3):224–225, September 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

leDubarle:1952:DPD

[214] R. P. le Dubarle. D. Le principe de dualité en physique: est-il possible d’axiomatiser la physique?. (French) [D. the principle of duality in physics: is it possible to axiomatize physics?]. Dialectica: International Review of Philosophy of Knowledge, 6(3):226–253, September 1952. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1952:IJD

[215] Anonymous. Intervention de J.-L. Destouches à la communication du R. P. Dubarle. (French) [Response by J.-L. Destouches to the communication of R. P. Dubarle]. Dialectica: International Review of Philosophy of Knowledge, 6(3):254–257, September 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1952:DNR

[216] Anonymous. Diskussion nach dem referat Dubarle. (German) [Discussion on the reference Dubarle]. Dialectica: International Review of Philosophy of Knowledge, 6(3):258–259, September 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rossel:1952:TEF

[217] J. Rossel. Théorie et expérience. (French) [Theory and experience]. Dialectica: International Review of Philosophy of Knowledge, 6(3):260–263, September 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Paris:1952:ETP

[218] Carlos Paris. Expérience et théorie en physique. (French) [Experience and theory in physics]. Dialectica: International Review of Philosophy of Knowledge, 6(3):264–265, September 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Clay:1952:REE

[219] J. Clay. Le rapport entre l’expérience et la théorie. (French) [The relationship between experience and theory]. Dialectica: International Review of Philosophy of Knowledge, 6(3):266–269, September 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vanMiescher:1952:VBW

[220] K. van Miescher. Varia: die Bedeutung der Wertung in der Gegensätzlichkelt. (German) [Varia: the meaning of the valuation in the counter-terrorist]. Dialectica: International Review of Philosophy of Knowledge, 6(3):270–283, September 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Echarri:1952:ETN

[221] Jacques Echarri. Expérience et théorie niveaux d’experience. (French) [Experience and theory levels of experience]. Dialectica: International Review of Philosophy of Knowledge, 6(4):291–304, December 1952. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bouligand:1952:PPM

[222] G. Bouligand. La pensée prospective en mathématiques. (French) [Prospective thinking in mathematics]. Dialectica: International Review of Philosophy of Knowledge, 6(4):305–308, December 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Apeary:1952:MSE

[223] R. Apeary. Les mathématiques sont-elles une théorie pure?. (French) [Is mathematics a pure theory?]. Dialectica: International Review of Philosophy of Knowledge, 6(4):309–310, December 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hirsch:1952:TEM

[224] Guy Hirsch. Theorie et expérience en mathématiques. (French) [Theory and experience in mathematics]. Dialectica: International Review of Philosophy of Knowledge, 6(4):311–326, December 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vanReidemeister:1952:LLR

[225] Kurt van Reidemeister. Zur Logik der Lehre vom Raum. (German) [On the logic of the doctrine of space]. Dialectica: International Review of Philosophy of Knowledge, 6(4):327–342, December 1952. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vanDingler:1952:VEO

[226] Hugo van Dingler. Varia: Empirismus und Operationismus: die beiden Wissenschaftslehren e-Lehre und o-Lehre in ihrem Verhältnis. (German)

[Varia: Empiricism and operationism: the two doctrines of science edoctrine and o-doctrine in their relationship]. Dialectica: International Review of Philosophy of Knowledge, 6(4):343–376, December 1952. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vanAltwegg:1953:TEG

[227] M. van Altwegg. Theorie und Erfahrung. (German) [Theory and experience]. Dialectica: International Review of Philosophy of Knowledge, 7 (1):5–21, March 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Husson:1953:TES

[228] L. Husson. Théorie et expérience dans les sciences juridiques. (French) [Theory and experience in the legal sciences]. Dialectica: International Review of Philosophy of Knowledge, 7(1):22–31, March 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vanBernays:1953:DRH

[229] Paul van Bernays. Diskussionsbemerkung zum referat von herrn Husson. (German) [Discussions on the reference by Mr. Husson]. Dialectica: International Review of Philosophy of Knowledge, 7(1):32–33, March 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wittenberg:1953:QRD

[230] Alexandre Wittenberg. Quelques remarques sur la dualité théorie– expérience en morale. (French) [Some remarks on the theory–experience duality in morals]. Dialectica: International Review of Philosophy of Knowledge, 7(1):34–38, March 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1953:VRE

[231] F. Gonseth. Varia: remarques sur un exposé de H. Dingler. (French) [Varia: remarks on a presentation by H. Dingler]. Dialectica: International Review of Philosophy of Knowledge, 7(1):39–51, March 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gagnebin:1953:NBH

[232] S. Gagnebin. Notes bibliographiques: Histoire de la mécanique par René Dugas. Dialectica: International Review of Philosophy of Knowledge, 7 (1):52–54, March 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361

(electronic). URL http://www.jstor.org/stable/42968355.

Kroner:1953:BAE

[233] F. Kröner. Buchbesprechung: Albert Einstein, philosopher-scientist, Herausgegeben von P. A. Schilpp. The Library of Living Philosophers, Inc., Evanston, Illinois, 1949. Dialectica: International Review of Philosophy of Knowledge, 7(1):61–69, March 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1953:NGS

[234] G. H. Müller. Nelson Goodman: the structure of appearance. Dialectica: International Review of Philosophy of Knowledge, 7(1):70–77, March 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1953:KDP

[235] Gert Heinz Müller. Karl Dürr: the propositional logic of Boethius (with an appendix by Norman M. Martin). Dialectica: International Review of Philosophy of Knowledge, 7(1):78–83, March 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wittenberg:1953:HBC

[236] A. Wittenberg. Haskell B. Curry: outlines of a formalist philosophy of mathematics. Dialectica: International Review of Philosophy of Knowledge, 7(1):84–89, March 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1953:SSI

[237] Anonymous. Science et société — impact. (French) [Science and society — impact]. Dialectica: International Review of Philosophy of Knowledge, 7(1):90–93, March 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mebville:1953:CCD

[238] Mm. H. Mébville, D. Christoff, S. Gagnebin, and Et F. Gonseth. Cohérence et concordance: discussion: sur le principe formel de la pensée. (French) [Coherence and concordance: discussion: on the formal principle of thought]. Dialectica: International Review of Philosophy of Knowledge, 7(2):95–98, June 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mieville:1953:CFI

[239] Henri-L. Miéville. Des constituants formels de l’idée de vérité et de leur signification ontologique. (French) [Formal constituents of the idea of truth and their ontological meaning]. Dialectica: International Review of Philosophy of Knowledge, 7(2):99–140, June 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Christoff:1953:PFP

[240] D. Christoff. Pensée formelle et pensée dialecticienne. (French) [Formal thought and dialectical thought]. Dialectica: International Review of Philosophy of Knowledge, 7(2):141–151, June 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gagnebin:1953:BRP

[241] S. Gagnebin. Brèves réflexions sur le principe d’identité et la norme de concordance 1: résponse à M. Miéville. (French) [Brief reflections on the principle of identity and the standard of concordance 1: Response to M. Miéville]. Dialectica: International Review of Philosophy of Knowledge, 7(2):152–160, June 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1953:RMF

[242] Anonymous. Réponse de M. F. Gonseth à M. H. Miéville. (French) [Response of M. F. Gonseth to M. H. Miéville]. Dialectica: International Review of Philosophy of Knowledge, 7(2):161–175, June 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Curry:1953:TE

[243] Haskell B. Curry. Theory and experience. Dialectica: International Review of Philosophy of Knowledge, 7(2):176–178, June 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1953:ASM

[244] Gert Heinz Müller. Andreas Speiser: Die mathematische Denkweise. (German) [Andreas Speiser: Mathematical thinking]. Dialectica: International Review of Philosophy of Knowledge, 7(2):179–182, June 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1953:CIPa

[245] Anonymous. Congrés international de philosophie des sciences Zurich, automne 1954. (French) [International Congress of Philosophy of Science Zurich, Autumn 1954]. Dialectica: International Review of Philosophy of Knowledge, 7(2):183–184, June 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1953:IKPa

[246] Anonymous. Internationaler Kongress für Philosophie der Wissenschaften, Zürich, Herbst 1954. (German) [International Congress of Philosophy of Science, Zurich, Fall 1954]. Dialectica: International Review of Philosophy of Knowledge, 7(2):185–186, June 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1953:ICPa

[247] Anonymous. International Congress for the Philosophy of Science Zurich, Autumn 1954. Dialectica: International Review of Philosophy of Knowledge, 7(2):187–188, June 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

McKeon:1953:PF

[248] Richard McKeon. Process and function. Dialectica: International Review of Philosophy of Knowledge, 7(3):191–227, September 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

McKeon:1953:APF

[249] Richard McKeon. Abstract of process and function. Dialectica: International Review of Philosophy of Knowledge, 7(3):228–231, September 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wittenberg:1953:APM

[250] Alexandre Wittenberg. Über adäquate Problemstellung in der mathematischen Grundlagenforschung. (German) [On adequate problem solving in basic mathematical research]. Dialectica: International Review of Philosophy of Knowledge, 7(3):232–254, September 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See remarks [280, 281].

Anonymous:1953:RCLa

[251] Anonymous. Reviews — chronique — Literaturblatt: Zur Methodologie der Physik und zum Standpunkt von J.-L. Destouches. (German) [On the methodology of physics and on the standpoint of J.-L. Destouches]. Dialectica: International Review of Philosophy of Knowledge, 7(3):255– 275, September 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1953:FII

[252] Anonymous. Fondation de l’institut international de philosophie politique. (French) [Foundation of the International Institute of Political Philosophy]. Dialectica: International Review of Philosophy of Knowledge, 7(3):276–277, September 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1953:CIPb

[253] Anonymous. Congrès international de philosophie des sciences Zurich, automne 1954. (French) [International Congress of Philosophy of Science Zurich, Autumn 1954]. Dialectica: International Review of Philosophy of Knowledge, 7(3):278–279, September 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1953:IKPb

[254] Anonymous. Internationaler Kongress für Philosophie der Wissenschaften, Zürich, Herbst 1954. (German) [International Congress of Philosophy of Science, Zurich, Fall 1954]. Dialectica: International Review of Philosophy of Knowledge, 7(3):280–281, September 1953. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1953:ICPb

[255] Anonymous. International Congress for the Philosophy of Science, Zurich, Autumn 1954. Dialectica: International Review of Philosophy of Knowledge, 7(3):282–283, September 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vonSchottlaender:1953:SZM

[256] Prof. Rudolf von Schottlaender. Die spannung zwischen metaphysik und erfahrung meditation über die vormittage des brüsseler philosophenkongresses. (German) [The tension between metaphysics and experience meditation on the premiere of the Brussels Philosopher’s Congress]. Dialectica: International Review of Philosophy of Knowledge, 7(4):287–302, December 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1953:QEZ

[257] Prof. F. Gonseth. Quatrièmes entretiens de Zurich: Conférence d’ouverture. (French) [Fourth interviews of Zurich: Opening conference]. Dialectica: International Review of Philosophy of Knowledge, 7(4):303– 317, December 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vonBernays:1953:ER

[258] Prof. P. von Bernays. Einleitendes referat. Dialectica: International Review of Philosophy of Knowledge, 7(4):318–321, December 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vonJecklin:1953:ER

[259] Prof. H. von Jecklin. Einleitendes referat. Dialectica: International Review of Philosophy of Knowledge, 7(4):322–325, December 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vonNolfi:1953:ERG

[260] Dr. Padrot von Nolfi. Einleitendes referat. (German) [Introductory statement]. Dialectica: International Review of Philosophy of Knowledge, 7 (4):326–330, December 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vonErismann:1953:WIS

[261] Prof. Th. von Erismann. Wahrscheinlichkeit im Sein und Denken. (German) [Probability in being and thinking]. Dialectica: International Review of Philosophy of Knowledge, 7(4):331–346, December 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nolfi:1953:OUG

[262] Padrot Nolfi. Über ordnung und unordnung. (German) [Superordinate and disorder]. Dialectica: International Review of Philosophy of Knowledge, 7(4):347–348, December 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1953:RCLb

[263] Anonymous. Reviews — chronique — Literaturblatt: reviews chronique Literaturblatt. Dialectica: International Review of Philosophy of Knowledge, 7(4):349–370, December 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1953:CIPc

[264] Anonymous. Congrès international de philosophie des sciences, Zurich, août 1954 2e congrès international de U.I.P.S.. (French) [International Congress of Philosophy of Science, Zurich, August 1954 2nd International Congress of U.I.P.S.]. Dialectica: International Review of Philosophy of Knowledge, 7(4):371–373, December 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kongress-Vorstand:1953:IKP

[265] Der Kongress-Vorstand. Internationaler Kongress für Philosophie der Wissenschaften, Zürich, August 1954 zweiter internationaler Kongress der U.I.P.S.. (German) [International Congress of Philosophy of Science, Zurich, August, 1954, Second International Congress of U.I.P.S.]. Dialectica: International Review of Philosophy of Knowledge, 7(4):374–376, December 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1953:ICPc

[266] Anonymous. International Congress for the Philosophy of Science Zurich, August 1954 Second International Congress of the U.I.P.S. Dialectica: International Review of Philosophy of Knowledge, 7(4):377–379, December 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1953:RBR

[267] Anonymous. Received books: received books. Dialectica: International Review of Philosophy of Knowledge, 7(4):380–384, December 1953. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Piaget:1954:ITE

[268] Jean Piaget. Inconditionnés transcendantaux et epistémologie génétique. (French) [Transcendental unconditioned and genetic epistemology]. Dialectica: International Review of Philosophy of Knowledge, 8(1):5–13, March 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mieville:1954:RQC

[269] Henri-L. Miéville. Réponse à quelques critiques. (French) [Reply to some criticism]. Dialectica: International Review of Philosophy of Knowledge, 8(1):14–30, March 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1954:FDC

[270] F. Gonseth. Sur le fondement du calcul des probabilités. II: Zur Frage der Grundlagen der Wahrscheinlichkeitsrechnung II: remarques préliminaires. (French) [On the basis of the calculation of probabilities. II: preliminary remarks]. Dialectica: International Review of Philosophy of Knowledge, 8(1):31–36, March 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1954:AWG

[271] Anonymous. Zur Axiomatik der Wahrscheinlichkeitsrechnung. (German) [The axiomatics of probability calculation]. Dialectica: International Review of Philosophy of Knowledge, 8(1):37–47, March 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Richter:1954:BWG

[272] Prof. Dr. Hans Richter. Zur Begründung der Wahrscheinlichkeitsrechnung. (German) [On justification of probability calculation]. Dialectica: International Review of Philosophy of Knowledge, 8(1):48–77, March 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1954:RMB

[273] Gert Heinz Müller. Reviews: I. M. Bocheński: Ancient Formal Logic. Dialectica: International Review of Philosophy of Knowledge, 8(1):78–80, March 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1954:EMT

[274] Gert Heinz Müller. Ernest A. Moody: truth and consequence in medieval logic. Dialectica: International Review of Philosophy of Knowledge, 8(1): 81–83, March 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1954:HWP

[275] Gert H. Müller. Hermann Wein: das Problem des Relativismus. (Philosophie im Übergang zur Anthropologie). (German) [Hermann Wein: the problem of relativism. (Philosophy in transition to anthropology)]. Dialectica: International Review of Philosophy of Knowledge, 8(1):84–87, March 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sehvais:1954:EDJ

[276] W. Sehvais. Éditorial du journal de la société belge de professeurs de mathématiques. (French) [Editorial on the Journal of the Belgian Society of Mathematics Teachers]. Dialectica: International Review of Philosophy of Knowledge, 8(1):88–91, March 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mahalanobis:1954:FS

[277] P. C. Mahalanobis. The foundations of statistics. Dialectica: International Review of Philosophy of Knowledge, 8(2):95–111, June 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pauli:1954:WPG

[278] Wolfgang Pauli. Wahrscheinlichkeit und Physik. (German) [Probability and physics]. Dialectica: International Review of Philosophy of Knowledge, 8(2):112–118; Diskussion 118–124, June 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1954:APT

[279] Anonymous. Sur un aspect paradoxal de la theorie des probabilités. (French) [On a paradoxical aspect of probability theory]. Dialectica: International Review of Philosophy of Knowledge, 8(2):125–144, June 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Guggenheimer:1954:QRC

[280] H. Guggenheimer. Quelques remarques concernant l’article de M. Alexandre Wittenberg: Über adäquate Problemstellung in der mathematischen Grundlagenforschung. (French) [A few remarks concerning M. Alexandre Wittenberg’s article]. Dialectica: International Review of Philosophy of Knowledge, 8(2):145–146, June 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See [250].

Bernays:1954:BBA

[281] P. Bernays. Bemerkungen zu der Betrachtung von Alexander Wittenberg: Über adäquate Problemstellung in der mathematischen Grundlagenforschung. (German) [Remarks on the consideration of Alexander Wittenberg: On Adequate Problems in Mathematical Fundamental Research]. Dialectica: International Review of Philosophy of Knowledge, 8 (2):147–151, June 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See [250].

Wittenberg:1954:APM

[282] Alexander Wittenberg. Über adäquate Problemstellung in der mathematischen Grundlagenforschung. Eine Antwort. (German) [On adequate problem solving in basic mathematical research. An answer]. Dialectica: International Review of Philosophy of Knowledge, 8(2):152–157, June 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1954:NBP

[283] Anonymous. Note bibliographique: philosophie néo-scolastique et philosophie ouverte. (French) [Bibliographical note: Neo-scholastic philosophy and open philosophy]. Dialectica: International Review of Philosophy of Knowledge, 8(2):158–167, June 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deLHorbe:1954:EPN

[284] F. de L’Horbe. Événement philosophique: un nouveau discours de la méthode. (French) [Philosophical event: a new discourse of the method].

Dialectica: International Review of Philosophy of Knowledge, 8(2):168– 172, June 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kroner:1954:WPE

[285] Franz Kröner. W. Pauli: Der Einfluss archetypischer Vorstellungen auf die Bildung naturwissenschaftlicher Theorien bei Kepler. (German) [W. Pauli: The influence of archetypal ideas on the formation of scientific theories of Kepler]. Dialectica: International Review of Philosophy of Knowledge, 8(2):173–180, June 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1954:CPS

[286] F. Gonseth. Connaître par la science. (French) [Knowledge through science]. Dialectica: International Review of Philosophy of Knowledge, 8 (3):183–198, September 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lande:1954:QIC

[287] Alfred Landé. Quantum indeterminacy, a consequence of cause-effect continuity. Dialectica: International Review of Philosophy of Knowledge, 8(3):199–209, September 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Turbayne:1954:BRS

[288] Colin Murray Turbayne. Berkeley and Russell on space. Dialectica: International Review of Philosophy of Knowledge, 8(3):210–227, September 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Curry:1954:RDN

[289] H. B. Curry. Remarks on the definition and nature of mathematics. Dialectica: International Review of Philosophy of Knowledge, 8(3):228– 233, September 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Specker:1954:AMG

[290] E. Specker. Die Antinomien der Mengenlehre. (German) [The antinomies of quantum theory]. Dialectica: International Review of Philosophy of Knowledge, 8(3):234–244, September 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Johnstone:1954:SAP

[291] Henry W. Johnstone. Some aspects of philosophical disagreement. Dialectica: International Review of Philosophy of Knowledge, 8(3):245–257, September 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kroner:1954:HDS

[292] F. Kröner. Herbert Dingle: the scientific adventure. Dialectica: International Review of Philosophy of Knowledge, 8(3):258–267, September 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1954:KMG

[293] Gert Heinz Müller. Karl Menger: géométrie générale. (French) [Karl Menger: general geometry]. Dialectica: International Review of Philosophy of Knowledge, 8(3):267–269, September 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1954:AJE

[294] Gert Heinz Müller. Arnaud den Joy: l’énumération transfinie. (French) [Arnaud den Joy: the transfinite enumeration]. Dialectica: International Review of Philosophy of Knowledge, 8(3):270–273, September 1954. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1954:OBJ

[295] Gert Heinz Müller. Oskar Becker und Jos. E. Hofmann: Geschichte der Mathematik. (German) [Oskar Becker and Jos. E. Hofmann: History of mathematics]. Dialectica: International Review of Philosophy of Knowledge, 8(3):274, September 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1954:BJP

[296] Anonymous. British journal for the philosophy of science. Dialectica: International Review of Philosophy of Knowledge, 8(3):275–280, September 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pauli:1954:NEA

[297] Wolfgang Pauli. Naturwissenschaftliche und erkenntnistheoretische Aspekte der Ideen vom Unbewussten. (German) [Natural sciences and epistemological aspects of the ideas of the unconscious]. Dialectica: International Review of Philosophy of Knowledge, 8(4):283–301, December 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Meier:1954:PUS

[298] C. A. Meier. Projection, übertragung und subjekt-objektrelation in der psychologie. (German) [Projection, transfer and subject-object-relation in psychology]. Dialectica: International Review of Philosophy of Knowledge, 8(4):302–321, December 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1954:ECG

[299] Anonymous. Les évolutionnismes contemporains de Gustave Mercier et du Père Teilhard de Chardin. (French) [The contemporary evolutionism of Gustave Mercier and Father Teilhard de Chardin]. Dialectica: International Review of Philosophy of Knowledge, 8(4):322–346, December 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cbahay:1954:LFE

[300] F. Cbahay. Logique formelle, empirisme logique et vérité. (French) [Formal logic, logical empiricism and truth]. Dialectica: International Review of Philosophy of Knowledge, 8(4):347–356, December 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gagnebin:1954:EMF

[301] S. Gagnebin. Esprit et méthode. (French) [Spirit and method]. Dialectica: International Review of Philosophy of Knowledge, 8(4):357–363, December 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kroner:1954:AKJ

[302] F. Kroner. Alfred Klemmt, John Locke: theoretische philosophie. (German) [Alfred Klemmt, John Locke: theoretical philosophy]. Dialectica: International Review of Philosophy of Knowledge, 8(4):364–365, December 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1954:KR

[303] Gert Heinz Müller. Kurt Reidemeister. Dialectica: International Review of Philosophy of Knowledge, 8(4):366–368, December 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pikler:1954:BRB

[304] Andrew G. Pikler. Book review: British Journal for the Philosophy of Science: abstracts of main articles in the February number. Dialectica: International Review of Philosophy of Knowledge, 8(4):369–370, December 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1954:LA

[305] Anonymous. Liste des auteurs. Dialectica: International Review of Philosophy of Knowledge, 8(4):371–374, December 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1954:TMF

[306] Anonymous. Table des matières. (French) [Table of contents]. Dialectica: International Review of Philosophy of Knowledge, 8(4):375–376, December 1954. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1955:LEP

[307] F. Gonseth. L’ouverture à l’expérience et les a priori. (French) [Openness to experience and a priori]. Dialectica: International Review of Philosophy of Knowledge, 9(1–2):5–22, June 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1955:FAKa

[308] Paul Bernays. Zur Frage der Anknüpfung an die kantische Erkenntnistheorie: eine kritische Erörterung. (German) [On the question of the connection to the Kantian theory of knowledge: a critical discussion]. Dialectica: International Review of Philosophy of Knowledge, 9(1–2):23–65, June 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Scholz:1955:TZI

[309] Heinrich Scholz. Eine Topologie der Zeit im kantischen Sinne. (German) [A topology of time in the Kantian sense]. Dialectica: International Review of Philosophy of Knowledge, 9(1–2):66–113, June 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lachieze-Rey:1955:KSF

[310] P. Lachiéze-Rey. Le kantisme et la science. (French) [Kantism and science]. Dialectica: International Review of Philosophy of Knowledge, 9 (1–2):114–122, June 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1955:CPS

[311] F. Gonseth. Connaître par la science (suite). (French) [Knowledge through science (end)]. Dialectica: International Review of Philosophy of Knowledge, 9(1–2):123–136, June 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1955:RMC

[312] F. Gonseth. Recherches méthodologiques: conclusions de l’ouvrage La géométrie et le problème de l’espace. (French) [Methodological research: conclusions of the book Geometry and the problem of space]. Dialectica: International Review of Philosophy of Knowledge, 9(1–2):137–185, June 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Freundlich:1955:BJE

[313] Rudolf Freundlich. Bela Juhos: die Erkenntnis und ihre Leistung die naturwissenschaftliche Methode. (German) [Bela Juhos: knowledge and its achievement: the natural science method]. Dialectica: International Review of Philosophy of Knowledge, 9(1–2):186–187, June 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1955:BJPa

[314] Anonymous. British journal for the philosophy of science. Dialectica: International Review of Philosophy of Knowledge, 9(1–2):188–190, June 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1955:FAKb

[315] Paul Bernays. Zur Frage der Anknüpfung an die kantische Erkenntnistheorie. (German) [On the question of the link to the Kantian theory of knowledge]. Dialectica: International Review of Philosophy of Knowledge, 9(3–4):195–221, December 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1955:MPF

[316] F. Gonseth. Des mathématiques à la philosophie. (French) [From mathematics to philosophy]. Dialectica: International Review of Philosophy of Knowledge, 9(3–4):222–243, December 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Moch:1955:ONP

[317] François Moch. Oui, non — peut-être. (French) [Yes, no — perhaps]. Dialectica: International Review of Philosophy of Knowledge, 9(3–4): 244–262, December 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rougier:1955:LER

[318] L. Rougier. L’univers est-il rationnel? Dialectica: International Review of Philosophy of Knowledge, 9(3–4):263–272, December 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vanDantzig:1955:FN

[319] D. van Dantzig. Is 10^{10^{1}0} a finite number? Dialectica: International Review of Philosophy of Knowledge, 9(3–4):273–277, December 1955. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1955:ADD

[320] Anonymous. Actes du deuxième congrès international de l’Union Internationale de Philosophie des Sciences. (French) [Proceedings of the Second International Congress of the International Union Philosophy of Science]. Dialectica: International Review of Philosophy of Knowledge, 9(3–4):278, December 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Feibleman:1955:KAS

[321] James K. Feibleman. Knowing about Semipalatinsk. Dialectica: International Review of Philosophy of Knowledge, 9(3–4):279–286, December 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bird:1955:DPI

[322] Otto Bird. Dialectic in philosophical inquiry. Dialectica: International Review of Philosophy of Knowledge, 9(3–4):287–304, December 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Miescher:1955:GWG

[323] K. Miescher. Gegensätzlichkeit und Wirklichkeit. (German) [Contradiction and reality]. Dialectica: International Review of Philosophy of Knowledge, 9(3–4):305–340, December 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dingler:1955:GWG

[324] Hugo Dingler. Geometrie und Wirklichkeit. (German) [Geometry and reality]. Dialectica: International Review of Philosophy of Knowledge, 9 (3–4):341–362, December 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Grize:1955:LNV

[325] Jean-Blaise Grize. L’implication et la négation vues au travers des méthodes de Gentzen et de Fitch. (French) [The implication and negation seen through the methods of Gentzen and Fitch]. Dialectica: International Review of Philosophy of Knowledge, 9(3–4):363–381, December 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1955:BJPb

[326] Anonymous. British Journal for the Philosophy of Science. Dialectica: International Review of Philosophy of Knowledge, 9(3–4):382–384, December 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1955:BHS

[327] Anonymous. Buchbesprechung: Heinricii Scholz: Eine Topologie der Zeit im kantischen Sinne. Dialectica: International Review of Philosophy of Knowledge, 9(3–4):385, December 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1955:RBL

[328] Anonymous. Received books — livres reçus — eingelaufene Bücher. Dialectica: International Review of Philosophy of Knowledge, 9(3–4): 386–389, December 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1955:LAL

[329] Anonymous. Liste des auteurs: liste des auteurs. Dialectica: International Review of Philosophy of Knowledge, 9(3–4):391–395, December 1955. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1956:DES

[330] F. Gonseth. Dialectica entre dans sa dixième année d’existence. (French) [Dialectica enters its tenth year of existence]. Dialectica: International Review of Philosophy of Knowledge, 10(1):5–8, March 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rotenstreich:1956:KSC

[331] Nathan Rotenstreich. Kant’s schematism in its context. Dialectica: International Review of Philosophy of Knowledge, 10(1):9–30, March 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Binder:1956:BNP

[332] H. Binder. Der Begriff des Normalen in der Psychiatrie. (German) [The concept of normal in psychiatry]. Dialectica: International Review of Philosophy of Knowledge, 10(1):31–40, March 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Simenauer:1956:BBG

[333] Erich Simenauer. Betrachtungen über die biologischen Grundlagen des Denkens und des Willens. (German) [Reflections on the biological foundations of thought and will]. Dialectica: International Review of Philosophy of Knowledge, 10(1):41–53, March 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gex:1956:CCN

[334] Maurice Gex. Chronique de caractérologie no. 3. (French) [Chronicle of caracterology no. 3]. Dialectica: International Review of Philosophy of Knowledge, 10(1):54–79, March 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dingler:1956:GWG

[335] Hugo Dingler. Geometrie und Wirklichkeit. (German) [Geometry and reality]. Dialectica: International Review of Philosophy of Knowledge, 10(1):80–93, March 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1956:BJPa

[336] Anonymous. The British Journal for the Philosophy of Science. Dialectica: International Review of Philosophy of Knowledge, 10(1):94–95, March 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1956:HTF

[337] F. Gonseth. De l’humanisation de la technique. (French) [From the humanization of technology]. Dialectica: International Review of Philosophy of Knowledge, 10(2):99–112, June 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Reenpaa:1956:VBE

[338] Yrjö Reenpää. Versuch über die Beobachungsgrundlagen der exakten empirischen Wissenschaften. (German) [Experiment on the observations of the exact empirical sciences]. Dialectica: International Review of Philosophy of Knowledge, 10(2):113–147, June 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pap:1956:LER

[339] Arthur Pap. Locical empiricism and rationalism. Dialectica: International Review of Philosophy of Knowledge, 10(2):148–166, June 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

DeLaHorbe:1956:SPF

[340] F. De La Horbe. Sociologie et philosophie. (French) [Sociology and philosophy]. Dialectica: International Review of Philosophy of Knowledge, 10(2):167–179, June 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1956:RCL

[341] Anonymous. Reviews — chronique — Literaturblatt: F. van Steenbergen, ontologie. (French) [F. van Steenbergen, ontology]. Dialectica: International Review of Philosophy of Knowledge, 10(2):180–183, June 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1956:BJPb

[342] Anonymous. British journal for the philosophy of science. Dialectica: International Review of Philosophy of Knowledge, 10(2):184–189, June 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Moch:1956:LA

[343] Franqois Moch. La logique des attitudes. Dialectica: International Review of Philosophy of Knowledge, 10(3):191–203, September 1956. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Polanyi:1956:PAS

[344] Michael Polanyi. Pure and applied science and their appropriate forms of organization. Dialectica: International Review of Philosophy of Knowledge, 10(3):231–242, September 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Conwaary:1956:PSP

[345] Joachim Werner Conwaary. Pholosophy, science of psychology metaphysics. Dialectica: International Review of Philosophy of Knowledge, 10(3):243–249, September 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Finsler:1956:PSM

[346] P. Finsler. Der platinische Standpunkt in der Mathematik. (German) [The Platonic viewpoint in mathematics]. Dialectica: International Review of Philosophy of Knowledge, 10(3):250–255, September 1956. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wittenberg:1956:WKP

[347] Alexander Wittenberg. Waryn kein platonisums? Dialectica: International Review of Philosophy of Knowledge, 10(3):256–261, September 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1956:DTP

[348] P. Bernays. Zur Diskussion des Themas der platonische Standpunkt in der Mathematik. (German) [On discussion of the subject of the Platonic position in mathematics]. Dialectica: International Review of Philosophy of Knowledge, 10(3):262–265, September 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Finsler:1956:UP

[349] Paul Finsler. Unddoch platonismus. Dialectica: International Review of Philosophy of Knowledge, 10(3):266–270, September 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lorenzer:1956:BZ

[350] P. Lorenzer and P. Finsler. Briefwechsel zwischen. Dialectica: International Review of Philosophy of Knowledge, 10(3):271–277, September 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1956:BJPc

[351] Anonymous. British journal for the philosophy of science. Dialectica: International Review of Philosophy of Knowledge, 10(3):278–280, September 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deLatil:1956:MCC

[352] P. de Latil. De la machine considérée comme un moyen de connaître l’homme. (French) [Of the machine considered as a means of knowing man]. Dialectica: International Review of Philosophy of Knowledge, 10 (4):287–291, December 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Doyon:1956:MCD

[353] A. Doyon and L. Liaigre. Méthodologie comparée du biomécanisme et de la mécanique comparée. (French) [Comparative methodology of biomechanism and comparative mechanics]. Dialectica: International Review of Philosophy of Knowledge, 10(4):292–323, December 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Russo:1956:D

[354] P. F. Russo. Discussion. Dialectica: International Review of Philosophy of Knowledge, 10(4):324–335, December 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Couffignal:1956:QRS

[355] L. Couffignal. Quelques réflexions et suggestions. (French) [Some thoughts and suggestions]. Dialectica: International Review of Philosophy of Knowledge, 10(4):336–339, December 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1956:D

[356] F. Gonseth and F. Bonsack. Discussion. Dialectica: International Review of Philosophy of Knowledge, 10(4):339–346, December 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vernotte:1956:DMC

[357] Pierre Vernotte. Le domaine de la machine et celui de l’homme. (French) [The domain of the machine and that of man]. Dialectica: International Review of Philosophy of Knowledge, 10(4):347–356, December 1956. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1956:D

[358] Anonymous. Discussion. Dialectica: International Review of Philosophy of Knowledge, 10(4):356–367, December 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hertz:1956:GTF

[359] Paul Hertz. Gibbs’ theory its foundations and applications. Dialectica: International Review of Philosophy of Knowledge, 10(4):368–383, December 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1956:RBR

[360] Anonymous. Received books: received books. Dialectica: International Review of Philosophy of Knowledge, 10(4):384–386, December 1956. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1956:LAL

[361] Anonymous. Liste des auteurs: liste des auteurs. Dialectica: International Review of Philosophy of Knowledge, 10(4):387–391, December 1956. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1957:E

[362] F. Gonseth. Editorial. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):5–8, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1957:GMM

[363] F. Gonseth. Generalités méthodologiques: la méthodologie des sciences peut-elle etre élevée au rang de discipline scientifique?. (French) [General methodology: can the methodology of science be elevated to scientific discipline?]. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):9–20, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Perelman:1957:EPF

[364] Ch. Perelman. Évidence et preuve. (French) [Evidence and proof]. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):21– 35, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pauli:1957:PPR

[365] Wolfgang Pauli. Phänomen und Physikalische Realität. (German) [Phenomena and physical reality]. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):36–48, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). Lecture of 24 August 1954 at the International Philosophy Congress in Zurich, Switzerland.

Rossel:1957:CTI

[366] J. Rossel. Caractéristiques, tendances et implications de la recherche atomique actuelle. (French) [Characteristics, trends and implications of current atomic research]. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):49–56, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Destouches:1957:ADN

[367] J.-L. Destouches. Aspect dialectique de la notion de système physique. (French) [Dialectical aspect of the physical system]. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):57–69, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Konig:1957:AGG

[368] H. König. Zur Analyse des Grössenbegriffs. (German) [Toward analysis of the concept of size]. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):70–87, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dubarle:1957:MLM

[369] R. P. Dubarle. Méthodologie de la logique et des mathématiques methodologie der logik und der mathematik: formalisations et théorèmes critiques. (French) [Methodology of logic and mathematics: Formalizations and critical theorems]. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):88–120, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bouligand:1957:LMS

[370] G. Bouligand. L’activité mathématique et son dualisme. (French) [Mathematical activity and its dualism]. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):121–139, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Guggenheimer:1957:DSF

[371] Heinrich Guggenheimer. Dialektische Synthese von Formalismus und Intuitionismus. (German) [Dialectical synthesis of formalism and intuitionism]. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):140–147, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dugue:1957:QRC

[372] Daniel Dugué. Quelques remarques sur le caractère provisoire de toute axiomatique. (French) [Some remarks on the provisional nature of any axiomatic]. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):148–153, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ruesch:1957:PHC

[373] Jurgen Ruesch. Principles of human communication. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):154–166, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Monnier:1957:POC

[374] Marcel Monnier. Plan d’organisation, continuité et stades du développement des activités électriques cérébrales. (French) [Plan of organization, continuity and stages of the development of brain electrical activities]. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):167–178, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Herrick:1957:AIN

[375] C. Judson Herrick. Analytic and integrative nervous functions. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):179–186, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Buhler:1957:AGS

[376] Charlotte Buhler. About the goal-structure of human life. theoretical considerations. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):187–205, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Servais:1957:PMO

[377] Willy Servais. Pédagogie mathèmatique ouverte. (French) [Teaching open mathematics]. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):206–215, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1957:BJP

[378] Anonymous. British journal for the philosophy of science: authors’ abstracts of main articles in the February number, 1957. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):216–218, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1957:BRP

[379] Anonymous. Book review: Philosophy of Science journal (U.S.A.): authors’ abstracts of main articles in the January number, 1957. Dialectica: International Review of Philosophy of Knowledge, 11(1–2):219–225, June 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gagnebin:1957:E

[380] S. Gagnebin. Editorial. Dialectica: International Review of Philosophy of Knowledge, 11(3–4):232, December 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1957:SSP

[381] Paul Bernays. Von der Syntax der Sprache zur Philosophie der Wissenschaften. (German) [From the syntax of language to the philosophy of science]. Dialectica: International Review of Philosophy of Knowledge, 11(3–4):233–246, December 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Blanc:1957:EID

[382] Alberto Carlo Blanc. De l’emploi inadéquat du terme · primitif ·: considérations sur l‘évolution et la systématique. (French) [From the inadequate use of the · primitive ·: considerations on evolution and systematics]. Dialectica: International Review of Philosophy of Knowledge, 11 (3–4):247–275, December 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mercier:1957:SPF

[383] André Mercier. Science et philosophie. (French) [Science and philosophy]. Dialectica: International Review of Philosophy of Knowledge, 11(3–4): 276–295, December 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bayer:1957:ERL

[384] Raymond Bayer. Engagement et raison: l‘évolution de l’intelligence et les formes modernes de la dialectique. (French) [Commitment and reason: the evolution of intelligence and modern forms of dialectics]. Dialectica: International Review of Philosophy of Knowledge, 11(3–4):296–305, December 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rougier:1957:ERL

[385] Louis Rougier. Engagement et raison: l‘évolution du concept de raison dans la pensée occidentale. (French) [Commitment and reason: The evolution of the concept of reason in Western thought]. Dialectica: International Review of Philosophy of Knowledge, 11(3–4):306–326, December 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Paris:1957:ERP

[386] Carlos Paris. Engagement et raison: programme et position historique d’un rationalisme humaniste. (French) [Commitment and reason: Program and historical position of a humanist rationalism]. Dialectica: International Review of Philosophy of Knowledge, 11(3–4):327–336, December 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Reidemeister:1957:ERS

[387] Kurt Reidemeister. Engagement et raison — Spekulation und Vernunft. (French) [Commitment and reason]. Dialectica: International Review of Philosophy of Knowledge, 11(3–4):337–344, December 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dupreel:1957:ERC

[388] E. Dupréel. Engagement et raison: consistance et valeurs. (French) [Commitment and reason: Consistency and values]. Dialectica: International Review of Philosophy of Knowledge, 11(3–4):345–353, December 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Popper:1957:KPP

[389] Karl R. Popper. Knowledge and probability: probability magic or knowledge out of ignorance. Dialectica: International Review of Philosophy of Knowledge, 11(3–4):354–374, December 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Moch:1957:KPR

[390] François Moch. Knowledge and probability: réflexions sur les probabilités. (French) [Knowledge and probability: reflections on probabilities]. Dialectica: International Review of Philosophy of Knowledge, 11 (3–4):375–391, December 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nolfi:1957:KPS

[391] Padrot Nolfi. Knowledge and probability: Spieltheorie und Willensfreiheit. Dialectica: International Review of Philosophy of Knowledge, 11 (3–4):392–404, December 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gagnebin:1957:VPG

[392] Samuel Gagnebin. Valeur philosophique de la géométrie: structure et substructure de la géométrie. (French) [Philosophical value of geometry: structure and substructure of geometry]. Dialectica: International Review of Philosophy of Knowledge, 11(3–4):405–433, December 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1957:LAL

[393] Anonymous. Liste des auteurs: liste des auteurs. Dialectica: International Review of Philosophy of Knowledge, 11(3–4):435–439, December 1957. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1958:RDD

[394] Anonymous. Reouverture dn dialogue. Dialectica: International Review of Philosophy of Knowledge, 12(1):5–6, March 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Freudenthal:1958:MSP

[395] Hans Freudenthal. Ist die mathematische Statistik Paradox?. (German) [Is mathematical statistics a paradox?]. Dialectica: International Review of Philosophy of Knowledge, 12(1):7–32, March 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1958:POD

[396] F. Gonseth. Pour ouvrir la discussion. (French) [To open the discussion]. Dialectica: International Review of Philosophy of Knowledge, 12 (1):33–36, March 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Favarger:1958:VSC

[397] Claude Favarger. Vérité scientifique et compréhension du vivant réflexions d’un naturaliste. (French) [Scientific truth and understanding of the living reflections of a naturalist]. Dialectica: International Review of Philosophy of Knowledge, 12(1):37–67, March 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Favarger:1958:DR

[398] M. C. Favarger and F. Gonseth. Deux remarques. Dialectica: International Review of Philosophy of Knowledge, 12(1):68–69, March 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Russo:1958:PFT

[399] F. Russo. Une perspective fondamentale de la théorie de l’information et de l’action technique: la dualité de l’action concrète et du signal. (French) [A fundamental perspective of the theory of information and technical action: the duality of concrete action and the signal]. Dialectica: International Review of Philosophy of Knowledge, 12(1):70–80, March 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rankin:1958:ICF

[400] H. D. Rankin. Immediate cognition of the forms in the Phædo? Dialectica: International Review of Philosophy of Knowledge, 12(1):81–86, March 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gex:1958:RCC

[401] Maurice Gex. Reviews: chronique de caractérologie n^{∘} 4. (French) [Chronicle of caracterology n^{∘} 4]. Dialectica: International Review of Philosophy of Knowledge, 12(1):87–92, March 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1958:BJP

[402] Anonymous. British journal for the philosophy of science. Dialectica: International Review of Philosophy of Knowledge, 12(1):93–95, March 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ayer:1958:PLO

[403] A. J. Ayer. Philosophie et langage ordinaire. (French) [Philosophy and ordinary language]. Dialectica: International Review of Philosophy of Knowledge, 12(2):99–129, June 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rey:1958:IMP

[404] André Rey. Les images mentales en psychophysiologie. (French) [Mental images in psychophysiology]. Dialectica: International Review of Philosophy of Knowledge, 12(2):130–145, June 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Husser:1958:GZG

[405] F. F. Husser. Zur genealogie der zahl. (German) [On the genealogy of the numbers]. Dialectica: International Review of Philosophy of Knowledge, 12(2):146–184, June 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Scherrer:1958:GEG

[406] Willy Scherrer. Geometrie und Erkenntnistheorie. (German) [Geometry and cognitive theory]. Dialectica: International Review of Philosophy of Knowledge, 12(2):185–196, June 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1958:RMF

[407] Anonymous. Réponse de M. F. Gonseth à M. W. Scherrer. (French) [Reply by M. F. Gonseth to M. W. Scherrer]. Dialectica: International Review of Philosophy of Knowledge, 12(2):197–199, June 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1958:PSJ

[408] Anonymous. Philosophy of Science journal (U.S.A.). Dialectica: International Review of Philosophy of Knowledge, 12(2):200–204, June 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ackermann:1958:BZS

[409] Wilhelm Ackermann. Über die Beziehung zwischen strikter und strenger Implikation. (German) [On the relationship between strict and strict implication]. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):213–222, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Beth:1958:FCE

[410] E. W. Beth. ^{≪}Cogito ergo sum^{≫}— raisonnement ou intuition?. (French) [^{≪}I think, therefore I exist^{≫}— reasoning or intuition?]. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):223–235, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Carnap:1958:BTS

[411] Rudolf Carnap. Beobachtungssprache und theoretische Sprache. (German) [Observation language and theoretical language]. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):236–248, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Curry:1958:CFS

[412] Haskell B. Curry. Calculuses and formal systems. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):249–273, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fraenkel:1958:PBB

[413] Abraham A. Fraenkel. Paul Bernays und die Begründung der Mengenlehre. (German) [Paul Bernays and the justification of quantum theory]. Dialectica: International Review of Philosophy of Knowledge, 12(3–4): 274–279, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Godel:1958:BNB

[414] Kurt Gödel. Über eine Bisher noch nicht benützte Erweiterung des finiten Standpunktes. (German) [On an as yet unused extension of the finite position]. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):280–287, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonsetu:1958:PDL

[415] F. Gonsetu. Le problème du langage et l’ouverture à l’expèrience. (French) [The problem of language and openness to experience]. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):288–295, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Goodstein:1958:NMS

[416] R. L. Goodstein. On the nature of mathematical systems. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):296–316, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hermes:1958:EWG

[417] Hans Hermes. Zum Einfachheitsprinzip in der Wahrscheinlichkeitsrechnung. (German) [Toward a simplicity principle in probability calculation]. Dialectica: International Review of Philosophy of Knowledge, 12 (3–4):317–331, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Heyting:1958:BIW

[418] A. Heyting. Blick von der intuitionistischen Warte. (German) [View from the intuitionist]. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):332–345, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kreisel:1958:HP

[419] Georg Kreisel. Hilbert’s programme. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):346–372, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Peter:1958:GRF

[420] Rózsa Péter. Graphschemata und rekursive Funktionen. (German) [Graph schemata and recursive functions]. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):373–393, December 1958. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Robinson:1958:RMC

[421] Abraham Robinson. Relative model-completeness and the elimination of quantifiers. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):394–407, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schmidt:1958:ENU

[422] H. Arnold Schmidt. Über einige neuere Untersuchungen zur Modalitätenlogik. (German) [Some recent studies on modality logic]. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):408–421, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sceutte:1958:AGF

[423] Kurt Sceütte. Aussagenlogische grundeigenschaften formaler systeme. (German) [Relativity of formal systems]. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):422–442, December 1958. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Skolem:1958:RAS

[424] Th. Skolem. Reduction of axiom systems with axiom schemes to systems with only simple axioms. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):443–450, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Specker:1958:DGD

[425] Ernst Specker. Dualität. (German) [Duality]. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):451–465, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wang:1958:EYF

[426] Hao Wang. Eighty years of foundational studies. Dialectica: International Review of Philosophy of Knowledge, 12(3–4):466–497, December 1958. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Piaget:1959:PAE

[427] Jean Piaget. Perception, apprentissage et empirisme. (French) [Perception, learning and empiricism]. Dialectica: International Review of Philosophy of Knowledge, 13(1):5–15, March 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1959:LEE

[428] F. Gonseth. L’ouverture à l’expérience en épistémologie. (French) [Openness has experience in epistemology]. Dialectica: International Review of Philosophy of Knowledge, 13(1):16–26, March 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Konig:1959:KMG

[429] H. König. Kulturbedeutung des Messens. (German) [Cultural meaning of measurement]. Dialectica: International Review of Philosophy of Knowledge, 13(1):27–41, March 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schrag:1959:WHP

[430] Calvin O. Schrag. Whitehead and Heidegger: process philosophy and existential philosophy. Dialectica: International Review of Philosophy of Knowledge, 13(1):42–56, March 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Munz:1959:IP

[431] Peter Munz. Investigations of philosophy. Dialectica: International Review of Philosophy of Knowledge, 13(1):57–80, March 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Scherrer:1959:BRM

[432] W. Scherrer. Bemerkungen zur Réponse de M. F. Gonseth à M. W. Scherrer. (German) [Remarks on the response of M. F. Gonseth to M. W. Scherrer]. Dialectica: International Review of Philosophy of Knowledge, 13(1):81–83, March 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1959:PSJa

[433] Anonymous. Philosophy of Science journal (USA). Dialectica: International Review of Philosophy of Knowledge, 13(1):84–94, March 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ruyer:1959:PSF

[434] Raymond Ruyer. La psychobiologie et la science. (French) [Psychobiology and science]. Dialectica: International Review of Philosophy of Knowledge, 13(2):103–122, June 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chauchard:1959:VLA

[435] Paul Chauchard. Valeur et limites de l’apport scientifique à la philosophie. (French) [Value and limits of scientific contribution to philosophy]. Dialectica: International Review of Philosophy of Knowledge, 13(2):123– 143, June 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tomlin:1959:PL

[436] E. W. F. Tomlin. The philosophy of life. Dialectica: International Review of Philosophy of Knowledge, 13(2):144–159, June 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gex:1959:PIS

[437] Maurice Gex. La philosophie d’inspiration scientifique. (French) [The philosophy of scientific inspiration]. Dialectica: International Review of Philosophy of Knowledge, 13(2):160–184, June 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Piguet:1959:NPI

[438] J.-Claude Piguet. Note sur la philosophie d’inspiration scientifique. (French) [Note on philosophy of scientific inspiration]. Dialectica: International Review of Philosophy of Knowledge, 13(3–4):191–207, December 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Feibleman:1959:LSS

[439] James K. Feibleman. The logical structure of the scientific method. Dialectica: International Review of Philosophy of Knowledge, 13(3–4):208– 225, December 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1959:QRA

[440] F. Gonseth. Quelques remarques sur l’article de M. James K. Feibleman, “The logical structure of the scientific method”. (French) [A few remarks on M. James K. Feibleman’s, “The logical structure of the scientific method”]. Dialectica: International Review of Philosophy of Knowledge, 13(3–4):226–234, December 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1959:DLP

[441] Gustav E. Müller. Dialectic-the logic of philosophy. Dialectica: International Review of Philosophy of Knowledge, 13(3–4):235–261, December 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1959:RPS

[442] Anonymous. Le robot peut-il servir à la connaissance de l’homme? (French) [Can the robot be used for the knowledge of man?]. Dialectica: International Review of Philosophy of Knowledge, 13(3–4):262–268, December 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sauvan:1959:FCF

[443] Jacques Sauvan. Les facteurs de connaissance. (French) [Knowledge factors]. Dialectica: International Review of Philosophy of Knowledge, 13 (3–4):269–280, December 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

David:1959:RPN

[444] Aurel David. Réflexions pour un nouveau schéma de l’homme: but et moyens. (French) [Reflections on a new schema of man: purpose and means]. Dialectica: International Review of Philosophy of Knowledge, 13 (3–4):281–320, December 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vernotte:1959:PCH

[445] Pierre Vernotte. À propos de la connaissance — H. (French) [About knowledge — H]. Dialectica: International Review of Philosophy of Knowledge, 13(3–4):321–330, December 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bonsack:1959:CVF

[446] F. Bonsack. Connaissance et vérité. (French) [Knowledge and truth]. Dialectica: International Review of Philosophy of Knowledge, 13(3–4): 331–342, December 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1959:CHP

[447] F. Gonseth. La connaissance de l’homme par l’intermédiaire du robot. (French) [The knowledge of man through the robot]. Dialectica: International Review of Philosophy of Knowledge, 13(3–4):343–349, December 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tiebout:1959:SWC

[448] Harry M. Tiebout. Subjectivity in Whitehead: a comment on “Whitehead and Heidegger”. Dialectica: International Review of Philosophy of Knowledge, 13(3–4):350–353, December 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1959:NIC

[449] Anonymous. Notice: 1960 international congress for logic, methodology and philosophy of science. Dialectica: International Review of Philosophy of Knowledge, 13(3–4):354, December 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1959:PSJb

[450] Anonymous. Philosophy of Science journal (U.S.A.). Dialectica: International Review of Philosophy of Knowledge, 13(3–4):355–360, December 1959. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vandel:1960:HS

[451] A. Vandel. Un humanisme scientifique. Dialectica: International Review of Philosophy of Knowledge, 14(1):5–20, March 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gex:1960:DPI

[452] Maurice Gex. Défense de la philosophie d’inspiration scientifique. (French) [Defense of the philosophy of scientific inspiration]. Dialectica: International Review of Philosophy of Knowledge, 14(1):21–31, March 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ruyeb:1960:RNM

[453] Raymond Ruyeb. Réponse à la note de M. J.-Claude Piguet. (French) [Reply to M. J.-Claude Piguet]. Dialectica: International Review of Philosophy of Knowledge, 14(1):32–36, March 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Portmann:1960:LHP

[454] Adolphe Portmann. L’ontogenése humaine et le probléme de l’origine. (French) [Human ontogenesis and the problem of origin]. Dialectica: International Review of Philosophy of Knowledge, 14(1):37–52, March 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Blair:1960:SSG

[455] George A. Blair. Science, sufficient ground, and the possibility of metaphysics. Dialectica: International Review of Philosophy of Knowledge, 14 (1):53–79, March 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Supek:1960:EPL

[456] Ivan Supek. Escape of philosophy into linguistic depths. Dialectica: International Review of Philosophy of Knowledge, 14(1):80–92, March 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1960:RCL

[457] Anonymous. Reviews — chronique — Literaturblatt: Philippe Wehrle: L’univers aléatoire. Preface de Ferdinand Gonseth. Dialectica: International Review of Philosophy of Knowledge, 14(1):93–99, March 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1960:BJP

[458] Anonymous. British journal for the philosophy of science. Dialectica: International Review of Philosophy of Knowledge, 14(1):100–103, March 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gagnebin:1960:GTQ

[459] S. Gagnebin. Gonseth tel que je le connais. (French) [Gonseth as I know hime]. Dialectica: International Review of Philosophy of Knowledge, 14 (2–3):109–120, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mehcier:1960:TTP

[460] André Mehcier. Tea for two (a piece of fiction). Dialectica: International Review of Philosophy of Knowledge, 14(2–3):121–126, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Aebi:1960:SGF

[461] M. Aebi. Die stellung von Gonseths ^{≪}loffener philosophie^{≫}im ganzen der philosophia perennis. (German) [The position of Gonseth’s philosophical philosophy in the whole of the philosophia perennis]. Dialectica: International Review of Philosophy of Knowledge, 14(2–3):127–150, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Behnays:1960:CPG

[462] Paul Behnays. Charakterzüge der philosophie Gonseths. (German) [Characteristics of the philosophy of Gonseth]. Dialectica: International Review of Philosophy of Knowledge, 14(2–3):151–156, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bonsack:1960:MSF

[463] François Bonsack. La métaphysique et la science. (French) [Metaphysics and science]. Dialectica: International Review of Philosophy of Knowledge, 14(2–3):157–166, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bouligand:1960:ETL

[464] G. Bouligand. Effort théorique et limites au pouvoir d’unifier. (French) [Theoretical effort and limits to the power to unify]. Dialectica: International Review of Philosophy of Knowledge, 14(2–3):167–175, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Christoff:1960:SOF

[465] Daniel Christoff. Situation et objectivité. (French) [Situation and objectivity]. Dialectica: International Review of Philosophy of Knowledge, 14 (2–3):176–182, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Guggenheimer:1960:GRN

[466] H. W. Guggenheimer. General relativity and nuclear reactions. Dialectica: International Review of Philosophy of Knowledge, 14(2–3):183–187, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hainard:1960:SAF

[467] Robert Hainard. Science et art. (French) [Science and art]. Dialectica: International Review of Philosophy of Knowledge, 14(2–3):188–196, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hirsch:1960:POC

[468] Guy Hirsch. Philosophie ouverte et connaissance probable. (French) [Open philosophy and probable knowledge]. Dialectica: International Review of Philosophy of Knowledge, 14(2–3):197–202, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lindeh:1960:BVZ

[469] Arthur Lindeh. Beobachtung, Versuch und Zufall. (German) [Observation, experiment and chance]. Dialectica: International Review of Philosophy of Knowledge, 14(2–3):203–214, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mercier:1960:RDM

[470] André Mercier. De la réalité et du mythe. (French) [From reality and the myth]. Dialectica: International Review of Philosophy of Knowledge, 14 (2–3):215–220, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nolfi:1960:DFB

[471] Padrot Nolfi. Deutung der Freiheit als eine Begegnung. (German) [Interpretation of freedom as an encounter]. Dialectica: International Review of Philosophy of Knowledge, 14(2–3):221–229, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Perelman:1960:RJF

[472] Ch. Perelman. La règle de justice. (French) [The rule of justice]. Dialectica: International Review of Philosophy of Knowledge, 14(2–3):230–238, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Specker:1960:LGE

[473] Ernst Specker. Die Logik nicht gleichzeitig entscheidbarer Aussagen. (German) [The logic of not simultaneously decidable statements]. Dialectica: International Review of Philosophy of Knowledge, 14(2–3):239– 246, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Veillon:1960:S

[474] Charles Veillon. Songe. Dialectica: International Review of Philosophy of Knowledge, 14(2–3):247–253, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Walteb:1960:WAP

[475] Emil J. Walteb. Wissenschaftstheoretische ansätze zum problem der überwindung der ideologischen sozialphilosophie durch empirische sozialforschung. (German) [Scientific theoretical approaches to the problem of overcoming ideological social philosophy by empirical social research]. Dialectica: International Review of Philosophy of Knowledge, 14(2–3): 254–266, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1960:MPF

[476] Anonymous. Miscellaneous: publications de Ferdinand Gonseth. (French) [Miscellaneous: publications by Ferdinand Gonseth]. Dialectica: International Review of Philosophy of Knowledge, 14(2–3):267–275, September 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Redacteur:1960:AFA

[477] Le Rédacteur. Avertissement. (French) [Advertisement]. Dialectica: International Review of Philosophy of Knowledge, 14(4):278, December 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ladriere:1960:LFL

[478] Jean Ladriere. Les limitations des formalismes et leur signification philosophique. (French) [The limitations of formalisms and their philosophical significance]. Dialectica: International Review of Philosophy of Knowledge, 14(4):279–328, December 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Breton:1960:POE

[479] Stanislas Breton. Phénoménologie et ouverture à l’experience. (French) [Phenomenology and openness to experience]. Dialectica: International Review of Philosophy of Knowledge, 14(4):329–364, December 1960. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gex:1960:RCL

[480] Maurice Gex. Reviews — chronique — Literaturblatt: chronique de caractérologie n^{∘} 5. (French) [Chronicle of characterology n^{∘} 5]. Dialectica: International Review of Philosophy of Knowledge, 14(4):365–392, December 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1960:APS

[481] Anonymous. Abstracts: Philosophy of science journal (U.S.A.). Dialectica: International Review of Philosophy of Knowledge, 14(4):393–395, December 1960. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1961:E

[482] Anonymous. Editorial. Dialectica: International Review of Philosophy of Knowledge, 15(1–2):9–11, June 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1961:GB

[483] Anonymous. Gaston Berger. Dialectica: International Review of Philosophy of Knowledge, 15(1–2):13–14, June 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1961:SIF

[484] Anonymous. Séance inaugurale. (French) [Inaugural meeting]. Dialectica: International Review of Philosophy of Knowledge, 15(1–2):15–24, June 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1961:LCC

[485] Anonymous. I. Limites et critères de la connaissance. Thème des entretiens. (French) [I. Limits and criteria of knowledge. Theme of interviews]. Dialectica: International Review of Philosophy of Knowledge, 15(1–2): 25–51, June 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1961:ICS

[486] Anonymous. II. La connaissance scientifique. (French) [II. Scientific knowledge]. Dialectica: International Review of Philosophy of Knowledge, 15(1–2):53–87, June 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1961:ICP

[487] Anonymous. III. Le concept philosophique de Monde. (French) [III. The philosophical concept of Monde]. Dialectica: International Review of Philosophy of Knowledge, 15(1–2):89–127, June 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1961:ICE

[488] Anonymous. IV. La connaissance de l’existence. (French) [IV. The knowledge of existence]. Dialectica: International Review of Philosophy of Knowledge, 15(1–2):129–181, June 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1961:VCL

[489] Anonymous. V. la conscience et lee valeurs. (French) [V. consciousness and values]. Dialectica: International Review of Philosophy of Knowledge, 15(1–2):183–208, June 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1961:VLN

[490] Anonymous. VI. Liberté et nécessité dans l’existence historique. (French) [VI. Freedom and necessity in historical existence]. Dialectica: International Review of Philosophy of Knowledge, 15(1–2):209–227, June 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1961:VNL

[491] Anonymous. Vii. nature, limites et conditions de la connaissance historique. (French) [Vii. nature, limits and conditions of historical knowledge]. Dialectica: International Review of Philosophy of Knowledge, 15 (1–2):229–261, June 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1961:VLC

[492] Anonymous. VIII. Limites et critères de la connaissance métaphysique. (French) [VIII. Limits and criteria of metaphysical knowledge]. Dialectica: International Review of Philosophy of Knowledge, 15(1–2):263–296, June 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1961:ISF

[493] Anonymous. IX. Synthése. (French) [IX. Synthesis]. Dialectica: International Review of Philosophy of Knowledge, 15(1–2):297–318, June 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1961:SCF

[494] Anonymous. Séance de clôture. (French) [Closing session]. Dialectica: International Review of Philosophy of Knowledge, 15(1–2):320–330, June 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1961:CLP

[495] Anonymous. Comités et liste des participants. (French) [Committees and list of participants]. Dialectica: International Review of Philosophy of Knowledge, 15(1–2):332, June 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Perelman:1961:AP

[496] Ch. Perelman. Avant-propos. Dialectica: International Review of Philosophy of Knowledge, 15(3–4):338, December 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dekkers:1961:FDPa

[497] R. Dekkers. Le fait et le droit: problèmes qu’ils posent. (French) [The facts and the law: problems they pose]. Dialectica: International Review of Philosophy of Knowledge, 15(3–4):339–346, December 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dekkers:1961:FDPb

[498] R. Dekkers. Le fait et le droit dans la procédure classique romaine. (French) [Fact and the law in the classical Roman procedure]. Dialectica: International Review of Philosophy of Knowledge, 15(3–4):347–358, December 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Engisch:1961:FDD

[499] Karl Engisch. Le fait et le droit en droit allemand. (French) [Facts and the law in German law]. Dialectica: International Review of Philosophy of Knowledge, 15(3–4):359–382, December 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Foriers:1961:DDF

[500] P. Foriers. La distinction du fait et du droit devant la cour de cassation de Belgique. (French) [The distinction between fact and law before the Supreme Court of Appeal of Belgium]. Dialectica: International Review of Philosophy of Knowledge, 15(3–4):383–409, December 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hallemans:1961:JDD

[501] W. Hallemans. Le juge devant la distinction du fait et du droit. (French) [The judge before the distinction of fact and law]. Dialectica: International Review of Philosophy of Knowledge, 15(3–4):410–420, December 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bayart:1961:DDF

[502] Arnould Bayart. La distinction du fait et du droit. Le point de vue de l’avocat. (French) [The distinction between fact and law. The point of view of the lawyer]. Dialectica: International Review of Philosophy of Knowledge, 15(3–4):421–444, December 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ascareixi:1961:FDD

[503] Tullio Ascareixi. Le fait et le droit devant la cour de cassation italienne. (French) [Facts and law before the Italian Supreme Court of Appeal]. Dialectica: International Review of Philosophy of Knowledge, 15(3–4): 445–461, December 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Riveho:1961:DDD

[504] Jean Riveho. La distinction du droit et du fait dans la jurisprudence du conseil d’état français. (French) [The distinction between law and fact in the jurisprudence of the French State Council]. Dialectica: International Review of Philosophy of Knowledge, 15(3–4):462–480, December 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Buch:1961:DFJ

[505] H. Buch. Le droit et le fait dans la jurisprudence du conseil d’état de Belgique. (French) [The law and the fact in the jurisprudence of the Council of State of Belgium]. Dialectica: International Review of Philosophy of Knowledge, 15(3–4):481–506, December 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Boland:1961:NUJ

[506] G. Boland. La notion d’urgence dans la jurisprudence du conseil d’état de Belgique. (French) [The notion of urgency in the case-law of the Council of State of Belgium]. Dialectica: International Review of Philosophy of Knowledge, 15(3–4):507–522, December 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dumont:1961:CEJ

[507] M. Dumont. Le conseil d’état, juge de l’intéréet général. (French) [The Council of State, a judge of the general interest]. Dialectica: International Review of Philosophy of Knowledge, 15(3–4):523–552, December 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Morgenthal:1961:NSF

[508] L. Morgenthal. La notion de subordination, le fait et le droit. (French) [The notion of subordination, the fact and the right]. Dialectica: International Review of Philosophy of Knowledge, 15(3–4):553–576, December 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Motte:1961:LNE

[509] M. Th. Motte. L’évolution de la notion d’état d’ivresse dans la repression de l’ivresse au volant. (French) [The evolution of the notion of drunkenness in the repression of drunk driving]. Dialectica: International Review of Philosophy of Knowledge, 15(3–4):577–600, December 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Perelman:1961:DDF

[510] Ch. Perelman. La distinction du fait et du droit le point de vue du logicien. (French) [The distinction between fact and law the point of view of the logician]. Dialectica: International Review of Philosophy of Knowledge, 15(3–4):601–610, December 1961. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Feibleman:1962:PES

[511] James K. Feibleman. Philosophical empiricism from the scientific standpoint. Dialectica: International Review of Philosophy of Knowledge, 16 (1):5–14, March 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wittenberg:1962:MPS

[512] A. Wittenberg. May philosophy of science preach empiricism and practice apriorism? Dialectica: International Review of Philosophy of Knowledge, 16(1):15–24, March 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nador:1962:LDH

[513] György Nádor. L’importance de la doctrine heuristique de Descartes dans l’histoire de la science. (French) [The importance of the Descartes heuristic doctrine in the history of science]. Dialectica: International Review of Philosophy of Knowledge, 16(1):25–38, March 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zeleznik:1962:DEL

[514] Carter Zeleznik. Dualistic expressions: learning and instinct. Dialectica: International Review of Philosophy of Knowledge, 16(1):39–44, March 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Moch:1962:PCF

[515] François Moch. Pratique et connaissance. (French) [Practice and knowledge]. Dialectica: International Review of Philosophy of Knowledge, 16 (1):45–55, March 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tummers:1962:GPE

[516] J. H. Tummers. La géométrie et le problème de l’espace. (French) [Geometry and the problem of space]. Dialectica: International Review of Philosophy of Knowledge, 16(1):56–60, March 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Evans:1962:RSC

[517] Melbourne G. Evans. The relativity of simultaneity: a critical analysis. Dialectica: International Review of Philosophy of Knowledge, 16(1): 61–82, March 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bonsack:1962:RME

[518] F. Bonsack. Réponse à M. Evans et à quelques autres. (French) [Answer to M. Evans and a few others]. Dialectica: International Review of Philosophy of Knowledge, 16(1):83–91, March 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mercier:1962:RR

[519] André Mercier. Reviews: reviews. Dialectica: International Review of Philosophy of Knowledge, 16(1):92–95, March 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1962:BJP

[520] Anonymous. British journal for the philosophy of science. Dialectica: International Review of Philosophy of Knowledge, 16(1):96–104, March 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1962:PSJ

[521] Anonymous. Philosophy of science journal: philosophy of science journal. Dialectica: International Review of Philosophy of Knowledge, 16(1):105– 107, March 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1962:RBR

[522] Anonymous. Received books: received books. Dialectica: International Review of Philosophy of Knowledge, 16(1):108–110, March 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Konig:1962:GSF

[523] René König. Grundlagenprobleme der soziologischen Forschungsmethoden. (German) [Basic problems of sociological research methods]. Dialectica: International Review of Philosophy of Knowledge, 16(2):115–142, June 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Baumert:1962:QQV

[524] Gerhart Baumert. Über quantitative und qualitative Verfahren in den Sozialwissenschaften. (German) [On quantitative and qualitative methods in social sciences]. Dialectica: International Review of Philosophy of Knowledge, 16(2):143–154, June 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Compton:1962:US

[525] John J. Compton. Understanding science. Dialectica: International Review of Philosophy of Knowledge, 16(2):155–176, June 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deViller:1962:LEF

[526] A. Bonnet de Viller. L’envers de l’expérience. (French) [The reverse of the experience]. Dialectica: International Review of Philosophy of Knowledge, 16(2):177–197, June 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Duvanel:1962:RCS

[527] Blaise Duvanel. Reviews: chronique de sociologie. (French) [Chronicle of sociology]. Dialectica: International Review of Philosophy of Knowledge, 16(2):199–201, June 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Duvanel:1962:RRB

[528] B. Duvanel. Reviews: Roger Bastide: Les religions africaines au Brésil. Dialectica: International Review of Philosophy of Knowledge, 16(2):201– 203, June 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Duvanel:1962:RTS

[529] B. Duvanel. Reviews: traité de sociologie. (French) [Treatise of sociology]. Dialectica: International Review of Philosophy of Knowledge, 16 (2):204–207, June 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Topitsch:1962:VZS

[530] Ernst Topitsch. Das Verhältnis zwischen Sozial- und Naturwissenschaften : eine methodologisch-ideologiekritische Untersuchung. (German) [The relationship between social and natural sciences: a methodological– ideological critical investigation]. Dialectica: International Review of Philosophy of Knowledge, 16(3):211–231, September 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lenk:1962:PI

[531] Kurt Lenk. Die positivistische ideologienlehre. Dialectica: International Review of Philosophy of Knowledge, 16(3):232–254, September 1962. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stern:1962:IMP

[532] Herold S. Stern. Implications of the methodology of the physical sciences for the social sciences. Dialectica: International Review of Philosophy of Knowledge, 16(3):255–274, September 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tatarkiewicz:1962:MSD

[533] Krzysztof Tatarkiewicz. Modèles et systèmes déductifs. (French) [Deductive models and systems]. Dialectica: International Review of Philosophy of Knowledge, 16(3):275–298, September 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1962:RVR

[534] Anonymous. Relativité de la vitesse relative : suite et fin de la discussion Evans–Bonsack. (French) [Relativity of relative velocity: continuation and end of Evans–Bonsack discussion]. Dialectica: International Review of Philosophy of Knowledge, 16(3):299–301, September 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1962:APS

[535] Anonymous. Abstracts from · Philosophy of Science ·. Dialectica: International Review of Philosophy of Knowledge, 16(3):302–303, September 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Noblle-Neumann:1962:MFU

[536] Dr. Elisabeth Noblle-Neumann. Über den methodischen Fortschritt in der Umfrageforschung. (German) [On methodical progress in survey research]. Dialectica: International Review of Philosophy of Knowledge, 16(4):307–328, December 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hofstatter:1962:ASD

[537] Peter R. Hofstätter. Analyse eines Stereotyps: die deutsche Frau. (German) [Analysis of a stereotype: the German woman]. Dialectica: International Review of Philosophy of Knowledge, 16(4):329–353, December 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Walter:1962:VCS

[538] Emile J. Walter. Vers une classification scientifique de la socioloigie. (French) [Towards a scientific classification of socioloigy]. Dialectica: International Review of Philosophy of Knowledge, 16(4):354–360, December 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Maruyama:1962:POM

[539] Magoroh Maruyama. Philosophy as an open meta-science of interdisciplinary cross-induction. Dialectica: International Review of Philosophy of Knowledge, 16(4):361–384, December 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bonsack:1962:PIO

[540] François Bonsack. Pour une interprétation objectiviste de la théorie de l’information. (French) [For an objectivist interpretation of the theory of information]. Dialectica: International Review of Philosophy of Knowledge, 16(4):385–395, December 1962. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pap:1963:NKW

[541] Arthur Pap. Neuere Kernfragen der Wahrscheinlichkeitstheorie. (German) [Recent core probabilities of probability theory]. Dialectica: International Review of Philosophy of Knowledge, 17(1):5–22, March 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Harris:1963:EP

[542] Errol E. Harris. The end of a phase. Dialectica: International Review of Philosophy of Knowledge, 17(1):23–48, March 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See remarks [543].

Bernays:1963:REP

[543] Paul Bernays. Remarks to “The end of a phase”. Dialectica: International Review of Philosophy of Knowledge, 17(1):49–50, March 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See [542].

Emery:1963:PFM

[544] Eric Emery. Propos sur les fondements de la musique. (French) [On the fundamentals of music]. Dialectica: International Review of Philosophy of Knowledge, 17(1):51–77, March 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Maruyama:1963:BEMa

[545] Magoroh Maruyama. Basic elements in misunderstandings. I. Dialectica: International Review of Philosophy of Knowledge, 17(1):78–92, March 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1963:ABJ

[546] Anonymous. Abstract: British Journal for the Philosophy of Science. Dialectica: International Review of Philosophy of Knowledge, 17(1):93–96, March 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Maruyama:1963:BEMb

[547] Magoroh Maruyama. Basic elements in misunderstandings. II. Dialectica: International Review of Philosophy of Knowledge, 17(2–3):99–109, September 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1963:LTQ

[548] Anonymous. L’analogie en tant que méthode de connaissance: introduction: extrait d’une lettre à Mm. Cuénod et Pun. (French) [Analogy as a method of knowledge: introduction: extract from a letter to Mm. Cuénod and Pun]. Dialectica: International Review of Philosophy of Knowledge, 17(2–3):111–115, September 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cuenod:1963:RMG

[549] M. Cuénod and L. Pun. Réponse à M. Gonseth. (French) [Reply to M. Gonseth]. Dialectica: International Review of Philosophy of Knowledge, 17(2–3):116–118, September 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1963:PCA

[550] F. Gonseth. I. Principes de la connaissance analogique: analogie et modéles mathématiques. (French) [I. Principles of analogue knowledge: analogy and mathematical models]. Dialectica: International Review of Philosophy of Knowledge, 17(2–3):119–150, September 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Loeb:1963:PQM

[551] J. Loeb. Par quel mécanisme l’analogie est-elle une source de connaissance?. (French) [By what mechanism is analogy a source of knowledge?]. Dialectica: International Review of Philosophy of Knowledge, 17(2–3): 151–158, September 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Henning:1963:ACP

[552] Edward Henning and Mihajlo D. Mesarovic. Analogy in the creative processes and the objects of creation in art and sciences. Dialectica: International Review of Philosophy of Knowledge, 17(2–3):159–166, September 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pask:1963:UAP

[553] Gordon Pask. The use of analogy and parable in cybernetics with emphasis upon analogies for learning and creativity. Dialectica: International Review of Philosophy of Knowledge, 17(2–3):167–203, September 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Raymond:1963:AIA

[554] F. H. Raymond. Analogies et intelligence artificielle. (French) [Analogies and artificial intelligence]. Dialectica: International Review of Philosophy of Knowledge, 17(2–3):203–215, September 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Naslin:1963:IMC

[555] P. Naslin. II. Méthodes de la connaissance analogique: analogies, homologies et modéles. (French) [II. Analogical methods: analogies, homologies and models]. Dialectica: International Review of Philosophy of Knowledge, 17(2–3):215–239, September 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pun:1963:OAF

[556] L. Pun. Optimum de l’analogie. (French) [Optimum of analogy]. Dialectica: International Review of Philosophy of Knowledge, 17(2–3):240–259, September 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zelbstein:1963:MAI

[557] U. Zelbstein. Méthodes analogiques et isomorphisme des chaines fonctionnelles. (French) [Analog methods and isomorphism of functional chains]. Dialectica: International Review of Philosophy of Knowledge, 17 (2–3):260–266, September 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cuenod:1963:IAM

[558] Michel Cuénod. III. Application des méthodes analogiques: analogie et notion de rigueur pour l’ingénieur. (French) [III. Application of analog methods: analogy and rigor for the engineer]. Dialectica: International Review of Philosophy of Knowledge, 17(2–3):267–283, September 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hesse:1963:ACT

[559] Mary Hesse. Analogy and confirmation theory. Dialectica: International Review of Philosophy of Knowledge, 17(2–3):284–292, September 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Edeline:1963:ALD

[560] F. Edeline. Aspects logiques du jeu de mots poétique. (French) [Logical aspects of the poetic play on words]. Dialectica: International Review of Philosophy of Knowledge, 17(4):297–306, December 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zafiropulo:1963:RAD

[561] Jean Zafiropulo. Le rôle de l’analogie dans le déchiffrement de l’écriture mycénienne linéaire B. (French) [The role of analogy in deciphering Mycenaean linear script B]. Dialectica: International Review of Philosophy of Knowledge, 17(4):307–327, December 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Solari:1963:CPM

[562] Luigi Solari. Contenu et portée des modèles économétriques. (French) [Content and scope of econometric models]. Dialectica: International Review of Philosophy of Knowledge, 17(4):328–352, December 1963. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

VonGrunebaum:1963:MWV

[563] G. E. Von Grunebaum. Muslim world view and Muslim science. Dialectica: International Review of Philosophy of Knowledge, 17(4):353–367, December 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chari:1963:ICT

[564] C. T. K. Chari. Information, cosmology and time. Dialectica: International Review of Philosophy of Knowledge, 17(4):368–380, December 1963. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Perelman:1964:AP

[565] Ch. Perelman. Avant-propos. Dialectica: International Review of Philosophy of Knowledge, 18(1–4):6, December 1964. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Malgaud:1964:ADP

[566] W. Malgaud. Les antinomies en droit, à propos de l’étude de G. Gavazzi. (French) [The antinomies in law, concerning the study of G. Gavazzi]. Dialectica: International Review of Philosophy of Knowledge, 18(1–4): 7–19, December 1964. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Foriers:1964:AD

[567] Paul Foriers. Les antinomies en droit. (French) [Antinomies in law]. Dialectica: International Review of Philosophy of Knowledge, 18(1–4):

20–38, December 1964. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Morgenthal:1964:ADS

[568] Léopold Morgenthal. Les antinomies en droit social. (French) [The antinomies in social law]. Dialectica: International Review of Philosophy of Knowledge, 18(1–4):39–62, December 1964. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Silance:1964:QEA

[569] Lue. Silance. Quelques exemples d’antinomies et essai de classement. (French) [Some examples of antinomies and ranking test]. Dialectica: International Review of Philosophy of Knowledge, 18(1–4):63–137, December 1964. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

VanderElst:1964:ADI

[570] Raymond Vander Elst. Antinomies en droit international privé. (French) [Antinomies in private international law]. Dialectica: International Review of Philosophy of Knowledge, 18(1–4):138–176, December 1964. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hoeffler:1964:ADP

[571] Jacques Hoeffler. Les antinomies en droit public. (French) [Antinomies in public law]. Dialectica: International Review of Philosophy of Knowledge, 18(1–4):177–182, December 1964. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Boland:1964:QPA

[572] Georges Boland. Quelques propos sur les antinomies et pseudoantinomies en particulier en droit administratif. (French) [Some remarks on antinomies and pseudo-antinomies in particular in administrative law]. Dialectica: International Review of Philosophy of Knowledge, 18 (1–4):183–203, December 1964. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Huberlant:1964:ARP

[573] Charles Huberlant. Antinomies et recours aux principes généraux. (French) [Antinomies and use of general principles]. Dialectica: International Review of Philosophy of Knowledge, 18(1–4):204–236, December 1964. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bobbio:1964:CPR

[574] N. Bobbio. Des critères pour résoudre les antinomies. (French) [Criteria for resolving antinomies]. Dialectica: International Review of Philosophy of Knowledge, 18(1–4):237–258, December 1964. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Miedzianogoba:1964:LNA

[575] J. Miedzianogoba. L’adage nemo auditur propriam turpitudinem allegans et le code napoléon. (French) [The adage nemo auditur propriam turpitudinem allegans and the Napoleon Code]. Dialectica: International Review of Philosophy of Knowledge, 18(1–4):259–284, December 1964. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Salmon:1964:ADI

[576] Jean J.-A. Salmon. Les antinomies en droit international public. (French) [Antinomies in international public law]. Dialectica: International Review of Philosophy of Knowledge, 18(1–4):285–319, December 1964. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Foriers:1964:AED

[577] Paul Foriers. Les antinomies entre dispositions de droit communautaire et dispositions de droit interne. (French) [The antinomies between provisions of community law and provisions of domestic law]. Dialectica: International Review of Philosophy of Knowledge, 18(1–4):320–336, December 1964. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tammelo:1964:TTT

[578] Ilmar Tammelo. Tensions and tenebrae in treaty interpretation: reconnaissance of a battlefield for antinomies. Dialectica: International Review of Philosophy of Knowledge, 18(1–4):337–350, December 1964. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Szabo:1964:CED

[579] Inure Szabó. Des contradictions entre le droit des différents systèmes sociaux. (French) [Contradictions between the laws of different social systems]. Dialectica: International Review of Philosophy of Knowledge, 18(1–4):351–371, December 1964. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Buch:1964:CDA

[580] H. Buch. Conception dialectique des antinomies juridiques. Dialectica: International Review of Philosophy of Knowledge, 18(1–4):372–391, December 1964. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Perelman:1964:ADE

[581] Ch. Perelman. Les antinomies en droit: essai de synthèse. (French) [Antinomies in law: a synthesis essay]. Dialectica: International Review of Philosophy of Knowledge, 18(1–4):392–404, December 1964. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wittenberg:1965:I

[582] A. Wittenberg and F. Gonseth. Introduction. Dialectica: International Review of Philosophy of Knowledge, 19(1–2):6–15, June 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wittenberg:1965:QEM

[583] A. Wittenberg. Questionnaire sur l’enseignement des mathématiques. (French) [Mathematics education questionnaire]. Dialectica: International Review of Philosophy of Knowledge, 19(1–2):16–26, June 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sorensen:1965:MRF

[584] W. Sörensen. Un modèle de réponse. (French) [A response model]. Dialectica: International Review of Philosophy of Knowledge, 19(1–2):27– 33, June 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sorensen:1965:PRE

[585] W. Sörensen. À propos de la réforme de l’enseignement des mathématiques. (French) [On reforming mathematics education]. Dialectica: International Review of Philosophy of Knowledge, 19(1–2):34–39, June 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Goneth:1965:LP

[586] F. Goneth. L’homo phenomenologicus. Dialectica: International Review of Philosophy of Knowledge, 19(1–2):40–69, June 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Konig:1965:GEE

[587] Hans König. Über die Grenzen der exakten Erfassung der Qualität — ≪Hell^{≫}. (German) [On the limits of the exact detection of quality — “Hell”]. Dialectica: International Review of Philosophy of Knowledge, 19(1–2):70–90, June 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deViller:1965:SPF

[588] A. Bonnet de Viller. Science et philosophie. (French) [Science and philosophy]. Dialectica: International Review of Philosophy of Knowledge, 19(1–2):91–135, June 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Guggenheimer:1965:PRN

[589] H. Guggenheimer. The psychology of real numbers. Dialectica: International Review of Philosophy of Knowledge, 19(1–2):136–143, June 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nador:1965:SMD

[590] György Nädor. Scientia mirabilis: Descartes et Leibniz. (French) [Wonderful science: Descartes and Leibniz]. Dialectica: International Review of Philosophy of Knowledge, 19(1–2):144–157, June 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Staal:1965:WB

[591] J. F. Staal. E. W. Beth (1908–1964). Dialectica: International Review of Philosophy of Knowledge, 19(1–2):158–184, June 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1965:RBRa

[592] Anonymous. Received books: received books. Dialectica: International Review of Philosophy of Knowledge, 19(1–2):181–184, June 1965. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1965:APSa

[593] Anonymous. Abstracts: philosophy of science. Dialectica: International Review of Philosophy of Knowledge, 19(1–2):185–192, June 1965. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1965:CFC

[594] Anonymous. Congrès. (French) [Congress]. Dialectica: International Review of Philosophy of Knowledge, 19(1–2):192, June 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bunge:1965:PR

[595] Mario Bunge. Physics and reality. Dialectica: International Review of Philosophy of Knowledge, 19(3–4):195–222, December 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fock:1965:PQI

[596] V. Fock. La physique quantique et les idéalisations classiques. (French) [Quantum physics and classical idealizations]. Dialectica: International Review of Philosophy of Knowledge, 19(3–4):223–245, December 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Park:1965:IPM

[597] David Park. The idea of a particle in mictrophysics. Dialectica: International Review of Philosophy of Knowledge, 19(3–4):246–258, December 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Emch:1965:SLM

[598] Gérard Emch and Josef Maria Jauch. Structures logiques et mathématiques en physique quantique. (French) [Logical and mathematical structures in quantum physics]. Dialectica: International Review of Philosophy of Knowledge, 19(3–4):259–279, December 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deBeauregard:1965:PCE

[599] O. Costa de Beauregard. Le · paradoxe · des correlations d’Einstein: et de Schrödinger et l’épaisseur temporelle de la transition quantique. (French) [The · paradox · of the correlations of Einstein: and Schrödinger and the temporal thickness of the quantum transition]. Dialectica: International Review of Philosophy of Knowledge, 19(3–4):280–289, December 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cochran:1965:LWP

[600] Andrew A. Cochran. Life and the wave properties of matter. Dialectica: International Review of Philosophy of Knowledge, 19(3–4):290–312, December 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1965:MDC

[601] Ferdinand Gonseth. Sur la méthodologie du calcul des probabilités. (French) [On the methodology of calculating probabilities]. Dialectica: International Review of Philosophy of Knowledge, 19(3–4):313–328, December 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bonsack:1965:QSP

[602] F. Bonsack. En quel sens les probabilités sont-elles subjectives?. (French) [In what sense are probabilities subjective?]. Dialectica: International Review of Philosophy of Knowledge, 19(3–4):329–344, December 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Destouches:1965:GMP

[603] J. L. Destouches. General mathematical physics and schemas, application to the theory of particles. Dialectica: International Review of Philosophy of Knowledge, 19(3–4):345–348, December 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lande:1965:NQF

[604] Alfred Landé. Non-quantal foundations of quantum mechanics. Dialectica: International Review of Philosophy of Knowledge, 19(3–4):349–357, December 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1965:RBRb

[605] Anonymous. Received books: received books. Dialectica: International Review of Philosophy of Knowledge, 19(3–4):358–361, December 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1965:APSb

[606] Anonymous. Abstracts from Philosophy of Science. Dialectica: International Review of Philosophy of Knowledge, 19(3–4):362–363, December 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1965:ABJ

[607] Anonymous. Abstracts from: British Journal of Philosophy of Science. Dialectica: International Review of Philosophy of Knowledge, 19(3–4): 364–365, December 1965. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1966:HAP

[608] Paul Emile Pilet. Hommage au professeur Ferdinand Gonseth. Fondateur et directeur de “Dialectica”. (French) [Homage to Professor Ferdinand Gonseth. Founder and director of “Dialectica”]. Dialectica: International Review of Philosophy of Knowledge, 20(1):3–4, March 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1966:E

[609] Ferdinand Gonseth. Editorial. Dialectica: International Review of Philosophy of Knowledge, 20(1):5–13, March 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bouligand:1966:PSU

[610] G. Bouligand. Le philosophe et le savant unis dans l’œuvre épistémologique. hommage au fondateur de Dialectica. (French) [The philosopher and the scientist united in the epistemological work. Tribute to the founder of Dialectica]. Dialectica: International Review of Philosophy of Knowledge, 20(1):14–26, March 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1966:GBB

[611] Paul Bernays. Gedanken zu dem Buch Bildung und Mathematik (Mathematik als exemplarisches Gymnasialfach) von Alexander Israel Wittenberg. (German) [Thoughts on the book Education and Mathematics (Mathematics as an Exemplary Gymnasium Subject) by Alexander Israel Wittenberg]. Dialectica: International Review of Philosophy of Knowledge, 20(1):27–42, March 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1966:LB

[612] Paul Emile Pilet. L’analogie en biologie. (French) [Analogy in biology]. Dialectica: International Review of Philosophy of Knowledge, 20(1):43– 50, March 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Konig:1966:EDK

[613] H. König. Über ein einfaches Drei-Komponenten-Modell für das menschliche Denken. (German) [About a simple three-component model for human thinking]. Dialectica: International Review of Philosophy of Knowledge, 20(1):51–66, March 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chari:1966:ITQ

[614] C. T. K. Chari. Information theory, quantum mechanics and‘linguistic duality’. Dialectica: International Review of Philosophy of Knowledge, 20(1):67–88, March 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1966:PDL

[615] F. Gonseth. Le problème du langage et la philosophic ouverte. (French) [The problem of language and the open philosophic]. Dialectica: International Review of Philosophy of Knowledge, 20(1):89–105, March 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1966:D

[616] Anonymous. Dialectica. Dialectica: International Review of Philosophy of Knowledge, 20(2):129, June 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1966:CFD

[617] F. Gonseth. Comment fonder une discipline exacte. (French) [How to build an exact discipline]. Dialectica: International Review of Philosophy of Knowledge, 20(2):131–136, June 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Destouches:1966:OPI

[618] J.-L. Destouches. Observation, prévision, invention, objectivité, réalité dans les sciences ayant acquis une forme théorique. (French) [Observation, forecasting, invention, objectivity, reality in the sciences having acquired a theoretical form]. Dialectica: International Review of Philosophy of Knowledge, 20(2):137–142, June 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Born:1966:SR

[619] M. Born. Symbol and reality. Dialectica: International Review of Philosophy of Knowledge, 20(2):143–157, June 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dubarle:1966:ORC

[620] D. Dubarle. Objectivité et réalité dans le cas de la physique probabiliste. (French) [Objectivity and reality in the case of probabilistic physics]. Dialectica: International Review of Philosophy of Knowledge, 20(2):158– 173, June 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bunge:1966:PR

[621] Mario Bunge. Physics and reality. Dialectica: International Review of Philosophy of Knowledge, 20(2):174–195, June 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wolvekamp:1966:COI

[622] H. P. Wolvekamp. The concept of the organism as an integrated whole. Dialectica: International Review of Philosophy of Knowledge, 20(2):196– 214, June 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ladriegre:1966:ORM

[623] Jean Ladriegre. Objectivité et réalité en mathématiques. (French) [Objectivity and reality in mathematics]. Dialectica: International Review of Philosophy of Knowledge, 20(2):215–241, June 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1966:RSE

[624] Anonymous. Reviewing studies — études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 20(2):242–243, June 1966. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1967:E

[625] Anonymous. Editorial. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):5–7, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1967:ADC

[626] Anonymous. Actes du colloque international sur la réforme et la coordination des enseignements de mathématique et de physique, Lausanne, janvier 1967. (French) [Proceedings of the International Symposium on the Reform and Co-ordination of Mathematics and Physics, Lausanne, January 1967]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):8–9, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Grivet:1967:I

[627] Th. Grivet. I. Introduction. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):10–13, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Knecht:1967:IRG

[628] W. Knecht. II. Rapport général. (French) [II. General report]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):14–25, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Knecht:1967:GRB

[629] W. Knecht. General report. the background of the symposium. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):26–37, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wittenberg:1967:IEP

[630] A. Wittenberg. III. Enquêtes préparatoires et réponses: 1. Enquête préalable sur l’enseignement des mathématiques. (French) [III. preliminary investigations and responses: 1. Preliminary survey of mathematics education]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):38–46, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wittenberg:1967:QTM

[631] A. Wittenberg. A. Questionnaire on the teaching of mathematics. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):46– 54, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chaux-de-Fonds:1967:BRE

[632] La Chaux de Fonds and E. Emery. B. Réponses à l’enquête sur l’enseignement des mathématiques. B. Replies on the teaching of mathematics — une contribution au débat. (French) [B. Responses to the survey on mathematics education. B. Replies on the teaching of mathematics — a contribution to the debate]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):54–69, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Servais:1967:PAE

[633] W. Servais. Problèmes actuels de l’enseignement de la mathématique. (French) [Current problems in mathematics education]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):70–88, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Guggenheimer:1967:CVR

[634] H. Guggenheimer. A conservative’s view of the reform of school mathematics. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):88–91, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Felix:1967:EAM

[635] Mlle L. Félix. Sur l’enseignement actuel des mathématiques élémentaires. (French) [On the current teaching of elementary mathematics]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):92–103, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Freudenthal:1967:RFRa

[636] H. Freudenthal. Réponse. (French) [Response]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):103–105, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Calame:1967:PRE

[637] Andréa Calame. À propos d’une réforme de l’enseignement des mathématiques à Neuchâtel. (French) [About a reform of mathematics education in Neuchâtel]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):105–109, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Suter:1967:REM

[638] H. Suter. La réforme de l’enseignement des mathématiques à Neuchâtel. (French) [The reform of teaching of mathematics at Neuchâtel]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):109–116, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Traupel:1967:BFG

[639] Prof. Dr. W. Traupel. Bemerkungen zum Fragebogen von Gonseth und Wittenberg über den Mathematikunterricht an Mittelschulen. (German) [Remarks on the questionnaire of Gonseth and Wittenberg on mathematics teaching in middle school]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):116–124, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zweifel:1967:BFW

[640] O. Zweifel. Beantwortung des Fragebogens von A. Wittenberg betreffend den Mathematikunterricht. (German) [Answer to A. Wittenberg’s questionnaire concerning mathematics teaching]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):125–131, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1967:ERE

[641] Anonymous. 2. Enquête sur la réforme de l’enseignement de la physique. 2. Preliminary enquiry on the teaching of physics. (French) [2. Inquiry on the reform of the teaching of physics. 2. Preliminary investigation of the teaching of physics]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):132–134, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chaundy:1967:BRE

[642] D. C. F. Chaundy. B. Réponses à l’enquête sur l’enseignement de la physique B. Replies on the teaching on physics: Dialectica and GIREP enquiry on the teaching of physics. (French) [B. Answers to the survey on the teaching of physics B. Replies on the teaching of physics: Dialectica GIREP]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):135–136, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hecht:1967:Aa

[643] Dr. K. Hecht. Antwort. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):137–138, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Little:1967:RQT

[644] Prof. R. N. Little. Response to questionnaire on the teaching of physics. Dialectica: International Review of Philosophy of Knowledge, 21(1–4): 138–140, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mathieu:1967:RAQ

[645] Prof. J. P. Mathieu. Réponses au questionnaire sur l’enseignement de la physique. (French) [Responses to the questionnaire on the teaching of physics]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):140–142, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Messel:1967:Ra

[646] H. Messel. Reply. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):142–143, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1967:QEP

[647] Prof. J. Muller. Questionnaire sur l’enseignement de la physique au niveau secondaire. (French) [Questionnaire on the teaching of physics at the secondary level]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):143–146, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Soren:1967:QTP

[648] Prof. Søren and Prof. S. Sikjaer. Questionnaire on the teaching of physics. Dialectica: International Review of Philosophy of Knowledge, 21(1–4): 146–147, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stollberg:1967:STT

[649] Robert Stollberg. Some thoughts on the teaching of secondary school physics concerning unification of secondary school physics. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):148–152, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

VanHove:1967:RAQ

[650] Prof. L. Van Hove. Réponses au questionnaire sur l’enseignement de la physique. (French) [Responses to the questionnaire on the teaching of physics]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):153–156, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Badziag:1967:BRE

[651] Kazimierz Badziag. B. Réponses de l’enquête sur l’enseignement de mathématique et de physique. B. Replies on the teaching of mathematics and physics reply to the questionnaire. (French) [B. Responses of the survey on mathematics and physics teaching. B. Replies on the teaching of mathematics and physics]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):157–158, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bass:1967:RFR

[652] J. Bass. Réponse. (French) [Response]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):158–160, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Baurmann:1967:RAQ

[653] Dr. Erwin Baurmann. Réponse au questionnaire. (French) [Response to the questionnaire]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):160–162, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ferretti:1967:RFR

[654] Maria Ferretti. Réponse. (French) [Response]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):162–164, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Freudenthal:1967:RFRb

[655] H. Freudenthal. Réponse. (French) [Response]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):164–165, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Friskopp:1967:AMP

[656] K. G. Friskopp. On the articulation of mathematics and physics teaching. Dialectica: International Review of Philosophy of Knowledge, 21(1–4): 166–167, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hecht:1967:Ab

[657] Dr. K. Hecht. Antwort. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):167–168, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jauch:1967:RFR

[658] J. M. Jauch. Réponse. (French) [Response]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):168–170, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jeger:1967:RAQ

[659] M. Jeger and M. Rueff. Réponse au questionnaire sur l’articulation des enseignements de mathématique et de physique. (French) [Response to the questionnaire on the articulation of mathematics and physics]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):170– 174, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kourganoff:1967:RQQ

[660] V. Kourganoff. Réponse à quelques questions. (French) [Response to some questions]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):175–179, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kranzer:1967:RQ

[661] Dr. Walter Kranzer. Reply to the questionnaire. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):179–181, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Krygowska:1967:RFR

[662] A. Z. Krygowska. Réponse. (French) [Response]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):181–185, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lichtenberg:1967:A

[663] Jonas Lichtenberg. Answer. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):185–188, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lohisse:1967:RFR

[664] J. Lohisse, M. Delhiere, and A. Festraets. Réponse. (French) [Response]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4): 188–190, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mercier:1967:RFR

[665] Professeur R. Mercier. Réponse. (French) [Response]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):190–191, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Messel:1967:Rb

[666] Professor H. Messel. Reply. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):192–193, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pjeriskin:1967:RAQ

[667] Prof. A. V. Pjeriskin. Réponses au questionnaire sur les méthodes d’enseignement de la physique. (French) [Responses to the questionnaire on methods of teaching physics]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):194–197, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pisot:1967:RFR

[668] Ch. Pisot. Réponse. (French) [Response]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):197–200, December 1967. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1967:RSM

[669] Anonymous. Reply of the School Mathematics Study Group School of Education (USA). Dialectica: International Review of Philosophy of Knowledge, 21(1–4):200–202, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

VanHove:1967:RFR

[670] L. Van Hove. Réponse. (French) [Response]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):202–203, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Delessert:1967:IRR

[671] M. A. Delessert. IV. Rapports sur les résultats des enquêtes et remarques. 1. enquête sur l’enseignement de la mathématique rapport sur les réponses au questionnaire concernant l’enseignement mathématique. (French) [IV. Reports on survey results and remarks. 1. inquiry on the teaching of mathematics report on the answers to the questionnaire concerning mathematical education]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):204–214, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1967:I

[672] du Professeur F. Gonseth. Intervention. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):215, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Emery:1967:I

[673] M. E. Emery. Intervention. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):216–217, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Krygowska:1967:I

[674] Mme A. Z. Krygowska. Intervention. I. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):217–219, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Krygowska:1967:II

[675] Mme A. Z. Krygowska. Intervention. II. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):220–222, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Freudenthal:1967:R

[676] H. Freudenthal. Remarques. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):222–223, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Knecht:1967:EEP

[677] M. W. Knecht. 2. Enquête sur l’enseignement de la physique rapport sur les réponses au questionnaire de physique. (French) [2. Inquiry on the teaching of physics report on physics questionnaire responses]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):224–241, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fuka:1967:R

[678] M. Fuka, Mme M. Ferretti, M. H. Messel, and M. J. Smolec. Remarques. Dialectica: International Review of Philosophy of Knowledge, 21(1–4): 241–244, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pisot:1967:EAE

[679] M. Ch. Pisot. 3. enquête sur l’articulation des enseignements de mathématique et de physique rapport sur les réponses au questionnaire sur l’articulation des enseignements de mathématique et de physique. (French) [3. inquiry on the articulation of mathematics and physics lessons on the answers to the questionnaire on the articulation of mathematics and physics]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):245–255, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Messel:1967:Rc

[680] M. Harry Messel and M. H. Freudenthal. Remarque. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):255–257, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1967:RPF

[681] Anonymous. Résolutions, propositions finales et remarques — Resolutions, final proposals and remarks: 1. Propositions finales des participant à propos de l’enseignement mathématique. (French) [1. Participants’ final proposals of mathematical education]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):258–266, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Servais:1967:R

[682] M. W. Servais. Remarques. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):267, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1967:VCF

[683] F. Gonseth. VI. Considérations finales. (French) [VI. Final considerations]. Dialectica: International Review of Philosophy of Knowledge, 21 (1–4):268–276, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Orear:1967:VCL

[684] Jay Orear. VII. Conférences — lectures: The new physics. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):277–291, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Messel:1967:RMH

[685] H. Messel. Relativity and magnetism for high school students. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):292–310, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Servais:1967:IIE

[686] W. Servais. Introduction de l’intégrale dans l’enseignement secondaire. (French) [Introduction of the integral in secondary education]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):311–357, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bunt:1967:NTI

[687] Lucas N. H. Bunt. Notes on the teaching of integral calculus. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):358–365, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bunt:1967:PSI

[688] Lucas N. H. Bunt. Probability and statistical inference in the secondary school. Dialectica: International Review of Philosophy of Knowledge, 21 (1–4):366–382, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Smolec:1967:MME

[689] Ignaci Smolec. Modèles mathématiques dans l’enseignement de la physique. (French) [Mathematical models in the teaching of physics]. Dialectica: International Review of Philosophy of Knowledge, 21(1–4): 383–391, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1967:VA

[690] Anonymous. VIII. appendices. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):392–396, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1967:RSE

[691] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):397–398, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1967:RBL

[692] Anonymous. Received books — livres reçus — Eingelaufene Bücher. Dialectica: International Review of Philosophy of Knowledge, 21(1–4): 398–399, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1967:ABJ

[693] Anonymous. Abstracts from the ‘British Journal for the Philosophy of Science’. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):400–403, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Farber:1967:PPR

[694] Marvin Farber. Philosophy and phenomenologigal research. Dialectica: International Review of Philosophy of Knowledge, 21(1–4):404, December 1967. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rosenthal:1968:CST

[695] Sandra B. Rosenthal and Sandra B. Rosenthal. The cognitive status of theoretical terms. Dialectica: International Review of Philosophy of Knowledge, 22(1):3–17, March 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1968:RPD

[696] P. Bernays. Remarks to the paper by Dr Sandra Rosenthal: ‘The Cognitive Status of Theoretical Terms’. Dialectica: International Review of Philosophy of Knowledge, 22(1):18–19, March 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gruender:1968:CF

[697] David Gruender. Constructs and fictions. Dialectica: International Review of Philosophy of Knowledge, 22(1):20–27, March 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lee:1968:SMS

[698] Donald S. Lee and Donald S. Lee. Scientific method as a stage process. Dialectica: International Review of Philosophy of Knowledge, 22 (1):28–44, March 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Beard:1968:PAI

[699] Robert W. Beard and Robert W. Beard. The paradoxes of analysis and identity. Dialectica: International Review of Philosophy of Knowledge, 22 (1):45–46, March 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1968:ERE

[700] Anonymous. Enquêtes sur la réforme des enseignements de mathématiques, de physique et de biologie. (French) [Inquiries on reform of mathematics, physics and biology]. Dialectica: International Review of Philosophy of Knowledge, 22(1):57, March 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wittenberg:1968:PRR

[701] Alexander Wittenberg and Alexander Wittenberg. Priorities and responsibilities in the reform of mathematical education: An essay in educational meta-theory. Dialectica: International Review of Philosophy of Knowledge, 22(1):58–74, March 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Papy:1968:MMF

[702] G. Papy. Mathématique moderne. (French) [Modern mathematics]. Dialectica: International Review of Philosophy of Knowledge, 22(1):75–78, March 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1968:RBLa

[703] Anonymous. Received books — livres Reçus — eingelaufene Bücher. Dialectica: International Review of Philosophy of Knowledge, 22(1):79–81, March 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1968:AFS

[704] Anonymous. Abstracts from‘Philosophy of science’. Dialectica: International Review of Philosophy of Knowledge, 22(1):82–87, March 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1968:ATJa

[705] Anonymous. Abstracts from the‘British journal for the philosophy of science’. Dialectica: International Review of Philosophy of Knowledge, 22(1):88, March 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1968:D

[706] P. Bernays, F. Gonseth, H. König, P. Nolfi, and E. Pilet. dialectica. Dialectica: International Review of Philosophy of Knowledge, 22(1):89, March 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1968:E

[707] Anonymous. Editorial. Dialectica: International Review of Philosophy of Knowledge, 22(2):91–95, June 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hubien:1968:PAL

[708] Hubert Hubien. Philosophic analytique et linguistique moderne. (French) [Modern analytical and linguistic philosophy]. Dialectica: International Review of Philosophy of Knowledge, 22(2):96–119, June 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Starobinski:1968:MDD

[709] J. Starobinski. Montaigne et la dénonciation du mensonge. (French) [Montaigne and the denunciation of the lie]. Dialectica: International Review of Philosophy of Knowledge, 22(2):120–131, June 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cantoee:1968:SPS

[710] Enrico Cantoee. Science and philosophy. Some reflections on Man’s unending quest for understanding. Dialectica: International Review of Philosophy of Knowledge, 22(2):132–166, June 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1968:ATJb

[711] Anonymous. Abstracts from the‘British journal for the philosophy of science’. Dialectica: International Review of Philosophy of Knowledge, 22(2):167–168, June 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1968:SFS

[712] F. Gonseth. Stratégic de fondement et stratégic d’engagement. (French) [Strategic foundation and strategic engagement]. Dialectica: International Review of Philosophy of Knowledge, 22(3–4):171–186, December 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deBeauregard:1968:TIL

[713] O. Costa de Beauregard. On time, information and life. Dialectica: International Review of Philosophy of Knowledge, 22(3–4):187–205, December 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1968:TSF

[714] F. Gonseth. Temps et syntaxe. (French) [Time and syntax]. Dialectica: International Review of Philosophy of Knowledge, 22(3–4):206–213, December 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Knecht:1968:TMC

[715] Yvonne Knecht. Temps et modes chez Spinoza. (French) [Time and mode for Spinoza]. Dialectica: International Review of Philosophy of Knowledge, 22(3–4):214–237, December 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chauchard:1968:ICP

[716] Paul Chauchard. Importance de la connaissance psychophysiologique du langage. (French) [Importance of psychophysiological knowledge of language]. Dialectica: International Review of Philosophy of Knowledge, 22(3–4):238–246, December 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sauvan:1968:SMS

[717] T. Sauvan. Simulation de mecanismes sémantiques. (French) [Simulation of semantic mechanisms]. Dialectica: International Review of Philosophy of Knowledge, 22(3–4):247–259, December 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pun:1968:DFC

[718] Lucas Pun. Sur le déclic. (French) [On the click]. Dialectica: International Review of Philosophy of Knowledge, 22(3–4):260–271, December 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tschumi:1968:LPL

[719] Raymond Tschumi. Le langage et la pensée : Leurs rapports et le dépassement de lews niveaux dans les sciences et dans les littératures. (French) [Language and thought: Their relationships and the overcoming of levels in science and literature]. Dialectica: International Review of Philosophy of Knowledge, 22(3–4):272–292, December 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Loreau:1968:ALF

[720] Max Loreau. Art et langage. (French) [Art and language]. Dialectica: International Review of Philosophy of Knowledge, 22(3–4):293–299, December 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chari:1968:CDL

[721] C. T. K. Chari. Culturally different languages and philosophies. Dialectica: International Review of Philosophy of Knowledge, 22(3–4):300–312, December 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Junod:1968:NML

[722] H. Ph. Junod. Notes sur le modèle linguistique saussurien. (French) [Notes on the Saussurian linguistic model]. Dialectica: International Review of Philosophy of Knowledge, 22(3–4):313–317, December 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rosenthal:1968:CSC

[723] Sandra B. Rosenthal. A comment on some comments. Dialectica: International Review of Philosophy of Knowledge, 22(3–4):318–320, December 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1968:ABJ

[724] Anonymous. Abstracts from the ‘British Journal for the Philosophy of Science’. Dialectica: International Review of Philosophy of Knowledge, 22(3–4):321–323, December 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1968:RSE

[725] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 22(3–4):324, December 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1968:RBLb

[726] Anonymous. Received books — livres reçus — eingelaufene Bücher. Dialectica: International Review of Philosophy of Knowledge, 22(3–4): 325–326, December 1968. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Redaction:1969:AFA

[727] La Rédaction. Avertissement. (French) [Advertisement]. Dialectica: International Review of Philosophy of Knowledge, 23(1):3, March 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1969:E

[728] Anonymous. Editorial. Dialectica: International Review of Philosophy of Knowledge, 23(1):5–6, March 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nolfi:1969:SS

[729] P. Nolfi. Strategische spiele. Dialectica: International Review of Philosophy of Knowledge, 23(1):7–23, March 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Morf:1969:PPS

[730] Albert Morf, Jean-Blaise Grize, and Laurent Pauli. Pour une pédagogie scientifique. (French) [For a scientific pedagogy]. Dialectica: International Review of Philosophy of Knowledge, 23(1):24–31, March 1969. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Emery:1969:PDPa

[731] Eric Emery. À propos du problème du temps. (French) [On the problem of time]. Dialectica: International Review of Philosophy of Knowledge, 23 (1):32–44, March 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Robinson:1969:FPV

[732] Abraham Robinson. From a formalist’s point of view. Dialectica: International Review of Philosophy of Knowledge, 23(1):45–49, March 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Goodstein:1969:EM

[733] R. L. Goodstein. Empiricism in mathematics. Dialectica: International Review of Philosophy of Knowledge, 23(1):50–57, March 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Engeler:1969:PFC

[734] Erwin Engeler and Helmut Röhrl. On the problem of foundations of category theory. Dialectica: International Review of Philosophy of Knowledge, 23(1):58–66, March 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vonFinsler:1969:UKG

[735] Paul von Finsler and Zürich. Über die Unabhängigkeit der Kontinuumhypothese. (German) [On the independence of the Continuum Hypothesis]. Dialectica: International Review of Philosophy of Knowledge, 23(1):67–78, March 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sellars:1969:NSF

[736] R. W. Sellars. Le naturalisme de Sellars. (French) [The naturalism of Sellars]. Dialectica: International Review of Philosophy of Knowledge, 23(1):79–80, March 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Blandino:1969:TNV

[737] G. Blandino. Théories sur la nature de la vie. (French) [Theories about the nature of life]. Dialectica: International Review of Philosophy of Knowledge, 23(1):80, March 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Curry:1969:MBF

[738] Haskell B. Curry. Modified basic functionality in combinatory logic. Dialectica: International Review of Philosophy of Knowledge, 23(2):83–92, June 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kreisel:1969:TNF

[739] G. Kreisel. Two notes on the foundations of set-theory. Dialectica: International Review of Philosophy of Knowledge, 23(2):93–114, June 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Emery:1969:PDPb

[740] E. Emery. À propos du problème du temps (suite). (French) [On the problem of time (end)]. Dialectica: International Review of Philosophy of Knowledge, 23(2):115–134, June 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Feibleman:1969:NOO

[741] J. K. Feibleman. Naturally — occurring ontologies. Dialectica: International Review of Philosophy of Knowledge, 23(2):135–150, June 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1969:RSE

[742] Anonymous. Reviewing studies — Études critiques — Betrachtungen für Literatur. Dialectica: International Review of Philosophy of Knowledge, 23(2):151–152, June 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1969:LRE

[743] Anonymous. Livres reçus — Eingegangene Bücher — books received. Dialectica: International Review of Philosophy of Knowledge, 23(2):154– 158, June 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1969:PMF

[744] F. Gonseth. Prospective et méthode. (French) [Prospective and methodology]. Dialectica: International Review of Philosophy of Knowledge, 23 (3–4):163–175, December 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gayer:1969:OHS

[745] Jan Gayer. An outline of hierarchical systems theory and its role in philosophy. Dialectica: International Review of Philosophy of Knowledge, 23(3–4):177–188, December 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Phipps:1969:RPS

[746] T. E. Phipps, Jr. The relativity of physical size. Dialectica: International Review of Philosophy of Knowledge, 23(3–4):189–215, December 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gauthier:1969:NTS

[747] Yvon Gauthier. La notion théorétique de structure. (French) [The theoretical notion of structure]. Dialectica: International Review of Philosophy of Knowledge, 23(3–4):217–227, December 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Maruyama:1969:ESS

[748] Magoroh Maruyama. Epistemology of social science research: exploration in inculture researchers. Dialectica: International Review of Philosophy of Knowledge, 23(3–4):229–280, December 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lawvere:1969:AF

[749] F. William Lawvere. Adjointness in foundations. Dialectica: International Review of Philosophy of Knowledge, 23(3–4):281–296, December 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1969:PO

[750] Ferdinand Gonseth. La philosophie ouverte. Dialectica: International Review of Philosophy of Knowledge, 23(3–4):297–314, December 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1969:APS

[751] Anonymous. Abstracts from philosophy of science. Dialectica: International Review of Philosophy of Knowledge, 23(3–4):315–320, December 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1969:CSI

[752] Anonymous. Contents sommaire inhalt. Dialectica: International Review of Philosophy of Knowledge, 23(3–4):321, December 1969. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1970:PFP

[753] Paul Emile Pilet. Préface. (French) [Preface]. Dialectica: International Review of Philosophy of Knowledge, 24(1–3):7–10, September 1970. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1970:FGH

[754] Paul Emile Pilet. Ferdinand Gonseth, l’homme et le dialogue. (French) [Ferdinand Gonseth, man and dialogue]. Dialectica: International Review of Philosophy of Knowledge, 24(1–3):11–21, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Guggenheimer:1970:SCG

[755] H. Guggenheimer. Some comments on Gonseth’s principles of a methodology of science. Dialectica: International Review of Philosophy of Knowledge, 24(1–3):23–28, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bertholet:1970:VOS

[756] Edmond Bertholet. Vers un organon de la science et de la philosophie. (French) [Towards an organon of science and philosophy]. Dialectica: International Review of Philosophy of Knowledge, 24(1–3):29–39, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Echarri:1970:MAF

[757] Por Jaime Echarri. Metodología abierta y fenomenología abierta. (Spanish) [Open methodology and open phenomenology]. Dialectica: International Review of Philosophy of Knowledge, 24(1–3):41–51, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1970:SKI

[758] Paul Bernays. Die schematische Korrespondenz und die idealisierten Strukturen. (German) [Schematic correspondence and idealized structures]. Dialectica: International Review of Philosophy of Knowledge, 24 (1–3):53–66, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Grize:1970:LLF

[759] Jean-Blaise Grize. Logique et langue. (French) [Logic and language]. Dialectica: International Review of Philosophy of Knowledge, 24(1–3): 67–76, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Engeler:1970:GL

[760] Erwin Engeler. Geometry and language. Dialectica: International Review of Philosophy of Knowledge, 24(1–3):77–85, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tschumi:1970:LTL

[761] Raymond Tschumi. Langage, temps et liberté. I. (French) [Language, time, and liberty. I]. Dialectica: International Review of Philosophy of Knowledge, 24(1–3):87–110, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Emery:1970:EMP

[762] Eric Emery. Esthétique musicale et philosophic ouverte. (French) [Aesthetics musical and philosophy open]. Dialectica: International Review of Philosophy of Knowledge, 24(1–3):111–121, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nolfi:1970:HST

[763] P. Nolfi. Holonspiele, Struktur und Tragweite. (German) [Holon games, structure and scope]. Dialectica: International Review of Philosophy of Knowledge, 24(1–3):123–138, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tonini:1970:IEF

[764] V. Tonini. Information et explication. (French) [Information and explanation]. Dialectica: International Review of Philosophy of Knowledge, 24 (1–3):139–155, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Janoska:1970:DA

[765] Georg Jánoska and Judith Jánoska-Bendl. Das dialektische apriori. Dialectica: International Review of Philosophy of Knowledge, 24(1–3):157– 163, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Piguet:1970:FDS

[766] J.-Claude Piguet. Un fait dialectique: la science newtonienne. (French) [A dialectic fact: Newtonian science]. Dialectica: International Review of Philosophy of Knowledge, 24(1–3):165–183, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Guggenheimer:1970:ASC

[767] Eva H. Guggenheimer. Acoustic symmetry in Catullus. Dialectica: International Review of Philosophy of Knowledge, 24(1–3):185–195, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rossel:1970:ORD

[768] J. Rossel. Objectivité et réalité, deux concepts en devenir. (French) [Objectivity and reality, two concepts in the making]. Dialectica: International Review of Philosophy of Knowledge, 24(1–3):197–204, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schlechta:1970:PPE

[769] Karl Schlechta. Sind Physik und Philosophie am Ende?. (German) [Are physics and philosophy at the end?]. Dialectica: International Review of Philosophy of Knowledge, 24(1–3):205–213, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cantore:1970:SHS

[770] E. Cantore. Science and humanism: the sapiential role of philosophy. Dialectica: International Review of Philosophy of Knowledge, 24(1–3): 215–241, September 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deBeauregard:1970:HAP

[771] O. Costa de Beauregard. Hommage au professeur Ferdinand Gonseth (suite du fasc. 1–3) autofondation. (French) [Homage to Professor Ferdinand Gonseth (continuation of fascicles 1–3) autofondation]. Dialectica: International Review of Philosophy of Knowledge, 24(4):247–254, December 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deLatil:1970:CPF

[772] P. de Latil. Causalité et processus. (French) [Causality and process]. Dialectica: International Review of Philosophy of Knowledge, 24(4):255– 260, December 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Smith:1970:EAO

[773] Robin Smith. On eliminating the art object. Dialectica: International Review of Philosophy of Knowledge, 24(4):261–266, December 1970. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pour-El:1970:RTV

[774] Marian Boykan Pour-El. A recursion-theoretic view of axiomatizable theories. Dialectica: International Review of Philosophy of Knowledge, 24(4):267–276, December 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Montague:1970:PIL

[775] Richard Montague. Pragmatics and intensional logic. Dialectica: International Review of Philosophy of Knowledge, 24(4):277–302, December 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wette:1970:UEG

[776] Eduard Wette. Vom Unendlichen zum Endlichen. (German) [From the infinite to the finite]. Dialectica: International Review of Philosophy of Knowledge, 24(4):303–323, December 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1970:RSE

[777] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 24(4):325–328, December 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1970:APS

[778] Anonymous. Abstracts from “Philosophy of science”. Dialectica: International Review of Philosophy of Knowledge, 24(4):329–337, December 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1970:CSI

[779] Anonymous. Contents sommaire inhalt. Dialectica: International Review of Philosophy of Knowledge, 24(4):338, December 1970. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1971:QAM

[780] Paul-Emile Pilet. Quelques aspects méthodologiques de information biologique. (French) [Some methodological aspects of biological information]. Dialectica: International Review of Philosophy of Knowledge, 25 (1):3–15, March 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ruse:1971:TBR

[781] Michael E. Ruse. Two biological revolutions. Dialectica: International Review of Philosophy of Knowledge, 25(1):17–38, March 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ruse:1971:RRM

[782] Michael E. Ruse. Reduction, replacement, and molecular biology. Dialectica: International Review of Philosophy of Knowledge, 25(1):39–72, March 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1971:RSNa

[783] Anonymous. Reviewing studies — notes critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 25(1):73–78, March 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1971:LMF

[784] F. Gonseth. Langue et méthode. (French) [Language and method]. Dialectica: International Review of Philosophy of Knowledge, 25(2):83–118, June 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Angenot:1971:CCL

[785] Marc Angenot. Condillac et le “Cours de linguistique générale”. (French) [Condillac and the “Course on general linguistics”]. Dialectica: International Review of Philosophy of Knowledge, 25(2):119–130, June 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tibbetts:1971:LED

[786] Paul Tibbetts. The “Levels of experience” doctrine in modern philosophy of mind. Dialectica: International Review of Philosophy of Knowledge, 25(2):131–151, June 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Reignier:1971:QCQ

[787] Jean Reignier. Qu’est-ce qu’une particule élémentaire ?. (French) [What is an elementary particle?]. Dialectica: International Review of Philosophy of Knowledge, 25(2):153–165, June 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1971:RSNb

[788] Anonymous. Reviewing studies — notes critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 25(2):166, June 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernays:1971:SGM

[789] Paul Bernays. Zum Symposium über die Grundlagen der Mathematik. (German) [On the symposium on the fundamentals of mathematics]. Dialectica: International Review of Philosophy of Knowledge, 25(3–4):171– 195, December 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deLatil:1971:NIF

[790] Pierre de Latil. La notion d’information. (French) [The idea of information]. Dialectica: International Review of Philosophy of Knowledge, 25 (3–4):197–205, December 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Arsac:1971:CII

[791] Jacques J. Arsac. Le concept d’information et informatique. (French) [The concept of information and computing]. Dialectica: International Review of Philosophy of Knowledge, 25(3–4):207–219, December 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cuenod:1971:ICF

[792] Michel Cuénod. Identification et connaissance. (French) [Identification and knowledge]. Dialectica: International Review of Philosophy of Knowledge, 25(3–4):221–238, December 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Robinet:1971:CIR

[793] André Robinet. Courte introduction aux relations de la philosophic avec informatique. (French) [Short introduction to the relations of philosophic with computer science]. Dialectica: International Review of Philosophy of Knowledge, 25(3–4):239–249, December 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wenin:1971:IIP

[794] Christian Wenin. Informatique et indexation de périodiques philosophiques. Intérêt et limites d’une entreprise. (French) [Computer science and indexing of philosophical periodicals. Interest and limits to a business]. Dialectica: International Review of Philosophy of Knowledge, 25(3–4): 251–259, December 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tombeur:1971:TAT

[795] Paul Tombeur. Le traitement automatique de textes philosophiques. (French) [Automatic processing of philosophical texts]. Dialectica: International Review of Philosophy of Knowledge, 25(3–4):261–274, December 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zelbstein:1971:RPA

[796] Uri Zelbstein. Réflexions pour une approche épistémologique de informatique. (French) [Reflections on an epistemological approach to computing]. Dialectica: International Review of Philosophy of Knowledge, 25 (3–4):275–291, December 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cantore:1971:SDH

[797] Enrico Cantore. Science as dialogical humanizing process: Highlights of a vocation. Dialectica: International Review of Philosophy of Knowledge, 25(3–4):293–316, December 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1971:RSNc

[798] Anonymous. Reviewing studies — notes critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 25(3–4):317–320, December 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1971:RPL

[799] Anonymous. Recent publications — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 25(3–4): 321–323, December 1971. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sauvan:1972:CIF

[800] Jacques Sauvan. Connaissance et informatique. (French) [Knowledge and informatics]. Dialectica: International Review of Philosophy of Knowledge, 26(1):3–10, March 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pun:1972:DAI

[801] Lucas Pun. Deux aspects de l’informatique. (French) [Two aspects of computing]. Dialectica: International Review of Philosophy of Knowledge, 26(1):11–50, March 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Maruyama:1972:NCI

[802] Magoroh Maruyama. Non-classificational information and noninformational communication. Dialectica: International Review of Philosophy of Knowledge, 26(1):51–59, March 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Robinet:1972:LAO

[803] André Robinet. L’information appliquée aux œuvres complètes de Malebranche. (French) [Information on the complete works of Malebranche]. Dialectica: International Review of Philosophy of Knowledge, 26(1):61– 76, March 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1972:RSE

[804] P. E. Pilet. Reviewing studies — Études critiques — Betrachtungen zur Literatur: Hypothèse et perception. (French) [Hypothesis and perception]. Dialectica: International Review of Philosophy of Knowledge, 26 (1):77–78, March 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Apostol:1972:ODD

[805] Pavel Apostol. An operational demarcation of the domain of methodology versus epistemology and logic. Dialectica: International Review of Philosophy of Knowledge, 26(2):83–92, June 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tonini:1972:SPF

[806] Valerio Tonini. Science et philosophic: la fondation d’une épistémologie ouverte. (French) [Science and philosophy: the foundation of an open epistemology]. Dialectica: International Review of Philosophy of Knowledge, 26(2):93–104, June 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Morf:1972:FCT

[807] A. Morf. La formation des connaissances et la théorie didactique. (French) [The formation of knowledge and didactic theory]. Dialectica: International Review of Philosophy of Knowledge, 26(2):103–114, June 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Landea:1972:EQG

[808] Alfred Landea. Einheit in der Quantenwelt (Gegen den Bohr– Heisenberg’schen Positivismus). (German) [Unity in the quantum world (through Bohr–Heisenberg positivism)]. Dialectica: International Review of Philosophy of Knowledge, 26(2):115–130, June 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zuber:1972:PQD

[809] Ryszard Zuber. À propos de la question dite générale. (French) [On the so-called general question]. Dialectica: International Review of Philosophy of Knowledge, 26(2):131–137, June 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Koehler:1972:BRT

[810] Conrad J. Koehler. Bertrand Russell’s Theory of Knowledge by Elizabeth Ramsden Eames. Dialectica: International Review of Philosophy of Knowledge, 26(2):139–145, June 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1972:RSE

[811] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 26(2):146–148, June 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Maruyama:1972:THF

[812] Magoroh Maruyama. Toward human futuristics: Trans-epistemological process. Dialectica: International Review of Philosophy of Knowledge, 26(3–4):155–183, December 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schwartz:1972:RLC

[813] Elisabeth Schwartz. Remarques sur “L’Espace des choses” de Wittgenstein et ses origines frégéennes. (French) [Remarks on “The space of things” by Wittgenstein and its origins in France]. Dialectica: International Review of Philosophy of Knowledge, 26(3–4):185–226, December 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Brown:1972:HLS

[814] Harold I. Brown. Harris on the logic of science. Dialectica: International Review of Philosophy of Knowledge, 26(3–4):227–246, December 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tibbetts:1972:FRF

[815] Paul Tibbetts. Feigl on raw feels, the brain, and knowledge claims: Some problems regarding theoretical concepts. Dialectica: International Review of Philosophy of Knowledge, 26(3–4):247–266, December 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schwartz:1972:RPE

[816] Bertrand Schwartz. Réflexions prospectives sur l’éducation permanente. (French) [Reflections on lifelong learning]. Dialectica: International Review of Philosophy of Knowledge, 26(3–4):267–292, December 1972. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rohatyn:1972:LP

[817] Dennis A. Rohatyn. The language of philosophy. Dialectica: International Review of Philosophy of Knowledge, 26(3–4):293–299, December 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jakubec:1972:RSE

[818] Joel Jakubec. Reviewing studies — Études critiques — Betrachtungen zur Literatur. L’alternative. (French) [The alternative]. Dialectica: International Review of Philosophy of Knowledge, 26(3–4):301–303, December 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nlfi:1973:HA

[819] Padrot Nlfi. Holonspiele als arbeitshypothese. Dialectica: International Review of Philosophy of Knowledge, 27(1):3–11, March 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jauch:1973:DCQ

[820] J. M. Jauch. Determinism in classical and quantal physics. Dialectica: International Review of Philosophy of Knowledge, 27(1):13–26, March 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tibbetts:1973:MTR

[821] Paul Tibbetts. Mead’s theory of reality and the knower-known transaction. Dialectica: International Review of Philosophy of Knowledge, 27 (1):27–41, March 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zemach:1973:NC

[822] Eddy M. Zemach. The nature of consciousness. Dialectica: International Review of Philosophy of Knowledge, 27(1):43–65, March 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1973:RSEa

[823] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 27(1):67–72, March 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1973:MCO

[824] Anonymous. Miscellaneous: Communication to our readers — communication à nos lecteurs — Mitteilung an unsere Leser. Dialectica: International Review of Philosophy of Knowledge, 27(1):73, March 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mehra:1973:QPI

[825] Jagdish Mehra. The quantum principle: Its interpretation and epistemology. Dialectica: International Review of Philosophy of Knowledge, 27 (2):75–157, June 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1973:HPN

[826] Ferdinand Gonseth. En hommage à Padrot Nolfi. (French) [A tribute to Padrot Nolfi]. Dialectica: International Review of Philosophy of Knowledge, 27(2):159–161, June 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1973:RSEb

[827] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 27(2):162–163, June 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lauener:1973:VGF

[828] H. Lauener. Vorwort. (German) [Foreword]. Dialectica: International Review of Philosophy of Knowledge, 27(3–4):167, December 1973. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Brulisauer:1973:WWP

[829] Bruno Brülisauer. Wieweit werden philosophische Probleme durch Sprachanalyse gelöst?. (German) [How are philosophical problems solved by speech analysis?]. Dialectica: International Review of Philosophy of Knowledge, 27(3–4):169–178, December 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1973:SSG

[830] Ferdinand Gonseth. Über die Sprache sprechen. (German) [Talk about language]. Dialectica: International Review of Philosophy of Knowledge, 27(3–4):179–217, December 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stroll:1973:LCP

[831] Avrum Stroll. Linguistic clusters and the problem of universals. Dialectica: International Review of Philosophy of Knowledge, 27(3–4):219–260, December 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Black:1973:PCE

[832] Max Black. Paradigm cases and evaluative words. Dialectica: International Review of Philosophy of Knowledge, 27(3–4):261–272, December 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stenius:1973:PAS

[833] Erik Stenius. The problem of AKRASIA as a subject of philosophical analysis. Dialectica: International Review of Philosophy of Knowledge, 27(3–4):273–288, December 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Foellesdal:1973:ITU

[834] Dagfinn Foellesdal. Indeterminacy of translation and under-determination of the theory of nature. Dialectica: International Review of Philosophy of Knowledge, 27(3–4):289–301, December 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bar-Hillel:1973:PTB

[835] Y. Bar-Hillel. Primary truth bearers. Dialectica: International Review of Philosophy of Knowledge, 27(3–4):303–312, December 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Davidson:1973:RI

[836] Donald Davidson. Radical interpretation. Dialectica: International Review of Philosophy of Knowledge, 27(3–4):313–328, December 1973. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hintikka:1973:QVQ

[837] Jaakko Hintikka. Quantifiers vs. quantification theory. Dialectica: International Review of Philosophy of Knowledge, 27(3–4):329–358, December 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See comments [886].

Bernays:1973:AFS

[838] Paul Bernays. Über die Auffassung formaler Systeme als Sprachen. (German) [On the conception of formal systems as languages]. Dialectica: International Review of Philosophy of Knowledge, 27(3–4):359–362, December 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Egli:1973:SRF

[839] Urs Egli. Semantische Repräsentation der Frage. (German) [Semantic representation of the question]. Dialectica: International Review of Philosophy of Knowledge, 27(3–4):363–370, December 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See comment [888].

Anonymous:1973:LLR

[840] Anonymous. Liste des livres reçus. Dialectica: International Review of Philosophy of Knowledge, 27(3–4):371–372, December 1973. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lauener:1974:PPB

[841] Henri Lauener. Zur philosophischen Problematik des Begriffes der Kausalität. (German) [The philosophical problem of the concept of causality]. Dialectica: International Review of Philosophy of Knowledge, 28(1–2):3–20, June 1974. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kaeser:1974:DKR

[842] E. Kaeser. Ein dogma des kritischen rationalismus. Dialectica: International Review of Philosophy of Knowledge, 28(1–2):21–42, June 1974. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tonoiu:1974:POP

[843] Vasile Tonoiu. La philosophie ouverte et le problème du commencement: Allez, et la foi vous viendra (d’Alembert). (French) [Open philosophy and the problem of the beginning: Go, and faith will come to you (d’Alembert)]. Dialectica: International Review of Philosophy of Knowledge, 28(1–2):43–68, June 1974. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gochet:1974:PDC

[844] Paul Gochet. Le problème du commencement et la philosophie analytique. (French) [The problem of beginning and analytic philosophy]. Dialectica: International Review of Philosophy of Knowledge, 28(1–2): 69–86, June 1974. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rohatyn:1974:ALS

[845] Dennis A. Rohatyn. Against the logicians: some informed polemics. Dialectica: International Review of Philosophy of Knowledge, 28(1–2): 87–102, June 1974. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lambros:1974:SDP

[846] Charles H. Lambros. Schlick’s doctrine of the a priori in Allgemeine Erkenntnislehre. Dialectica: International Review of Philosophy of Knowledge, 28(1–2):103–128, June 1974. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1974:RSEa

[847] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 28(1–2):129–131, June 1974. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Maruyama:1974:PAC

[848] Magoroh Maruyama. Paradigmatology and its application to crossdisciplinary, cross-professional and cross-cultural communication. Dialectica: International Review of Philosophy of Knowledge, 28(3–4):135– 196, December 1974. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Temple:1974:GGS

[849] Dennis Temple. Grue-green and some mistakes in confirmation theory. Dialectica: International Review of Philosophy of Knowledge, 28(3–4): 197–210, December 1974. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Capek:1974:QRP

[850] Milič Capek. Sur quelques résistances philosophiques à la physique du vingtième siècle. (French) [On some philosophical resistances to the physics of the twentieth century]. Dialectica: International Review of Philosophy of Knowledge, 28(3–4):211–222, December 1974. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Girill:1974:ICR

[851] T. R. Girill. Identity, causality, and the regressiveness of microexplanations. Dialectica: International Review of Philosophy of Knowledge, 28(3–4):223–238, December 1974. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Shotwell:1974:II

[852] David A. Shotwell. The identity of indiscernibles. Dialectica: International Review of Philosophy of Knowledge, 28(3–4):239–242, December 1974. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Saarnio:1974:GPI

[853] Uuno Saarnio. Die Grellingsche Paradoxie und ihre exakte Lösung. (German) [Grelling’s paradox and its exact solution]. Dialectica: International Review of Philosophy of Knowledge, 28(3–4):243–262, December 1974. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lake:1974:PM

[854] John Lake. A philosophy of mathematics? Dialectica: International Review of Philosophy of Knowledge, 28(3–4):263–270, December 1974. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1974:RSEb

[855] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 28(3–4):271–279, December 1974. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pier:1975:AP

[856] Jean-Paul Pier. Avant-propos. Dialectica: International Review of Philosophy of Knowledge, 29(1):3–4, March 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hirsch:1975:MRF

[857] G. Hirsch. Mathématisation et réalité. (French) [Mathematization and reality]. Dialectica: International Review of Philosophy of Knowledge, 29 (1):5–24, March 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1975:MRF

[858] Ferdinand Gonseth. Les mathématiques et la réalité. (French) [Mathematics and reality]. Dialectica: International Review of Philosophy of Knowledge, 29(1):25–38, March 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dieudonne:1975:LIM

[859] Jean Dieudonné. L’abstraction et l’intuition mathématique. (French) [Abstraction and mathematical intuition]. Dialectica: International Review of Philosophy of Knowledge, 29(1):39–54, March 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Meschkowski:1975:MRG

[860] Herbert Meschkowski. Mathematik und Realität bei Georg Cantor. (German) [Mathematics and Reality for Georg Cantor]. Dialectica: International Review of Philosophy of Knowledge, 29(1):55–70, March 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Thom:1975:MIF

[861] René Thom. Les mathématiques et l’intelligible. (French) [Mathematics and the intelligible]. Dialectica: International Review of Philosophy of Knowledge, 29(1):71–80, March 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pier:1975:CRS

[862] Jean-Paul Pier. Compte rendu succinct de la Table Ronde. (French) [Summary of the Round Table]. Dialectica: International Review of Philosophy of Knowledge, 29(1):81–84, March 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gonseth:1975:HJM

[863] F. Gonseth. Hommage à J.-M. Jauch, membre du Comité, directeur de Dialectica en 1973 et 1974. (French) [Tribute to J.-M. Jauch, member of the Committee, director of Dialectica in 1973 and 1974]. Dialectica: International Review of Philosophy of Knowledge, 29(1):85, March 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Feibleman:1975:AOL

[864] James K. Feibleman. Assumptions of operational logic. Dialectica: International Review of Philosophy of Knowledge, 29(2–3):91–104, September 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Agazzi:1975:TET

[865] Evandro Agazzi. De la théorie électroélastique à la théorie électromagnétique du champ. (French) [From electroelastic theory to electromagnetic field theory]. Dialectica: International Review of Philosophy of Knowledge, 29 (2–3):105–126, September 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Folse:1975:FOQ

[866] Henry J. Folse. The formal objectivity of quantum mechanical systems. Dialectica: International Review of Philosophy of Knowledge, 29(2–3): 127–143, September 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tibbetts:1975:FHK

[867] Paul Tibbetts. ^{≪}Hanson and Kuhn on observation reports and knowledge claims^{≫}. Dialectica: International Review of Philosophy of Knowledge, 29(2–3):145–155, September 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Weiss:1975:HST

[868] Raymond L. Weiss. Historicism and science: Thoughts on Quine. Dialectica: International Review of Philosophy of Knowledge, 29(2–3):157–165, September 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See comments [869].

Bernays:1975:ARW

[869] Paul Bernays. Zu der Abhandlung von Raymond L. Weiss ^{≪}Historicism and Science: Thoughts on Quine^{≫}. (German) [On the treatise of Raymond L. Weiss “Historicism and Science: Thoughts on Quine”]. Dialectica: International Review of Philosophy of Knowledge, 29(2–3):167–172, September 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See [868].

Bartlett:1975:PI

[870] Steven Bartlett. Phenomenology of the implicit. Dialectica: International Review of Philosophy of Knowledge, 29(2–3):173–188, September 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Legrand:1975:HPH

[871] Michel Legrand. Hypothèses pour une histoire de la psychanalyse. (French) [Hypotheses for a history of psychoanalysis]. Dialectica: International Review of Philosophy of Knowledge, 29(2–3):189–207, September 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1975:AP

[872] Paul-Emile Pilet. Avant-propos. Dialectica: International Review of Philosophy of Knowledge, 29(4):211–212, December 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Plamondon:1975:CRM

[873] Ann Plamondon. The contemporary reconciliation of mechanism and organicism. Dialectica: International Review of Philosophy of Knowledge, 29(4):213–221, December 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zaslawsky:1975:QRM

[874] Denis Zaslawsky. Quelques rapports méthodologiques entre les sciences biologiques et la philosophie. (French) [Some methodological reports between the biological sciences and philosophy]. Dialectica: International Review of Philosophy of Knowledge, 29(4):223–235, December 1975. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Burstrom:1975:EER

[875] H. G. Burström. Ethics of experimental research. Dialectica: International Review of Philosophy of Knowledge, 29(4):237–247, December 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1975:LEB

[876] Paul-Emile Pilet. L’enseignement expérimental de la biologie: quelques problèmes épistémologiques. (French) [Experimental teaching of biology: some epistemological problems]. Dialectica: International Review of Philosophy of Knowledge, 29(4):249–257, December 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rohner:1975:CEA

[877] Georges Rohner. Comment enseigner authentiquement la biologie au niveau secondaire. (French) [How to truly teach biology at the secondary level]. Dialectica: International Review of Philosophy of Knowledge, 29 (4):259–278, December 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1975:RSE

[878] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 29(4):279–280, December 1975. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1976:FG

[879] Anonymous. Ferdinand Gonseth. Dialectica: International Review of Philosophy of Knowledge, 30(1):3, March 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Herausgeber:1976:V

[880] Der Herausgeber and Henri Lauener. Vorwort. Dialectica: International Review of Philosophy of Knowledge, 30(1):5–6, March 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hintikka:1976:QVP

[881] Jaakko Hintikka. Quine vs. Peirce? Dialectica: International Review of Philosophy of Knowledge, 30(1):7–8, March 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Graybeal:1976:UDT

[882] David P. Graybeal. The in- and under-determinacy of translation. Dialectica: International Review of Philosophy of Knowledge, 30(1):9–15, March 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stroll:1976:TBP

[883] Avrum Stroll. Truth-bearers, propositions and the problem of universals. Dialectica: International Review of Philosophy of Knowledge, 30 (1):17–34, March 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stenius:1976:CDD

[884] Erik Stenius. Comments on Donald Davidson’s paper “Radical interpretation”. Dialectica: International Review of Philosophy of Knowledge, 30 (1):35–60, March 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hintikka:1976:PC

[885] Jaakko Hintikka. The prospects for convention T. Dialectica: International Review of Philosophy of Knowledge, 30(1):61–66, March 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stenius:1976:CJH

[886] Erik Stenius. Comments on Jaakko Hintikka’s paper “Quantifiers vs. quantification theory”. Dialectica: International Review of Philosophy of Knowledge, 30(1):67–88, March 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See [837].

Hintikka:1976:POQ

[887] Jaakko Hintikka. Partially ordered quantifiers vs. partially ordered ideas. Dialectica: International Review of Philosophy of Knowledge, 30(1):89– 99, March 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hintikka:1976:QCU

[888] Jaakko Hintikka. The question of?: A comment on Urs Egli. Dialectica: International Review of Philosophy of Knowledge, 30(1):101–103, March 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See

Anonymous:1976:RSEa

[889] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 30(1):105–106, March 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Temple:1976:SLP

[890] Dennis Temple. Some linguistic puzzles related to formal logic. Dialectica: International Review of Philosophy of Knowledge, 30(2–3):111–116, September 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

White:1976:SRS

[891] Michael J. White. A suggestion regarding the semantical analysis of performatives. Dialectica: International Review of Philosophy of Knowledge, 30(2–3):117–134, September 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tallet:1976:PC

[892] Jorge Tallet. Probability and credibility. Dialectica: International Review of Philosophy of Knowledge, 30(2–3):135–143, September 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Grieder:1976:PZR

[893] Alfons Grieder. Protophysik der Zeit und Relativitätstheorie. (German) [Protophysics of time and relativity theory]. Dialectica: International Review of Philosophy of Knowledge, 30(2–3):145–159, September 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Levy-Leblond:1976:TPQ

[894] Jean-Marc Lévy-Leblond. Towards a proper quantum theory (hints for a recasting). Dialectica: International Review of Philosophy of Knowledge, 30(2–3):161–192, September 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zipkin:1976:TZ

[895] Harry J. Zipkin. Theoretical zipperdynamics. Dialectica: International Review of Philosophy of Knowledge, 30(2–3):192–196, September 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Conn:1976:ODS

[896] Walter E. Conn. Objectivity — a developmental and structural analysis: The epistemologies of Jean Piaget and Bernard Lonergan. Dialectica: International Review of Philosophy of Knowledge, 30(2–3):197–221, September 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Allison:1976:KRR

[897] Henry E. Allison. Kant’s refutation of realism. Dialectica: International Review of Philosophy of Knowledge, 30(2–3):223–253, September 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Weiss:1976:RB

[898] Raymond L. Weiss. Response to Bernays. Dialectica: International Review of Philosophy of Knowledge, 30(2–3):255–259, September 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Morel:1976:SRS

[899] Bernard Morel. Société, religion et spiritualité. (French) [Society, religion and spirituality]. Dialectica: International Review of Philosophy of Knowledge, 30(4):263–265, December 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rudhardt:1976:RSR

[900] Jean Rudhardt. Religion sociale et religion personnelle dans la Grèce antique. (French) [Social and personal religion in Ancient Greece]. Dialectica: International Review of Philosophy of Knowledge, 30(4):267– 276, December 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Keller:1976:FSD

[901] Carl-A. Keller. La fonction sociale du saint. (French) [The social function of the saint]. Dialectica: International Review of Philosophy of Knowledge, 30(4):277–284, December 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

May:1976:AHI

[902] Jacques May. Aspects historiques de l’insertion du bouddhisme dans quelques sociétés d’Asie. (French) [Historical aspects of the integration of Buddhism in some societies of Asia]. Dialectica: International Review of Philosophy of Knowledge, 30(4):285–298, December 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Python:1976:MCF

[903] P. Python. Mystique et communauté (le fait religieux sous son aspect théologique). (French) [Mysticism and community (the religious fact in its theological aspect)]. Dialectica: International Review of Philosophy of Knowledge, 30(4):299–304, December 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

DEpinay:1976:RSS

[904] Christian Lalive D’Epinay. Religion, spiritualité et société de l’étude sociologique du pentecôtisme latino-américain. (French) [Religion, spirituality and society of the sociological study of Latin–American Pentecostalism]. Dialectica: International Review of Philosophy of Knowledge, 30(4):305–313, December 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vallin:1976:MTT

[905] G. Vallin. Modèles théoriques des traditions asiatiques et préjugés occidentaux. (French) [Theoretical models of Asian traditions and Western prejudices]. Dialectica: International Review of Philosophy of Knowledge, 30(4):315–319, December 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Morel:1976:CMF

[906] Bernard Morel. Conclusions méthodologiques. (French) [Methodological conclusions]. Dialectica: International Review of Philosophy of Knowledge, 30(4):321–322, December 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1976:RSEb

[907] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 30(4):323–324, December 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1976:RBL

[908] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 30(4):325– 326, December 1976. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Directeur:1977:AP

[909] Le Comité Directeur. Avant-propos. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):4, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1977:EFI

[910] Anonymous. Editorial of the first issue (1947). Dialectica: International Review of Philosophy of Knowledge, 31(1–2):5–6, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1977:EDP

[911] Anonymous. Éditorial du premier numéro (1947). (French) [Editorial from the first issue (1947)]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):7–8, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1977:GGE

[912] Anonymous. Geleitwort: Geleitwort zur ersten Nummer (1947). (German) [Command: First commandment (1947)]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):9–11, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gagnebin:1977:DEF

[913] Samuel Gagnebin. Le dialogue engagé. (French) [Engaged dialogue]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):13– 18, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Piaget:1977:SFG

[914] Jean Piaget. Souvenirs sur Ferdinand Gonseth. (French) [Memories of Ferdinand Gonseth]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):19–22, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1977:FGS

[915] Paul-Emile Pilet. Ferdinand Gonseth — sa vie, son œuvre. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):23–33, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dockx:1977:NHC

[916] Stanislas Dockx. Note historique concernant la fondation de l’Union internationale de Philosophie des Sciences. (French) [Historical note concerning the founding of the International Union of Philosophy of Science]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2): 35–38, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Joel:1977:GPF

[917] Joel and Doris Jakubec. Gonseth, un protestant. (French) [Gonseth, a Protestant]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):39–43, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Morel:1977:LES

[918] Bernard Morel. L’ouverture á l’expérience spirituelle. (French) [The opening to the spiritual experience]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):45–55, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tschumi:1977:EFR

[919] Raymond Tschumi. Les esprits fins à la recherche de l’honnête homme. (French) [Fine minds in search of the honest man]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):57–64, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Emery:1977:PFG

[920] Eric Emery. La pensée de F. Gonseth et le temps dans l’art musical. (French) [The thought of F. Gonseth and time in musical art]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):65–97, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bertholet:1977:QEC

[921] Edmond Bertholet. La question de l’Être chez Heidegger et chez Gonseth. (French) [The question of the being in Heidegger and Gonseth]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2): 99–105, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bonsack:1977:DSD

[922] François Bonsack. La dialectique de la source et du filtre. (French) [The dialectic of the source and the filter]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):107–112, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lauener:1977:FG

[923] Henri Lauener. Ferdinand Gonseth 1890–1975. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):113–118, June 1977. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

VonBernays:1977:UFG

[924] Paul Von Bernays. Überlegungen zu Ferdinand Gonseths Philosophie. (German) [Reflections on Ferdinand Gonseth’s philosophy]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):119–128, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Grize:1977:DCA

[925] Jean-Biaise Grize. Discours, connaissance et actions. (French) [Speech, knowledge and actions]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):129–139, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Soerensen:1977:FGE

[926] Werner Soerensen. Ferdinand Gonseth et l’enseignement des mathématiques. (French) [Ferdinand Gonseth and the teaching of mathematics]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):141–144, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Destouches:1977:MSC

[927] Jean-Louis Destouches. Sur la méthodologie scientifique et les conceptions de Ferdinand Gonseth. (French) [On the scientific methodology and conceptions of Ferdinand Gonseth]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):145–163, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Christoff:1977:RSI

[928] Daniel Christoff. Raison suffisante et idonéisme, fermeture et ouverture. (French) [Reason enough and idoneism, closure and opening]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):165–176, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Popper:1977:SRP

[929] Karl R. Popper. Some remarks on panpsychism and epiphenomenalism. Dialectica: International Review of Philosophy of Knowledge, 31(1–2): 177–186, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Guggenheimer:1977:ABE

[930] Heinrich Guggenheimer and Ferdinand Gonseth. The axioms of betweenness in Euclid. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):187–192, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Guggenheimer:1977:MT

[931] Eva H. Guggenheimer and F. Gonseth. Magic textiles. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):193–198, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1977:NLP

[932] Paul-Emile Pilet. Notes de lecture: Pour une ouverture biologique à la psychologie de l’intelligence. (French) [Reading notes: For a biological opening to the psychology of intelligence]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):199–202, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1977:RSEa

[933] P. E. Pilet. Reviewing studies — Études critiques — Betrachtungen zur Literatur: Le comportement, moteur de l’évolution. (French) [Behavior, the driver of evolution]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):203–204, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1977:RSEb

[934] P. E. Pilet. Reviewing studies — Études critiques — Betrachtungen zur Literatur: Évolution et conscience. (French) [Evolution and conscience]. Dialectica: International Review of Philosophy of Knowledge, 31(1–2): 204, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1977:RSEc

[935] P. E. Pilet. Reviewing studies — Études critiques — Betrachtungen zur Literatur: l’œuvre de F. L. G. Frege. (French) [The works of F. L. G. Frege]. Dialectica: International Review of Philosophy of Knowledge, 31 (1–2):205, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1977:RSEd

[936] P. E. Pilet. Reviewing studies — Études critiques — Betrachtungen zur Literatur: Science et métaphysique. (French) [Science and metaphysics].

Dialectica: International Review of Philosophy of Knowledge, 31(1–2): 205, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1977:RSEe

[937] P. E. Pilet. Reviewing studies — Études critiques — Betrachtungen zur Literatur: The problem of knowledge. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):206, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1977:RSEf

[938] P. E. Pilet. Reviewing studies — Études critiques — Betrachtungen zur Literatur: Patterns of aging. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):206, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1977:ASB

[939] Anonymous. Announcement: Sixth Biennial Meeting of the Philosophy of Science Association. Dialectica: International Review of Philosophy of Knowledge, 31(1–2):207, June 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Comite:1977:PB

[940] Le Comité. Paul Bernays. Dialectica: International Review of Philosophy of Knowledge, 31(3–4):211, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1977:V

[941] Anonymous. Vorwort. Dialectica: International Review of Philosophy of Knowledge, 31(3–4):215, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vonBernays:1977:SWI

[942] Paul von Bernays. Sprache und Wirklichkeit: Im Vorfeld der Fragestellungen. (German) [Language and reality: Before the question]. Dialectica: International Review of Philosophy of Knowledge, 31(3–4):217–223, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hintikka:1977:LG

[943] Jaakko Hintikka. Language-games. Dialectica: International Review of Philosophy of Knowledge, 31(3–4):225–245, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Davidson:1977:RR

[944] Donald Davidson. Reality without reference. Dialectica: International Review of Philosophy of Knowledge, 31(3–4):247–258, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Moravcsik:1977:STB

[945] J. M. E. Moravcsik. Singular terms, belief, and reality. Dialectica: International Review of Philosophy of Knowledge, 31(3–4):259–272, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kaeser:1977:PLP

[946] E. Kaeser. Physical laws, physical entities and ontology. Dialectica: International Review of Philosophy of Knowledge, 31(3–4):273–299, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vonJanich:1977:SPW

[947] Peter von Janich. Die Sprache der Physik und die Wirklichkeit der Naturwissenschaften. (German) [The language of physics and the reality of natural science]. Dialectica: International Review of Philosophy of Knowledge, 31(3–4):301–312, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bonsack:1977:ICR

[948] F. Bonsack. Invariance as a criterion of reality. Dialectica: International Review of Philosophy of Knowledge, 31(3–4):313–331, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vonLauener:1977:QOS

[949] Henri von Lauener. Quine über Ontologie und substitutionelle Quantifikation. (German) [Quine on ontology and substitutional quantification]. Dialectica: International Review of Philosophy of Knowledge, 31 (3–4):333–357, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Graeser:1977:LTR

[950] Andreas Graeser. On language, thought, and reality in ancient Greek philosophy. Dialectica: International Review of Philosophy of Knowledge, 31(3–4):359–388, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gochet:1977:MTP

[951] Paul Gochet. Model theory and the pragmatics of indexicals. Dialectica: International Review of Philosophy of Knowledge, 31(3–4):389–408, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stroll:1977:TAT

[952] Avrum Stroll and Robert Foelber. Talk about talk about surfaces. Dialectica: International Review of Philosophy of Knowledge, 31(3–4):409– 430, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Black:1977:MAM

[953] Max Black. More about metaphor. Dialectica: International Review of Philosophy of Knowledge, 31(3–4):431–457, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1977:E

[954] Anonymous. Errata. Dialectica: International Review of Philosophy of Knowledge, 31(3–4):459, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1977:RBL

[955] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 31(3–4): 461–462, December 1977. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hattiangadi:1978:RPS

[956] J. N. Hattiangadi. Rationality and the problem of scientific traditions. Dialectica: International Review of Philosophy of Knowledge, 32(1):3– 28, March 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Binns:1978:SAB

[957] Peter Binns. The supposed asymmetry between falsification and verification. Dialectica: International Review of Philosophy of Knowledge, 32 (1):29–40, March 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Allison:1978:TTN

[958] Henri E. Allison. Things in themselves, noumena, and the transcendental object. Dialectica: International Review of Philosophy of Knowledge, 32 (1):41–76, March 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Blum:1978:NDN

[959] Alex Blum. Notes and discussions — notes et discussions — notizen und diskussionen: Convention T and natural languages. Dialectica: International Review of Philosophy of Knowledge, 32(1):77–80, March 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1978:RSE

[960] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 32(1):81–84, March 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Woods:1978:FAI

[961] John Woods and Douglas Walton. The fallacy of ‘ad ignorantiam’. Dialectica: International Review of Philosophy of Knowledge, 32(2):87–99, June 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Black:1978:SQA

[962] Max Black. Some questions about “Practical reasoning”. Dialectica: International Review of Philosophy of Knowledge, 32(2):101–114, June 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lafrance:1978:DPJ

[963] Guy Lafrance. Les deux principes de la justice selon Rawls. (French) [The two principles of justice according to Rawls]. Dialectica: International Review of Philosophy of Knowledge, 32(2):115–123, June 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See comments [964].

Nielsen:1978:EJC

[964] Kai Nielsen. Equality, justice and class: Comments on “Les Deux Principes de la Justice selon Rawls”. Dialectica: International Review of Philosophy of Knowledge, 32(2):125–133, June 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See [963].

Duchesneau:1978:TDP

[965] François Duchesneau. Téléologie et détermination positive de l’ordre biologique. (French) [Teleology and positive determination of the biological order]. Dialectica: International Review of Philosophy of Knowledge, 32 (2):135–153, June 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Duchesne:1978:RNP

[966] Jules Duchesne. Reflections on a new perspective of the evolution of living organisms. Dialectica: International Review of Philosophy of Knowledge, 32(2):155–163, June 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Brown:1978:DSJ

[967] Richard H. Brown. Dialectic and structure in Jean-Paul Sartre and Claude Levi-Strauss. Dialectica: International Review of Philosophy of Knowledge, 32(2):165–184, June 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Engeler:1978:LWP

[968] Erwin Engeler. Zum logischen Werk von Paul Bernays. (German) [On the logical work of Paul Bernays]. Dialectica: International Review of Philosophy of Knowledge, 32(3–4):191–200, December 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Raggio:1978:RAB

[969] Andrès R. Raggio. Die Rolle der Analogie in Bernays’ Philosophie der Mathematik. (German) [The role of analogy in Bernays’ philosophy of mathematics]. Dialectica: International Review of Philosophy of Knowledge, 32(3–4):201–207, December 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hasenjaeger:1978:PZG

[970] G. Hasenjaeger. Prädikatenvariablen in der Zahlentheorie. (German) [Predicate variables in number theory]. Dialectica: International Review of Philosophy of Knowledge, 32(3–4):209–220, December 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lorenzen:1978:KAG

[971] P. Lorenzen. Konstruktive Analysis und das geometrische Kontinuum. (German) [Constructive analysis and the geometric continuum]. Dialectica: International Review of Philosophy of Knowledge, 32(3–4):221–227, December 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fraisse:1978:ASE

[972] Roland Fraïssé. Les axiomatiques sont-elles un jeu?. (French) [Are axioms a game?]. Dialectica: International Review of Philosophy of Knowledge, 32(3–4):229–244, December 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ghose:1978:DNC

[973] Amitabha Ghose. Different notions of constructivity — their ontology. Dialectica: International Review of Philosophy of Knowledge, 32(3–4): 245–253, December 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stenius:1978:FMA

[974] Erik Stenius. Foundations of mathematics: Ancient Greek and modern. Dialectica: International Review of Philosophy of Knowledge, 32(3–4): 255–290, December 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vanDalen:1978:BGH

[975] Dirk van Dalen. Brouwer: The Genesis of his intuitionism. Dialectica: International Review of Philosophy of Knowledge, 32(3–4):291–303, December 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Friedman:1978:QLC

[976] Michael Friedman and Hilary Putnam. Quantum logic, conditional probability, and interference. Dialectica: International Review of Philosophy of Knowledge, 32(3–4):305–315, December 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wermus:1978:EMA

[977] H. Wermus. Esquisse d’un modèle des activités cognitives. (French) [Outline of a model of cognitive activities]. Dialectica: International Review of Philosophy of Knowledge, 32(3–4):317–338, December 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Popper:1978:NSE

[978] Karl Popper. Natural selection and the emergence of mind. Dialectica: International Review of Philosophy of Knowledge, 32(3–4):339–355, December 1978. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vatrdy:1979:SRR

[979] P. Vatrdy and L. Fleischhacker. Some remarks on the relationship between Russell’s vicious-circle principle and Russell’s paradox. Dialectica: International Review of Philosophy of Knowledge, 33(1):3–19, March 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1979:N

[980] Anonymous. Nachwort. Dialectica: International Review of Philosophy of Knowledge, 33(1):20–22, March 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fleischhacker:1979:RVC

[981] L. E. Fleischhacker. Is Russell’s vicious circle principle false or meaningless? Dialectica: International Review of Philosophy of Knowledge, 33 (1):23–35, March 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Richards:1979:CPC

[982] John Richards and Ernst von Glasersfeld. The control of perception and the construction of reality: Epistemological aspects of the feedbackcontrol system. Dialectica: International Review of Philosophy of Knowledge, 33(1):37–58, March 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Doppelt:1979:AMA

[983] Gerald Doppelt. The Austin–Malcolm argument for the incorrigibility of perceptual reports. Dialectica: International Review of Philosophy of Knowledge, 33(1):59–75, March 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1979:RSE

[984] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 33(1):77–81, March 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Blazek:1979:CEP

[985] Bohuslaw Blažek. Can epistemology as a philosophical discipline develop into a science? Dialectica: International Review of Philosophy of Knowledge, 33(2):87–108, June 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ohe:1979:NHW

[986] Seizo Ohe. The natural and human world and the creative flow of our consciousness. Dialectica: International Review of Philosophy of Knowledge, 33(2):109–121, June 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stalker:1979:WHP

[987] Douglas F. Stalker. What’s happening in philosophy of art today? Dialectica: International Review of Philosophy of Knowledge, 33(2):123– 130, June 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Athanasiades-Mercanton:1979:VEP

[988] Marguerite Athanasiadès-Mercanton. La vestale et étinceile: À propos de pédagogie universitaire. (French) [The vestal and spark: About academic pedagogy]. Dialectica: International Review of Philosophy of Knowledge, 33(2):133–149, June 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Borel:1979:RSE

[989] Marie-Jeanne Borel. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 33(2):151, June 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1979:E

[990] Anonymous. Errata. Dialectica: International Review of Philosophy of Knowledge, 33(2):153, June 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1979:RBL

[991] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 33(2):155– 156, June 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tranoy:1979:V

[992] K. E. Tranoy. Vorwort. Dialectica: International Review of Philosophy of Knowledge, 33(3–4):159–163, December 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rorty:1979:EHF

[993] Richard Rorty. De l’épistémologie à l’herméneutique. (French) [From epistemology to hermeneutics]. Dialectica: International Review of Philosophy of Knowledge, 33(3–4):165–188, December 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Thurnher:1979:EGS

[994] Rainer Thurnher. ‘Erklärung’ in genetischen und systematischen Zusammenhängen. (German) [’Explanation’ in genetic and systematic contexts]. Dialectica: International Review of Philosophy of Knowledge, 33 (3–4):189–200, December 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Moravcsik:1979:U

[995] J. M. Moravcsik. Understanding. Dialectica: International Review of Philosophy of Knowledge, 33(3–4):201–216, December 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kung:1979:URJ

[996] Guido Kung. Understanding and its rational justification. Dialectica: International Review of Philosophy of Knowledge, 33(3–4):217–232, December 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chisholm:1979:VEQ

[997] Roderick M. Chisholm. Verstehen: The epistemological question. Dialectica: International Review of Philosophy of Knowledge, 33(3–4):233–246, December 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Frey:1979:KIG

[998] Gerhard Frey. Über die Konstruktion von Interpretationsschemata. (German) [On the construction of interpretive schemata]. Dialectica: International Review of Philosophy of Knowledge, 33(3–4):247–261, December 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jacques:1979:CCC

[999] Francis Jacques. Croyance commune et croyance communiquée. (French) [Common belief and belief communicated]. Dialectica: International Review of Philosophy of Knowledge, 33(3–4):263–280, December 1979. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zaslawsky:1979:MEI

[1000] Denis Zaslawsky. La méthode de explication informelle en philosophie logique et en linguistique. (French) [The informal method of explanation in logical philosophy and linguistics]. Dialectica: International Review of Philosophy of Knowledge, 33(3–4):281–295, December 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Parret:1979:SU

[1001] Herman Parret. Significance and understanding. Dialectica: International Review of Philosophy of Knowledge, 33(3–4):297–317, December 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Follesdal:1979:HHD

[1002] Dagfinn Føllesdal. Hermeneutics and the hypothetico-deductive method. Dialectica: International Review of Philosophy of Knowledge, 33(3–4): 319–336, December 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Black:1979:WLG

[1003] Max Black. Wittgenstein’s language-games. Dialectica: International Review of Philosophy of Knowledge, 33(3–4):337–353, December 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Weingartner:1979:AAS

[1004] P. Weingartner. Analogy among systems. Dialectica: International Review of Philosophy of Knowledge, 33(3–4):355–378, December 1979. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stroll:1979:MPE

[1005] Avrum Stroll. Moore’s proof of an external world. Dialectica: International Review of Philosophy of Knowledge, 33(3–4):379–397, December 1979. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Winters:1980:RB

[1006] Barbara Winters. Reasonable believing. Dialectica: International Review of Philosophy of Knowledge, 34(1):3–16, March 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gram:1980:KDP

[1007] Moltke S. Gram. Kant’s duplication problem. Dialectica: International Review of Philosophy of Knowledge, 34(1):17–59, March 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hottois:1980:LDG

[1008] G. Hottois. L’itinéaire déontique de G. H. von Wright. (French) [The deontic itinerary of G. H. von Wright]. Dialectica: International Review of Philosophy of Knowledge, 34(1):61–88, March 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Swiggers:1980:NDA

[1009] P. Swiggers. Notes and discussions: ‘analytically true’: De dicto and de dicto dicti. Dialectica: International Review of Philosophy of Knowledge, 34(1):89–90, March 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See remarks [1042].

Pitt:1980:HVS

[1010] Joseph C. Pitt. Hempel versus Sellars on explanation. Dialectica: International Review of Philosophy of Knowledge, 34(2):95–120, June 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rothbart:1980:PAI

[1011] Daniel Rothbart. Popper against inductivism. Dialectica: International Review of Philosophy of Knowledge, 34(2):121–128, June 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gray:1980:PAP

[1012] Bennison Gray. Popper and the 7th approximation: The problem of taxonomy. Dialectica: International Review of Philosophy of Knowledge, 34(2):129–154, June 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tibbetts:1980:NDF

[1013] Paul Tibbetts. Notes and discussions: ^{≪}the subjective element in scientific discovery: Popper versus ‘traditional epistemology’^{≫}. Dialectica: International Review of Philosophy of Knowledge, 34(2):155–160, June 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1980:C

[1014] Anonymous. Corrigendum. Dialectica: International Review of Philosophy of Knowledge, 34(2):161, June 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Grieder:1980:LR

[1015] Alfons Grieder. On the logic of relations. Dialectica: International Review of Philosophy of Knowledge, 34(3):167–182, September 1980. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Goldstick:1980:CS

[1016] D. Goldstick. Cognitive synonymy. Dialectica: International Review of Philosophy of Knowledge, 34(3):183–203, September 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Esposito:1980:CIN

[1017] Joseph L. Esposito. Convention, invention, and necessity. Dialectica: International Review of Philosophy of Knowledge, 34(3):205–210, September 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Glouberman:1980:CCH

[1018] M. Glouberman. Causation, cognition, and historical typology. Dialectica: International Review of Philosophy of Knowledge, 34(3):211–227, September 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Richmond:1980:DST

[1019] Sheldon Richmond. A discussion of some theories of pictorial representation. Dialectica: International Review of Philosophy of Knowledge, 34 (3):229–240, September 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Baker:1980:MSR

[1020] W. Baker. Mathematical structures as representations of intellectual structures. Dialectica: International Review of Philosophy of Knowledge, 34(4):247–262, December 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Byrne:1980:SEU

[1021] Patrick H. Byrne. The significance of Einstein’s use of the history of science. Dialectica: International Review of Philosophy of Knowledge, 34(4):263–276, December 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wuketits:1980:NTC

[1022] Franz M. Wuketits. On the notion of teleology in contemporary life sciences. Dialectica: International Review of Philosophy of Knowledge, 34(4):277–290, December 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1980:RSE

[1023] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 34(4):291–298, December 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1980:RBL

[1024] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 34(4):299– 302, December 1980. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Meyer:1981:TDC

[1025] Michel Meyer. The transcendental deduction of the categories: Its impact on German idealism and neo-positivism. Dialectica: International Review of Philosophy of Knowledge, 35(1):7–20, June 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vuillemin:1981:TPI

[1026] Jules Vuillemin. Trois philosophes intuitionnistes: Epicure, Descartes et Kant. (French) [Three intuitionist philosophers: Epicure, Descartes and Kant]. Dialectica: International Review of Philosophy of Knowledge, 35 (1):21–41, June 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Beck:1981:RNK

[1027] Lewis White Beck. Über die Regelmässigkeit der Natur bei Kant. (German) [On the regularity of nature in Kant]. Dialectica: International Review of Philosophy of Knowledge, 35(1):43–56, June 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Allison:1981:TSP

[1028] Henry E. Allison. Transcendental schematism and the problem of the synthetic a priori. Dialectica: International Review of Philosophy of Knowledge, 35(1):57–83, June 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Howell:1981:KFC

[1029] Robert Howell. Kant’s first- critique theory of the transcendental object. Dialectica: International Review of Philosophy of Knowledge, 35 (1):85–125, June 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hintikka:1981:KEP

[1030] Jaakko Hintikka. Kant on existence, predication, and the ontological argument. Dialectica: International Review of Philosophy of Knowledge, 35(1):127–146, June 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stenius:1981:KAM

[1031] Erik Stenius. Kant and the apriority of mathematics. Dialectica: International Review of Philosophy of Knowledge, 35(1):147–166, June 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Silber:1981:KMR

[1032] John R. Silber. Kant and the mythic roots of morality. Dialectica: International Review of Philosophy of Knowledge, 35(1):167–193, June 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hoffe:1981:TVE

[1033] Otfried Höffe. Transzendentale oder vernunftkritische Ethik (Kant)? : Zur Methodenkomplexität einer sachgerechten Moralphilosophie. (German) [Transcendental or critical criticism (Kant)? : The method complexity of a proper moral philosophy]. Dialectica: International Review of Philosophy of Knowledge, 35(1):195–221, June 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rod:1981:KAK

[1034] Wolfgang Rod. Kants Annahme einer Kausalität aus Freiheit und die Idee einer transzendentalen Ethik. (German) [Kant’s assumption of causality from freedom and the idea of transcendental ethics]. Dialectica: International Review of Philosophy of Knowledge, 35(1):223–241, June 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lauener:1981:SSG

[1035] Henri Lauener. Der systematische Stellenwert des Gefühls der Achtung in Kants Ethik. (German) [The systematic importance of the feeling of respect in Kant’s ethics]. Dialectica: International Review of Philosophy of Knowledge, 35(1):243–264, June 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Prauss:1981:KTA

[1036] Gerold Prauss. Kants Theorie der ästhetischen Einstellung. (German) [Kant’s theory of aesthetic attitude]. Dialectica: International Review of Philosophy of Knowledge, 35(1):265–281, June 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Broughton:1981:QQS

[1037] Lynne M. Broughton. Quine’s ‘quality space’. Dialectica: International Review of Philosophy of Knowledge, 35(3):291–302, September 1981. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Swiggers:1981:SC

[1038] P. Swiggers. The supermaxim of conversation. Dialectica: International Review of Philosophy of Knowledge, 35(3):303–306, September 1981. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bache:1981:PPM

[1039] Christopher M. Bache. Paraphrase and paraphrasing metaphors. Dialectica: International Review of Philosophy of Knowledge, 35(3):307–326, September 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Newman:1981:EII

[1040] Jay Newman. Epistemic inference and illative judgment. Dialectica: International Review of Philosophy of Knowledge, 35(3):327–339, September 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chisholm:1981:EP

[1041] Roderick M. Chisholm. Epistemic principles. Dialectica: International Review of Philosophy of Knowledge, 35(3):341–345, September 1981. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Auroux:1981:NDS

[1042] S. Auroux. Notes et discussions: Some remarks on Swiggers’ analyticity ‘de dicto’ and ‘de dicto dicti’. Dialectica: International Review of Philosophy of Knowledge, 35(3):347–349, September 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See [1009].

Stiffler:1981:ARV

[1043] Eric Stiffler. “On an argument for the relational view of belief”. Dialectica: International Review of Philosophy of Knowledge, 35(3):351–355, September 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1981:RSEa

[1044] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 35(3):357–363, September 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Malherbe:1981:KPC

[1045] J.-F. Malherbe. Karl Popper et Claude Bernard. (French) [Karl Popper and Claude Bernard]. Dialectica: International Review of Philosophy of Knowledge, 35(4):373–388, December 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jones:1981:KPT

[1046] Gary E. Jones. Kuhn, Popper, and theory comparison. Dialectica: International Review of Philosophy of Knowledge, 35(4):389–397, December 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

McHoul:1981:RMC

[1047] A. W. McHoul. Remarks on Moore’s conception and defence of common sense. Dialectica: International Review of Philosophy of Knowledge, 35 (4):399–413, December 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Badriyeh:1981:ATW

[1048] Katherine Badriyeh. An antiskeptical theory of when and how we know. Dialectica: International Review of Philosophy of Knowledge, 35(4):415– 432, December 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1981:RSEb

[1049] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 35(4):433–434, December 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1981:RBL

[1050] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 35(4):435– 436, December 1981. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1982:E

[1051] Paul Emile Pilet. Editorial. Dialectica: International Review of Philosophy of Knowledge, 36(1):3–5, March 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Duchesne:1982:PVU

[1052] Jules Duchesne. De la physique à la vie dans l’univers. (French) [From physics to life in the Universe]. Dialectica: International Review of Philosophy of Knowledge, 36(1):7–14, March 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1982:BBB

[1053] Paul-Emile Pilet. Biologie, biophysique et biochimie: complémentarite et discordance. (French) [Biology, biophysics and biochemistry: complementarity and discordance]. Dialectica: International Review of Philosophy of Knowledge, 36(1):15–22, March 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rivier:1982:NTD

[1054] Laurent Rivier. Une nouvelle technique de dosage biochimique: la GC/MS, progrès et quelques réflexions critiques. (French) [A new biochemical assay technique: the GC/MS, progress and some critical reflections]. Dialectica: International Review of Philosophy of Knowledge, 36 (1):23–28, March 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Boiler:1982:IEZ

[1055] Thomas Boiler. Isolation und Eigenschaften von Zellorganellen. (German) [Isolation and properties of cell organelles]. Dialectica: International Review of Philosophy of Knowledge, 36(1):29–35, March 1982. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Beffa:1982:LMA

[1056] Roland Beffa. Les limites d’un modèle d’analyse en phytobiologie expérimentale: quelques remarques critiques. (French) [The limitations of a model of analysis in experimental phytobiology: some critical remarks]. Dialectica: International Review of Philosophy of Knowledge, 36 (1):37–41, March 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zaslawsky:1982:NEC

[1057] Denis Zaslawsky. Note sur explication causale en biologie. (French) [Note on causal explanation in biology]. Dialectica: International Review of Philosophy of Knowledge, 36(1):43–50, March 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pernet:1982:BLS

[1058] Jean-Jacques Pernet. Buts et limites de la statistique en biologie expérimentale. (French) [Aims and limits of statistics in experimental biology]. Dialectica: International Review of Philosophy of Knowledge, 36(1):51–54, March 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Heller:1982:LRP

[1059] René Heller. L’emploi des radioisotopes dans les problèmes d’absorption. (French) [The use of radioisotopes in absorption problems]. Dialectica: International Review of Philosophy of Knowledge, 36(1):55–61, March 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Turian:1982:MMF

[1060] Gilbert Turian. Méthodes en micromycologie. (French) [Methods in micromycology]. Dialectica: International Review of Philosophy of Knowledge, 36(1):63–69, March 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wahli:1982:ACE

[1061] Walter Wahli. Analyses critiques de expression génétique. (French) [Critical expression of gene expression]. Dialectica: International Review of Philosophy of Knowledge, 36(1):71–81, March 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jolles:1982:APD

[1062] Pierre Jollés and Jacqueline Jollés. Autopsie d’une protéine: determination de sa séquence. (French) [Autopsy of a protein: determination of its sequence]. Dialectica: International Review of Philosophy of Knowledge, 36(1):83–89, March 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Widmer:1982:ERE

[1063] François Widmer. Les enzymes: réalité expérimentale et implications physiologiques. (French) [Enzymes: experimental reality and physiological implications]. Dialectica: International Review of Philosophy of Knowledge, 36(1):91–98, March 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Guignard:1982:NPG

[1064] Roland Guignard. NAD(P)-glycohydrolase et différentiation sporale chez Neurospora crassa: analyse de la fonction d’une enzyme. (French) [NAD (P)-glycohydrolase and sporadic differentiation in Neurospora crassa: analysis of the function of an enzyme]. Dialectica: International Review of Philosophy of Knowledge, 36(1):99–104, March 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Steck:1982:SRC

[1065] Paul Steck. Situation de la recherche à caractère appliqué. (French) [Status of applied research]. Dialectica: International Review of Philosophy of Knowledge, 36(1):105–116, March 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1982:VV

[1066] Anonymous. Vorwort: Vorwort. Dialectica: International Review of Philosophy of Knowledge, 36(2–3):125, September 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Agassi:1982:IT

[1067] Joseph Agassi. Irrationalism today. Dialectica: International Review of Philosophy of Knowledge, 36(2–3):127–146, September 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Black:1982:WSR

[1068] Max Black and Oscar Wilde. “Why should I be rational?”. Dialectica: International Review of Philosophy of Knowledge, 36(2–3):147–168, September 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chisholm:1982:ER

[1069] Roderick M. Chisholm. Epistemic reasoning. Dialectica: International Review of Philosophy of Knowledge, 36(2–3):169–177, September 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stroll:1982:PKR

[1070] Avrum Stroll. Primordial knowledge and rationality. Dialectica: International Review of Philosophy of Knowledge, 36(2–3):179–201, September 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bornet:1982:MAN

[1071] Gérard Bornet. Modes of analyticity — a note to A. Stroll’s “Primordial Knowledge and Rationality”. Dialectica: International Review of Philosophy of Knowledge, 36(2–3):203–205, September 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Moravcsik:1982:UE

[1072] J. M. E. Moravcsik. Understanding and the emotions. Dialectica: International Review of Philosophy of Knowledge, 36(2–3):207–224, September 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Harding:1982:GVC

[1073] Sandra Harding. Is gender a variable in conceptions of rationality? a survey of issues. Dialectica: International Review of Philosophy of Knowledge, 36(2–3):225–242, September 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Weingartner:1982:CRC

[1074] Paul Weingartner. Conditions of rationality for the concepts belief, knowledge, and assumption. Dialectica: International Review of Philosophy of Knowledge, 36(2–3):243–263, September 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tennant:1982:PP

[1075] Neil Tennant. Proof and paradox. Dialectica: International Review of Philosophy of Knowledge, 36(2–3):265–296, September 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Follesdal:1982:SRA

[1076] Dagfinn Follesdal. The status of rationality assumptions in interpretation and in the explanation of action. Dialectica: International Review of Philosophy of Knowledge, 36(4):301–316, December 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Davidson:1982:RA

[1077] Donald Davidson. Rational animals. Dialectica: International Review of Philosophy of Knowledge, 36(4):317–327, December 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cooke:1982:RAR

[1078] Roger M. Cooke. Risk assessment and rational decision theory. Dialectica: International Review of Philosophy of Knowledge, 36(4):329–351, December 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lauener:1982:MPL

[1079] Henri Lauener. Method in philosophy and logic. Dialectica: International Review of Philosophy of Knowledge, 36(4):353–371, December 1982. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1982:RBL

[1080] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 36(4):373– 374, December 1982. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Thibaud:1983:NPI

[1081] Pierre Thibaud. La notion peircéenne d’interprétant. (French) [The Peircean notion of interpreter]. Dialectica: International Review of Philosophy of Knowledge, 37(1):3–33, March 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dau:1983:WHT

[1082] Paolo Dau. What historical theorists haven’t shown. Dialectica: International Review of Philosophy of Knowledge, 37(1):35–56, March 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1983:RSEa

[1083] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 37(1):57–70, March 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schmid:1983:CIE

[1084] Anne-Françoise Schmid. La correspondance inédite entre Bertrand Russell et Louis Couturat. (French) [The unpublished correspondence between Bertrand Russell and Louis Couturat]. Dialectica: International Review of Philosophy of Knowledge, 37(2):75–109, June 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vernant:1983:TLE

[1085] Denis Vernant. Le traitement logique de l’existence et les présupposés de l’ontologie. (French) [The logical treatment of existence and the assumptions of ontology]. Dialectica: International Review of Philosophy of Knowledge, 37(2):111–132, June 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

LePore:1983:CMR

[1086] Ernest Le Pore. The concept of meaning and its role in understanding language: Notes and discussions. Dialectica: International Review of Philosophy of Knowledge, 37(2):133–139, June 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Titze:1983:DWB

[1087] Hans Titze. Dialektik und Widerspruch Bemerkungen zu einem neuen ungarischen Buch. (German) [Dialectics and opposition remarks on a new Hungarian book]. Dialectica: International Review of Philosophy of Knowledge, 37(2):141–151, June 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wegener:1983:QWV

[1088] Ursula Wegener. Quantenmechanik und Wahrscheinlichkeit : Ein Vergleich der von von Weizsäcker und Drieschner bzw. von Ludwig vorgeschlagenen Interpretationen der Quantenmechanik. (German) [Quantum mechanics and probability: A comparison of the interpretations of quantum mechanics proposed by Weizsäcker and Drieschner or by Ludwig]. Dialectica: International Review of Philosophy of Knowledge, 37(3):155– 167, September 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wassermann:1983:HBB

[1089] G. D. Wassermann. Human behaviour and biology. Dialectica: International Review of Philosophy of Knowledge, 37(3):169–184, September 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Paty:1983:CRC

[1090] Michel Paty. La critique rationaliste de la création au 18éme siècle. (French) [The rationalist critique of creation in the 18th century]. Dialectica: International Review of Philosophy of Knowledge, 37(3):185– 200, September 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Woods:1983:KTS

[1091] M. Woods. Kant’s transcendental schematism. Dialectica: International Review of Philosophy of Knowledge, 37(3):201–219, September 1983. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Moser:1983:NDR

[1092] Paul K. Moser. Notes et discussions: : Rationality without surprises: Davidson on rational belief. Dialectica: International Review of Philosophy of Knowledge, 37(3):221–226, September 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Beckermann:1983:RSE

[1093] Ansgar Beckermann. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 37(3):227–232, September 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vuillemin:1983:CCM

[1094] Jules Vuillemin. Le carré Chrysippéen des modalités. (French) [The Chrysippean square of the modalities]. Dialectica: International Review of Philosophy of Knowledge, 37(4):235–247, December 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Couture:1983:CPM

[1095] Jocelyne Couture. Les classes dans les Principia Mathematica sontelles des expressions incomplétes?. (French) [Are the classes in Principia Mathematica incomplete expressions?]. Dialectica: International Review of Philosophy of Knowledge, 37(4):249–267, December 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marti-Huang:1983:PRM

[1096] Duen Marti-Huang. Poppers Rückzug in die Metaphysik. (German) [Popper’s retreat into metaphysics]. Dialectica: International Review of Philosophy of Knowledge, 37(4):269–283, December 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Maloney:1983:NMM

[1097] J. Christopher Maloney. A new model for metaphor. Dialectica: International Review of Philosophy of Knowledge, 37(4):285–301, December 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1983:RSEb

[1098] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 37(4):303–314, December 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1983:RBL

[1099] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 37(4):315– 316, December 1983. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Shapiro:1984:ASE

[1100] Stewart Shapiro. Arithmetic sinn and effectiveness. Dialectica: International Review of Philosophy of Knowledge, 38(1):3–16, March 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Slezak:1984:MMS

[1101] Peter Slezak. Minds, machines and self-reference. Dialectica: International Review of Philosophy of Knowledge, 38(1):17–34, March 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Molino:1984:NDP

[1102] Jean Molino. Notes et discussions: À propos de être ou être selon D. Zaslawsky. (French) [Notes and discussions: To be or be according to D. Zaslawsky]. Dialectica: International Review of Philosophy of Knowledge, 38(1):35–55, March 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See response [1103].

Zaslawsky:1984:RJM

[1103] Denis Zaslawsky. Réponse à J. Molino. (French) [Response to J. Molino]. Dialectica: International Review of Philosophy of Knowledge, 38(1):56– 66, March 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See [1102].

Khatchadourian:1984:MI

[1104] Haig Khatchadourian. Metaphor II. Dialectica: International Review of Philosophy of Knowledge, 38(1):67–76, March 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1984:RSEa

[1105] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 38(1):77–80, March 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1984:MM

[1106] Anonymous. Meetings: Meetings. Dialectica: International Review of Philosophy of Knowledge, 38(1):80, March 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Searle:1984:IPN

[1107] John R. Searle. Intentionality and its place in nature. Dialectica: International Review of Philosophy of Knowledge, 38(2–3):87–99, September 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Davidson:1984:FPA

[1108] Donald Davidson. First person authority. Dialectica: International Review of Philosophy of Knowledge, 38(2–3):101–111, September 1984. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sokolowski:1984:IAN

[1109] Robert Sokolowski. Intentional analysis and the noema. Dialectica: International Review of Philosophy of Knowledge, 38(2–3):113–129, September 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chisholm:1984:WPO

[1110] Roderick M. Chisholm. What is the problem of objective reference? Dialectica: International Review of Philosophy of Knowledge, 38(2–3): 131–142, September 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kung:1984:IRO

[1111] Guido Küng. The intentional and the real object. Dialectica: International Review of Philosophy of Knowledge, 38(2–3):143–156, September 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Smith:1984:AFR

[1112] Barry Smith. Acta cum fundamentis in Re. Dialectica: International Review of Philosophy of Knowledge, 38(2–3):157–178, September 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Haller:1984:OAA

[1113] Rudolf Haller. Objects, acts, and attitudes. Dialectica: International Review of Philosophy of Knowledge, 38(2–3):179–190, September 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stroker:1984:IKW

[1114] Elisabeth Ströker. Intentionalität und Konstitution. Wandlungen des Intentionalitätskonzepts in der Philosophie Husserls. (German) [Intentality and constitution. Changes of the intentionality concept in Husserl’s philosophy]. Dialectica: International Review of Philosophy of Knowledge, 38(2–3):191–208, September 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marbach:1984:UIE

[1115] Eduard Marbach. On using intentionality in empirical phenomenology: the problem of‘Mental images. Dialectica: International Review of Philosophy of Knowledge, 38(2–3):209–229, September 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Aebli:1984:WIW

[1116] Hans Aebli. What is intentionality and who has intentions in a structuralist model of knowledge, action and thought. Dialectica: International Review of Philosophy of Knowledge, 38(2–3):231–242, September 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Christoff:1984:LPS

[1117] Daniel Christoff. L’intentionnalité et le problème de la signification. (French) [Intentionality and the problem of meaning]. Dialectica: International Review of Philosophy of Knowledge, 38(2–3):245–262, September 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1984:E

[1118] Anonymous. Erratum. Dialectica: International Review of Philosophy of Knowledge, 38(4):285, December 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Robert:1984:ANA

[1119] A. Robert. Une approche naïve de analyse non-standard. (French) [A nonstandard analysis approach]. Dialectica: International Review of Philosophy of Knowledge, 38(4):287–296, December 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Farkas:1984:PNP

[1120] E. J. Farkas and M. E. Szabo. “On the plausibility of nonstandard proofs in analysis”. Dialectica: International Review of Philosophy of Knowledge, 38(4):297–310, December 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Szafarz:1984:RML

[1121] Ariane Szafarz. Richard von Mises: L’échec d’une axiomatique. (French) [Richard von Mises: The failure of an axiomatic]. Dialectica: International Review of Philosophy of Knowledge, 38(4):311–317, December 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Parrochia:1984:OMC

[1122] Daniel Parrochia. Optique, mécanique et calcul des chances chez Huygens et Spinoza: (sur quelques paradigmes possibles du discours philosophique). (French) [Optics, mechanics and calculus of chances in Huygens and Spinoza: (on some possible paradigms of philosophical discourse)]. Dialectica: International Review of Philosophy of Knowledge, 38(4):319–345, December 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schock:1984:NDR

[1123] Rolf Schock. Notes and discussions: Ravens, grue, and material implication. Dialectica: International Review of Philosophy of Knowledge, 38 (4):347–350, December 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1984:RSEb

[1124] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 38(4):351–357, December 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1984:RBL

[1125] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 38(4):359– 360, December 1984. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Soerensen:1985:AP

[1126] W. Soerensen. Avant-propos. Dialectica: International Review of Philosophy of Knowledge, 39(1):4, March 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lauener:1985:GBF

[1127] Henri Lauener. Gaston Bachelard et Ferdinand Gonseth, philosophes de la dialectique scientifique. (French) [Gaston Bachelard and Ferdinand Gonseth, philosophers of scientific dialectics]. Dialectica: International Review of Philosophy of Knowledge, 39(1):5–18, March 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Black:1985:MIC

[1128] Max Black. Making intelligent choices: How useful is decision theory? Dialectica: International Review of Philosophy of Knowledge, 39(1):19– 34, March 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Millar:1985:WUM

[1129] Alan Millar. Where’s the use in meaning? Dialectica: International Review of Philosophy of Knowledge, 39(1):35–51, March 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jason:1985:TPI

[1130] Gary J. Jason. Two problems of induction. Dialectica: International Review of Philosophy of Knowledge, 39(1):53–74, March 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1985:NHS

[1131] Anonymous. A new heading: Summaries of doctoral theses. Dialectica: International Review of Philosophy of Knowledge, 39(1):75, March 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Shimony:1985:I

[1132] Abner Shimony. Introduction. Dialectica: International Review of Philosophy of Knowledge, 39(2):83–84, June 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1985:ELB

[1133] Anonymous. An exchange on local beables. Dialectica: International Review of Philosophy of Knowledge, 39(2):85–96, June 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bonsack:1985:IBD

[1134] François Bonsack. L;inégalité de Bell: démonstration intuitive et commentaires. (French) [Bell inequality: intuitive demonstration and comments]. Dialectica: International Review of Philosophy of Knowledge, 39 (2):111–125, June 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Balzer:1985:NDT

[1135] W. Balzer. On a new definition of theoreticity. Dialectica: International Review of Philosophy of Knowledge, 39(2):127–145, June 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wreen:1985:MOV

[1136] Michael Wreen. Mackie on the objectivity of values. Dialectica: International Review of Philosophy of Knowledge, 39(2):147–156, June 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gasser:1985:RSE

[1137] James Gasser. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 39(2):157, June 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lauener:1985:V

[1138] H. Lauener. Vorwort. Dialectica: International Review of Philosophy of Knowledge, 39(3):163–164, September 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mieville:1985:ACE

[1139] Denis Miéville. Un aperçu des caractéristiques et de esprit des systèmes logiques de Stanisław Leniewski. (French) [An overview of the characteristics and spirit of the logical systems of Stanisław Leniewski]. Dialectica: International Review of Philosophy of Knowledge, 39(3):165–179, September 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rickey:1985:ILO

[1140] V. Frederick Rickey. Interpretations of Leśniewski’s ontology. Dialectica: International Review of Philosophy of Knowledge, 39(3):181–192, September 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Simons:1985:SO

[1141] Peter M. Simons. A semantics for ontology. Dialectica: International Review of Philosophy of Knowledge, 39(3):193–216, September 1985. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lejewski:1985:AIN

[1142] Czestaw Lejewski. Accommodating the informal notion of class within the framework of Lesaniewski’s ontology. Dialectica: International Review of Philosophy of Knowledge, 39(3):217–241, September 1985. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kung:1985:LEE

[1143] Guido Küng. La logique est-elle une discipline des mathématiques ou faitelle partie de ontologie?. (French) [Is logic a discipline of mathematics or is it part of ontology?]. Dialectica: International Review of Philosophy of Knowledge, 39(3):243–258, September 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Black:1985:EOF

[1144] Max Black. Entretiens d’Oxford. (French) [Interviews at Oxford]. Dialectica: International Review of Philosophy of Knowledge, 39(4):263, December 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ayer:1985:SR

[1145] Alfred J. Ayer. The scope of reason. Dialectica: International Review of Philosophy of Knowledge, 39(4):265–277, December 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kolakowski:1985:IP

[1146] Leszek Kolakowski. Irrationality in politics. Dialectica: International Review of Philosophy of Knowledge, 39(4):279–290, December 1985. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ladriere:1985:RIP

[1147] Jean Ladrière. Rationalité et irrationalité en politique. (French) [Rationality and irrationality in politics]. Dialectica: International Review of Philosophy of Knowledge, 39(4):291–296, December 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ricoeur:1985:IPP

[1148] Paul Ricoeur. Irrationality and the plurality of philosophical systems. Dialectica: International Review of Philosophy of Knowledge, 39(4):297– 319, December 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marcus:1985:TIE

[1149] Ruth Barcan Marcus. Is there irrationality in the existence of a plurality of philosophical theories. Dialectica: International Review of Philosophy of Knowledge, 39(4):321–328, December 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marx:1985:EMR

[1150] Werner Marx. Ethos and mortality: Reflections on nonrational elements in the formation of personal virtues. Dialectica: International Review of Philosophy of Knowledge, 39(4):329–338, December 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pears:1985:RPM

[1151] David Pears. Reply to Professor Marx’s paper. Dialectica: International Review of Philosophy of Knowledge, 39(4):339–344, December 1985. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Davidson:1985:II

[1152] Donald Davidson. Incoherence and irrationality. Dialectica: International Review of Philosophy of Knowledge, 39(4):345–354, December 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Granger:1985:DNR

[1153] Gilles G. Granger. Les deux niveaux de la rationalité. (French) [The two levels of rationality]. Dialectica: International Review of Philosophy of Knowledge, 39(4):355–363, December 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1985:RBL

[1154] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 39(4):365– 366, December 1985. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1986:RSI

[1155] Anonymous. Reviewing studies: 11. internationale Wittgenstein symposium. Dialectica: International Review of Philosophy of Knowledge, 40 (1):i, March 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1986:RSP

[1156] Anonymous. Reviewing studies: Philosophica: An international journal of philosophical inquiry on epistemological, axiological and social-political issues. Dialectica: International Review of Philosophy of Knowledge, 40 (1):ii, March 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1986:RSM

[1157] Anonymous. Reviewing studies: MANUSCRITO. Dialectica: International Review of Philosophy of Knowledge, 40(1):iii, March 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1986:RSG

[1158] Anonymous. Reviewing studies: Grazer Philosophische Studien. Dialectica: International Review of Philosophy of Knowledge, 40(1):iv, March 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stroud-Drinkwater:1986:SFS

[1159] Clive Stroud-Drinkwater. Seeing and following some rules. Dialectica: International Review of Philosophy of Knowledge, 40(1):3–18, March 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Thibaud:1986:NPO

[1160] Pierre Thibaud. La notion peircéenne d’objet d’un signe. (French) [The Peircean notion of the object of a sign]. Dialectica: International Review of Philosophy of Knowledge, 40(1):19–43, March 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jason:1986:NDE

[1161] G. James Jason. Notes and discussions: Epistemologies and apologies. Dialectica: International Review of Philosophy of Knowledge, 40(1):45– 58, March 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Granger:1986:NDN

[1162] Gilles Granger. Notes and discussions: Nécéssité ou contingence. (French) [Necessity or contingency]. Dialectica: International Review of Philosophy of Knowledge, 40(1):59–70, March 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hirsch:1986:NDA

[1163] Guy Hirsch. Notes and discussions: Abraham Robinson-selected papers. Dialectica: International Review of Philosophy of Knowledge, 40 (1):71–78, March 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1986:RSE

[1164] Paul Emile Pilet. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 40(1):79–80, March 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Laurier:1986:NCP

[1165] Daniel Laurier. Nouvelles catégories pour l’analyse du sens du locuteur. (French) [New categories for the analysis of the speaker’s meaning]. Dialectica: International Review of Philosophy of Knowledge, 40(2):87–106, June 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Qvarnstrom:1986:QTO

[1166] Bengt-Olof Qvarnström. Quine’s theory of observation sentence understanding and his inscrutability thesis. Dialectica: International Review of Philosophy of Knowledge, 40(2):107–120, June 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Palmquist:1986:SPO

[1167] S. R. Palmquist. Six perspectives on the object in Kant’s theory of knowledge. Dialectica: International Review of Philosophy of Knowledge, 40(2):121–151, June 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Holly:1986:NDD

[1168] W. J. Holly. Notes et discussions: On Donald Davidson’s first person authority. Dialectica: International Review of Philosophy of Knowledge, 40(2):153–156, June 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Loh:1986:FKA

[1169] Werner Loh. Fehldeutungen der klassischen Aussagenlogik. (German) [Misconceptions of classical propositional logic]. Dialectica: International Review of Philosophy of Knowledge, 40(2):157–162, June 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Weir:1986:RB

[1170] Alan Weir. Realism and behaviourism. Dialectica: International Review of Philosophy of Knowledge, 40(3):167–200, September 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Swiggers:1986:FDD

[1171] P. Swiggers. Les fondements définitionnels du réalisme constructionnel de Nelson Goodman. (French) [The definitive basis of the constructional realism of Nelson Goodman]. Dialectica: International Review of Philosophy of Knowledge, 40(3):201–211, September 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stroud-Drinkwater:1986:SSR

[1172] Clive Stroud-Drinkwater. Seeing and speaking of rabbits. Dialectica: International Review of Philosophy of Knowledge, 40(3):213–227, September 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sesardic:1986:PPC

[1173] Neven Sesardić. Psychology without principle of charity. Dialectica: International Review of Philosophy of Knowledge, 40(3):229–240, September 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Laurier:1986:LPL

[1174] Daniel Laurier. La langue d’une population: le lien entre la sémantique et la pragmatique. (French) [The language of a population: the link between semantics and pragmatics]. Dialectica: International Review of Philosophy of Knowledge, 40(4):251–272, December 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vernant:1986:QSC

[1175] Denis Vernant. Quantification substitutionnelle, contextes intensionnels et question d’existence. (French) [Substitutional quantification, intensional contexts and the question of existence]. Dialectica: International Review of Philosophy of Knowledge, 40(4):273–296, December 1986. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dubucs:1986:RAC

[1176] Jacques Dubucs. Réalisme et antimécanisme chez K. Gödel. (French) [Realism and antimechanism in K. Gödel]. Dialectica: International Review of Philosophy of Knowledge, 40(4):297–308, December 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Grene:1986:EMD

[1177] Marjorie Grene. Die Einheit des Menschen: Descartes unter den Scholastikern. (German) [The unity of man: Descartes among scholastics]. Dialectica: International Review of Philosophy of Knowledge, 40 (4):309–322, December 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1986:RSE

[1178] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 40(4):323–325, December 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1986:RBL

[1179] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 40(4):326, December 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1986:AAA

[1180] Anonymous. Advices to authors: Advices to authors. Dialectica: International Review of Philosophy of Knowledge, 40(4):329, December 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1986:CAC

[1181] Anonymous. Conseils aux auteurs: Conseils aux auteurs. Dialectica: International Review of Philosophy of Knowledge, 40(4):329–330, December 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1986:AAG

[1182] Anonymous. Anweisungen Für Autoren. (German) [Instructions for authors]. Dialectica: International Review of Philosophy of Knowledge, 40 (4):330, December 1986. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1987:VGF

[1183] Anonymous. Vorwort. (German) [Foreword]. Dialectica: International Review of Philosophy of Knowledge, 41(1–2):5–6, June 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stroll:1987:N

[1184] Avrum Stroll. Norms. Dialectica: International Review of Philosophy of Knowledge, 41(1–2):7–22, June 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lauener:1987:PNT

[1185] Henri Lauener. Philosophie als normative Tätigkeit (offener Transzendentalismus versus Naturalismus). (German) [Philosophy as a normative activity (open transcendentalism versus naturalism)]. Dialectica: International Review of Philosophy of Knowledge, 41(1–2):23–38, June 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Granger:1987:CNA

[1186] Gilles Gaston Granger. Conventions, normes, axiomes dans la connaissance des faits humains. (French) [Conventions, norms, axioms in the knowledge of human facts]. Dialectica: International Review of Philosophy of Knowledge, 41(1–2):39–67, June 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Putnam:1987:TCD

[1187] Hilary Putnam. Truth and convention: On Davidson’s refutation of conceptual Relativism. Dialectica: International Review of Philosophy of Knowledge, 41(1–2):69–77, June 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tennant:1987:CNC

[1188] Neil Tennant. Conventional necessity and the contingency of convention. Dialectica: International Review of Philosophy of Knowledge, 41(1–2): 79–95, June 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Essler:1987:SKG

[1189] Wilhelm K. Essler. Sprache und Konvention. (German) [Language and convention]. Dialectica: International Review of Philosophy of Knowledge, 41(1–2):97–116, June 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Haller:1987:RUG

[1190] Rudolf Haller. Regelbrauch und Übereinkunft. (German) [Common use and agreement]. Dialectica: International Review of Philosophy of Knowledge, 41(1–2):117–128, June 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lorenz:1987:OR

[1191] Kuno Lorenz. Is and ought revisited. Dialectica: International Review of Philosophy of Knowledge, 41(1–2):129–144, June 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marti-Huang:1987:OC

[1192] Duen Marti-Huang. The “Is” and “Ought” convention. Dialectica: International Review of Philosophy of Knowledge, 41(1–2):145–153, June 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vuillemin:1987:JPC

[1193] Jules Vuillemin. La justice par convention: signification philosophique de la doctrine de Rawls. (French) [Justice by convention: philosophical significance of the doctrine of Rawls]. Dialectica: International Review of Philosophy of Knowledge, 41(1–2):155–166, June 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Boos:1987:TRS

[1194] William Boos. Theory-relative skepticism. Dialectica: International Review of Philosophy of Knowledge, 41(3):175–207, September 1987. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kuyper:1987:KIE

[1195] Gerard Kuyper. Is knowledge an internal or an external relation between a person and a proposition? Dialectica: International Review of Philosophy of Knowledge, 41(3):209–220, September 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lugg:1987:BFF

[1196] Andrew Lugg. Bunkum, flim-flam and quackery: Pseudoscience as a philosophical problem. Dialectica: International Review of Philosophy of Knowledge, 41(3):221–230, September 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Larson:1987:CTD

[1197] David T. Larson. Correspondence and the third dogma. Dialectica: International Review of Philosophy of Knowledge, 41(3):231–236, September 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Holcomb:1987:LAP

[1198] Harmon Holcomb. Logicism and Achinstein’s pragmatic theory of scientific explanation. Dialectica: International Review of Philosophy of Knowledge, 41(3):239–248, September 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1987:RSE

[1199] Paul-Emile Pilet and Paul-Emile Pilet. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 41(3):249–250, September 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stekeler-Weithofer:1987:HSB

[1200] Pirmin Stekeler-Weithofer. Handlung, Sprache und Bewusstsein Zum Szientismus in Sprach- und Erkenntnistheorien. (German) [Action, language, and consciousness of science in linguistic and cognitive theory]. Dialectica: International Review of Philosophy of Knowledge, 41(4):255– 272, December 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nola:1987:SPC

[1201] Robert Nola. Some problems concerning the theory-ladenness of observations. Dialectica: International Review of Philosophy of Knowledge, 41(4):273–292, December 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jacquette:1987:KMB

[1202] Dale Jacquette. Kripke and the mind-body problem. Dialectica: International Review of Philosophy of Knowledge, 41(4):293–300, December 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Neale:1987:MGI

[1203] Stephen Neale. Meaning, grammar, and indeterminacy. Dialectica: International Review of Philosophy of Knowledge, 41(4):301–319, December 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Leroux:1987:SVT

[1204] Jean Leroux. The structuralist view of theories and empiricism. Dialectica: International Review of Philosophy of Knowledge, 41(4):321–326, December 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1987:RBL

[1205] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 41(4):327– 328, December 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1987:AA

[1206] Anonymous. Advices to authors. Dialectica: International Review of Philosophy of Knowledge, 41(4):347–348, December 1987. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Larson:1988:TDT

[1207] David Larson. Tarski, Davidson, and theories of truth. Dialectica: International Review of Philosophy of Knowledge, 42(1):3–16, March 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Malpas:1988:NIC

[1208] J. E. Malpas. The nature of interpretative charity. Dialectica: International Review of Philosophy of Knowledge, 42(1):17–36, March 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Geurts:1988:NDI

[1209] J. P. M. Geurts and J. Van Brakel. Notes et discussions: Internal realism, truth and charity. Dialectica: International Review of Philosophy of Knowledge, 42(1):37–44, March 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Callaway:1988:ATC

[1210] H. G. Callaway. Ascent to truth: a critical examination of Quine’s philosophy. Dialectica: International Review of Philosophy of Knowledge, 42 (1):45–58, March 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bertolet:1988:CSM

[1211] Rod Bertolet. Critical study of Michael Devitt, realism and truth. Dialectica: International Review of Philosophy of Knowledge, 42(1):59–72, March 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Biro:1988:RSE

[1212] John Biro. Reviewing studies — Études critiques — Betrachtungen zur Literatur. contemporary philosophy: A new survey. Dialectica: International Review of Philosophy of Knowledge, 42(1):73–76, March 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dilworth:1988:IEE

[1213] Craig Dilworth. Identity, equality and equivalence. Dialectica: International Review of Philosophy of Knowledge, 42(2):83–92, June 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stemmer:1988:HTA

[1214] Nathan Stemmer. Hume’s two assumptions. Dialectica: International Review of Philosophy of Knowledge, 42(2):93–104, June 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Henderson:1988:WDA

[1215] David Henderson. Wittgenstein’s descriptivist approach to understanding: Is there a place for explanation in interpretive accounts? Dialectica: International Review of Philosophy of Knowledge, 42(2):105–115, June 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vuillemin:1988:MMC

[1216] M. Jules Vuillemin. Le Misanthrope, mythe de la comédie. (French) [The Misanthrope, a myth of comedy]. Dialectica: International Review of Philosophy of Knowledge, 42(2):117–127, June 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Elugardo:1988:NDA

[1217] Reinaldo Elugardo. Notes et discussions: Against weak psychophysical supervenience. Dialectica: International Review of Philosophy of Knowledge, 42(2):129–143, June 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

OMeara:1988:RSP

[1218] Dominic J. O’Meara. Reviewing studies: Die Philosophie der Antike 3 Stoa, Epikureismus und Skepsis. Dialectica: International Review of Philosophy of Knowledge, 42(2):145, June 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Specker:1988:PMA

[1219] Ernst Specker. Postmoderne mathematik: Abschied vom paradies? Dialectica: International Review of Philosophy of Knowledge, 42(3):163– 170, September 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dorato:1988:WWQ

[1220] Mauro Dorato and Galileo. The world of the worms and the quest for reality. Dialectica: International Review of Philosophy of Knowledge, 42 (3):171–182, September 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Humberstone:1988:SEC

[1221] I. L. Humberstone. Some epistemic capacities. Dialectica: International Review of Philosophy of Knowledge, 42(3):183–200, September 1988. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Needham:1988:CRC

[1222] Paul Needham. Causation: Relation or connective? Dialectica: International Review of Philosophy of Knowledge, 42(3):201–220, September 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schwerin:1988:NDA

[1223] Alan Schwerin. Notes et discussions: An analysis of two accounts on the sense of singular terms. Dialectica: International Review of Philosophy of Knowledge, 42(3):221–232, September 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Grize:1988:RSE

[1224] Jean-Blaise Grize. Reviewing studies — Études critiques — Betrachtungen zur Literatur — Epistémologie et sciences humaines. (French) [Epistemology and social sciences]. Dialectica: International Review of Philosophy of Knowledge, 42(3):233–235, September 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

White:1988:EVH

[1225] Morton White. Epistemologie vor dem hintergrund politischer wertungen. Dialectica: International Review of Philosophy of Knowledge, 42 (4):243–264, December 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

DeQueiroz:1988:PTA

[1226] Ruy J. G. B. De Queiroz. A proof-theoretic account of programming and the role of reduction rules. Dialectica: International Review of Philosophy of Knowledge, 42(4):265–282, December 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Genova:1988:QDU

[1227] A. C. Genova. Quine’s dilemma of underdetermination. Dialectica: International Review of Philosophy of Knowledge, 42(4):283–294, December 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Duran:1988:NDR

[1228] Jane Duran. Notes et discussions: Reductionism and the naturalization of epistemology. Dialectica: International Review of Philosophy of Knowledge, 42(4):295–306, December 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Blum:1988:VFQ

[1229] Alex Blum. A version of the Frege — Quine argument. Dialectica: International Review of Philosophy of Knowledge, 42(4):307–312, December 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1988:RSE

[1230] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur: Pragmatik. (German) [Pragmatics]. Dialectica: International Review of Philosophy of Knowledge, 42(4):313–318, December 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1988:RBL

[1231] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 42(4):319– 320, December 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1988:CA

[1232] Anonymous. Conseils aux auteurs. Dialectica: International Review of Philosophy of Knowledge, 42(4):323–324, December 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1988:AA

[1233] Anonymous. Advices to authors. Dialectica: International Review of Philosophy of Knowledge, 42(4):323, December 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1988:AAG

[1234] Anonymous. Anweisungen für Autoren. (German) [Instructions for authors]. Dialectica: International Review of Philosophy of Knowledge, 42 (4):324, December 1988. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Boyd:1989:WRI

[1235] Richard Boyd. What realism implies and what it does not. Dialectica: International Review of Philosophy of Knowledge, 43(1–2):5–29, June 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Newton-Smith:1989:TR

[1236] W. H. Newton-Smith. The truth in realism. Dialectica: International Review of Philosophy of Knowledge, 43(1–2):31–45, June 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stein:1989:YBS

[1237] Howard Stein. Yes, but… some skeptical remarks on realism and antirealism. Dialectica: International Review of Philosophy of Knowledge, 43(1–2):47–65, June 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bonsack:1989:PRE

[1238] François Bonsack. Prolegomena to a realist epistemology. Dialectica: International Review of Philosophy of Knowledge, 43(1–2):67–81, June 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Agazzi:1989:NRN

[1239] Evandro Agazzi. Naive realism and naive antirealism. Dialectica: International Review of Philosophy of Knowledge, 43(1–2):83–98, June 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Worrall:1989:SRB

[1240] John Worrall. Structural realism: The best of both worlds? Dialectica: International Review of Philosophy of Knowledge, 43(1–2):99–124, June 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ladriere:1989:PRP

[1241] Jean Ladriére. Physical reality. a phenomenological approach. Dialectica: International Review of Philosophy of Knowledge, 43(1–2):125–139, June 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kantscheider:1989:RBP

[1242] Bernulf Kantscheider. Realism from a biological point of view. Dialectica: International Review of Philosophy of Knowledge, 43(1–2):141–156, June 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

DEspagnat:1989:RPF

[1243] Bernard D’Espagnat. Réalité et physique. (French) [Reality and physics]. Dialectica: International Review of Philosophy of Knowledge, 43(1–2): 157–172, June 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

McDowell:1989:MPD

[1244] John McDowell. Mathematical Platonism and Dummettian anti-realism. Dialectica: International Review of Philosophy of Knowledge, 43(1–2): 173–192, June 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gumanski:1989:NAR

[1245] Leon Gumanski. A new approach to realistic epistemology. Dialectica: International Review of Philosophy of Knowledge, 43(1–2):193–210, June 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lepage:1989:KT

[1246] François Lepage. Knowledge and truth. Dialectica: International Review of Philosophy of Knowledge, 43(3):215–229, September 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Steinhoff:1989:PEO

[1247] Gordon Steinhoff. Putnam on “Empirical objects”. Dialectica: International Review of Philosophy of Knowledge, 43(3):231–248, September 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Daniel:1989:BER

[1248] Wojciech Daniel. Bohr, Einstein and realism. Dialectica: International Review of Philosophy of Knowledge, 43(3):249–261, September 1989. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Brenner-Golomb:1989:NDP

[1249] N. Brenner-Golomb and J. van Brakel. Notes et discussions: Putnam on Davidson on conceptual schemes. Dialectica: International Review of Philosophy of Knowledge, 43(3):263–269, September 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ward:1989:RIG

[1250] Andrew Ward. Radical interpretation and the Gunderson Game. Dialectica: International Review of Philosophy of Knowledge, 43(3):271–280, September 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Widerker:1989:DDI

[1251] David Widerker. In defense of Davidson’s identity thesis regarding action individuation. Dialectica: International Review of Philosophy of Knowledge, 43(3):281–288, September 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hugly:1989:CTP

[1252] Philip Hugly and Charles Sayward. Can there be a proof that some unprovable arithmetic sentence is true? Dialectica: International Review of Philosophy of Knowledge, 43(3):289–292, September 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1989:RSEa

[1253] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 43(3):293–294, September 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wreen:1989:LDI

[1254] Michael Wreen. Light from darkness, from ignorance knowledge. Dialectica: International Review of Philosophy of Knowledge, 43(4):299–314, December 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Khlentzos:1989:ARU

[1255] Drew Khlentzos. Anti-realism under mind? Dialectica: International Review of Philosophy of Knowledge, 43(4):315–328, December 1989. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gauthier:1989:FAI

[1256] Yvon Gauthier. Finite arithmetic with infinite descent. Dialectica: International Review of Philosophy of Knowledge, 43(4):329–337, December 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Loeck:1989:DLM

[1257] Gisela Loeck. Descartes’ logic of magnitudes. Dialectica: International Review of Philosophy of Knowledge, 43(4):339–372, December 1989. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Thibaud:1989:NPI

[1258] P. Thibaud. Nom propre et individuation chez Peirce. (French) [Proper name and individuation in Peirce]. Dialectica: International Review of Philosophy of Knowledge, 43(4):373–386, December 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rogerson:1989:NDC

[1259] Kenneth F. Rogerson. Notes et discussions: Causal hermits. Dialectica: International Review of Philosophy of Knowledge, 43(4):387–396, December 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1989:RSEb

[1260] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 43(4):397–398, December 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1989:RBL

[1261] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 43(4):399, December 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1989:CVT

[1262] Anonymous. Contents of vol. 43, 1989 — table annuelle, vol. 43 — Inhalt Band 43. Dialectica: International Review of Philosophy of Knowledge, 43(4):401–402, December 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1989:AAC

[1263] Anonymous. Advices to authors conseils aux auteurs. Dialectica: International Review of Philosophy of Knowledge, 43(4):403–404, December 1989. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lauener:1990:VHG

[1264] Henri Lauener. Vorwort des Herausgebers. (German) [Editors’ foreword]. Dialectica: International Review of Philosophy of Knowledge, 44(1–2): 3–4, June 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marcus:1990:MB

[1265] Ruth Barcan Marcus. Max Black (1909–1988). Dialectica: International Review of Philosophy of Knowledge, 44(1–2):5–8, June 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Granger:1990:VMF

[1266] G. G. Granger and André Weil. Sur le vague en mathématiques. (French) [On the wave in mathematics]. Dialectica: International Review of Philosophy of Knowledge, 44(1–2):9–22, June 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sanford:1990:ISI

[1267] David H. Sanford. The inductive support of inductive rules: Themes from max black. Dialectica: International Review of Philosophy of Knowledge, 44(1–2):23–41, June 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ziff:1990:TP

[1268] Paul Ziff. Time preference. Dialectica: International Review of Philosophy of Knowledge, 44(1–2):43–54, June 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hintikka:1990:MVL

[1269] Jaakko Hintikka and Gabriel Sandu. Metaphor and the varieties of lexical meaning: Dedicated to the memory of Max Black. Dialectica: International Review of Philosophy of Knowledge, 44(1–2):55–78, June 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Martinich:1990:MIB

[1270] A. P. Martinich. Meaning and intention: Black versus Grice. Dialectica: International Review of Philosophy of Knowledge, 44(1–2):79–98, June 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Donnellan:1990:GNK

[1271] Keith S. Donnellan. Genuine names and knowledge by acquaintance. Dialectica: International Review of Philosophy of Knowledge, 44(1–2): 99–112, June 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wettstein:1990:FRS

[1272] Howard Wettstein. Frege-Russell semantics? Dialectica: International Review of Philosophy of Knowledge, 44(1–2):113–135, June 1990. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lambert:1990:RTD

[1273] Karel Lambert. Russell’s theory of definite descriptions. Dialectica: International Review of Philosophy of Knowledge, 44(1–2):137–152, June 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stroll:1990:MM

[1274] Avrum Stroll. Max on Moore. Dialectica: International Review of Philosophy of Knowledge, 44(1–2):153–163, June 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pears:1990:WH

[1275] David Pears. Wittgenstein’s holism. Dialectica: International Review of Philosophy of Knowledge, 44(1–2):165–173, June 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Garver:1990:FLW

[1276] Newton Garver. Form of life in Wittgenstein’s later work. Dialectica: International Review of Philosophy of Knowledge, 44(1–2):175–201, June 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sachs:1990:PRF

[1277] David Sachs. A philosophical resistance to Freud. Dialectica: International Review of Philosophy of Knowledge, 44(1–2):204–214, June 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1990:E

[1278] P. E. Pilet. Editorial. Dialectica: International Review of Philosophy of Knowledge, 44(3–4):i, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vuillemin:1990:MFG

[1279] Jules Vuillemin. Sur la méthode de Ferdinand Gonseth. (French) [On the method of Ferdinand Gonseth]. Dialectica: International Review of Philosophy of Knowledge, 44(3–4):225–228, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ory:1990:CDM

[1280] André Ory. Aux confins du monde physique et du monde psychique: Essai sur le thème du réferentiel. (French) [On the confines of the physical world and the psychic world: Essay on the theme of the referential]. Dialectica: International Review of Philosophy of Knowledge, 44(3–4): 229–242, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bonsack:1990:VEG

[1281] François Bonsack. Une version de l’épistémologie gonséthienne: le béton et les cailloux. (French) [A version of Gonsethian epistemology: concrete and pebbles]. Dialectica: International Review of Philosophy of Knowledge, 44(3–4):243–253, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tonoiu:1990:VPD

[1282] Vasile Tonoiu. Vers une pédagogie du dialogue. (French) [Towards a pedagogy of dialogue]. Dialectica: International Review of Philosophy of Knowledge, 44(3–4):255–278, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Heinzmann:1990:LMG

[1283] Gerhard Heinzmann. L’épistémologie mathématique de Gonseth dans la perspective du pragmatisme de Peirce. (French) [The mathematical epistemology of Gonseth in the perspective of the pragmatism of Peirce]. Dialectica: International Review of Philosophy of Knowledge, 44(3–4): 279–286, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lauener:1990:POF

[1284] Henri Lauener. La philosophic ouverte de F. Gonseth aboutit-elle à une conception réaliste ou relativiste des théories scientifiques?. (French) [Does the open philosophy of F. Gonseth lead to a realistic or relativist conception of scientific theories?]. Dialectica: International Review of Philosophy of Knowledge, 44(3–4):287–293, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jorge:1990:CQF

[1285] Maria Manuel Araújo Jorge. Ce que Ferdinand Gonseth à d’important á dire à l’épistémologie contemporaine. (French) [What Ferdinand Gonseth has to say about contemporary epistemology]. Dialectica: International Review of Philosophy of Knowledge, 44(3–4):295–311, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Witkowski:1990:PME

[1286] Lech Witkowski. On the phenomenon of marginality in epistemology: Gonseth and his tradition. Dialectica: International Review of Philosophy of Knowledge, 44(3–4):313–322, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cohen-Tannoudji:1990:PDC

[1287] Gilles Cohen-Tannoudji. La pertinence du concept horizon de réalité en physique théorique contemporaine. (French) [The relevance of the horizon concept of reality in contemporary theoretical physics]. Dialectica: International Review of Philosophy of Knowledge, 44(3–4):323–332, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1990:LDV

[1288] Paul-Emile Pilet. L’étude du vivant et la méthodologie ouverte. (French) [The study of life and the open methodology]. Dialectica: International Review of Philosophy of Knowledge, 44(3–4):333–341, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Peguiron:1990:PGM

[1289] Nicolas Peguiron. La pensée gonséthienne et le mouvement systémique. (French) [Gonsethian thought and the systemic movement]. Dialectica: International Review of Philosophy of Knowledge, 44(3–4):343–351, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Morel:1990:GDT

[1290] Bernard Morel. Gonseth et le discours théologique. (French) [Gonseth and theological discourse]. Dialectica: International Review of Philosophy of Knowledge, 44(3–4):353–361, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1990:ECB

[1291] Anonymous. Études critiques — Betrachtungen zur Literatur — reviewing studies. Dialectica: International Review of Philosophy of Knowledge, 44(3–4):363–366, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1990:RBL

[1292] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 44(3–4): 367–369, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1990:CVT

[1293] Anonymous. Contents of vol. 44, 1990 — table annuelle, vol. 44 inhalt band 44. Dialectica: International Review of Philosophy of Knowledge, 44(3–4):371–372, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1990:AAC

[1294] Anonymous. Advices to authors conseils aux auteurs. Dialectica: International Review of Philosophy of Knowledge, 44(3–4):373–375, December 1990. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Herman:1991:IKT

[1295] David J. Herman. The incoherence of Kant’s transcendental dialectic: Specifying the minimal conditions for dialectical error. Dialectica: International Review of Philosophy of Knowledge, 45(1):3–29, March 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Glouberman:1991:TID

[1296] M. Glouberman. Transcendental idealism: The dialectical dimension. Dialectica: International Review of Philosophy of Knowledge, 45(1):31–45, March 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Malpas:1991:HI

[1297] Jeff Malpas. Holism and indeterminacy. Dialectica: International Review of Philosophy of Knowledge, 45(1):47–58, March 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gibson:1991:MQD

[1298] Roger F. Gibson. More on Quine’s dilemma of underdetermination. Dialectica: International Review of Philosophy of Knowledge, 45(1):59–66, March 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Green:1991:DO

[1299] Karen Green. Dummett’s ought from is. Dialectica: International Review of Philosophy of Knowledge, 45(1):67–82, March 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deQueiroz:1991:NDM

[1300] Ruy J. G. B. de Queiroz. Notes et discussions: Meaning as grammar plus consequences. (French) [Notes and discussions: Meaning as grammar plus consequences]. Dialectica: International Review of Philosophy of Knowledge, 45(1):83–86, March 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1991:ECB

[1301] Anonymous. Études critiques — Betrachtungen zur Literatur — reviewing studies. Dialectica: International Review of Philosophy of Knowledge, 45(1):87–89, March 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Brunner:1991:MRC

[1302] Fernand Brunner. Mysticisme et rationalité chez maître Eckhart. (French) [Mysticism and rationality in master Eckhart]. Dialectica: International Review of Philosophy of Knowledge, 45(2–3):99–115, September 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rod:1991:KRN

[1303] Wolfgang Röd. Kants Reine Naturwissenschaft als kritische Metaphysik. (German) [Kant’s natural science as critical metaphysics]. Dialectica: International Review of Philosophy of Knowledge, 45(2–3):117–131, September 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Essler:1991:KKE

[1304] Wilhelm K. Essler. Kant und kein Ende. (German) [Kant and no end]. Dialectica: International Review of Philosophy of Knowledge, 45(2–3): 133–163, September 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stroll:1991:OH

[1305] Avrum Stroll. Observation and the hidden. Dialectica: International Review of Philosophy of Knowledge, 45(2–3):165–179, September 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schulthess:1991:EPS

[1306] Daniel Schulthess. Eléments et parties des systémes: Note sur l’interprétation temporelle des deux relations. (French) [Elements and parts of systems: Note on the temporal interpretation of the two relations]. Dialectica: International Review of Philosophy of Knowledge, 45 (2–3):181–189, September 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Davidson:1991:EE

[1307] Donald Davidson. Epistemology externalized. Dialectica: International Review of Philosophy of Knowledge, 45(2–3):191–202, September 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hintikka:1991:OOM

[1308] Jaakko Hintikka. Overcoming “Overcoming metaphysics through logical analysis of language” through logical analysis of language. Dialectica: International Review of Philosophy of Knowledge, 45(2–3):203–218, September 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Quine:1991:IV

[1309] W. V. Quine. Immanence and validity. Dialectica: International Review of Philosophy of Knowledge, 45(2–3):219–230, September 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vuillemin:1991:ESA

[1310] Jules Vuillemin. Etre sensible, acte, puissance. Dialectica: International Review of Philosophy of Knowledge, 45(2–3):231–248, September 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Giannoli:1991:MS

[1311] Giovanni I. Giannoli and Alberto Gianquinto. On meaningful signals. Dialectica: International Review of Philosophy of Knowledge, 45(4):255– 271, December 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nowak:1991:TFP

[1312] Leszek Nowak. Thoughts are facts in possible worlds, truths are facts of a given world. Dialectica: International Review of Philosophy of Knowledge, 45(4):273–288, December 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Duran:1991:BN

[1313] Jane Duran. Boydian nightmares. Dialectica: International Review of Philosophy of Knowledge, 45(4):289–299, December 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hermann:1991:PEP

[1314] P. Hermann. Pour une épistémologie psychanalytique. Sur trois mots de Sigmund Freud. (French) [For a psychoanalytic epistemology. On three words from Sigmund Freud]. Dialectica: International Review of Philosophy of Knowledge, 45(4):301–315, December 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1991:RSE

[1315] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 45(4):317–327, December 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1991:RBL

[1316] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 45(4):328– 329, December 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1991:CVT

[1317] Anonymous. Contents of vol. 45, 1991 –table annuelle, vol. 45 inhalt band 45. Dialectica: International Review of Philosophy of Knowledge, 45(4): 330, December 1991. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Moser:1992:AE

[1318] Paul K. Moser. Analyticity and epistemology. Dialectica: International Review of Philosophy of Knowledge, 46(1):3–19, March 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Seymour:1992:EBC

[1319] Michel Seymour. Les expériences de Burge et les contenus de pensée. (French) [The experiences of Burge and the contents of thought]. Dialectica: International Review of Philosophy of Knowledge, 46(1):21–39, March 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Callaway:1992:MHS

[1320] H. G. Callaway. Meaning holism and semantic realism. Dialectica: International Review of Philosophy of Knowledge, 46(1):41–59, March 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wiggins:1992:MTC

[1321] David Wiggins. Meaning, truth-conditions, proposition: Frege’s doctrine of sense retrieved, resumed and redeployed in the light of certain recent criticisms. Dialectica: International Review of Philosophy of Knowledge, 46(1):61–90, March 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schwartz:1992:JRF

[1322] Elisabeth Schwartz. Le jugement de recognition fregéen et la supposition de determination complète. (French) [The judgment of Frege recognition and the assumption of complete determination]. Dialectica: International Review of Philosophy of Knowledge, 46(1):91–114, March 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1992:RSEa

[1323] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 46(1):115–117, March 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gottlob:1992:HSC

[1324] Rainer Gottlob. How scientists confirm universal propositions (laws of nature): The minor significance of “Positive instances”. Dialectica: International Review of Philosophy of Knowledge, 46(2):123–139, June 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Majer:1992:WKD

[1325] Ulrich Majer. Weyls Kritik an Dedekinds Zahlbegriff. (German) [Weyl’s criticism of Dedekind’s number concept]. Dialectica: International Review of Philosophy of Knowledge, 46(2):141–149, June 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ibarra:1992:LTQ

[1326] Andoni Ibarra and Thomas Mormann. L’explication en tant que généralisation théorique. (French) [The explanation as theoretical generalization]. Dialectica: International Review of Philosophy of Knowledge, 46(2):151–168, June 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Haaparanta:1992:ATT

[1327] Leila Haaparanta. The analogy theory of thinking. Dialectica: International Review of Philosophy of Knowledge, 46(2):169–183, June 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1992:RSEb

[1328] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 46(2):184–185, June 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lorenz:1992:SPD

[1329] Kuno Lorenz. Die Selbstorganisation der Philosophie durch das dialogische Prinzip. (German) [The self-organization of philosophy by the dialogic principle]. Dialectica: International Review of Philosophy of Knowledge, 46(3–4):191–199, September 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dilworth:1992:LTS

[1330] Craig Dilworth. The linguistic turn: Shortcut or detour? Dialectica: International Review of Philosophy of Knowledge, 46(3–4):201–214, September 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Burri:1992:LWR

[1331] Alex Burri. The linguistic winding road. Dialectica: International Review of Philosophy of Knowledge, 46(3–4):215–224, September 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Callaway:1992:DLD

[1332] H. G. Callaway. Does language determine our scientific ideas? Dialectica: International Review of Philosophy of Knowledge, 46(3–4):225–242, September 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vanBrakel:1992:NKM

[1333] J. van Brakel. Natural kinds and manifest forms of life. Dialectica: International Review of Philosophy of Knowledge, 46(3–4):243–261, September 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Simons:1992:WTS

[1334] Perer Simons and Ernst Mayr. Was trägt die Sprachanalyse zur Philosophie der Biologie bei — und Umgekehrt?. (German) [What does linguistics contribute to the philosophy of biology — and vice versa?]. Dialectica: International Review of Philosophy of Knowledge, 46(3–4):263–280, September 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Shapere:1992:TTA

[1335] Dudley Shapere. Talking and thinking about nature roots, evolution, and future prospects. Dialectica: International Review of Philosophy of Knowledge, 46(3–4):281–296, September 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Blau:1992:CNG

[1336] U. Blau. Cantor und Nigarjuna. (German) [Cantor and Nigarjuna]. Dialectica: International Review of Philosophy of Knowledge, 46(3–4):297– 311, September 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1992:RSEc

[1337] Anonymous. Reviewing studies — Études critiques — Betrachtungen zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 46(3–4):313–315, September 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1992:RBL

[1338] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 46(3–4): 317, September 1992. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zibakalam:1993:ERS

[1339] Saeid Zibakalam. Emergence of a radical sociology of scientific knowledge: The strong programme in the early writings of Barry Barnes. Dialectica: International Review of Philosophy of Knowledge, 47(1):3–25, March 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Havas:1993:DWT

[1340] Katalin G. Havas. Do we tolerate inconsistencies? Dialectica: International Review of Philosophy of Knowledge, 47(1):27–35, March 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Slater:1993:FO

[1341] B. H. Slater. Farewell to opacity. Dialectica: International Review of Philosophy of Knowledge, 47(1):37–53, March 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chokr:1993:NDN

[1342] Nader N. Chokr. Notes et discussions: Nelson Goodman on truth, Relativism, and criteria of rightness or why we should dispense with truth and adopt rightness?. (French) [Notes and discussions: Nelson Goodman on truth, Relativism, and criteria of rightness or why we should dispense with truth and adopt rightness?]. Dialectica: International Review of Philosophy of Knowledge, 47(1):55–73, March 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1993:RSE

[1343] Anonymous. Reviewing studies — Études critiques — Betrachtung zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 47(1):75–79, March 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schulthess:1993:AP

[1344] Daniel Schulthess. Avant-propos. Dialectica: International Review of Philosophy of Knowledge, 47(2–3):83–84, June 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lauener:1993:FB

[1345] Henri Lauener. Fernand Brunner (1920–1991). Dialectica: International Review of Philosophy of Knowledge, 47(2–3):85–87, June 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rousseau:1993:FBP

[1346] Nicolas Rousseau. Fernand Brunner, professeur. Dialectica: International Review of Philosophy of Knowledge, 47(2–3):89–96, June 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Brunner:1993:TRF

[1347] Fernand Brunner. Travaux représentatifs. (French) [Representative works]. Dialectica: International Review of Philosophy of Knowledge, 47(2–3):97–98, June 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ladriere:1993:MMF

[1348] Jean Ladrière. Métaphysique et mystique. (French) [Metaphysics and mystical]. Dialectica: International Review of Philosophy of Knowledge, 47(2–3):99–121, June 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deMuralt:1993:CAS

[1349] André de Muralt. La causalité aristotélicienne et la structure de pensée scotiste. (French) [Aristotelian causality and the scotistic thought structure]. Dialectica: International Review of Philosophy of Knowledge, 47 (2–3):121–141, June 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gillet:1993:LCP

[1350] Eric Gillet and Paul Gochet. La logique de la connaissance: Le problème de l’omniscience logique. (French) [The logic of knowledge: The problem of logical omniscience]. Dialectica: International Review of Philosophy of Knowledge, 47(2–3):143–171, June 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Follesdal:1993:NIC

[1351] Dagfinn Føllesdal. La notion d’intentionalité chez Husserl. (French) [The notion of intentionality in Husserl]. Dialectica: International Review of Philosophy of Knowledge, 47(2–3):173–187, June 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lauener:1993:BDG

[1352] Henri Lauener. Bemerkungen zur Deutung von Gödels Unentscheidbarkeitstheorem. (German) [Remarks on the interpretation of Gødel’s Incompleteness Theorem]. Dialectica: International Review of Philosophy of Knowledge, 47(2–3):189–197, June 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Boss:1993:SPP

[1353] Gilbert Boss. Le songe d’une poétique philosophique (les rêves de Descartes). (French) [The dream of a philosophical poetic (Descartes’ dreams)]. Dialectica: International Review of Philosophy of Knowledge, 47(2–3):199–215, June 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pena:1993:CDP

[1354] Lorenzo Peña. Le choix de Dieu et le principe du meilleur. (French) [The choice of God and the principle of the best]. Dialectica: International Review of Philosophy of Knowledge, 47(2–3):217–254, June 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schulthess:1993:PEC

[1355] Daniel Schulthess. Psychologie et épistémologie de la croyance (belief, assent) selon Hume. (French) [Psychology and epistemology of belief (belief, assent) according to Hume]. Dialectica: International Review of Philosophy of Knowledge, 47(2–3):255–267, June 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rivier:1993:LTP

[1356] Dominique Rivier. L’espace et le temps de la physique contemporaine. (French) [The space and time of contemporary physics]. Dialectica: International Review of Philosophy of Knowledge, 47(2–3):269–285, June 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stachowiak:1993:PKP

[1357] Herbert Stachowiak. Zur pragmatik kognitiver prozesse. Dialectica: International Review of Philosophy of Knowledge, 47(4):291–304, December 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Szamosi:1993:NCI

[1358] G. Szamosi. Naturalizing the Copenhagen interpretation: (an evolutionary approach to quantum-measurement theory.). Dialectica: International Review of Philosophy of Knowledge, 47(4):305–325, December 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Duran:1993:NDN

[1359] Jane Duran and Ruth Doell. Notes et discussions: Naturalized epistemology, connectionism and biology. Dialectica: International Review of Philosophy of Knowledge, 47(4):327–336, December 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1993:RBL

[1360] Anonymous. Received books — livres reçus — Einegange Bücher. Dialectica: International Review of Philosophy of Knowledge, 47(4):337– 338, December 1993. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Liston:1994:HAO

[1361] Michael Liston. How abstract objects strike us. Dialectica: International Review of Philosophy of Knowledge, 48(1):3–27, March 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hernandez-Iglesias:1994:ID

[1362] Manuel Hernández-Iglesias. Incommensurability without dogmas. Dialectica: International Review of Philosophy of Knowledge, 48(1):29–45, March 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nielsen:1994:BCC

[1363] Kai Nielsen. Broad’s conception of critical and speculative philosophy. Dialectica: International Review of Philosophy of Knowledge, 48(1):47– 64, March 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vahid:1994:NDO

[1364] Hamid Vahid. Notes et discussions: Origin, subsequent history and necessity. Dialectica: International Review of Philosophy of Knowledge, 48 (1):65–72, March 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1994:RSEa

[1365] Anonymous. Reviewing studies — Études critiques — Betrachtung zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 48(1):73–78, March 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deQueiroz:1994:NLG

[1366] Ruy J. G. B. de Queiroz. Normalisation and language-games. Dialectica: International Review of Philosophy of Knowledge, 48(2):83–123, June 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cortois:1994:QAD

[1367] Paul Cortois. Quelques aspects du programme épistémologique de Cavaillés. (French) [Some aspects of the Cavaillés epistemological program]. Dialectica: International Review of Philosophy of Knowledge, 48 (2):125–141, June 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Emery-Hellwig:1994:POF

[1368] Eric Emery-Hellwig. Philosophie ouverte de F. Gonseth et philosophie analytique. (French) [Open philosophy of F. Gonseth and analytical philosophy]. Dialectica: International Review of Philosophy of Knowledge, 48(2):143–155, June 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Thom:1994:NDR

[1369] René Thom. Notes et discussions: Le Référential selon Ferdinand Gonseth, et ses aspects philosophiques. (French) [Notes and discussions: Ferdinand Gonseth’s Referentia, and its philosophical aspects]. Dialectica: International Review of Philosophy of Knowledge, 48(2):157–168, June 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1994:RSEb

[1370] Anonymous. Reviewing studies — Études critiques — Betrachtung zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 48(2):169–170, June 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Glauser:1994:PFP

[1371] Richard Glauser. Présentation. (French) [Presentation]. Dialectica: International Review of Philosophy of Knowledge, 48(3–4):177–183, September 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bruns:1994:ODP

[1372] Manfred Bruns and Gianfranco Soldati. Object-dependent and property dependent contents. Dialectica: International Review of Philosophy of Knowledge, 48(3–4):185–208, September 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Recanati:1994:HNN

[1373] François Recanati. How narrow is narrow content? Dialectica: International Review of Philosophy of Knowledge, 48(3–4):209–229, September 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cuypers:1994:LNP

[1374] Stefaan E. Cuypers. L’impossible naturalisme de la psychosémantique de Fodor. (French) [The impossible naturalism of the psycho-semantics of Fodor]. Dialectica: International Review of Philosophy of Knowledge, 48 (3–4):231–248, September 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Proust:1994:DDE

[1375] Joëlle Proust. Descripteurs distaux et externalisme. (French) [Distal descriptors and externalism]. Dialectica: International Review of Philosophy of Knowledge, 48(3–4):249–265, September 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Buekens:1994:ECC

[1376] Filip Buekens. Externalism, content, and causal histories. Dialectica: International Review of Philosophy of Knowledge, 48(3–4):267–286, September 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Guttenplan:1994:BKO

[1377] Samuel Guttenplan. Belief, knowledge and the origins of content. Dialectica: International Review of Philosophy of Knowledge, 48(3–4):287–305, September 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Corazza:1994:PTP

[1378] Eros Corazza. Perspectival thoughts and psychological generalizations. Dialectica: International Review of Philosophy of Knowledge, 48(3–4): 307–336, September 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Burri:1994:CSD

[1379] Alex Burri. Le caractère social du langage et de la pensée. (French) [The social character of language and thought]. Dialectica: International Review of Philosophy of Knowledge, 48(3–4):337–352, September 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1994:RBL

[1380] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 48(3–4): 353, September 1994. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Douven:1995:SRE

[1381] Igor Douven and Jaap van Brakel. Is scientific realism an empirical hypothesis? Dialectica: International Review of Philosophy of Knowledge, 49(1):3–14, March 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Psilos:1995:SRB

[1382] Stathis Psilos. Is structural realism the best of both worlds? Dialectica: International Review of Philosophy of Knowledge, 49(1):15–46, March 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Combourieu:1995:CAD

[1383] Marie-Christine Combourieu. Controverse autour de la définition de la réalité physique. Le paradoxe d’Einstein–Podolsky–Rosen (1935) et la non-séparabilité quantique. (French) [Controversy around the definition of physical reality. The paradox of Einstein–Podolsky–Rosen (1935) and the quantum non-separability]. Dialectica: International Review of Philosophy of Knowledge, 49(1):47–74, March 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Davidson:1995:IRT

[1384] Donald Davidson. Interview: Relations and transitions — an interview with. Dialectica: International Review of Philosophy of Knowledge, 49 (1):75–86, March 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Furrer:1995:V

[1385] Stephan Furrer and Stepha Hottinger. Vorwort. Dialectica: International Review of Philosophy of Knowledge, 49(2–4):93–94, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stroll:1995:TNO

[1386] Arum Stroll. Transcendentalism or naturalism: The only alternatives? Dialectica: International Review of Philosophy of Knowledge, 49(2–4): 95–112, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marcus:1995:ANS

[1387] Ruth Barcan Marcus. The anti-naturalism of some language centered accounts of belief. Dialectica: International Review of Philosophy of Knowledge, 49(2–4):113–130, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Simons:1995:MO

[1388] Peter Simons. Mind and opacity. Dialectica: International Review of Philosophy of Knowledge, 49(2–4):131–146, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Callaway:1995:INC

[1389] H. G. Callaway. Intentionality naturalized: Continuity, reconstruction, and instrumentalism. Dialectica: International Review of Philosophy of Knowledge, 49(2–4):147–168, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bau:1995:EPT

[1390] Jong Bau and Ulrich Blau. Epistemic Paradoxe, Teil 1. (German) [Epistemic paradoxes, part 1]. Dialectica: International Review of Philosophy of Knowledge, 49(2–4):169–194, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lambert:1995:GTP

[1391] Karel Lambert. General terms, predicates and extensionality. Dialectica: International Review of Philosophy of Knowledge, 49(2–4):195–202, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1995:TTM

[1392] Benito Müller. Toward a transcendental model-theoretic semantics for scientific languages. Dialectica: International Review of Philosophy of Knowledge, 49(2–4):203–228, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hintikka:1995:GLG

[1393] Jaakko Hintikka. The games of logic and the games of inquiry. Dialectica: International Review of Philosophy of Knowledge, 49(2–4):229–250, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Quine:1995:NLW

[1394] W. V. Quine. Naturalism; or, living within One’s means. Dialectica: International Review of Philosophy of Knowledge, 49(2–4):251–263, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Davidson:1995:LC

[1395] Donald Davidson. Laws and cause. Dialectica: International Review of Philosophy of Knowledge, 49(2–4):263–280, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Essler:1995:WWE

[1396] Wilhelm K. Essler. Was ist und zu welchem Ende treibt man Metaphysik?. (German) [What is and what is the end of metaphysics?]. Dialectica: International Review of Philosophy of Knowledge, 49(2–4):281–317, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rod:1995:KHT

[1397] Wolfgang Röd. Kant und Hume: Die Transzendentalphilosophie als Alternative zum Naturalismus. (German) [Kant and Hume: Transcendental philosophy as an alternative to naturalism]. Dialectica: International Review of Philosophy of Knowledge, 49(2–4):317–334, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Allison:1995:NKT

[1398] Henry E. Allison. On naturalizing Kant’s transcendental psychology. Dialectica: International Review of Philosophy of Knowledge, 49(2–4): 335–356, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1995:PHL

[1399] Anonymous. Publikationen von Henri Lauener. (German) [Publications by Henri Lauener]. Dialectica: International Review of Philosophy of Knowledge, 49(2–4):353–356, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1995:RBL

[1400] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 49(2–4): 357, June 1995. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Plumer:1996:TCT

[1401] Gilbert Plumer. Truth and collective truth. Dialectica: International Review of Philosophy of Knowledge, 50(1):3–24, March 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Moreno:1996:UWF

[1402] Luis Fernández Moreno. Die Undefinierbarkeit der Wahrheit bei Frege. (German) [The undefinability of truth in Frege]. Dialectica: International Review of Philosophy of Knowledge, 50(1):25–36, March 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tibbetts:1996:RDC

[1403] Paul Tibbetts. Residual dualism in computational theories of mind. Dialectica: International Review of Philosophy of Knowledge, 50(1):37–52, March 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pilet:1996:LAF

[1404] Paul-Emile Pilet and Bernard Morel. L’Idonéisme, aujourd’hui. (French) [Idoneism, today]. Dialectica: International Review of Philosophy of Knowledge, 50(1):53–62, March 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1996:RSEa

[1405] Anonymous. Reviewing studies — Études critiques — Betrachtung zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 50(1):63–72, March 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1996:ENEa

[1406] Anonymous. ESAP news: ESAP news. Dialectica: International Review of Philosophy of Knowledge, 50(1):86–88, March 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Place:1996:IMD

[1407] Ullin T. Place. Intentionality as the mark of the dispositional. Dialectica: International Review of Philosophy of Knowledge, 50(2):91–120, June 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Folse:1996:OCU

[1408] Henry J. Folse. Ontological constraints and understanding quantum phenomena. Dialectica: International Review of Philosophy of Knowledge, 50(2):121–136, June 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rodriguez-Alcazar:1996:NEO

[1409] Javier Rodríguez-Alcazar. Naturalized epistemology and the is/ought gap. Dialectica: International Review of Philosophy of Knowledge, 50 (2):137–152, June 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1996:RSEb

[1410] Anonymous. Reviewing studies — Études critiques — Betrachtung zur Literatur. Dialectica: International Review of Philosophy of Knowledge, 50(2):153–161, June 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1996:ENEb

[1411] Anonymous. ESAP news: ESAP news. Dialectica: International Review of Philosophy of Knowledge, 50(2):163, June 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Blau:1996:EPT

[1412] Jong Blau/Ulrich Blau. Epistemische Paradoxien, Teil 2. (German) [Epistemic paradoxes, part 2]. Dialectica: International Review of Philosophy of Knowledge, 50(3):167–182, September 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Weber:1996:CCE

[1413] Erik Weber. Comment construit-on une explication déductivenomologique?. (French) [How do we construct a deductive–nomological explanation?]. Dialectica: International Review of Philosophy of Knowledge, 50(3):183–204, September 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Glock:1996:AU

[1414] H. J. Glock. Abusing use. Dialectica: International Review of Philosophy of Knowledge, 50(3):205–224, September 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Weintraub:1996:SL

[1415] Ruth Weintraub. The sceptical life. Dialectica: International Review of Philosophy of Knowledge, 50(3):225–234, September 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rosner:1996:QGS

[1416] Jennifer A. Rosner. Quine’s global structuralism. Dialectica: International Review of Philosophy of Knowledge, 50(3):235–242, September 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1996:ENEc

[1417] Anonymous. ESAP news: ESAP news. Dialectica: International Review of Philosophy of Knowledge, 50(3):243, September 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Landsberg:1996:ITA

[1418] Peter T. Landsberg. Irreversibility and Time’s arrow. Dialectica: International Review of Philosophy of Knowledge, 50(4):247–258, December 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Formi:1996:IRM

[1419] Gábor Formi. Internal realism, metaphysical realism, and brains in a vat. Dialectica: International Review of Philosophy of Knowledge, 50(4):259– 274, December 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pihlstrom:1996:SRW

[1420] Sami Pihlström. A solipsist in a real world. Dialectica: International Review of Philosophy of Knowledge, 50(4):275–290, December 1996. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1996:REC

[1421] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 50(4):291– 304, December 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1996:ENEd

[1422] Anonymous. ESAP news: ESAP news. Dialectica: International Review of Philosophy of Knowledge, 50(4):305–314, December 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1996:RBL

[1423] Anonymous. Received books — livres reçus — Eingegangene Bücher. Dialectica: International Review of Philosophy of Knowledge, 50(4):315, December 1996. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1997:I

[1424] Benito Müller. Introduction. Dialectica: International Review of Philosophy of Knowledge, 51(1):5–15, March 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Harre:1997:TBO

[1425] Rom Harré. Is there a basic ontology for the physical sciences? Dialectica: International Review of Philosophy of Knowledge, 51(1):17–34, March 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Aronson:1997:DFF

[1426] Jerrold L. Aronson. Dispositions as the foundation for Feynman’s formulation of quantum mechanics. Dialectica: International Review of Philosophy of Knowledge, 51(1):35–64, March 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cartwright:1997:WDL

[1427] Nancy Cartwright. Where do laws of nature come from? Dialectica: International Review of Philosophy of Knowledge, 51(1):65–78, March 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Swinburne:1997:IC

[1428] Richard Swinburne. The irreducibility of causation. Dialectica: International Review of Philosophy of Knowledge, 51(1):79–92, March 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Couture:1997:ALA

[1429] Jocelyne Couture. Analyse logique et analyticité; de Carnap à Gödel. (French) [Logical analysis and analyticity; from Carnap to Gödel]. Dialectica: International Review of Philosophy of Knowledge, 51(2):95–117, June 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Broach:1997:NRQ

[1430] Ronald J. Broach. A noncognitivist reading of Quine’s ethics. Dialectica: International Review of Philosophy of Knowledge, 51(2):119–134, June 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Virvidakis:1997:APG

[1431] Stelios Virvidakis. Analytic philosophy in Greece. Dialectica: International Review of Philosophy of Knowledge, 51(2):135–142, June 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1997:RECa

[1432] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 51(2):143–153, June 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1997:ENEa

[1433] Anonymous. ESAP news: ESAP news. Dialectica: International Review of Philosophy of Knowledge, 51(2):155–165, June 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vassallo:1997:DEF

[1434] Nicla Vassallo. On Dummett’s early Frege and analytical philosophy. Dialectica: International Review of Philosophy of Knowledge, 51(3):171– 187, September 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:1997:IO

[1435] Benito Müller. Identity and ontology. Dialectica: International Review of Philosophy of Knowledge, 51(3):189–223, September 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Weinert:1997:SSL

[1436] Friedel Weinert. On the status of social laws. Dialectica: International Review of Philosophy of Knowledge, 51(3):225–242, September 1997. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1997:ENEb

[1437] Anonymous. ESAP news: ESAP news. Dialectica: International Review of Philosophy of Knowledge, 51(3):243–249, September 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Simons:1997:HOQ

[1438] Peter Simons. Higher-order quantification and ontological commitment. Dialectica: International Review of Philosophy of Knowledge, 51(4):255– 271, December 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Giaretta:1997:ALF

[1439] Pierdaniele Giaretta. Analysis and logical form in Russell: The 1913 paradigm. Dialectica: International Review of Philosophy of Knowledge, 51(4):273–293, December 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1997:RECb

[1440] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 51(4):295– 312, December 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1997:ENEc

[1441] Anonymous. ESAP news: ESAP news. Dialectica: International Review of Philosophy of Knowledge, 51(4):313–319, December 1997. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bordes:1998:APF

[1442] Montse Bordes. Abstract particulars in a four-dimensional ontological frame. Dialectica: International Review of Philosophy of Knowledge, 52 (1):3–12, March 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Maurin:1998:DP

[1443] Anna-Sofia Maurin. Davidson on properties. Dialectica: International Review of Philosophy of Knowledge, 52(1):13–22, March 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Campbell:1998:AMC

[1444] Neil Campbell. Anomalous monism and the charge of epiphenomenalism. Dialectica: International Review of Philosophy of Knowledge, 52(1): 23–39, March 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Berg:1998:FPA

[1445] Jonathan Berg. First person authority, externalism, and wh-knowledge. Dialectica: International Review of Philosophy of Knowledge, 52(1):41– 44, March 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hum:1998:DIC

[1446] Daniel D. Hum. Davidson’s identity crisis. Dialectica: International Review of Philosophy of Knowledge, 52(1):45–61, March 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1998:RECa

[1447] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 52(1):63–73, March 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1998:ENEa

[1448] Anonymous. ESAP news: ESAP news. Dialectica: International Review of Philosophy of Knowledge, 52(1):74–80, March 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mumford:1998:LNO

[1449] Stephen Mumford. Laws of nature outlawed. Dialectica: International Review of Philosophy of Knowledge, 52(2):83–101, June 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Atkinson:1998:LQM

[1450] D. Atkinson. The light of quantum mechanics. Dialectica: International Review of Philosophy of Knowledge, 52(2):103–126, June 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Douven:1998:TES

[1451] Igor Douven. Truly empiricist semantics. Dialectica: International Review of Philosophy of Knowledge, 52(2):127–150, June 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lambert:1998:NDF

[1452] Karel Lambert. Notes et discussions: Fixing Quine’s theory of predication. Dialectica: International Review of Philosophy of Knowledge, 52(2): 153–160, June 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1998:RECb

[1453] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 52(2):161–164, June 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1998:ENEb

[1454] Anonymous. ESAP news: ESAP news. Dialectica: International Review of Philosophy of Knowledge, 52(2):165–166, June 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Blau:1998:PAC

[1455] U. Blau. Ein platonistisches Argument für Cantors Kontinuumshypothese. (German) [A platonic argument for Cantor’s Continuum Hypothesis]. Dialectica: International Review of Philosophy of Knowledge, 52(3):175–202, September 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

MacBride:1998:WPU

[1456] Fraser MacBride. Where are particulars and universals? Dialectica: International Review of Philosophy of Knowledge, 52(3):203–227, September 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vahid:1998:IEC

[1457] Hamid Vahid. The Internalism/Externalism controversy: The epistemization of an older debate. Dialectica: International Review of Philosophy of Knowledge, 52(3):229–246, September 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rothbart:1998:EPE

[1458] Daniel Rothbart. Extending Popper’s epistemology to the lab. Dialectica: International Review of Philosophy of Knowledge, 52(3):247–254, September 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1998:RECc

[1459] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 52(3):255– 261, September 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1998:ENEc

[1460] Anonymous. ESAP news: ESAP news. Dialectica: International Review of Philosophy of Knowledge, 52(3):263–269, September 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Clark:1998:ASQ

[1461] Chalmers C. Clark. The art of science: Quine and the speculative reach of philosophy in natural science. Dialectica: International Review of Philosophy of Knowledge, 52(4):275–290, December 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wilburn:1998:SOA

[1462] Ron Wilburn. Skepticism, objectivity and the aspirations of immanence. Dialectica: International Review of Philosophy of Knowledge, 52(4):291– 318, December 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Shockley:1998:QED

[1463] Kenneth Eric Shockley. Quine’s ethical dilemma. Dialectica: International Review of Philosophy of Knowledge, 52(4):319–338, December 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Casalegno:1998:RLC

[1464] Paolo Casalegno. The referential and the logical component in Fodor’s semantics. Dialectica: International Review of Philosophy of Knowledge, 52(4):339–363, December 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1998:ENEd

[1465] Anonymous. ESAP news: ESAP news. Dialectica: International Review of Philosophy of Knowledge, 52(4):365–372, December 1998. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vanBrakel:1999:SAM

[1466] J. van Brakel. Supervenience and anomalous monism. Dialectica: International Review of Philosophy of Knowledge, 53(1):3–24, March 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kistler:1999:CEF

[1467] Max Kistler. Causes as events and facts. Dialectica: International Review of Philosophy of Knowledge, 53(1):25–46, March 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Adams:1999:ODT

[1468] Max F. Adams, R. Stecker, and G. Fuller. Object dependent thoughts, perspectival thoughts, and psychological generalization. Dialectica: International Review of Philosophy of Knowledge, 53(1):47–59, March 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marton:1999:SVD

[1469] Peter Marton. Skeptics versus dogmatics: The battle over the criterion. Dialectica: International Review of Philosophy of Knowledge, 53 (1):61–71, March 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Barnes:1999:REC

[1470] J. Barnes. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 53(1):73–74, March 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1999:ENEa

[1471] Anonymous. ESAP news: ESAP news. Dialectica: International Review of Philosophy of Knowledge, 53(1):75–81, March 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Roth:1999:EEN

[1472] Paul Roth. The epistemology of “Epistemology naturalized”. Dialectica: International Review of Philosophy of Knowledge, 53(2):87–110, June 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Thalos:1999:TDN

[1473] Mariam Thalos. Two dogmas of naturalized epistemology. Dialectica: International Review of Philosophy of Knowledge, 53(2):111–138, June 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Weberman:1999:CCR

[1474] David Weberman. Cambridge changes revisited: Why certain relational changes are indispensable. Dialectica: International Review of Philosophy of Knowledge, 53(2):139–149, June 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1999:REC

[1475] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 53(2):151–159, June 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:1999:ENEb

[1476] Anonymous. ESAP news: ESAP news. Dialectica: International Review of Philosophy of Knowledge, 53(2):160–166, June 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lauener:1999:E

[1477] Henri Lauener. Einleitung. Dialectica: International Review of Philosophy of Knowledge, 53(3–4):173–177, September 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Smith:1999:MAT

[1478] Quentin Smith. Marcus’ 1961 article as a termination of Russellian dominance in analytic philosophy. Dialectica: International Review of Philosophy of Knowledge, 53(3–4):179–210, September 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Engel:1999:DBA

[1479] Pascal Engel. Dispositional belief, assent, and acceptance. Dialectica: International Review of Philosophy of Knowledge, 53(3–4):211–226, September 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sosa:1999:EP

[1480] Ernest Sosa. The essentials of persons. Dialectica: International Review of Philosophy of Knowledge, 53(3–4):227–241, September 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gochet:1999:QML

[1481] Paul Gochet and Eric Gillet. Quantified modal logic, dynamic semantics and S5. Dialectica: International Review of Philosophy of Knowledge, 53 (3–4):243–251, September 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Williamson:1999:TCB

[1482] Timothy Williamson. Truthmakers and the converse Barcan formula. Dialectica: International Review of Philosophy of Knowledge, 53(3–4): 253–270, September 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lambert:1999:LPD

[1483] Karel Lambert. Logically proper definite descriptions: An essay in honor of Ruth Marcus. Dialectica: International Review of Philosophy of Knowledge, 53(3–4):271–282, September 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hintikka:1999:ACL

[1484] Jaako Hintikka. Is the axiom of choice a logical or set-theoretical principle? Dialectica: International Review of Philosophy of Knowledge, 53 (3–4):283–290, September 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Belnap:1999:TA

[1485] Nuel Belnap. Truth by ascent. Dialectica: International Review of Philosophy of Knowledge, 53(3–4):291–306, September 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Essler:1999:MAM

[1486] Wilhelm K. Essler, Joachim Labude, and Stefanie Ucsnay. Métaphysische Anfangsgründe der Mechanik. (German) [Metaphysical beginnings of mechanics]. Dialectica: International Review of Philosophy of Knowledge, 53(3–4):307–344, September 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vuillemin:1999:IEP

[1487] J. Vuillemin. Induction et existence physique. (French) [Induction and physical existence]. Dialectica: International Review of Philosophy of Knowledge, 53(3–4):345–383, September 1999. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Humberstone:2000:WFS

[1488] Lloyd Humberstone. What Fa says about a. Dialectica: International Review of Philosophy of Knowledge, 54(1):3–28, March 2000. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Butler:2000:PSE

[1489] Keith Butler. Problems for semantic externalism and a priori refutations of skeptical arguments. Dialectica: International Review of Philosophy of Knowledge, 54(1):29–49, March 2000. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Voltolini:2000:IML

[1490] Alberto Voltolini. Is it merely loose talk?: A ‘bizarre’ solution to the opacity puzzle. Dialectica: International Review of Philosophy of Knowledge, 54(1):51–72, March 2000. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Batitsky:2000:MCL

[1491] Vadim Batitsky. Measurement in Carnap’s late philosophy of science. Dialectica: International Review of Philosophy of Knowledge, 54(2):87–108, June 2000. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Needham:2000:DQ

[1492] Paul Needham. Duhem and Quine. Dialectica: International Review of Philosophy of Knowledge, 54(2):109–132, June 2000. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

White:2000:LCP

[1493] Graham White. Lewis, causality, and possible worlds. Dialectica: International Review of Philosophy of Knowledge, 54(2):133–137, June 2000. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2000:RECa

[1494] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 54(2):139–142, June 2000. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Orilia:2000:PTP

[1495] Francesco Orilia. The property-theoretical, performative-nominalistic theory of proper names. Dialectica: International Review of Philosophy of Knowledge, 54(3):155–176, September 2000. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pagin:2000:ST

[1496] Peter Pagin. Sensation terms. Dialectica: International Review of Philosophy of Knowledge, 54(3):177–199, September 2000. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Buchanan:2000:TWD

[1497] Reid Buchanan. The tension in Wittgenstein’s diagnosis of scepticism. Dialectica: International Review of Philosophy of Knowledge, 54(3):201– 225, September 2000. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2000:RECb

[1498] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 54(3):225– 231, September 2000. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Laurier:2000:QSS

[1499] Daniel Laurier. Que sera sera. Dialectica: International Review of Philosophy of Knowledge, 54(4):247–264, December 2000. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Harre:2000:DSA

[1500] Rom Harré. Defending science from all of its enemies and some of its friends. Dialectica: International Review of Philosophy of Knowledge, 54(4):265–281, December 2000. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Varzi:2000:MC

[1501] Achille C. Varzi. Mereological commitments. Dialectica: International Review of Philosophy of Knowledge, 54(4):283–305, December 2000. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hetherington:2000:RBV

[1502] Stephen Hetherington. Re: Brains in a vat. Dialectica: International Review of Philosophy of Knowledge, 54(4):307–312, December 2000. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2000:RECc

[1503] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 54(4):313– 328, December 2000. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vidal-Rosset:2001:JVF

[1504] Joseph Vidal-Rosset and Karel Lambert. Jules Vuillemin : (15th February 1920–16th January 2001). Dialectica: International Review of Philosophy of Knowledge, 55(1):3–7, March 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Otero:2001:FAM

[1505] Manuel Ppérez Otero. A fallacy about the modal status of logic. Dialectica: International Review of Philosophy of Knowledge, 55(1):9–27, March 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Child:2001:TDR

[1506] William Child. Triangulation: Davidson, realism and natural kinds. Dialectica: International Review of Philosophy of Knowledge, 55(1):29–50, March 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Castle:2001:SVE

[1507] David G. A. Castle. A semantic view of ecological theories. Dialectica: International Review of Philosophy of Knowledge, 55(1):51–66, March 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2001:RECa

[1508] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 55(1):67–72, March 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Follesdal:2001:QPP

[1509] Dagfinn Føllesdal. Quine — the philosophers’ philosopher. Dialectica: International Review of Philosophy of Knowledge, 55(2):99–103, June 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Katzoff:2001:EVE

[1510] Charlotte Katzoff. Epistemic virtue and epistemic responsibility. Dialectica: International Review of Philosophy of Knowledge, 55(2):105–118, June 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chrudzimski:2001:TIM

[1511] Arkadiusz Chrudzimski. Die Theorie der Intentionalität Meinongs. (German) [The theory of meinong’s intentionality]. Dialectica: International Review of Philosophy of Knowledge, 55(2):119–143, June 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Feleppa:2001:QDN

[1512] Robert Feleppa. Quine, Davidson, and the naturalization of metaethics. Dialectica: International Review of Philosophy of Knowledge, 55(2):145– 166, June 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Raatikainen:2001:PLW

[1513] Panu Raatikainen. Putnam, languages and worlds. Dialectica: International Review of Philosophy of Knowledge, 55(2):167–174, June 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2001:RECb

[1514] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 55(2):175–184, June 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2001:ESAa

[1515] Anonymous. European Society for Analytic Philosophy Newsletter N^{∘} 17. Dialectica: International Review of Philosophy of Knowledge, 55(2): 185–192, June 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Soldati:2001:E

[1516] Gianfranco Soldati. Editorial. Dialectica: International Review of Philosophy of Knowledge, 55(3):195–198, September 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Esfeld:2001:GQ

[1517] Michael Esfeld. Gonseth and Quine. Dialectica: International Review of Philosophy of Knowledge, 55(3):199–219, September 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Haji:2001:LFW

[1518] Ishtiyaque Haji and Stefaan E. Cuypers. Libertarian free will and CNC manipulation. Dialectica: International Review of Philosophy of Knowledge, 55(3):221–239, September 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Green:2001:DDC

[1519] Karen Green. Davidson’s derangement: Of the conceptual priority of language. Dialectica: International Review of Philosophy of Knowledge, 55(3):239–258, September 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Burscheid:2001:ERN

[1520] Hans Joachim Burscheid and Horst Struve. Zur Entwicklung und Rechtfertigung normativer Theorien: das Beispiel der Gerechtigkeit von Glücksspielen. (German) [On the development and justification of normative theories: the example of justice in gambling]. Dialectica: International Review of Philosophy of Knowledge, 55(3):259–282, September 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2001:RECc

[1521] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 55(3):282– 288, September 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bonk:2001:HRS

[1522] Thomas Bonk. How Reichenbach solved the quantum measurement problem. Dialectica: International Review of Philosophy of Knowledge, 55(4): 291–314, December 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jackman:2001:OLC

[1523] Henry Jackman. Ordinary language, conventionalism and a priori knowledge. Dialectica: International Review of Philosophy of Knowledge, 55 (4):315–325, December 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pritchard:2001:CSP

[1524] Duncan Pritchard. Contextualism, scepticism, and the problem of epistemic descent. Dialectica: International Review of Philosophy of Knowledge, 55(4):327–349, December 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vahi:2001:KVE

[1525] Hamid Vahi. Knowledge and varieties of epistemic luck. Dialectica: International Review of Philosophy of Knowledge, 55(4):351–362, December 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Eklund:2001:SVS

[1526] Matti Eklund. Supervaluationism, vagueifiers, and semantic overdetermination. Dialectica: International Review of Philosophy of Knowledge, 55(4):363–378, December 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2001:ESAb

[1527] Anonymous. European Society for Analytic Philosophy Newsletter N^{∘} 18. Dialectica: International Review of Philosophy of Knowledge, 55(4):379– 385, December 2001. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vallicella:2002:RMV

[1528] William F. Vallicella. Relations, monism, and the vindication of Bradley’s regress. Dialectica: International Review of Philosophy of Knowledge, 56(1):3–35, March 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Okasha:2002:HSQ

[1529] Samir Okasha. How to be a selective Quinean. Dialectica: International Review of Philosophy of Knowledge, 56(1):37–47, March 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dresner:2002:WBP

[1530] Eli Dresner. Wittgenstein’s builders and Perry’s objection to sentence priority. Dialectica: International Review of Philosophy of Knowledge, 56 (1):49–63, March 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Barnes:2002:WB

[1531] Jonathan Barnes. What is a begriffsschrift? Dialectica: International Review of Philosophy of Knowledge, 56(1):65–80, March 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2002:RECa

[1532] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 56(1):81–85, March 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2002:ESAa

[1533] Anonymous. European Society for Analytic Philosophy Newsletter N^{∘} 19. Dialectica: International Review of Philosophy of Knowledge, 56(1): 89–96, March 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Patterson:2002:TAA

[1534] Douglas Patterson. Two arguments against disquotationalism. Dialectica: International Review of Philosophy of Knowledge, 56(2):99–108, June 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hauser:2002:WNA

[1535] Kai Hauser. What new axioms could not be. Dialectica: International Review of Philosophy of Knowledge, 56(2):109–124, June 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Collins:2002:VIS

[1536] John Collins. On the very idea of a science forming faculty. Dialectica: International Review of Philosophy of Knowledge, 56(2):125–151, June 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Moretti:2002:BAI

[1537] Luca Moretti. For a Bayesian account of indirect confirmation. Dialectica: International Review of Philosophy of Knowledge, 56(2):153–173, June 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2002:RECb

[1538] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 56(2):175–188, June 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2002:ESAb

[1539] Anonymous. European Society for Analytic Philosophy Newsletter N^{∘} 20. Dialectica: International Review of Philosophy of Knowledge, 56(2): 189–192, June 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Balzer:2002:SSI

[1540] Wolfgang Balzer. Searle on social institutions: A critique. Dialectica: International Review of Philosophy of Knowledge, 56(3):195–211, September 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Romerales:2002:PEI

[1541] Enrique Romerales. Precise entities but irredeemably vague concepts? Dialectica: International Review of Philosophy of Knowledge, 56(3):213– 233, September 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Iacona:2002:ER

[1542] Andrea Iacona. The expressing relation. Dialectica: International Review of Philosophy of Knowledge, 56(3):235–260, September 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Predelli:2002:HHM

[1543] Stefano Predelli. ‘Holmes’ and Holmes — a Millian analysis of names from fiction. Dialectica: International Review of Philosophy of Knowledge, 56(3):261–279, September 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2002:RECc

[1544] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 56(3):281– 287, September 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Soldati:2002:E

[1545] Gianfranco Soldati and Philipp Keller. Editorial. Dialectica: International Review of Philosophy of Knowledge, 56(4):293–294, December 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lambert:2002:HLV

[1546] Karel Lambert. Henri lauener: A very personal remembrance. Dialectica: International Review of Philosophy of Knowledge, 56(4):295–296, December 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gochet:2002:PRH

[1547] Paul Gochet. Personal reminiscence to Henri Lauener (1933–2002). Dialectica: International Review of Philosophy of Knowledge, 56(4):297– 298, December 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Geirsson:2002:FOD

[1548] Heimir Geirsson. Frege and object dependent propositions. Dialectica: International Review of Philosophy of Knowledge, 56(4):299–314, December 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Montminy:2002:NID

[1549] Martin Montminy. Normativité et irréductibilité du mental. (French) [Normativity and irreducibility of the mind]. Dialectica: International Review of Philosophy of Knowledge, 56(4):315–333, December 2002. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nagasawa:2002:EMA

[1550] Yujin Nagasawa. Externalism and the memory argument. Dialectica: International Review of Philosophy of Knowledge, 56(4):335–346, December 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Niebergall:2002:HPG

[1551] Karl-Georg Niebergall and Matthias Schirn. Hilbert’s programme and Gödel’s theorems. Dialectica: International Review of Philosophy of Knowledge, 56(4):347–370, December 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2002:RECd

[1552] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 56(4):371– 373, December 2002. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Garcia-Carpintero:2003:EPC

[1553] Manuel Garcia-Carpintero. Editorial: Philosophy and cognitive neuroscience. Dialectica: International Review of Philosophy of Knowledge, 57 (1):3–6, March 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Orilia:2003:DTS

[1554] Francesco Orilia. A description theory of singular reference. Dialectica: International Review of Philosophy of Knowledge, 57(1):7–40, March 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Garcia-Encinas:2003:PNS

[1555] M. J. García-Encinas. A posteriori necessity in singular causation and the Humean argument. Dialectica: International Review of Philosophy of Knowledge, 57(1):41–55, March 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Coliva:2003:AFG

[1556] Annalisa Coliva. The argument from the finer-grained content of colour experiences a redefinition of its role within the debate between McDowell and non-conceptual theorists. Dialectica: International Review of Philosophy of Knowledge, 57(1):57–70, March 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mcarthur:2003:PPR

[1557] Dan Mcarthur. Is Pickering’s “Pragmatic realism” viable? Dialectica: International Review of Philosophy of Knowledge, 57(1):71–88, March 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2003:RECa

[1558] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 57(1):89–94, March 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Voltolini:2003:I

[1559] Alberto Voltolini. Introduction. Dialectica: International Review of Philosophy of Knowledge, 57(2):99–102, June 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bonomi:2003:PFT

[1560] Andrea Bonomi and Sandro Zucchi. A pragmatic framework for truth in fiction. Dialectica: International Review of Philosophy of Knowledge, 57(2):103–120, June 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Corazza:2003:IFF

[1561] Eros Corazza and Mark Whitsey. Indexicals, fictions, and ficta. Dialectica: International Review of Philosophy of Knowledge, 57(2):121–136, June 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Courrle:2003:CC

[1562] Gregory Courrle. Characters and contingency. Dialectica: International Review of Philosophy of Knowledge, 57(2):137–148, June 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kroon:2003:QNE

[1563] Frederick Kroon. Quantified negative existentials. Dialectica: International Review of Philosophy of Knowledge, 57(2):149–164, June 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Leonardi:2003:NI

[1564] Paolo Leonardi. Names and illusions. Dialectica: International Review of Philosophy of Knowledge, 57(2):165–176, June 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Orenstein:2003:FPA

[1565] Alex Orenstein. Fiction, prepositional attitudes, and some truths about falsehood. Dialectica: International Review of Philosophy of Knowledge, 57(2):177–190, June 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pelletier:2003:VD

[1566] Jérôme Pelletier. “Vergil and Dido”. Dialectica: International Review of Philosophy of Knowledge, 57(2):191–203, June 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Thomasson:2003:SFC

[1567] Amie L. Thomasson. Speaking of fictional characters. Dialectica: International Review of Philosophy of Knowledge, 57(2):205–223, June 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Voltolini:2003:HFW

[1568] Alberto Voltolini. How fictional works are related to fictional entities. Dialectica: International Review of Philosophy of Knowledge, 57(2):225– 238, June 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Walton:2003:RQN

[1569] Kendall L. Walton. Restricted quantification, negative existentials, and fiction. Dialectica: International Review of Philosophy of Knowledge, 57 (2):239–242, June 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zalta:2003:RFC

[1570] Edward N. Zalta. Referring to fictional characters. Dialectica: International Review of Philosophy of Knowledge, 57(2):243–254, June 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Santambrogio:2003:EHS

[1571] Marco Santambrogio. Editorial: How should we read the classics of philosophy? Dialectica: International Review of Philosophy of Knowledge, 57(3):257–259, September 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Greimann:2003:WFJ

[1572] Dirk Greimann. What is Frege’s Julius Caesar problem? Dialectica: International Review of Philosophy of Knowledge, 57(3):261–278, September 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rodych:2003:MGN

[1573] Victor Rodych. Misunderstanding Gödel: new arguments about Wittgenstein and new remarks by Wittgenstein. Dialectica: International Review of Philosophy of Knowledge, 57(3):279–313, September 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sayward:2003:NDD

[1574] Charles Sayward. Notes and discussions: A defense of Mill on other minds. Dialectica: International Review of Philosophy of Knowledge, 57 (3):315–322, September 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Slater:2003:MCC

[1575] Matthew H. Slater. Minimalism and coincidence: Comments on Varzi. Dialectica: International Review of Philosophy of Knowledge, 57(3):323– 329, September 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dilworth:2003:RGT

[1576] John Dilworth. A refutation of Goodman’s type-token theory of notation. Dialectica: International Review of Philosophy of Knowledge, 57(3):330– 336, September 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2003:RECb

[1577] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 57(3):337– 350, September 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dokic:2003:EPF

[1578] Jérôme Dokic. Editorial: On the philosophical foundations of situated representing. Dialectica: International Review of Philosophy of Knowledge, 57(4):353–355, December 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Medina:2003:DTC

[1579] José Medina. Deflationism and the true colours of necessity in Wittgenstein’s Tractatus. Dialectica: International Review of Philosophy of Knowledge, 57(4):357–385, December 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Varzi:2003:NS

[1580] Achille C. Varzi. Naming the stages. Dialectica: International Review of Philosophy of Knowledge, 57(4):387–412, December 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nimtz:2003:NDP

[1581] Christian Nimtz and Michael Schütte. Notes and discussions: On physicalism, physical properties, and panpsychism. Dialectica: International Review of Philosophy of Knowledge, 57(4):413–422, December 2003. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Blum:2003:CCA

[1582] Alex Blum. The core of the consequence argument: Dedicated to the memory of my son Binyamin. Dialectica: International Review of Philosophy of Knowledge, 57(4):423–429, December 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

McCullagh:2003:DIR

[1583] Mark McCullagh. Do inferential roles compose? Dialectica: International Review of Philosophy of Knowledge, 57(4):431–438, December 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2003:RECc

[1584] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 57(4):439– 444, December 2003. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bodrozic:2004:I

[1585] Davor Bodrožić. Introduction. Dialectica: International Review of Philosophy of Knowledge, 58(1):3–10, March 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bealer:2004:OME

[1586] George Bealer. The origins of modal error. Dialectica: International Review of Philosophy of Knowledge, 58(1):11–42, March 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Miscevic:2004:EI

[1587] Nenad Miščević. The explainability of intuitions. Dialectica: International Review of Philosophy of Knowledge, 58(1):43–70, March 2004. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pust:2004:EKN

[1588] Joel Pust. On explaining knowledge of necessity. Dialectica: International Review of Philosophy of Knowledge, 58(1):71–87, March 2004. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Brendel:2004:IPP

[1589] Elke Brendel. Intuition pumps and the proper use of thought experiments. Dialectica: International Review of Philosophy of Knowledge, 58 (1):89–108, March 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Williamson:2004:PIS

[1590] Timothy Williamson. Philosphical ‘intuitions’ and scepticism about judgement. Dialectica: International Review of Philosophy of Knowledge, 58(1):109–153, March 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wright:2004:IEE

[1591] Crispin Wright. Intuition, entitlement and the epistemology of logical laws. Dialectica: International Review of Philosophy of Knowledge, 58 (1):155–175, March 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Brink:2004:JAT

[1592] Ingar Brink. Joint attention, triangulation and radical interpretation: A problem and its solution. Dialectica: International Review of Philosophy of Knowledge, 58(2):179–206, June 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tappolet:2004:NDT

[1593] Christine Tappolet. Notes and discussions: Through thick and thin: good and its determinates. Dialectica: International Review of Philosophy of Knowledge, 58(2):207–221, June 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Carrara:2004:NDM

[1594] Massimiliano Carrara and Pierdaniele Giaretta. Notes and discussions: The many facets of identity criteria. Dialectica: International Review of Philosophy of Knowledge, 58(2):221–232, June 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2004:RECa

[1595] Anonymous. Review — Étude critique — Buchbesprechung. Dialectica: International Review of Philosophy of Knowledge, 58(2):233–237, June 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Recanati:2004:BSO

[1596] François Recanati, Jonathan Barnes, and Marco Santambrogio. Book symposium “Oratio Obliqua, Oratio Recta”. Dialectica: International Review of Philosophy of Knowledge, 58(2):237–247, June 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Barnes:2004:II

[1597] Jonathan Barnes. Innocence and iconicity. Dialectica: International Review of Philosophy of Knowledge, 58(2):248–264, June 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Santambrogio:2004:SIR

[1598] Marco Santambrogio. Semantic innocence recovered? Dialectica: International Review of Philosophy of Knowledge, 58(2):264–272, June 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2004:RECb

[1599] Anonymous. Reviews — Études critiques — Buchbesprechungen. Dialectica: International Review of Philosophy of Knowledge, 58(2):273–274, June 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Correia:2004:I

[1600] Fabrice Correia and Philipp Keller. Introduction. Dialectica: International Review of Philosophy of Knowledge, 58(3):275–278, September 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Smith:2004:CFO

[1601] Barry Smith and Pierre Grenon. The cornucopia of formal-ontological relations. Dialectica: International Review of Philosophy of Knowledge, 58(3):279–296, September 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lowe:2004:SFO

[1602] E. J. Lowe. Some formal ontological relations. Dialectica: International Review of Philosophy of Knowledge, 58(3):297–316, September 2004. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

MacBride:2004:PMU

[1603] Fraser MacBride. Particulars, modes and universals: An examination of E. J. Lowe’s four-fold ontology. Dialectica: International Review of Philosophy of Knowledge, 58(3):317–333, September 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lowe:2004:PUD

[1604] E. J. Lowe. The particular–universal distinction: A reply to MacBride. Dialectica: International Review of Philosophy of Knowledge, 58(3):335– 340, September 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Simons:2004:L

[1605] Peter Simons. Location. Dialectica: International Review of Philosophy of Knowledge, 58(3):341–347, September 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Correia:2004:HF

[1606] Fabrice Correia. Husserl on foundation. Dialectica: International Review of Philosophy of Knowledge, 58(3):349–367, September 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Frascolla:2004:NTO

[1607] Pasquale Frascolla. On the nature of Tractatus objects. Dialectica: International Review of Philosophy of Knowledge, 58(3):369–382, September 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Casati:2004:OCF

[1608] Roberto Casati. Is the object concept formal? Dialectica: International Review of Philosophy of Knowledge, 58(3):383–394, September 2004. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Casalegno:2004:LCL

[1609] Paolo Casalegno. Logical concepts and logical inferences. Dialectica: International Review of Philosophy of Knowledge, 58(3):395–411, September 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hyttinen:2004:DAT

[1610] Tapani Hyttinen and Gabriel Sandu. Deflationism and arithmetical truth. Dialectica: International Review of Philosophy of Knowledge, 58 (3):413–426, September 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wolenski:2004:WFF

[1611] Jan Woleński. What is formal in formal semantics? Dialectica: International Review of Philosophy of Knowledge, 58(3):427–436, September 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rizzi:2004:ICO

[1612] Luigi Rizzi. From inventories to computations: Open /closed class items and substantive /functional heads. Dialectica: International Review of Philosophy of Knowledge, 58(3):437–451, September 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2004:S

[1613] Anonymous. Statistics. Dialectica: International Review of Philosophy of Knowledge, 58(3):453–454, September 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dokic:2004:I

[1614] Jérôme Dokic and Pascal Engel. Introduction. Dialectica: International Review of Philosophy of Knowledge, 58(4):459, December 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Levi:2004:LCL

[1615] Isaac Levi. The logic of consistency and the logic of truth. Dialectica: International Review of Philosophy of Knowledge, 58(4):461–482, December 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bradley:2004:RRT

[1616] Richard Bradley. Ramsey’s representation theorem. Dialectica: International Review of Philosophy of Knowledge, 58(4):483–497, December 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vickers:2004:RJT

[1617] John M. Vickers. Ramsey on judgment: The theory of “Facts and propositions”. Dialectica: International Review of Philosophy of Knowledge, 58(4):499–516, December 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rivenc:2004:CDV

[1618] François Rivenc. Contre la déflation de la vérité. (French) [Against deflation of truth]. Dialectica: International Review of Philosophy of Knowledge, 58(4):517–528, December 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tiercelin:2004:RP

[1619] Claudine Tiercelin. Ramsey’s pragmatism. Dialectica: International Review of Philosophy of Knowledge, 58(4):529–547, December 2004. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Olsson:2004:FPR

[1620] Erik J. Olsson. F. P. Ramsey on knowledge and fallibilism. Dialectica: International Review of Philosophy of Knowledge, 58(4):549–557, December 2004. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Soldati:2005:E

[1621] Gianfranco Soldati. Editorial. Dialectica: International Review of Philosophy of Knowledge, 59(1):3–4, March 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Esfeld:2005:MCM

[1622] Michael Esfeld. Mental causation and mental properties. Dialectica: International Review of Philosophy of Knowledge, 59(1):5–18, March 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

McEvoy:2005:BIP

[1623] Mark McEvoy. Belief-independent processes and the generality problem for reliabilism. Dialectica: International Review of Philosophy of Knowledge, 59(1):19–35, March 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Griesmaier:2005:KSU

[1624] Franz-Peter Griesmaier. Kitcher-style unificationism and explanatory relevance. Dialectica: International Review of Philosophy of Knowledge, 59(1):37–50, March 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Levy:2005:NDC

[1625] Neil Levy. Notes and discussions: Contrastive explanations: A dilemma for libertarians. Dialectica: International Review of Philosophy of Knowledge, 59(1):51–61, March 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pineda:2005:NDC

[1626] David Pineda. Notes and discussions: Causal exclusion and causal homogeneity. Dialectica: International Review of Philosophy of Knowledge, 59 (1):63–66, March 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Borghini:2005:NDC

[1627] Andrea Borghini. Notes and discussions: Counterpart theory vindicated: A reply to Merricks. Dialectica: International Review of Philosophy of Knowledge, 59(1):67–73, March 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Casetta:2005:BRL

[1628] Elena Casetta and Achille C. Varzi. Book review: On Location: Aristotle’s Concept of Place. Dialectica: International Review of Philosophy of Knowledge, 59(1):75–81, March 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

MacBride:2005:BRI

[1629] Fraser MacBride. Book review: Introducing Analytic Philosophy: Its Sense and its Nonsense 1879–2002. Dialectica: International Review of Philosophy of Knowledge, 59(1):81–87, March 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Garcia-Carpintero:2005:BST

[1630] Manuel García-Carpintero. Book symposium “Tractatus Logico-Philosophicus. Introduzione alla lettura”: Frascolla on tractarian logical pictures of facts. Dialectica: International Review of Philosophy of Knowledge, 59(1):87–97, March 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marconi:2005:BST

[1631] Diego Marconi. Book symposium “Tractatus logico-philosophicus. Introduzione alla lettura”: Frascolla on logic in the Tractatus. Dialectica: International Review of Philosophy of Knowledge, 59(1):97–107, March 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Frascolla:2005:BSTa

[1632] Pasquale Frascolla. Book symposium “Tractatus logico-philosophicus. Introduzione alla lettura”: Rejoinder to Carpintero. Dialectica: International Review of Philosophy of Knowledge, 59(1):107–109, March 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Frascolla:2005:BSTb

[1633] Pasquale Frascolla. Book symposium “Tractatus logico-philosophicus. Introduzione alla lettura”: Rejoinder to Marconi. Dialectica: International Review of Philosophy of Knowledge, 59(1):109–112, March 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Macia:2005:II

[1634] Josep Macià. Introduction: Introduction. Dialectica: International Review of Philosophy of Knowledge, 59(2):115–119, June 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cook:2005:AI

[1635] Roy T. Cook and Philip A. Ebert. Abstraction and identity. Dialectica: International Review of Philosophy of Knowledge, 59(2):121–139, June 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Demopoulos:2005:OKN

[1636] William Demopoulos. Our knowledge of numbers as self-subsistent objects. Dialectica: International Review of Philosophy of Knowledge, 59 (2):141–159, June 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Heck:2005:JCB

[1637] Richard G. Heck, Jr. Julius Caesar and basic law V. Dialectica: International Review of Philosophy of Knowledge, 59(2):161–178, June 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kemp:2005:CFP

[1638] Gary Kemp. Caesar from Frege’s perspective. Dialectica: International Review of Philosophy of Knowledge, 59(2):179–199, June 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Linnebo:2005:F

[1639] Øystein Linnebo. To be is to be an F. Dialectica: International Review of Philosophy of Knowledge, 59(2):201–222, June 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

MacBride:2005:JCP

[1640] Fraser MacBride. The Julio César problem. Dialectica: International Review of Philosophy of Knowledge, 59(2):223–236, June 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tappenden:2005:CPH

[1641] Jamie Tappenden. The Caesar Problem in its historical context: Mathematical background. Dialectica: International Review of Philosophy of Knowledge, 59(2):237–264, June 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Uzquiano:2005:SN

[1642] Gabriel Uzquiano. Semantic nominalism. Dialectica: International Review of Philosophy of Knowledge, 59(2):265–282, June 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jespersen:2005:EIA

[1643] Bjørn Jespersen. Explicit intensionalization, anti-actualism, and how Smith’s murderer might not have murdered Smith. Dialectica: International Review of Philosophy of Knowledge, 59(3):285–314, September 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Douven:2005:DCI

[1644] Igor Douven and Frank Hindriks. Deflating the correspondence intuition. Dialectica: International Review of Philosophy of Knowledge, 59(3):315– 329, September 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Katzav:2005:WPC

[1645] Joel Katzav. On what powers cannot do. Dialectica: International Review of Philosophy of Knowledge, 59(3):331–345, September 2005. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bates:2005:OPI

[1646] Jared Bates. The old problem of induction and the new reflective equilibrium. Dialectica: International Review of Philosophy of Knowledge, 59(3):347–356, September 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Philie:2005:MSA

[1647] Patrice Philie. Meaning-scepticism and analyticity. Dialectica: International Review of Philosophy of Knowledge, 59(3):357–365, September 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Einheuser:2005:NDT

[1648] Iris Einheuser. Notes and discussions: Two types of rigid designation. Dialectica: International Review of Philosophy of Knowledge, 59(3):367– 374, September 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gochet:2005:BRW

[1649] Paul Gochet. Book review: Wissenschaft und Empfindung: Die Immanuel Kant Lectures. Dialectica: International Review of Philosophy of Knowledge, 59(3):375–378, September 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bottani:2005:I

[1650] Andrea Bottani. Introduction. Dialectica: International Review of Philosophy of Knowledge, 59(4):381–400, December 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zimmerman:2005:TTB

[1651] Dean W. Zimmerman. The A-theory of time, the B-theory of time, and ‘taking tense seriously’. Dialectica: International Review of Philosophy of Knowledge, 59(4):401–457, December 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Paganini:2005:CZ

[1652] Elisa Paganini. Comments on Zimmerman. Dialectica: International Review of Philosophy of Knowledge, 59(4):459–462, December 2005. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hughes:2005:MFA

[1653] Christopher Hughes. More fuss about formulation: Sider (and me) on three- and four-dimensionalism. Dialectica: International Review of Philosophy of Knowledge, 59(4):463–480, December 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Carrara:2005:CH

[1654] Massimiliano Carrara. Comments on Hughes. Dialectica: International Review of Philosophy of Knowledge, 59(4):481–484, December 2005. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Varzi:2005:CTP

[1655] Achille C. Varzi. Change, temporal parts, and the argument from vagueness. Dialectica: International Review of Philosophy of Knowledge, 59 (4):485–498, December 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Correia:2005:CV

[1656] Fabrice Correia. Comments on Varzi. Dialectica: International Review of Philosophy of Knowledge, 59(4):499–502, December 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mugnai:2005:LSC

[1657] Massimo Mugnai. Leibniz on substance and changing properties. Dialectica: International Review of Philosophy of Knowledge, 59(4):503–516, December 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Morena:2005:CM

[1658] Luca Morena. Comments on mugnai. Dialectica: International Review of Philosophy of Knowledge, 59(4):517–520, December 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2005:VCI

[1659] Anonymous. Volume contents: Index for volume 59. Dialectica: International Review of Philosophy of Knowledge, 59(4):521–522, December 2005. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Engel:2006:ETB

[1660] Pascal Engel. Editorial: Top down and bottom up. Dialectica: International Review of Philosophy of Knowledge, 60(1):3–4, March 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schneider:2006:TFO

[1661] Christina Schneider. Towards a field ontology. Dialectica: International Review of Philosophy of Knowledge, 60(1):5–27, March 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Deonna:2006:EPP

[1662] Julien A. Deonna. Emotion, perception and perspective. Dialectica: International Review of Philosophy of Knowledge, 60(1):29–46, March 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rosenkranz:2006:MAL

[1663] Sven Rosenkranz. Metaethics, agnosticism, and logic. Dialectica: International Review of Philosophy of Knowledge, 60(1):47–61, March 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Janvid:2006:CSS

[1664] Mikael Janvid. Contextualism and the structure of skeptical arguments. Dialectica: International Review of Philosophy of Knowledge, 60(1):63– 77, March 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Becker:2006:NDC

[1665] Kelly Becker. Notes and discussions: Is counterfactual reliabilism compatible with higher-level knowledge? Dialectica: International Review of Philosophy of Knowledge, 60(1):79–84, March 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gemes:2006:NDC

[1666] Ken Gemes. Notes and discussions: Content & Watkins’s account of natural axiomatizations. Dialectica: International Review of Philosophy of Knowledge, 60(1):85–92, March 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anstey:2006:BRS

[1667] Peter Anstey. Book review: La structure du monde: objets, propriétés, états de choses. Dialectica: International Review of Philosophy of Knowledge, 60(1):93–96, March 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cameron:2006:TNC

[1668] Ross Cameron. Tropes, necessary connections, and non-transferability. Dialectica: International Review of Philosophy of Knowledge, 60(2):99– 113, June 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wilholt:2006:TAC

[1669] Torsten Wilholt. Think about the consequences! nominalism and the argument from the philosophy of logic. Dialectica: International Review of Philosophy of Knowledge, 60(2):115–133, June 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gorman:2006:TAI

[1670] Michael Gorman. Talking about intentional objects. Dialectica: International Review of Philosophy of Knowledge, 60(2):135–144, June 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Paseau:2006:SA

[1671] Alexander Paseau. The subtraction argument(s). Dialectica: International Review of Philosophy of Knowledge, 60(2):145–156, June 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bagnoli:2006:BTS

[1672] Carla Bagnoli. Breaking ties: The significance of choice in symmetrical moral dilemmas. Dialectica: International Review of Philosophy of Knowledge, 60(2):157–170, June 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ellis:2006:CEE

[1673] Jonathan Ellis. Color, error, and explanatory power. Dialectica: International Review of Philosophy of Knowledge, 60(2):171–179, June 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gaudet:2006:QNF

[1674] Eve Gaudet. Quine’s notion of fact of the matter. Dialectica: International Review of Philosophy of Knowledge, 60(2):181–193, June 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Murphy:2006:RCB

[1675] Peter Murphy. Reliability connections between conceivability and inconceivability. Dialectica: International Review of Philosophy of Knowledge, 60(2):195–205, June 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mizrahi:2006:II

[1676] Vivian Mizrahi and Martine Nida-Rümelin. Introduction: Introduction. Dialectica: International Review of Philosophy of Knowledge, 60(3):209– 222, September 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Byrne:2006:CMB

[1677] Alex Byrne. Color and the mind-body problem. Dialectica: International Review of Philosophy of Knowledge, 60(3):223–244, September 2006. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Maund:2006:ITC

[1678] Barry Maund. The illusory theory of colours: An anti-realist theory. Dialectica: International Review of Philosophy of Knowledge, 60(3):245– 268, September 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Levine:2006:CCE

[1679] Joseph Levine. Color and color experience: Colors as ways of appearing. Dialectica: International Review of Philosophy of Knowledge, 60(3):269– 282, September 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mizrahi:2006:COC

[1680] Vivian Mizrahi. Color objectivism and color pluralism. Dialectica: International Review of Philosophy of Knowledge, 60(3):283–306, September 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cohen:2006:CPV

[1681] Jonathan Cohen. Color and perceptual variation revisited: Unknown facts, alien modalities, and perfect psychosemantics. Dialectica: International Review of Philosophy of Knowledge, 60(3):307–319, September 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nida-Rumelin:2006:PAC

[1682] Martine Nida-Rümelin. A puzzle about colors. Dialectica: International Review of Philosophy of Knowledge, 60(3):321–336, September 2006. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Byrne:2006:CC

[1683] Alex Byrne. Comments: Comments. Dialectica: International Review of Philosophy of Knowledge, 60(3):337–340, September 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hardin:2006:CC

[1684] Larry Hardin. Comments: Comments. Dialectica: International Review of Philosophy of Knowledge, 60(3):341–345, September 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Maund:2006:CC

[1685] Barry Maund. Comments: Comments. Dialectica: International Review of Philosophy of Knowledge, 60(3):347–353, September 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dougherty:2006:CUI

[1686] Randall Dougherty and Jan Mycielski. Canonical universes and intuitions about probabilities. Dialectica: International Review of Philosophy of Knowledge, 60(4):357–368, December 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Potter:2006:DRG

[1687] Dennis Potter. Diagrammatic representation in geometry. Dialectica: International Review of Philosophy of Knowledge, 60(4):369–382, December 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Montero:2006:WDC

[1688] Barbara Montero. What does the conservation of energy have to do with physicalism? Dialectica: International Review of Philosophy of Knowledge, 60(4):383–396, December 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sa:2006:EP

[1689] Dan Lópezde Sa. Is ‘everything’ precise? Dialectica: International Review of Philosophy of Knowledge, 60(4):397–409, December 2006. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Textor:2006:BSN

[1690] Mark Textor. Brentano (and some neo-Brentanians) on inner consciousness. Dialectica: International Review of Philosophy of Knowledge, 60 (4):411–432, December 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lasonen-Aarnio:2006:EPK

[1691] Maria Lasonen-Aarnio. Externalism and a priori knowledge of the world: Why privileged access is not the issue. Dialectica: International Review of Philosophy of Knowledge, 60(4):433–445, December 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

McKinsey:2006:DRL

[1692] Michael McKinsey. Direct reference and logical truth: a reply to Lasonen-Aarnio. Dialectica: International Review of Philosophy of Knowledge, 60 (4):447–451, December 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Levy:2006:DD

[1693] Neil Levy. Determinist deliberations. Dialectica: International Review of Philosophy of Knowledge, 60(4):453–459, December 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Contessa:2006:STA

[1694] Gabriele Contessa. On the supposed temporal asymmetry of counterfactual dependence; or: It wouldn’t have taken a miracle! Dialectica: International Review of Philosophy of Knowledge, 60(4):461–473, December 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Antony:2006:PVP

[1695] Michael V. Antony. Papineau on the vagueness of phenomenal concepts. Dialectica: International Review of Philosophy of Knowledge, 60(4):475– 483, December 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bigelow:2006:RWA

[1696] John Bigelow and Robert Pargetter. Real work for aggregates. Dialectica: International Review of Philosophy of Knowledge, 60(4):485–503, December 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Urkia:2006:BRL

[1697] Igor Aristegi Urkia. Book review: L’Intentionnalité. Problèmes de philosophie de l’esprit — Pierre Jacob. Dialectica: International Review of Philosophy of Knowledge, 60(4):505–508, December 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sachse:2006:BRP

[1698] Christian Sachse. Book review: Physicalism, or Something Near Enough — Jaegwon Kim. Dialectica: International Review of Philosophy of Knowledge, 60(4):508–512, December 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Johansson:2006:BRF

[1699] Ingvar Johansson. Book review: The Four-Category Ontology. A Metaphysical Foundation for Natural Science — E. Jonathan Lowe. Dialectica: International Review of Philosophy of Knowledge, 60(4):513–518, December 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2006:VCI

[1700] Anonymous. Volume contents: Index for volume 60. Dialectica: International Review of Philosophy of Knowledge, 60(4):519–520, December 2006. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fine:2007:SIPa

[1701] Kit Fine. Special issue: The philosophy of Kit Fine Guest Editor: Kevin Mulligan: Introduction. Dialectica: International Review of Philosophy of Knowledge, 61(1):3, March 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Horwich:2007:SIP

[1702] Paul Horwich. Special issue: The philosophy of Kit Fine Guest Editor: Kevin Mulligan: The quest for reality. Dialectica: International Review of Philosophy of Knowledge, 61(1):5–16, March 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fine:2007:SIPb

[1703] Kit Fine. Special issue: The philosophy of Kit Fine Guest Editor: Kevin Mulligan: Response to Paul Horwich. Dialectica: International Review of Philosophy of Knowledge, 61(1):17–23, March 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

MacBride:2007:SIP

[1704] Fraser MacBride. Special issue: The philosophy of Kit Fine Guest Editor: Kevin Mulligan: Neutral relations revisited. Dialectica: International Review of Philosophy of Knowledge, 61(1):25–56, March 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fine:2007:SIPc

[1705] Kit Fine. Special issue: The philosophy of Kit Fine Guest Editor: Kevin Mulligan: Response to Fraser MacBride. Dialectica: International Review of Philosophy of Knowledge, 61(1):57–62, March 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Correia:2007:SIP

[1706] Fabrice Correia. Special issue: The philosophy of Kit Fine Guest Editor: Kevin mulligan: (Finean) essence and (Priorean) modality. Dialectica: International Review of Philosophy of Knowledge, 61(1):63–84, March 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fine:2007:SIPd

[1707] Kit Fine. Special issue: The philosophy of Kit Fine Guest Editor: Kevin Mulligan: Response to Fabrice Correia. Dialectica: International Review of Philosophy of Knowledge, 61(1):85–88, March 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Weir:2007:SIP

[1708] Alan Weir. Special issue: The philosophy of Kit Fine Guest Editor: Kevin Mulligan: Honest toil or sheer magic? Dialectica: International Review of Philosophy of Knowledge, 61(1):89–115, March 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fine:2007:SIPe

[1709] Kit Fine. Special issue: The philosophy of Kit Fine Guest Editor: Kevin Mulligan: Response to Alan Weir. Dialectica: International Review of Philosophy of Knowledge, 61(1):117–125, March 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Koslicki:2007:SIP

[1710] Kathrin Koslicki. Special issue: The philosophy of Kit Fine Guest Editor: Kevin Mulligan: Towards a neo-Aristotelian mereology. Dialectica: International Review of Philosophy of Knowledge, 61(1):127–159, March 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fine:2007:SIPf

[1711] Kit Fine. Special issue: The philosophy of Kit Fine Guest Editor: Kevin Mulligan: Response to Kathrin Koslicki. Dialectica: International Review of Philosophy of Knowledge, 61(1):161–166, March 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Garcia-Carpintero:2007:SIP

[1712] Manuel García-Carpintero. Special issue: The philosophy of Kit Fine Guest Editor: Kevin Mulligan: Bivalence and what is said. Dialectica: International Review of Philosophy of Knowledge, 61(1):167–190, March 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fine:2007:RMG

[1713] Kit Fine. Response to Manuel García–Carpintero. Dialectica: International Review of Philosophy of Knowledge, 61(1):191–194, March 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2007:SIP

[1714] Anonymous. Special issue: The philosophy of Kit Fine Guest Editor: Kevin Mulligan: Bibliography of Kit Fine. Dialectica: International Review of Philosophy of Knowledge, 61(1):195–200, March 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lillehammer:2007:DVO

[1715] Hallvard Lillehammer. Davidson on value and objectivity. Dialectica: International Review of Philosophy of Knowledge, 61(2):203–217, June 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Aguilar:2007:IIB

[1716] Jesús H. Aguilar. Interpersonal interactions and the bounds of agency. Dialectica: International Review of Philosophy of Knowledge, 61(2):219– 234, June 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Berto:2007:DIR

[1717] Francesco Berto. Is dialetheism an idealism? The Russellian fallacy and the Dialetheist’s dilemma. Dialectica: International Review of Philosophy of Knowledge, 61(2):235–263, June 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hansson:2007:PCR

[1718] Tobias Hansson. The problem(s) of change revisited. Dialectica: International Review of Philosophy of Knowledge, 61(2):265–274, June 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gluer:2007:CND

[1719] Kathrin Glüer. Critical notice: Donald Davidson’s collected essays. Dialectica: International Review of Philosophy of Knowledge, 61(2):275– 284, June 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ziv:2007:BRB

[1720] Anita Konzelmann Ziv. Book review: Bolzanos Metaphysik — By Andrej Krause. Dialectica: International Review of Philosophy of Knowledge, 61 (2):285–289, June 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lando:2007:BRT

[1721] Giorgio Lando. Book review: Teorie dell’assurdo. I rivali del Principio di Non-Contraddizione — By Francesco Berto. Dialectica: International Review of Philosophy of Knowledge, 61(2):291–295, June 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Volpe:2007:BRF

[1722] Giorgio Volpe. Book review: Fattori di verità — By Stefano Caputo. Dialectica: International Review of Philosophy of Knowledge, 61(2):297– 302, June 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sparber:2007:BRO

[1723] Georg Sparber. Book review: From an Ontological Point of View — By John Heil. Dialectica: International Review of Philosophy of Knowledge, 61(2):303–307, June 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Deonna:2007:SIA

[1724] Julien Deonna, Christine Clavien, and Ivo Wallimann. Special issue: Affective Intentionality and Practical Rationality. Guest Editors: Julien Deonna, Christine Clavien and Ivo Wallimann: Affective intentionality and practical rationality. Dialectica: International Review of Philosophy of Knowledge, 61(3):311–322, September 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

DeSousa:2007:SIA

[1725] Ronald De Sousa. Special issue: Affective intentionality and practical rationality: Guest Editors: Julien Deonna, Christine Clavien and Ivo Wallimann: Truth, authenticity, and rationality. Dialectica: International Review of Philosophy of Knowledge, 61(3):323–345, September 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Goldie:2007:SIA

[1726] Peter Goldie. Special issue: Affective intentionality and practical rationality: Guest Editors: Julien Deonna, Christine Clavien and Ivo Wallimann: Seeing what is the kind thing to do: Perception and emotion in morality. Dialectica: International Review of Philosophy of Knowledge, 61(3):347–361, September 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Doring:2007:SIA

[1727] Sabine A. Döring. Special issue: Affective intentionality and practical rationality: Guest Editors: Julien Deonna, Christine Clavien and Ivo Wallimann: Seeing what to do: Affective perception and rational motivation. Dialectica: International Review of Philosophy of Knowledge, 61(3):363–394, September 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Teroni:2007:SIA

[1728] Fabrice Teroni. Special issue: Affective intentionality and practical rationality: Guest Editors: Julien Deonna, Christine Clavien and Ivo Wallimann: Emotions and formal objects. Dialectica: International Review of Philosophy of Knowledge, 61(3):395–415, September 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dorsch:2007:SIA

[1729] Fabian Dorsch. Special issue: Affective intentionality and practical rationality: Guest Editors: Julien Deonna, Christine Clavien and Ivo Wallimann: Sentimentalism and the intersubjectivity of aesthetic evaluations. Dialectica: International Review of Philosophy of Knowledge, 61(3):417– 446, September 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Betzler:2007:SIA

[1730] Monika Betzler. Special issue: Affective intentionality and practical rationality: Guest Editors: Julien Deonna, Christine Clavien and Ivo Wallimann: Making sense of actions expressing emotions. Dialectica: International Review of Philosophy of Knowledge, 61(3):447–466, September 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ziv:2007:SIA

[1731] Anita Konzelmann Ziv. Special issue: Affective intentionality and practical rationality: Guest Editors: Julien Deonna, Christine Clavien and Ivo Wallimann: Collective guilt feeling revisited. Dialectica: International Review of Philosophy of Knowledge, 61(3):467–493, September 2007. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pettit:2007:RRG

[1732] Philip Pettit. Rationality, reasoning and group agency. Dialectica: International Review of Philosophy of Knowledge, 61(4):495–519, December 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Weiss:2007:TEK

[1733] Bernhard Weiss. Truth and the enigma of knowability. Dialectica: International Review of Philosophy of Knowledge, 61(4):521–537, December 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gibb:2007:PIA

[1734] Sophie Gibb. Is the partial identity account of property resemblance logically incoherent?: Winner of the 2006 Dialectica Essay Prize. Dialectica: International Review of Philosophy of Knowledge, 61(4):539–558, December 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Goddu:2007:BPT

[1735] G. C. Goddu. Banana peels and time travel. Dialectica: International Review of Philosophy of Knowledge, 61(4):559–572, December 2007. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Koksvik:2007:NDC

[1736] Ole Koksvik. Notes and discussions: Conservation of energy is relevant to physicalism. Dialectica: International Review of Philosophy of Knowledge, 61(4):573–582, December 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Morgan:2007:BRM

[1737] Daniel Morgan. Book review: I: The Meaning of the First Person Term — By Robert Maximilian de Gaynesford. Dialectica: International Review of Philosophy of Knowledge, 61(4):583–587, December 2007. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schnieder:2007:BRE

[1738] Benjamin Schnieder. Book review: Existential Dependence and Cognate Notions — By Fabrice Correia. Dialectica: International Review of Philosophy of Knowledge, 61(4):589–594, December 2007. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Meinertsen:2008:RUS

[1739] Bo Meinertsen. A relation as the unifier of states of affairs. Dialectica: International Review of Philosophy of Knowledge, 62(1):1–19, March 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Borghini:2008:DTP

[1740] Andrea Borghini and Neil E. Williams. A dispositional theory of possibility. Dialectica: International Review of Philosophy of Knowledge, 62 (1):21–41, March 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marti:2008:DRD

[1741] Genoveva Marti. Direct reference and definite descriptions. Dialectica: International Review of Philosophy of Knowledge, 62(1):43–57, March 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rosenkranz:2008:KCA

[1742] Sven Rosenkranz. Knowability, closure, and anti-realism. Dialectica: International Review of Philosophy of Knowledge, 62(1):59–75, March 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cox:2008:CBR

[1743] Edward T. Cox. Crimson brain, red mind: Yablo on mental causation. Dialectica: International Review of Philosophy of Knowledge, 62 (1):77–99, March 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Douven:2008:KPR

[1744] Igor Douven. Knowledge and practical reasoning. Dialectica: International Review of Philosophy of Knowledge, 62(1):101–118, March 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sylvand:2008:BRL

[1745] Benjamin Sylvand. Book review: L’institution du langage — By Michel Seymour. Dialectica: International Review of Philosophy of Knowledge, 62(1):119–123, March 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Beyer:2008:BRR

[1746] Christian Beyer. Book review: Rediscovering Empathy — Agency, Folk Psychology, and the Human Sciences — By Karsten R. Stueber. Dialectica: International Review of Philosophy of Knowledge, 62(1):123–128, March 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Esfeld:2008:BRS

[1747] Michael Esfeld and Michael Sollberger. Book review: Strukturale Repräsentation — By Andreas Bartels; Subjektivität, Intersubjektivität, Personalität. Ein Beitrag zur Philosophie der Person — By Christian Beyer; Bilder im Geiste. Die Imagery-Debatte — By Verena Gottschling; Der Blick von Innen. Zur transtemporalen Identität bewusstseinsfähiger Wesen — By Martine Nida–Rümelin; Illusion Freiheit? Mögliche und unmögliche Konsequenzen der Hirnforschung — By Michael Pauen; Willensfreiheit und Hirnforschung. Das Freiheitsmodell des epistemischen Libertarismus — By Bettina Walde; Der mentale Zugang zur Welt. Realismus, Skeptizismus und Intentionalität By Marcus Willaschek. Dialectica: International Review of Philosophy of Knowledge, 62(1):128–135, March 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bonardi:2008:BRR

[1748] Paolo Bonardi. Book review: Reflecting the Mind. Indexicality and Quasi-Indexicality — By Eros Corazza. Dialectica: International Review of Philosophy of Knowledge, 62(1):135–141, March 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Strahm:2008:II

[1749] Thomas Strahm. Introduction: Introduction. Dialectica: International Review of Philosophy of Knowledge, 62(2):145–147, June 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Diller:2008:FIC

[1750] Justus Diller. Functional interpretations of constructive set theory in all finite types. Dialectica: International Review of Philosophy of Knowledge, 62(2):149–177, June 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Feferman:2008:LHB

[1751] Solomon Feferman. Lieber Herr Bernays!, Lieber Herr Gödel! Gödel on finitism, constructivity and Hilbert’s program. Dialectica: International Review of Philosophy of Knowledge, 62(2):179–203, June 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ferreira:2008:MAP

[1752] Fernando Ferreira. A most artistic package of a jumble of ideas. Dialectica: International Review of Philosophy of Knowledge, 62(2):205–222, June 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kohlenbach:2008:GFI

[1753] Ulrich Kohlenbach. Gödel’s functional interpretation and its use in current mathematics. Dialectica: International Review of Philosophy of Knowledge, 62(2):223–267, June 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Oliva:2008:AGD

[1754] Paulo Oliva. An analysis of Gödel’s dialectica: Interpretation via linear logic. Dialectica: International Review of Philosophy of Knowledge, 62 (2):269–290, June 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Leitgeb:2008:IRS

[1755] Hannes Leitgeb. An impossibility result on semantic resemblance. Dialectica: International Review of Philosophy of Knowledge, 62(3):293– 306, September 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stoneham:2008:NAP

[1756] Tom Stoneham. A neglected account of perception. Dialectica: International Review of Philosophy of Knowledge, 62(3):307–322, September 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Brown:2008:IPR

[1757] Derek Brown. Indirect perceptual realism and multiple reference. Dialectica: International Review of Philosophy of Knowledge, 62(3):323– 334, September 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Blumson:2008:DC

[1758] Ben Blumson. Depiction and convention. Dialectica: International Review of Philosophy of Knowledge, 62(3):335–348, September 2008. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gross:2008:SSW

[1759] Steven Gross. Sincerely saying what you don’t believe again. Dialectica: International Review of Philosophy of Knowledge, 62(3):349–354, September 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kunne:2008:BSP

[1760] Wolfgang Künne. Book symposium: Précis of conceptions of truth. Dialectica: International Review of Philosophy of Knowledge, 62(3):355– 357, September 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sundholm:2008:BSS

[1761] Göran Sundholm. Book symposium: Summa de Veritate Hamburgensis: Truth according to Wolfgang Künne. Dialectica: International Review of Philosophy of Knowledge, 62(3):359–371, September 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wolenski:2008:BSK

[1762] Jan Woleński. Book symposium: Künne on truth and its conceptions. Dialectica: International Review of Philosophy of Knowledge, 62(3):373– 384, September 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kunne:2008:BSR

[1763] Wolfgang Künne. Book symposium: Replies to commentators. Dialectica: International Review of Philosophy of Knowledge, 62(3):385–401, September 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Debus:2008:EPR

[1764] Dorothea Debus. Experiencing the past: A relational account of recollective memory. Dialectica: International Review of Philosophy of Knowledge, 62(4):405–432, December 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Elder:2008:AUM

[1765] Crawford Elder. Against universal mereological composition. Dialectica: International Review of Philosophy of Knowledge, 62(4):433–454, December 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Herzberg:2008:CBE

[1766] Larry A. Herzberg. Constitutivism, belief, and emotion. Dialectica: International Review of Philosophy of Knowledge, 62(4):455–482, December 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Efird:2008:WPR

[1767] David Efird and Tom Stoneham. What is the principle of recombination? Dialectica: International Review of Philosophy of Knowledge, 62(4):483– 494, December 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Harrison:2008:MLC

[1768] Gerald K. Harrison. Modest libertarianism and clandestine control. Dialectica: International Review of Philosophy of Knowledge, 62(4):495– 507, December 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wieland:2008:RSR

[1769] Jan Willem Wieland and Arianna Betti. Relata-specific relations: A response to Vallicella. Dialectica: International Review of Philosophy of Knowledge, 62(4):509–524, December 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hansson:2008:DWN

[1770] Sven Ove Hansson. Do we need second-order probabilities? Dialectica: International Review of Philosophy of Knowledge, 62(4):525–533, December 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lasonen-Aarnio:2008:WEB

[1771] Maria Lasonen-Aarnio. Why the externalist is better off without free logic: A reply to McKinsey. Dialectica: International Review of Philosophy of Knowledge, 62(4):535–540, December 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Steglich-Petersen:2008:DDT

[1772] Asbjørn Steglich-Petersen. Does doxastic transparency support evidentialism? Dialectica: International Review of Philosophy of Knowledge, 62(4):541–547, December 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Oppy:2008:BRD

[1773] Graham Oppy. Book review: Developmental Theism: From Pure Will to Unbounded Love — By Peter Forrest. Dialectica: International Review of Philosophy of Knowledge, 62(4):549–553, December 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Broncano:2008:BRM

[1774] Fernando Broncano. Book review: Moral Responsibility. The Ways of Scepticism — By Carlos Moya. Dialectica: International Review of Philosophy of Knowledge, 62(4):553–557, December 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Voltolini:2008:BRT

[1775] Alberto Voltolini. Book review: Towards Non-being. The Logic And Metaphysics of Intentionality — By G. Priest. Dialectica: International Review of Philosophy of Knowledge, 62(4):557–561, December 2008. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Iacona:2008:BRP

[1776] Andrea Iacona. Book review: Per la verità: relativismo e filosofia — By Diego Marconi. Dialectica: International Review of Philosophy of Knowledge, 62(4):561–564, December 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Paternoster:2008:BRL

[1777] Alfredo Paternoster. Book review: Langage et cognition humaine — By Anne Reboul. (French) [Book review: Language and Human Cognition — By Anne Reboul]. Dialectica: International Review of Philosophy of Knowledge, 62(4):564–568, December 2008. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Engel:2009:EMM

[1778] Pascal Engel. Editorial: Measure for measure. Dialectica: International Review of Philosophy of Knowledge, 63(1):1–5, March 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rychter:2009:TNP

[1779] Pablo Rychter. There is no puzzle about change. Dialectica: International Review of Philosophy of Knowledge, 63(1):7–22, March 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Miller:2009:DCM

[1780] Kristie Miller. Defending contingentism in metaphysics. Dialectica: International Review of Philosophy of Knowledge, 63(1):23–49, March 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Voltolini:2009:BSP

[1781] Alberto Voltolini. Book symposium: Précis of how ficta follow fiction. Dialectica: International Review of Philosophy of Knowledge, 63(1):51–55, March 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Garcia-Carpintero:2009:BSV

[1782] Manuel García-Carpintero. Book symposium: Voltolini’s ficta. Dialectica: International Review of Philosophy of Knowledge, 63(1):57–66, March 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Corazza:2009:BSF

[1783] Eros Corazza. Book symposium: Fictions without ficta. Dialectica: International Review of Philosophy of Knowledge, 63(1):67–74, March 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Voltolini:2009:BSH

[1784] Alberto Voltolini. Book symposium: How ficta follow fiction: Replies to commentators. Dialectica: International Review of Philosophy of Knowledge, 63(1):75–84, March 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Meylan:2009:BRV

[1785] Anne Meylan. Book review: A Virtue Epistemology. Apt Belief and Reflective Knowledge — By E. Sosa. Dialectica: International Review of Philosophy of Knowledge, 63(1):85–89, March 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Giaretta:2009:BRM

[1786] Pierdaniele Giaretta. Book review: Medicine & Philosophy. A Twenty-First Century Introduction — By Ingvar Johansson and Niels Lynøe. Dialectica: International Review of Philosophy of Knowledge, 63(1):89–94, March 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nuccetelli:2009:BRT

[1787] Susana Nuccetelli. Book review: Two-Dimensional Semantics — Edited by Manuel García–Carpintero and Josep Maciá. Dialectica: International Review of Philosophy of Knowledge, 63(1):94–99, March 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kistler:2009:BRN

[1788] Max Kistler. Book review: Naturphilosophie als Metaphysik der Natur — By Michael Esfeld. Dialectica: International Review of Philosophy of Knowledge, 63(1):99–103, March 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kasabova:2009:BRN

[1789] Anita Kasabova. Book review: Neuer Anti-Kant und Atomenlehre des seligen Bolzano — By Franz Příhonský. Dialectica: International Review of Philosophy of Knowledge, 63(1):103–105, March 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Caputo:2009:BRT

[1790] Stefano Caputo. Book review: Teorie della Verità- — By Giorgio Volpe. Dialectica: International Review of Philosophy of Knowledge, 63(1):106– 110, March 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marconi:2009:OPC

[1791] Diego Marconi. Obituary of Paolo Casalegno. Dialectica: International Review of Philosophy of Knowledge, 63(2):115–116, June 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gerken:2009:CEE

[1792] Mikkel Gerken. Conceptual equivocation and epistemic relevance. Dialectica: International Review of Philosophy of Knowledge, 63(2):117– 132, June 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wall:2009:TPD

[1793] David Wall. Are there passive desires? Dialectica: International Review of Philosophy of Knowledge, 63(2):133–155, June 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rauti:2009:CWD

[1794] Antonio Rauti. Can we derive the principle of compositionality (if we deflate understanding)? Dialectica: International Review of Philosophy of Knowledge, 63(2):157–174, June 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Baumgartner:2009:ICI

[1795] Michael Baumgartner. Interdefining causation and intervention. Dialectica: International Review of Philosophy of Knowledge, 63(2):175–194, June 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Darby:2009:LWR

[1796] George Darby. Lewis’s worldmate relation and the apparent failure of Humean supervenience. Dialectica: International Review of Philosophy of Knowledge, 63(2):195–204, June 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schrenk:2009:CPE

[1797] Markus Schrenk. Can physics ever be complete if there is no fundamental level in nature? Dialectica: International Review of Philosophy of Knowledge, 63(2):205–208, June 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Turri:2009:IAD

[1798] John Turri. An infinitist account of doxastic justification. Dialectica: International Review of Philosophy of Knowledge, 63(2):209–218, June 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gomes:2009:OMP

[1799] Anil Gomes. Other minds and perceived identity. Dialectica: International Review of Philosophy of Knowledge, 63(2):219–230, June 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Casalegno:2009:RBA

[1800] Paolo Casalegno. Reasons to believe and assertion. Dialectica: International Review of Philosophy of Knowledge, 63(3):231–248, September 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Recanati:2009:RS

[1801] François Recanati. De re and De se. Dialectica: International Review of Philosophy of Knowledge, 63(3):249–269, September 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Morganti:2009:OPF

[1802] Matteo Morganti. Ontological priority, fundamentality and monism: Winner of the 2008 Dialectica Essay Prize. Dialectica: International Review of Philosophy of Knowledge, 63(3):271–288, September 2009. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Goff:2009:WPD

[1803] Philip Goff. Why panpsychism doesn’t help us explain consciousness. Dialectica: International Review of Philosophy of Knowledge, 63(3):289– 311, September 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Eklund:2009:SRC

[1804] Matti Eklund. On some recent criticisms of the ‘linguistic’ approach to ontology. Dialectica: International Review of Philosophy of Knowledge, 63(3):313–323, September 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Buleandra:2009:DTP

[1805] Andrei Buleandra. Doxastic transparency and prescriptivity. Dialectica: International Review of Philosophy of Knowledge, 63(3):325–332, September 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Orilia:2009:BRU

[1806] Francesco Orilia. Bradley’s regress and ungrounded dependence chains: A reply to Cameron. Dialectica: International Review of Philosophy of Knowledge, 63(3):333–341, September 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Berto:2009:ECN

[1807] Francesco Berto and Massimiliano Carrara. To exist and to count: a note on the minimalist view. Dialectica: International Review of Philosophy of Knowledge, 63(3):343–356, September 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Callaway:2009:BRF

[1808] H. G. Callaway. Book review: Fear of Knowledge, Against Relativism and Constructivism — By Paul Boghossian. Dialectica: International Review of Philosophy of Knowledge, 63(3):357–360, September 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Woodward:2009:BRM

[1809] Richard Woodward. Book review: Metaphysics and the Representational Fallacy — By Heather Dyke. Dialectica: International Review of Philosophy of Knowledge, 63(3):361–365, September 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

DelMar:2009:BRM

[1810] Maksymilian Del Mar. Book review: The Mystery of Capital and the Construction of Social Reality — Edited by Barry Smith, David M. Mark and Isaac Ehrlich. Dialectica: International Review of Philosophy of Knowledge, 63(3):365–368, September 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Leuenberger:2009:IPV

[1811] Stephan Leuenberger and Philipp Keller. Introduction: The philosophy of vectors. Dialectica: International Review of Philosophy of Knowledge, 63(4):369–380, December 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Simons:2009:VBG

[1812] Peter Simons. Vectors and beyond: Geometric algebra and its philosophical significance. Dialectica: International Review of Philosophy of Knowledge, 63(4):381–395, December 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lange:2009:TTV

[1813] Marc Lange. A tale of two vectors. Dialectica: International Review of Philosophy of Knowledge, 63(4):397–431, December 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Johansson:2009:MVP

[1814] Ingvar Johansson. Mathematical vectors and physical vectors. Dialectica: International Review of Philosophy of Knowledge, 63(4):433–447, December 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Busse:2009:HSV

[1815] Ralf Busse. Humean supervenience, vectorial fields, and the spinning sphere. Dialectica: International Review of Philosophy of Knowledge, 63 (4):449–489, December 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Forrest:2009:VCS

[1816] Peter Forrest. Vectors on curved space. Dialectica: International Review of Philosophy of Knowledge, 63(4):491–501, December 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Beisbart:2009:HFD

[1817] Claus Beisbart. How to fix directions or are assignments of vector characteristics attributions of intrinsic properties? Dialectica: International Review of Philosophy of Knowledge, 63(4):503–524, December 2009. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wilson:2009:CAA

[1818] Jessica Wilson. The causal argument against component forces. Dialectica: International Review of Philosophy of Knowledge, 63(4):525–554, December 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Massin:2009:MF

[1819] Olivier Massin. The metaphysics of forces. Dialectica: International Review of Philosophy of Knowledge, 63(4):555–589, December 2009. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wilson:2009:DMR

[1820] Alastair Wilson. Disposition-manifestations and reference-frames. Dialectica: International Review of Philosophy of Knowledge, 63(4):591– 601, December 2009. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lalumera:2010:II

[1821] Elisabetta Lalumera. Introduction: Introduction. Dialectica: International Review of Philosophy of Knowledge, 64(1):1–9, March 2010. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Glock:2010:CAJ

[1822] Hans-Johann Glock. Can animals judge? Dialectica: International Review of Philosophy of Knowledge, 64(1):11–33, March 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Peacocke:2010:RBT

[1823] Christopher Peacocke. Relation-based thought, objectivity and disagreement. Dialectica: International Review of Philosophy of Knowledge, 64 (1):35–56, March 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Martinez-Manrique:2010:DBS

[1824] Fernando Martínez-Manrique. On the distinction between semantic and conceptual representation. Dialectica: International Review of Philosophy of Knowledge, 64(1):57–78, March 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vicente:2010:CSL

[1825] Agustín Vicente. Clusters: On the structure of lexical concepts. Dialectica: International Review of Philosophy of Knowledge, 64(1):79–106, March 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Weiskopf:2010:CMT

[1826] Daniel A. Weiskopf. Concepts and the modularity of thought. Dialectica: International Review of Philosophy of Knowledge, 64(1):107–130, March 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dickie:2010:SCR

[1827] Imogen Dickie and Gurpreet Rattan. Sense, communication, and rational engagement. Dialectica: International Review of Philosophy of Knowledge, 64(2):131–151, June 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stephanou:2010:MA

[1828] Yannis Stephanou. The meaning of ‘actually’. Dialectica: International Review of Philosophy of Knowledge, 64(2):153–185, June 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jagnow:2010:SEP

[1829] René Jagnow. Shadow-experiences and the phenomenal structure of colors. Dialectica: International Review of Philosophy of Knowledge, 64(2): 187–212, June 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Whiting:2010:SBT

[1830] Daniel Whiting. Should I believe the truth? Dialectica: International Review of Philosophy of Knowledge, 64(2):213–224, June 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kellenberg:2010:AV

[1831] Aleksandar Kellenberg. The antinomy of the variable. Dialectica: International Review of Philosophy of Knowledge, 64(2):225–236, June 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Einheuser:2010:MTA

[1832] Iris Einheuser. The model-theoretic argument against quantifying over everything. Dialectica: International Review of Philosophy of Knowledge, 64(2):237–246, June 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bradley:2010:CBS

[1833] Darren Bradley. Conditionalization and belief De Se. Dialectica: International Review of Philosophy of Knowledge, 64(2):247–250, June 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gustafsson:2010:MPA

[1834] Johan E. Gustafsson. A money-pump for acyclic intransitive preferences. Dialectica: International Review of Philosophy of Knowledge, 64(2):251– 257, June 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gaskin:2010:PUP

[1835] Richard Gaskin. Précis of the unity of the proposition. Dialectica: International Review of Philosophy of Knowledge, 64(2):259–264, June 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vallicella:2010:GUP

[1836] William F. Vallicella. Gaskin on the unity of the proposition. Dialectica: International Review of Philosophy of Knowledge, 64(2):265–277, June 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Garcia-Carpintero:2010:GIU

[1837] Manuel García-Carpintero. Gaskin’s ideal unity. Dialectica: International Review of Philosophy of Knowledge, 64(2):279–288, June 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schnieder:2010:PUG

[1838] Benjamin Schnieder. Propositions united: Gaskin on Bradley’s regress and the unity of the proposition. Dialectica: International Review of Philosophy of Knowledge, 64(2):289–301, June 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gaskin:2010:UPR

[1839] Richard Gaskin. The unity of the proposition: Replies to Vallicella, Schnieder, and García–Carpintero. Dialectica: International Review of Philosophy of Knowledge, 64(2):303–311, June 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Henden:2010:DI

[1840] Edmund Henden. Deliberation incompatibilism. Dialectica: International Review of Philosophy of Knowledge, 64(3):313–333, September 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Roca-Royes:2010:MEM

[1841] Sonia Roca-Royes. Modal epistemology, modal concepts and the integration challenge. Dialectica: International Review of Philosophy of Knowledge, 64(3):335–361, September 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gibb:2010:CPL

[1842] Sophie Gibb. Closure principles and the laws of conservation of energy and momentum. Dialectica: International Review of Philosophy of Knowledge, 64(3):363–384, September 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cain:2010:LPS

[1843] M. J. Cain. Linguistics, psychology and the scientific study of language. Dialectica: International Review of Philosophy of Knowledge, 64(3):385– 404, September 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Evers:2010:ERT

[1844] Daan Evers. The end-relational theory of ‘ought’ and the weight of reasons. Dialectica: International Review of Philosophy of Knowledge, 64(3):405–417, September 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sabbarton-Leary:2010:DRM

[1845] Nigel Sabbarton-Leary. Descriptivist reference from metaphysical essence. Dialectica: International Review of Philosophy of Knowledge, 64(3):419–433, September 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Darby:2010:LPR

[1846] George Darby and Duncan Watson. Lewis’s principle of recombination: Reply to Efird and Stoneham. Dialectica: International Review of Philosophy of Knowledge, 64(3):435–445, September 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hochberg:2010:BRP

[1847] Herbert Hochberg. Book review: Phenomenological Realism Versus Scientific Realism: Reinhardt Grossmann–David M. Armstrong Metaphysical Correspondence — Edited by Javier Cumpa and Erwin Tegtmeier. Dialectica: International Review of Philosophy of Knowledge, 64(3):447– 451, September 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marconi:2010:BRP

[1848] Diego Marconi. Book review: Philosophie du langage (et de l’esprit) — By François Recanati. Dialectica: International Review of Philosophy of Knowledge, 64(3):452–458, September 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Frascolla:2010:BRW

[1849] Pasquale Frascolla. Book review: Wittgenstein’s Notes on Logic — By Michael Potter. Dialectica: International Review of Philosophy of Knowledge, 64(3):458–463, September 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Spitzley:2010:E

[1850] Thomas Spitzley. Editorial. Dialectica: International Review of Philosophy of Knowledge, 64(4):465, December 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Beckermann:2010:DWI

[1851] Ansgar Beckermann. Darwin — what if man is only an animal, after all? Dialectica: International Review of Philosophy of Knowledge, 64(4):467– 482, December 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sainsbury:2010:PV

[1852] R. Mark Sainsbury. Paderewski variations. Dialectica: International Review of Philosophy of Knowledge, 64(4):483–502, December 2010. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wallace:2010:RVA

[1853] R. Jay Wallace. Reasons, values and agent-relativity. Dialectica: International Review of Philosophy of Knowledge, 64(4):503–528, December 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kunne:2010:SRH

[1854] Wolfgang Künne. Sense, reference and hybridity: Reflections on Kripke’s recent reading of Frege. Dialectica: International Review of Philosophy of Knowledge, 64(4):529–551, December 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Boghossian:2010:OGC

[1855] Paul Boghossian. Our grasp of the concept of truth: Reflections on Künne. Dialectica: International Review of Philosophy of Knowledge, 64(4):553–563, December 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mulligan:2010:TPV

[1856] Kevin Mulligan. The truth predicate vs. the truth connective. On taking connectives seriously. Dialectica: International Review of Philosophy of Knowledge, 64(4):565–584, December 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kunne:2010:RPB

[1857] Wolfgang Künne. Replies to Paul Boghossian and Kevin Mulligan. Dialectica: International Review of Philosophy of Knowledge, 64(4):585– 615, December 2010. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Salmela:2011:CEM

[1858] Mikko Salmela. Can emotion be modelled on perception? Dialectica: International Review of Philosophy of Knowledge, 65(1):1–29, March 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lemos:2011:WWT

[1859] John Lemos. Wanting, willing, trying and Kane’s theory of free will. Dialectica: International Review of Philosophy of Knowledge, 65(1):31–48, March 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kappel:2011:EED

[1860] Klemens Kappel. Is epistemic expressivism dialectically incoherent? Dialectica: International Review of Philosophy of Knowledge, 65(1):49–69, March 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hossack:2011:PMK

[1861] Keith Hossack. Précis of the metaphysics of knowledge. Dialectica: International Review of Philosophy of Knowledge, 65(1):71–73, March 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Textor:2011:KF

[1862] Mark Textor. Knowing the facts. Dialectica: International Review of Philosophy of Knowledge, 65(1):75–86, March 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

MacBride:2011:EMH

[1863] Fraser MacBride. Extreme metaphysics: Hossack on logical objects, facts, propositions and universals. Dialectica: International Review of Philosophy of Knowledge, 65(1):87–101, March 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Leech:2011:MR

[1864] Jessica Leech. Modal rationalism. Dialectica: International Review of Philosophy of Knowledge, 65(1):103–115, March 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Correia:2011:GRM

[1865] Fabrice Correia, Jessica Leech, and Mollie Molyneaux. Genevan ruminations on the metaphysics of knowledge. Dialectica: International Review of Philosophy of Knowledge, 65(1):117–123, March 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hossack:2011:RC

[1866] Keith Hossack. Replies to comments. Dialectica: International Review of Philosophy of Knowledge, 65(1):125–151, March 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Allen:2011:RNC

[1867] Keith Allen. Revelation and the nature of colour. Dialectica: International Review of Philosophy of Knowledge, 65(2):153–176, June 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kriegel:2011:TDC

[1868] Uriah Kriegel. Two defenses of common-sense ontology. Dialectica: International Review of Philosophy of Knowledge, 65(2):177–204, June 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Saka:2011:QCL

[1869] Paul Saka. Quotation and conceptions of language. Dialectica: International Review of Philosophy of Knowledge, 65(2):205–220, June 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hunter:2011:AB

[1870] David Hunter. Alienated belief. Dialectica: International Review of Philosophy of Knowledge, 65(2):221–240, June 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Shagrir:2011:SAC

[1871] Oron Shagrir. Supervenience and anomalism are compatible. Dialectica: International Review of Philosophy of Knowledge, 65(2):241–266, June 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Coffman:2011:CDC

[1872] E. J. Coffman. Clarke’s defense of the contrast argument. Dialectica: International Review of Philosophy of Knowledge, 65(2):267–275, June 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Suarez:2011:BRE

[1873] Juan Suarez. Book review: Esbozo de la Filosofía de Kripke — By Manuel Pérez Otero. Dialectica: International Review of Philosophy of Knowledge, 65(2):277–281, June 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lando:2011:BRF

[1874] Giorgio Lando. Book review: Fostering the Ontological Turn. Gustav Bergmann (1906–1987) — Edited by Rosaria Egidi, Guido Bonino. Dialectica: International Review of Philosophy of Knowledge, 65(2):281– 285, June 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vonWachter:2011:BRS

[1875] Daniel von Wachter. Book review: Saving God: Religion after Idolatry — By Mark Johnston. Dialectica: International Review of Philosophy of Knowledge, 65(2):286–292, June 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Plourde:2011:BSA

[1876] Jimmy Plourde. Book symposium: The Arrow and the Point. Russell and Wittgenstein’s — By Guido Bonino. Dialectica: International Review of Philosophy of Knowledge, 65(2):293–315, June 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bonino:2011:BSR

[1877] Guido Bonino. Book symposium: Reply to Plourde. Dialectica: International Review of Philosophy of Knowledge, 65(2):317–326, June 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Keefe:2011:PSP

[1878] Rosanna Keefe. Phenomenal Sorites paradoxes and looking the same. Dialectica: International Review of Philosophy of Knowledge, 65(3):327– 344, September 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Toribio:2011:WWD

[1879] Josefa Toribio. What we do when we judge. Dialectica: International Review of Philosophy of Knowledge, 65(3):345–367, September 2011. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mchugh:2011:WDW

[1880] Conor Mchugh. What do we aim at when we believe? Dialectica: International Review of Philosophy of Knowledge, 65(3):369–392, September 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Peacock:2011:TPA

[1881] Howard Peacock. Is there a problem about propositional unity? Dialectica: International Review of Philosophy of Knowledge, 65(3):393–418, September 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Keinanen:2011:TBC

[1882] Markku Keinänen. Tropes — the basic constituents of powerful particulars? Dialectica: International Review of Philosophy of Knowledge, 65 (3):419–450, September 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Effingham:2011:UC

[1883] Nikk Effingham. Universalism and classes. Dialectica: International Review of Philosophy of Knowledge, 65(3):451–472, September 2011. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Carrara:2011:BRM

[1884] Massimiliano Carrara and Giuseppe Spolaore. Book review: Il mondo messo a fuoco. Storie di allucinazioni e miopie filosofiche (The World in Focus. Stories of Philosophical Hallucinations and Myopias) — By Achille Varzi. Dialectica: International Review of Philosophy of Knowledge, 65(3):473–477, September 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Plebani:2011:BRL

[1885] Matteo Plebani. Book review: Logic and Philosophy of Mathematics in the Early Husserl — By Stefania Centrone. Dialectica: International Review of Philosophy of Knowledge, 65(3):477–482, September 2011. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Meylan:2011:IPB

[1886] Anne Meylan. Introduction: The providential bad luck of justification. Dialectica: International Review of Philosophy of Knowledge, 65(4):483– 491, December 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Brown:2011:TEI

[1887] Jessica Brown. Thought experiments, intuitions and philosophical evidence. Dialectica: International Review of Philosophy of Knowledge, 65 (4):493–516, December 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

DeCruz:2011:EAE

[1888] Helen De Cruz, Maarten Boudry, Johan De Smedt, and Stefaan Blancke. Evolutionary approaches to epistemic justification. Dialectica: International Review of Philosophy of Knowledge, 65(4):517–535, December 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Steup:2011:BVI

[1889] Matthias Steup. Belief, voluntariness and intentionality. Dialectica: International Review of Philosophy of Knowledge, 65(4):537–559, December 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chudnoff:2011:WST

[1890] Elijah Chudnoff. What should a theory of knowledge do? Dialectica: International Review of Philosophy of Knowledge, 65(4):561–579, December 2011. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kelp:2011:J

[1891] Christoph Kelp. Not without justification. Dialectica: International Review of Philosophy of Knowledge, 65(4):581–595, December 2011. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Engel:2012:BAP

[1892] Pascal Engel. Bad analytic philosophy. Dialectica: International Review of Philosophy of Knowledge, 66(1):1–4, March 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kallestrup:2012:ARD

[1893] Jesper Kallestrup. Actually-rigidified descriptivism revisited. Dialectica: International Review of Philosophy of Knowledge, 66(1):5–21, March 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lowenstein:2012:DSA

[1894] David Löwenstein. Davidsonian semantics and anaphoric deflationism. Dialectica: International Review of Philosophy of Knowledge, 66(1):23– 44, March 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lemaire:2012:GBA

[1895] Stéphane Lemaire. A gate-based account of intentions. Dialectica: International Review of Philosophy of Knowledge, 66(1):45–67, March 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Damnjanovic:2012:RP

[1896] Nic Damnjanovic. Revelation and physicalism. Dialectica: International Review of Philosophy of Knowledge, 66(1):69–91, March 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hindriks:2012:WU

[1897] Frank Hindriks. But where is the University? Dialectica: International Review of Philosophy of Knowledge, 66(1):93–113, March 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fink:2012:FNP

[1898] Julian Fink. The function of normative process-requirements. Dialectica: International Review of Philosophy of Knowledge, 66(1):115–136, March 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Coleman:2012:MCC

[1899] Sam Coleman. Mental chemistry: Combination for panpsychists. Dialectica: International Review of Philosophy of Knowledge, 66(1):137–166, March 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gaskin:2012:NR

[1900] Richard Gaskin and Daniel J. Hill. On neutral relations. Dialectica: International Review of Philosophy of Knowledge, 66(1):167–186, March 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Luzzi:2012:NDC

[1901] Federico Luzzi. Notes and discussions: Contextualism and counterclosure. Dialectica: International Review of Philosophy of Knowledge, 66(1):187–199, March 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Edwards:2012:NDA

[1902] Douglas Edwards. Notes and discussions: On alethic disjunctivism. Dialectica: International Review of Philosophy of Knowledge, 66(1):200– 214, March 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Paganini:2012:IPC

[1903] Elisa Paganini. Introduction: Paolo Casalegno’s good points. Dialectica: International Review of Philosophy of Knowledge, 66(2):215–219, June 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Boghossian:2012:IEL

[1904] Paul Boghossian. Inferentialism and the epistemology of logic: Reflections on Casalegno and Williamson. Dialectica: International Review of Philosophy of Knowledge, 66(2):221–236, June 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Williamson:2012:BCU

[1905] Timothy Williamson. Boghossian and Casalegno on understanding and inference. Dialectica: International Review of Philosophy of Knowledge, 66(2):237–247, June 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wright:2012:MAS

[1906] Crispin Wright. Meaning and assertibility: Some reflections on Paolo Casalegno’s ‘The Problem of Non-conclusiveness’. Dialectica: International Review of Philosophy of Knowledge, 66(2):249–266, June 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Williamson:2012:WCM

[1907] Timothy Williamson. Wright and Casalegno on meaning and assertibility. Dialectica: International Review of Philosophy of Knowledge, 66(2): 267–271, June 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marconi:2012:SND

[1908] Diego Marconi. Semantic normativity, deference and reference. Dialectica: International Review of Philosophy of Knowledge, 66(2):273–287, June 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Douven:2012:DRC

[1909] Igor Douven and Christoph Kelp. In defense of the rational credibility account: A reply to Casalegno. Dialectica: International Review of Philosophy of Knowledge, 66(2):289–297, June 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Iacona:2012:BRB

[1910] Andrea Iacona. Book review: Brevissima introduzione alla filosofia del linguaggio — By Paolo Casalegno. Dialectica: International Review of Philosophy of Knowledge, 66(2):299–300, June 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bianchi:2012:BRV

[1911] Andrea Bianchi and Marco Santambrogio. Book review: Verità e significato: Scritti di filosofia del linguaggio — By Paolo Casalegno. Dialectica: International Review of Philosophy of Knowledge, 66(2):300–303, June 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ebert:2012:IOB

[1912] Philip A. Ebert and Martin Smith. Introduction: Outright belief and degrees of belief. Dialectica: International Review of Philosophy of Knowledge, 66(3):305–308, September 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wedgwood:2012:OB

[1913] Ralph Wedgwood. Outright belief. Dialectica: International Review of Philosophy of Knowledge, 66(3):309–329, September 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Milne:2012:BDB

[1914] Peter Milne. Belief, degrees of belief, and assertion. Dialectica: International Review of Philosophy of Knowledge, 66(3):331–349, September 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Douven:2012:LPP

[1915] Igor Douven. The lottery paradox and the pragmatics of belief. Dialectica: International Review of Philosophy of Knowledge, 66(3):351–373, September 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zardini:2012:LV

[1916] Elia Zardini. Luminosity and vagueness. Dialectica: International Review of Philosophy of Knowledge, 66(3):375–410, September 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hajek:2012:SCC

[1917] Alan Hájek. Is strict coherence coherent? Dialectica: International Review of Philosophy of Knowledge, 66(3):411–424, September 2012. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Easwaran:2012:EWA

[1918] Kenny Easwaran and Branden Fitelson. An ‘evidentialist’ worry about Joyce’s argument for probabilism. Dialectica: International Review of Philosophy of Knowledge, 66(3):425–433, September 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hoefer:2012:CBT

[1919] Carl Hoefer. Calibration: Being in tune with frequencies. Dialectica: International Review of Philosophy of Knowledge, 66(3):435–452, September 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Podlaskowski:2012:STA

[1920] Adam C. Podlaskowski. Simple tasks, abstractions, and semantic dispositionalism. Dialectica: International Review of Philosophy of Knowledge, 66(4):453–470, December 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Frapolli:2012:ME

[1921] María José Frápolli and Neftalí Villanueva. Minimal expressivism. Dialectica: International Review of Philosophy of Knowledge, 66(4):471– 487, December 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lerman:2012:DCE

[1922] Hemdat Lerman. Demonstrative content and the experience of properties. Dialectica: International Review of Philosophy of Knowledge, 66 (4):489–515, December 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Poston:2012:TE

[1923] Ted Poston. Is there an ‘i’ in epistemology? Dialectica: International Review of Philosophy of Knowledge, 66(4):517–541, December 2012. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

vanRiel:2012:PPP

[1924] Raphael van Riel. Pains, pills and properties — functionalism and the First-Order/Second-Order distinction. Dialectica: International Review of Philosophy of Knowledge, 66(4):543–562, December 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fallis:2012:LVG

[1925] Don Fallis. Lying as a violation of Grice’s first Maxim of quality. Dialectica: International Review of Philosophy of Knowledge, 66(4):563–581, December 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Montminy:2012:EMI

[1926] Martin Montminy. Epistemic modals and indirect weak suggestives. Dialectica: International Review of Philosophy of Knowledge, 66(4):583– 606, December 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Levy:2012:CCR

[1927] Neil Levy. Capacities and counterfactuals: A reply to Haji and McKenna. Dialectica: International Review of Philosophy of Knowledge, 66(4):607– 620, December 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vignolo:2012:RAE

[1928] Massimiliano Vignolo. Referential/Attributive: The explanatory gap of the contextualist theory. Dialectica: International Review of Philosophy of Knowledge, 66(4):621–633, December 2012. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Baumgartner:2013:RIN

[1929] Michael Baumgartner. Rendering interventionism and non-reductive physicalism compatible. Dialectica: International Review of Philosophy of Knowledge, 67(1):1–27, March 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Almang:2013:CSR

[1930] Jan Almäng. The causal self-referential theory of perception revisited. Dialectica: International Review of Philosophy of Knowledge, 67(1):29– 53, March 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ingthorsson:2013:PQP

[1931] R. D. Ingthorsson. Properties: Qualities, powers, or both? Dialectica: International Review of Philosophy of Knowledge, 67(1):55–80, March 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ben-Yami:2013:HLP

[1932] Hanoch Ben-Yami. Higher-level plurals versus articulated reference, and an elaboration of Salva Veritate. Dialectica: International Review of Philosophy of Knowledge, 67(1):81–102, March 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bates:2013:DSP

[1933] Jared Bates. “Damming the swamping problem, reliably”. Dialectica: International Review of Philosophy of Knowledge, 67(1):103–116, March 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mehta:2013:HEE

[1934] Neil Mehta. How to explain the explanatory gap. Dialectica: International Review of Philosophy of Knowledge, 67(2):117–135, June 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Liebesman:2013:CI

[1935] David Liebesman. Converse and identity. Dialectica: International Review of Philosophy of Knowledge, 67(2):137–155, June 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Feldman:2013:ASK

[1936] Simon D. Feldman and Allan Hazlett. Authenticity and self-knowledge. Dialectica: International Review of Philosophy of Knowledge, 67(2):157– 181, June 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Divers:2013:APE

[1937] John Divers. The analysis of possibility and the extent of possibility. Dialectica: International Review of Philosophy of Knowledge, 67(2):183– 200, June 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gaskin:2013:RPM

[1938] Richard Gaskin and Daniel J. Hill. Reach’s puzzle and mention. Dialectica: International Review of Philosophy of Knowledge, 67(2):201–222, June 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cohnitz:2013:USD

[1939] Daniel Cohnitz and Jaan Kangilaski. Understanding a sentence does not entail knowing its truth-conditions: Why the epistemological determination argument fails. Dialectica: International Review of Philosophy of Knowledge, 67(2):223–242, June 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Clapp:2013:BRA

[1940] Lenny Clapp. Book review: Asking Questions: Using Meaningful Structures to Imply Ignorance — By Robert Fiengo. Dialectica: International Review of Philosophy of Knowledge, 67(2):243–247, June 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Brand:2013:BRM

[1941] Cordula Brand. Book review: Mental Representation and Self-Consciousness. From Basic Self-Representation to Self-Related Cognition — By Gottfried Vosgerau. Dialectica: International Review of Philosophy of Knowledge, 67(2):248–252, June 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wright:2013:BRE

[1942] Stephen Wright. Book review: Epistemological Disjunctivism — By Duncan Pritchard. Dialectica: International Review of Philosophy of Knowledge, 67(2):252–257, June 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rossi:2013:BRE

[1943] Mauro Rossi. Book review: Ethical Naturalism. Current Debates — Edited by Susana Nuccetelli and Gary Seay. Dialectica: International Review of Philosophy of Knowledge, 67(2):257–264, June 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Archer:2013:NAS

[1944] Sophie Archer. Nondoxasticism about self-deception. Dialectica: International Review of Philosophy of Knowledge, 67(3):265–282, September 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Salis:2013:FNP

[1945] Fiora Salis. Fictional names and the problem of intersubjective identification. Dialectica: International Review of Philosophy of Knowledge, 67(3):283–301, September 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Morvarid:2013:RFI

[1946] Mahmoud Morvarid. Reference failure, illusion of thought and selfknowledge. Dialectica: International Review of Philosophy of Knowledge, 67(3):303–323, September 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Frederick:2013:STP

[1947] Danny Frederick. Singular terms, predicates and the spurious ‘is’ of identity. Dialectica: International Review of Philosophy of Knowledge, 67(3):325–343, September 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zhang:2013:NDC

[1948] Jiji Zhang. Notes and discussions: Can the incompatibilist get past the no past objection? Dialectica: International Review of Philosophy of Knowledge, 67(3):345–352, September 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Parent:2013:NDI

[1949] T. Parent. Notes and discussions: Infallibility naturalized: Reply to Hoffmann. Dialectica: International Review of Philosophy of Knowledge, 67(3):353–358, September 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Seel:2013:BRN

[1950] Gerhard Seel. Book review: Normenlogik. Grundlagen-Systeme-Anwendungen — By Edgar Morscher. Dialectica: International Review of Philosophy of Knowledge, 67(3):359–367, September 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Egg:2013:BRE

[1951] Matthias Egg. Book review: Explaining Science’s Success. Understanding How Scientific Knowledge Works — By John Wright. Dialectica: International Review of Philosophy of Knowledge, 67(3):367–372, September 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Reboul:2013:BRC

[1952] Anne Reboul. Book review: Consciousness in Interaction. The Role of the Natural and Social Context in Shaping Consciousness — Edited by Fabio Paglieri. Dialectica: International Review of Philosophy of Knowledge, 67(3):372–378, September 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tegtmeier:2013:BRA

[1953] Erwin Tegtmeier. Book review: Abstrakte Gegenstände. Semantik und Ontologie. 2., um einen Anhang erweiterte Auflage — By Wolfgang Künne. Dialectica: International Review of Philosophy of Knowledge, 67(3):378–382, September 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Moller:2013:DRO

[1954] Niklas Möller. Direct reference and the open question argument. Dialectica: International Review of Philosophy of Knowledge, 67(4):383–402, December 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Aakerman:2013:FMS

[1955] Jonas Åkerman. Forced-march Sorites arguments and linguistic competence. Dialectica: International Review of Philosophy of Knowledge, 67 (4):403–426, December 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Martinez:2013:TI

[1956] Manolo Martínez. Teleosemantics and indeterminacy. Dialectica: International Review of Philosophy of Knowledge, 67(4):427–453, December 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Frigerio:2013:RIC

[1957] Aldo Frigerio, Alessandro Giordani, and Luca Mari. On representing information: a characterization of the analog/digital distinction. Dialectica: International Review of Philosophy of Knowledge, 67(4):455–483, December 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Shani:2013:TIS

[1958] Itay Shani and Sungho Choi. Type-identity statements and the explanatory gap: An argument for compatibility. Dialectica: International Review of Philosophy of Knowledge, 67(4):485–502, December 2013. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Asay:2013:PT

[1959] Jamin Asay. Primitive truth. Dialectica: International Review of Philosophy of Knowledge, 67(4):503–519, December 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Strollo:2013:DIP

[1960] Andrea Strollo. Deflationism and the invisible power of truth. Dialectica: International Review of Philosophy of Knowledge, 67(4):521–543, December 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ludwig:2013:NWP

[1961] David Ludwig. New wave pluralism. Dialectica: International Review of Philosophy of Knowledge, 67(4):545–560, December 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schroer:2013:DPS

[1962] Robert Schroer. Do the primary and secondary intensions of phenomenal concepts coincide in all worlds? Dialectica: International Review of Philosophy of Knowledge, 67(4):561–577, December 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pettigrew:2013:AE

[1963] Richard Pettigrew. Accuracy and evidence. Dialectica: International Review of Philosophy of Knowledge, 67(4):579–596, December 2013. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Peterson:2013:GPP

[1964] Martin Peterson. A generalization of the Pasadena puzzle. Dialectica: International Review of Philosophy of Knowledge, 67(4):597–603, December 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Daly:2013:AD

[1965] Chris Daly and David Liggins. Animalism and deferentialism. Dialectica: International Review of Philosophy of Knowledge, 67(4):605–609, December 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bottani:2013:BRP

[1966] Andrea Bottani. Book review: Personal Identity: Complex or Simple? — Edited by Georg Gasser and Matthias Stefan. Dialectica: International Review of Philosophy of Knowledge, 67(4):611–617, December 2013. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Weber:2013:BRP

[1967] Marc-André Weber. Book review: On Privacy — By Annabelle Lever. Dialectica: International Review of Philosophy of Knowledge, 67(4):618– 621, December 2013. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schiaparelli:2013:BRA

[1968] Annamaria Schiaparelli. Book review: Aristotelian Ethics in Contemporary Perspective — Edited by Julia Peters. Dialectica: International Review of Philosophy of Knowledge, 67(4):621–625, December 2013. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stokes:2014:CPP

[1969] Dustin Stokes. Cognitive penetration and the perception of art: Winner of the dialectica essay competition on cognitive penetration. Dialectica: International Review of Philosophy of Knowledge, 68(1):1–34, March 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tucker:2014:IDP

[1970] Chris Tucker. If dogmatists have a problem with cognitive penetration, you do too. Dialectica: International Review of Philosophy of Knowledge, 68(1):35–62, March 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cecchi:2014:CPP

[1971] Ariel S. Cecchi. Cognitive penetration, perceptual learning and neural plasticity. Dialectica: International Review of Philosophy of Knowledge, 68(1):63–95, March 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fazekas:2014:PNP

[1972] Peter Fazekas. Pursuing natural piety: Understanding ontological emergence and distinguishing it from physicalism. Dialectica: International Review of Philosophy of Knowledge, 68(1):97–119, March 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Calosi:2014:EMP

[1973] Claudio Calosi. Extensionality, multilocation, persistence. Dialectica: International Review of Philosophy of Knowledge, 68(1):121–139, March 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bortolotti:2014:BRN

[1974] Lisa Bortolotti and Ema Sullivan-Bissett. Book review: New Essays on Belief: Constitution, Content and Structure — Edited by Nikolaj Nottelmann. Dialectica: International Review of Philosophy of Knowledge, 68(1):141–146, March 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Leuenberger:2014:BRM

[1975] Stephan Leuenberger. Book review: Metaphysical Grounding. Understanding the Structure of Reality — Edited by Fabrice Correia and Benjamin Schnieder. Dialectica: International Review of Philosophy of Knowledge, 68(1):147–151, March 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Garcia-Carpintero:2014:BRM

[1976] Manuel García-Carpintero. Book review: Morality, Self-Knowledge and Human Suffering: An Essay on the Loss of Confidence in the World — By Josep Corbí. Dialectica: International Review of Philosophy of Knowledge, 68(1):151–161, March 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

MacBride:2014:IHH

[1977] Fraser MacBride. Introduction: How Hochberg helped us take the ontological turn: An introduction. Dialectica: International Review of Philosophy of Knowledge, 68(2):163–169, June 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hochberg:2014:STR

[1978] Herbert Hochberg. Some things recalled. Dialectica: International Review of Philosophy of Knowledge, 68(2):171–182, June 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mantegani:2014:DIQ

[1979] Nicholas Mantegani. Difference, identity and quantification. Dialectica: International Review of Philosophy of Knowledge, 68(2):183–207, June 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pickel:2014:CPM

[1980] Bryan Pickel. Complex predication and the metaphysics of properties. Dialectica: International Review of Philosophy of Knowledge, 68(2):209– 230, June 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Levine:2014:RPR

[1981] James Levine. Russell, particularized relations and Bradley’s dilemma. Dialectica: International Review of Philosophy of Knowledge, 68(2):231– 261, June 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Leo:2014:TCF

[1982] Joop Leo. Thinking in a coordinate-free way about relations. Dialectica: International Review of Philosophy of Knowledge, 68(2):263–282, June 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Orilia:2014:POR

[1983] Francesco Orilia. Positions, ordering relations and O-roles. Dialectica: International Review of Philosophy of Knowledge, 68(2):283–303, June 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Simons:2014:RIS

[1984] Peter Simons. Relations and idealism: On some arguments of Hochberg against trope nominalism. Dialectica: International Review of Philosophy of Knowledge, 68(2):305–315, June 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fine:2014:PPW

[1985] Kit Fine. Permission and possible worlds. Dialectica: International Review of Philosophy of Knowledge, 68(3):317–336, September 2014. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cray:2014:IC

[1986] Wesley D. Cray. Inconstancy and content. Dialectica: International Review of Philosophy of Knowledge, 68(3):337–353, September 2014. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wallace:2014:AVM

[1987] Meg Wallace. The argument from vagueness for modal parts. Dialectica: International Review of Philosophy of Knowledge, 68(3):355–373, September 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Backmann:2014:BBS

[1988] Marius Backmann and Alexander Reutlinger. Better best systems — too good to be true. Dialectica: International Review of Philosophy of Knowledge, 68(3):375–390, September 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sander:2014:FGS

[1989] Thorsten Sander. A Frege-geach style objection to cognitivist judgment internalism. Dialectica: International Review of Philosophy of Knowledge, 68(3):391–408, September 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Harth:2014:RTR

[1990] Manfred Harth. Is relative truth really truth? Dialectica: International Review of Philosophy of Knowledge, 68(3):409–428, September 2014. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bernstein:2014:NDT

[1991] Sara Bernstein. Notes and discussions: Two problems for proportionality about omissions. Dialectica: International Review of Philosophy of Knowledge, 68(3):429–441, September 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cantwell:2014:NDU

[1992] John Cantwell. Notes and discussions: Unity and autonomy in expressivist logic. Dialectica: International Review of Philosophy of Knowledge, 68(3):443–457, September 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Doring:2014:BRE

[1993] Sabine A. Döring and Anika Lutz. Book review: The Emotions. A Philosophical Introduction — By Julien A. Deonna and Fabrice Teroni. Dialectica: International Review of Philosophy of Knowledge, 68(3):459– 463, September 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ludwig:2014:BRD

[1994] Kirk Ludwig. Book review: Donald Davidson: A Short Introduction — By Kathrin Glüer. Dialectica: International Review of Philosophy of Knowledge, 68(3):464–473, September 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Deroy:2014:BRW

[1995] Ophelia Deroy. Book review: Why Red Doesn’t Sound Like a Bell. Understanding the Feel of Consciousness — By Kevin O’Regan. Dialectica: International Review of Philosophy of Knowledge, 68(3):473–476, September 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gaultier:2014:ASE

[1996] Benoit Gaultier. Achievements, safety and environmental epistemic luck. Dialectica: International Review of Philosophy of Knowledge, 68(4):477– 497, December 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Davies:2014:TRO

[1997] Alex Davies. Off-target responses to occasion-sensitivity. Dialectica: International Review of Philosophy of Knowledge, 68(4):499–523, December 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Wahlberg:2014:IOR

[1998] Tobias Hansson Wahlberg. Institutional objects, reductionism and theories of persistence. Dialectica: International Review of Philosophy of Knowledge, 68(4):525–562, December 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sweeney:2014:UCR

[1999] Paula Sweeney. The utility of content-relativism. Dialectica: International Review of Philosophy of Knowledge, 68(4):563–579, December 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Felappi:2014:DS

[2000] Giulia Felappi. ‘In defence of sententialism’. Dialectica: International Review of Philosophy of Knowledge, 68(4):581–603, December 2014. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Baierle:2014:BRU

[2001] Emmanuel Baierlé. Book review: Understanding Human Agency — By Erasmus Mayr. Dialectica: International Review of Philosophy of Knowledge, 68(4):605–609, December 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ronnow-Rasmussen:2014:BRD

[2002] Toni Rønnow-Rasmussen. Book review: In Defense of Shame: The Faces of an Emotion — By Julien A. Deonna, Raffaele Rodogno and Fabrice Teroni. Dialectica: International Review of Philosophy of Knowledge, 68 (4):609–614, December 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cross:2014:BRG

[2003] Troy Cross. Book review: Getting Causes from Powers — By Stephen Mumford and Rani Lill Anjum. Dialectica: International Review of Philosophy of Knowledge, 68(4):614–619, December 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bader:2014:BRD

[2004] Ralf M. Bader. Book review: The Dimensions of Consequentialism: Ethics, Equality and Risk — By M. Peterson. Dialectica: International Review of Philosophy of Knowledge, 68(4):620–625, December 2014. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See reply [2005].

Peterson:2014:RRB

[2005] Martin Peterson. Reply to Ralf Bader. Dialectica: International Review of Philosophy of Knowledge, 68(4):625–629, December 2014. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See [2004].

Szigeti:2015:SMD

[2006] András Szigeti. Sentimentalism and moral dilemmas. Dialectica: International Review of Philosophy of Knowledge, 69(1):1–22, March 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Baave:2015:DET

[2007] Arvid Båve. A deflationist error theory of properties. Dialectica: International Review of Philosophy of Knowledge, 69(1):23–59, March 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cieslinski:2015:IT

[2008] Cezary Cieśliński. The innocence of truth. Dialectica: International Review of Philosophy of Knowledge, 69(1):61–85, March 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Varga:2015:IVC

[2009] Somogy Varga. Identifications, volitions and the case of successful psychopaths. Dialectica: International Review of Philosophy of Knowledge, 69(1):87–106, March 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Borgoni:2015:DMP

[2010] Cristina Borgoni. Dissonance and Moorean propositions. Dialectica: International Review of Philosophy of Knowledge, 69(1):107–127, March 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Nasim:2015:BRG

[2011] Omar W. Nasim. Book review: G. F. Stout and the Psychological Origins of Analytic Philosophy — By Maria van der Schaar. Dialectica: International Review of Philosophy of Knowledge, 69(1):129–134, March 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dokic:2015:BRP

[2012] Jérôme Dokic. Book review: Perception, Realism and the Problem of Reference — Edited by Athanasios Raftopoulos and Peter Machame. Dialectica: International Review of Philosophy of Knowledge, 69(1):134– 138, March 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Correia:2015:BRC

[2013] Fabrice Correia. Book review: Categories of Being: Essays on Metaphysics and Logic — Edited by L. Haaparanta and H. J. Koskinen. Dialectica: International Review of Philosophy of Knowledge, 69(1):138– 143, March 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

deSaussure:2015:BRT

[2014] Louis de Saussure. Book review: Telicity, Change, and State: A Cross-Categorial View of Event Structure — Edited by Violeta Demonte and Louise McNally. Dialectica: International Review of Philosophy of Knowledge, 69(1):143–148, March 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Merlo:2015:BRN

[2015] Giovanni Merlo. Book review: New Papers on the Present: Focus on Presentism — Edited by Roberto Ciuni, Kristie Miller and Giuliano Torrengo. Dialectica: International Review of Philosophy of Knowledge, 69 (1):148–153, March 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zeman:2015:RBP

[2016] Dan Zeman. Relativism and bound predicates of personal taste: An answer to Schaffer’s argument from binding. Dialectica: International Review of Philosophy of Knowledge, 69(2):155–183, June 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vlerick:2015:EEJ

[2017] Michael Vlerick and Alex Broadbent. Evolution and epistemic justification. Dialectica: International Review of Philosophy of Knowledge, 69(2): 185–203, June 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Andow:2015:NDE

[2018] James Andow. Notes and discussions: Expecting moral philosophers to be reliable. Dialectica: International Review of Philosophy of Knowledge, 69(2):205–220, June 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Massin:2015:BRP

[2019] Olivier Massin. Book review: Personal Value — By Rønnow-Rasmussen. Dialectica: International Review of Philosophy of Knowledge, 69(2):221– 231, June 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tassone:2015:BRT

[2020] Biagio G. Tassone. Book review: Themes From Brentano, Studien zur Österreichischen Philosophie — Edited by Denis Fisette and Guillaume Fréchette. Dialectica: International Review of Philosophy of Knowledge, 69(2):231–248, June 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Raz:2015:BRO

[2021] Tim Räz. Book review: The Outer Limits of Reason. What Science, Mathematics, and Logic Cannot Tell Us — By Noson S. Yanofsky. Dialectica: International Review of Philosophy of Knowledge, 69(2):248– 254, June 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Benovsky:2015:BRR

[2022] Jiri Benovsky. Book review: Resemblance and Representation: An Essay in the Philosophy of Pictures — By Ben Blumson. Dialectica: International Review of Philosophy of Knowledge, 69(2):254–258, June 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Doring:2015:IBP

[2023] Sabine A. Döring and Anika Lutz. Introduction: Beyond perceptualism: Introduction to the special issue. Dialectica: International Review of Philosophy of Knowledge, 69(3):259–270, September 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dokic:2015:EEM

[2024] Jérôme Dokic and Stéphane Lemaire. Are emotions evaluative modes? Dialectica: International Review of Philosophy of Knowledge, 69(3):271– 292, September 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Deonna:2015:EA

[2025] Julien A. Deonna and Fabrice Teroni. Emotions as attitudes. Dialectica: International Review of Philosophy of Knowledge, 69(3):293–311, September 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lutz:2015:PCE

[2026] Anika Lutz. The phenomenal character of emotional experience: A look at perception theory. Dialectica: International Review of Philosophy of Knowledge, 69(3):313–334, September 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lacewing:2015:EPS

[2027] Michael Lacewing. Emotion, perception, and the self in moral epistemology. Dialectica: International Review of Philosophy of Knowledge, 69(3):335–355, September 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schroeter:2015:DER

[2028] Laura Schroeter, François Schroeter, and Karen Jones. Do emotions represent values? Dialectica: International Review of Philosophy of Knowledge, 69(3):357–380, September 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Doring:2015:WWR

[2029] Sabine A. Döring. What’s wrong with recalcitrant emotions? From irrationality to challenge of agential identity. Dialectica: International Review of Philosophy of Knowledge, 69(3):381–402, September 2015. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Brady:2015:FBS

[2030] Michael S. Brady. Feeling bad and seeing bad. Dialectica: International Review of Philosophy of Knowledge, 69(3):403–416, September 2015. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Helm:2015:ERR

[2031] Bennett W. Helm. Emotions and recalcitrance: Reevaluating the perceptual model. Dialectica: International Review of Philosophy of Knowledge, 69(3):417–433, September 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lange:2015:EEN

[2032] Marc Lange. Explanation, existence and natural properties in mathematics — a case study: Desargues’ Theorem (winner of the 2014 Dialectica Prize). Dialectica: International Review of Philosophy of Knowledge, 69 (4):435–472, December 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hofweber:2015:PSM

[2033] Thomas Hofweber. The place of subjects in the metaphysics of material objects. Dialectica: International Review of Philosophy of Knowledge, 69(4):473–490, December 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pineda:2015:MAA

[2034] David Pineda. The modal argument against materialism and intertheoretic identities. Dialectica: International Review of Philosophy of Knowledge, 69(4):491–515, December 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Clapp:2015:NAA

[2035] Lenny Clapp. A non-alethic approach to faultless disagreement. Dialectica: International Review of Philosophy of Knowledge, 69(4):517–550, December 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rosenkranz:2015:PDC

[2036] Sven Rosenkranz and Moritz Schulz. Peer disagreement: A call for the revision of prior probabilities. Dialectica: International Review of Philosophy of Knowledge, 69(4):551–586, December 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Novaes:2015:DMA

[2037] Catarina Dutilh Novaes. A dialogical, multi-agent account of the normativity of logic. Dialectica: International Review of Philosophy of Knowledge, 69(4):587–609, December 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Voltolini:2015:BRN

[2038] Alberto Voltolini. Book review: The Nonexistent — By Anthony Everett. Dialectica: International Review of Philosophy of Knowledge, 69(4):611– 620, December 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dutant:2015:BRT

[2039] Julien Dutant. Book review: Tacit Knowledge — By Neil Gascoigne and Tim Thornton. Dialectica: International Review of Philosophy of Knowledge, 69(4):621–623, December 2015. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kail:2015:BRH

[2040] P. J. E. Kail. Book review: Hume’s Radical Scepticism and the Fate of Naturalized Epistemology — By Kevin Meeker. Dialectica: International Review of Philosophy of Knowledge, 69(4):623–630, December 2015. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2016:IIT

[2041] Anonymous. Issue information — TOC. Dialectica: International Review of Philosophy of Knowledge, 70(1):1, March 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Onofri:2016:TCT

[2042] Andrea Onofri. Two constraints on a theory of concepts. Dialectica: International Review of Philosophy of Knowledge, 70(1):3–27, March 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Belleri:2016:USE

[2043] Delia Belleri. The underdeterminacy of sentences and the expressibility of our thoughts. Dialectica: International Review of Philosophy of Knowledge, 70(1):29–48, March 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Susnik:2016:DIM

[2044] Matej Sušnik. Does it matter who is driving the trolley? Dialectica: International Review of Philosophy of Knowledge, 70(1):49–63, March 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Baker:2016:IAD

[2045] Derek Baker. Intuitions about disagreement do not support the normativity of meaning. Dialectica: International Review of Philosophy of Knowledge, 70(1):65–84, March 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Goldschmidt:2016:DCP

[2046] Tyron Goldschmidt. A demonstration of the causal power of absences. Dialectica: International Review of Philosophy of Knowledge, 70(1):85, March 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

McCarthy:2016:GTI

[2047] Timothy McCarthy. Gödel’s Third Incompleteness Theorem. Dialectica: International Review of Philosophy of Knowledge, 70(1):87–112, March 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lynch:2016:BRS

[2048] Kevin Lynch. Book review: Self-Knowledge for Humans — By Quassim Cassam. Dialectica: International Review of Philosophy of Knowledge, 70(1):113–119, March 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Ferguson:2016:BRL

[2049] Thomas Macaulay Ferguson. Book review: Logics and Falsifications. A New Perspective on Constructivist Semantics — By Andreas Kapsner. Dialectica: International Review of Philosophy of Knowledge, 70(1):119– 124, March 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Seron:2016:BRW

[2050] Denis Seron. Book review: Wittgenstein et la philosophie austroallemande — By K. Mulligan. (French) [Book review: booktitle Wittgenstein and Austro-German philosophy — By K. Mulligan]. Dialectica: International Review of Philosophy of Knowledge, 70(1):125–128, March 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Butterfield:2016:BRM

[2051] Paul Butterfield. Book review: Moral Failure: On The Impossible Demands of Morality — By Lisa Tessman. Dialectica: International Review of Philosophy of Knowledge, 70(1):129–136, March 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2016:IIa

[2052] Anonymous. Issue information. Dialectica: International Review of Philosophy of Knowledge, 70(2):137, June 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Barz:2016:IAP

[2053] Wolfgang Barz. Infallibility, acquaintance, and phenomenal concepts. Dialectica: International Review of Philosophy of Knowledge, 70(2):139– 168, June 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Otero:2016:HIC

[2054] Manuel Pérez Otero. How indeterminism could help incompatibilism on free action. Dialectica: International Review of Philosophy of Knowledge, 70(2):169–184, June 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Freitag:2016:DRG

[2055] Wolfgang Freitag. The disjunctive riddle and the grue-paradox. Dialectica: International Review of Philosophy of Knowledge, 70(2):185–200, June 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hoeltje:2016:NDM

[2056] Miguel Hoeltje. Notes and discussions: ‘meaning and truth’ and ‘truth and meaning’. Dialectica: International Review of Philosophy of Knowledge, 70(2):201–215, June 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sgaravatti:2016:NDK

[2057] Daniele Sgaravatti. Notes and discussions: Is knowledge of essence required for thinking about something? Dialectica: International Review of Philosophy of Knowledge, 70(2):217–228, June 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Becker:2016:NDR

[2058] Sebastian Becker. Notes and discussions: Resolving Turri’s puzzle about withholding. Dialectica: International Review of Philosophy of Knowledge, 70(2):229–243, June 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Voltolini:2016:BRO

[2059] Alberto Voltolini. Book review: The Objects of Thought — By Tim Crane. Dialectica: International Review of Philosophy of Knowledge, 70 (2):245–252, June 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Scarpati:2016:BRA

[2060] Maria Scarpati. Book review: Amore e altri inganni — By José A. Diez and Andrea Iacona. Dialectica: International Review of Philosophy of Knowledge, 70(2):252–256, June 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lacewing:2016:BRT

[2061] Michael Lacewing. Book review: True Emotions — By Mikko Salmela. Dialectica: International Review of Philosophy of Knowledge, 70(2):257– 265, June 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2016:IIb

[2062] Anonymous. Issue information. Dialectica: International Review of Philosophy of Knowledge, 70(3):267, September 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pashby:2016:HDT

[2063] Thomas Pashby. How do things persist? Location relations in physics and the metaphysics of persistence. Dialectica: International Review of Philosophy of Knowledge, 70(3):269–309, September 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Merlo:2016:SM

[2064] Giovanni Merlo. Subjectivism and the mental. Dialectica: International Review of Philosophy of Knowledge, 70(3):311–342, September 2016. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Green:2016:CBR

[2065] Sara Green and Nicholaos Jones. Constraint-based reasoning for search and explanation: Strategies for understanding variation and patterns in biology. Dialectica: International Review of Philosophy of Knowledge, 70(3):343–374, September 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marcus:2016:BKY

[2066] Eric Marcus. To believe is to know that you believe. Dialectica: International Review of Philosophy of Knowledge, 70(3):375–405, September 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Artiga:2016:LRD

[2067] Marc Artiga. Liberal representationalism: A deflationist defense. Dialectica: International Review of Philosophy of Knowledge, 70(3):407– 430, September 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See reply [2190].

Koziolek:2016:EIC

[2068] Nicholas Koziolek. Extensionality, indirect contexts and Frege’s hierarchy. Dialectica: International Review of Philosophy of Knowledge, 70 (3):431–462, September 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Duguid:2016:NDG

[2069] Callum Duguid. Notes and discussions: Graph theory and the identity of indiscernibles. Dialectica: International Review of Philosophy of Knowledge, 70(3):463–474, September 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Simkulet:2016:BRA

[2070] William Simkulet. Book review: Art and Ethics in a Material World: Kant’s Pragmatist Legacy — By Jennifer A. McMahon. Dialectica: International Review of Philosophy of Knowledge, 70(3):475–477, September 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rodogno:2016:BRV

[2071] Raffaele Rodogno. Book review: The View from Here: On Affirmation, Attachment, and the Limits of Regret — By R. Jay Wallace. Dialectica: International Review of Philosophy of Knowledge, 70(3):477–482, September 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Klenk:2016:BRR

[2072] Michael Klenk. Book review: Robust Ethics. The Metaphysics and Epistemology of Godless Normative Realism — By Eric Wielenberg. Dialectica: International Review of Philosophy of Knowledge, 70(3):482–488, September 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2016:IIc

[2073] Anonymous. Issue information. Dialectica: International Review of Philosophy of Knowledge, 70(4):489, December 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hansson:2016:SDG

[2074] Sven Ove Hansson, Karin Edvardsson Björnberg, and John Cantwell. Self-defeating goals. Dialectica: International Review of Philosophy of Knowledge, 70(4):491–512, December 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Paganini:2016:IIC

[2075] Elisa Paganini. Informative identities: A challenge for Frege’s puzzle. Dialectica: International Review of Philosophy of Knowledge, 70(4):513– 530, December 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jespersen:2016:LSH

[2076] Bjørn Jespersen. Left subsectivity: How to infer that a round peg is round. Dialectica: International Review of Philosophy of Knowledge, 70 (4):531–547, December 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Felline:2016:MPS

[2077] Laura Felline. It’s a matter of principle: Scientific explanation in information-theoretic reconstructions of quantum theory. Dialectica: International Review of Philosophy of Knowledge, 70(4):549–575, December 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schaffer:2016:IBL

[2078] Jonathan Schaffer. It is the business of laws to govern. Dialectica: International Review of Philosophy of Knowledge, 70(4):577–588, December 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rosa:2016:RDD

[2079] Luis Rosa. Rational doxastic dispositions and the epistemic regress problem. Dialectica: International Review of Philosophy of Knowledge, 70(4): 589–607, December 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dumsday:2016:NKE

[2080] Travis Dumsday. Natural-kind essentialism, substance ontology, and the unity problem: Two dispositionalist solutions. Dialectica: International Review of Philosophy of Knowledge, 70(4):609–626, December 2016. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Butterfield:2016:BRP

[2081] Paul Butterfield. Book review: The Philosophy of Conspiracy Theories — By Matthew Dentith. Dialectica: International Review of Philosophy of Knowledge, 70(4):627–632, December 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fish:2016:BRD

[2082] William Fish. Book review: Disjunctivism: Disjunctive Accounts in Epistemology and in the Philosophy of Perception — Edited by Marcus Willaschek. Dialectica: International Review of Philosophy of Knowledge, 70(4):632–635, December 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Lavers:2016:BRF

[2083] Gregory Lavers. Book review: Functions and Generality of Logic: Reflections on Dedekind’s and Frege’s Logicisms — By Hourya Benis-Sinaceur, Marco Panza and Gabriel Sandu. Dialectica: International Review of Philosophy of Knowledge, 70(4):636–640, December 2016. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2017:IIa

[2084] Anonymous. Issue information. Dialectica: International Review of Philosophy of Knowledge, 71(1):1, March 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Roelofs:2017:RAS

[2085] Luke Roelofs. Rational agency without self-knowledge: Could ‘we’ replace ‘i’? Dialectica: International Review of Philosophy of Knowledge, 71(1):3–33, March 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Salje:2017:CWA

[2086] Léa Salje. Crossed wires about crossed wires: Somatosensation and immunity to error through misidentification. Dialectica: International Review of Philosophy of Knowledge, 71(1):35–56, March 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mitchell:2017:EEE

[2087] Jonathan Mitchell. The epistemology of emotional experience. Dialectica: International Review of Philosophy of Knowledge, 71(1):57–84, March 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Boult:2017:CNC

[2088] Cameron Boult. Categorical norms and convention-Relativism about epistemic discourse. Dialectica: International Review of Philosophy of Knowledge, 71(1):85–99, March 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vlerick:2017:PCD

[2089] Michael Vlerick and Maarten Boudry. Psychological closure does not entail cognitive closure. Dialectica: International Review of Philosophy of Knowledge, 71(1):101–115, March 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Dellsen:2017:BRR

[2090] Finnur Dellsén. Book review: Reason and Explanation: A Defense of Explanatory Coherentism — By Ted Poston. Dialectica: International Review of Philosophy of Knowledge, 71(1):117–124, March 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Busch:2017:BRC

[2091] Kevin R. Busch. Book review: The Continuum Companion to Hume — Edited by Alan Bailey and Dan O’Brien. Dialectica: International Review of Philosophy of Knowledge, 71(1):124–140, March 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Chandler:2017:BRL

[2092] Jake Chandler. Book review: The Laws of Belief: Ranking Theory and its Philosophical Implications — By Wolfgang Spohn. Dialectica: International Review of Philosophy of Knowledge, 71(1):141–146, March 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jutten:2017:BRA

[2093] Timo Jütten. Book review: Axel Honneth — By Christopher F. Zurn. Dialectica: International Review of Philosophy of Knowledge, 71(1):146– 151, March 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2017:IIb

[2094] Anonymous. Issue information. Dialectica: International Review of Philosophy of Knowledge, 71(2):153, June 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hershenov:2017:PIP

[2095] David B. Hershenov and Adam P. Taylor. Personal identity and the possibility of autonomy. Dialectica: International Review of Philosophy of Knowledge, 71(2):155–179, June 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Monteleone:2017:CSE

[2096] John M. Monteleone. Concerns and the seriousness of emotion. Dialectica: International Review of Philosophy of Knowledge, 71(2):181–207, June 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sinclair:2017:RIF

[2097] Neil Sinclair. Reasons internalism and the function of normative reasons. Dialectica: International Review of Philosophy of Knowledge, 71 (2):209–229, June 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

McKenzie:2017:ABF

[2098] Kerry McKenzie. Against brute fundamentalism. Dialectica: International Review of Philosophy of Knowledge, 71(2):231–261, June 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Saad:2017:SRP

[2099] Bradford Saad. Should reductive physicalists reject the causal argument? Dialectica: International Review of Philosophy of Knowledge, 71 (2):263–279, June 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Muller:2017:HTE

[2100] Jean Moritz Müller. How (not) to think of emotions as evaluative attitudes. Dialectica: International Review of Philosophy of Knowledge, 71 (2):281–308, June 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Pappas:2017:BRH

[2101] George Pappas. Book review: How To Know: A Practicalist Conception of Knowledge — By Stephen Hetherington. Dialectica: International Review of Philosophy of Knowledge, 71(2):309–311, June 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tassone:2017:BRA

[2102] Biagio Gerard Tassone. Book review: Approaching Infinity — By Michael Huemer. Dialectica: International Review of Philosophy of Knowledge, 71(2):312–322, June 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kyriacou:2017:BRA

[2103] Christos Kyriacou. Book review: Assessment Sensitivity: Relative Truth and its Applications — By John MacFarlane. Dialectica: International Review of Philosophy of Knowledge, 71(2):322–332, June 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2017:IIc

[2104] Anonymous. Issue information. Dialectica: International Review of Philosophy of Knowledge, 71(3):333, September 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Carston:2017:IIS

[2105] Robyn Carston, Tim Pritchard, and Mark Textor. Introduction: Introduction to the special issue ‘Word Meaning — What it is and What it is not’. Dialectica: International Review of Philosophy of Knowledge, 71 (3):335–336, September 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Travis:2017:VMF

[2106] Charles Travis. Views of my fellows thinking (counting thoughts). Dialectica: International Review of Philosophy of Knowledge, 71(3):337– 378, September 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Recanati:2017:CP

[2107] François Recanati. Contextualism and polysemy. Dialectica: International Review of Philosophy of Knowledge, 71(3):379–397, September 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Allott:2017:LMC

[2108] Nicholas Allott and Mark Textor. Lexical modulation without concepts. Dialectica: International Review of Philosophy of Knowledge, 71(3):399– 424, September 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Cann:2017:WDW

[2109] Ronnie Cann and Ruth Kempson. What do words do for us? Dialectica: International Review of Philosophy of Knowledge, 71(3):425–460, September 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2017:IId

[2110] Anonymous. Issue information. Dialectica: International Review of Philosophy of Knowledge, 71(4):461, December 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1111/1746-8361.12157.

Yu:2017:MAP

[2111] Andy Demfree Yu. A modal account of propositions. Dialectica: International Review of Philosophy of Knowledge, 71(4):463–488, December 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1111/
1746-8361.12193.

Carrara:2017:CRC

[2112] Massimiliano Carrara and Giorgio Lando. Composition and relative counting. Dialectica: International Review of Philosophy of Knowledge, 71(4):489–529, December 2017. CODEN ???? ISSN 0012-2017 (print),

1746-8361 (electronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1111/1746-8361.12198.

Roberts:2017:FPL

[2113] Alan Roberts. Funny punny logic. Dialectica: International Review of Philosophy of Knowledge, 71(4):531–539, December 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1111/1746-8361.12200.

Miller:2017:FPC

[2114] Gregory Miller. Forming a positive concept of the phenomenal bonding relation for constitutive panpsychism. Dialectica: International Review of Philosophy of Knowledge, 71(4):541–562, December 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1111/1746-8361.12207.

Bourne:2017:EQM

[2115] Craig Bourne and Emily Caddick Bourne. Explanation and quasimiracles in narrative understanding: The case of poetic justice. Dialectica: International Review of Philosophy of Knowledge, 71(4):563–579, December 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1111/
1746-8361.12201.

Miller:2017:MSD

[2116] Alexander Miller. Moral supervenience: A defence of Blackburn’s argument. Dialectica: International Review of Philosophy of Knowledge, 71(4):581–601, December 2017. CODEN ???? ISSN 0012-2017 (print),

1746-8361 (electronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1111/1746-8361.12203.

Connolly:2017:IHN

[2117] Niall Connolly. I’m here now, but I won’t be here when you get this message. Dialectica: International Review of Philosophy of Knowledge, 71(4):603–622, December 2017. CODEN ???? ISSN 0012-2017 (print),

1746-8361 (electronic). URL https://onlinelibrary.wiley.com/doi/
abs/10.1111/1746-8361.12208.

Miller:2017:BRM

[2118] J. T. M. Miller. Book review: Meta-metaphysics: On Metaphysical Equivalence, Primitiveness, and Theory Choice — By Jiri Benovsky. Dialectica: International Review of Philosophy of Knowledge, 71(4):623– 628, December 2017. CODEN ???? ISSN 0012-2017 (print), 1746-

8361 (electronic). URL https://onlinelibrary.wiley.com/doi/abs/
10.1111/1746-8361.12195.

Torrengo:2017:BRT

[2119] Giuliano Torrengo and Cristian Mariani. Book review: Time: A Philosophical Introduction — By James Harrington. Dialectica: International Review of Philosophy of Knowledge, 71(4):628–631, December 2017. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/1746-8361.12196.

Baratella:2017:BRA

[2120] Riccardo Baratella and Claudio Calosi. Book review: Against Facts — By Arianna Betti. Dialectica: International Review of Philosophy of Knowledge, 71(4):631–640, December 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1111/1746-8361.12194.

Walker:2017:BRI

[2121] Jan Walker. Book review: Intuitionism vs. Classicism: A Mathematical Attack on Classical Logic — By Nick Haverkamp. Dialectica: International Review of Philosophy of Knowledge, 71(4):641–647, December 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1111/
1746-8361.12197.

Johnson:2017:BRI

[2122] Casey Rebecca Johnson. Book review: Intellectual Humility: An Introduction to the Philosophy and Science — By Ian M. Church & Peter L. Samuelson. Dialectica: International Review of Philosophy of Knowledge, 71(4):647–653, December 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1111/1746-8361.12204.

Schneider:2017:BRO

[2123] Luc Schneider. Book review: Ontology Made Easy — By Amie Thomasson. Dialectica: International Review of Philosophy of Knowledge, 71(4): 653–660, December 2017. CODEN ???? ISSN 0012-2017 (print), 1746-

8361 (electronic). URL https://onlinelibrary.wiley.com/doi/abs/
10.1111/1746-8361.12205.

Wurgler:2017:BRA

[2124] Judith Würgler. Book review: Against Authenticity: Why You Shouldn’t Be Yourself — By Simon Feldman. Dialectica: International Review of Philosophy of Knowledge, 71(4):660–669, December 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1111/1746-8361.12206.

Simons:2017:BRB

[2125] Peter Simons. Book review: Bolzano’s Conception of Grounding — By Stefan Roski. Dialectica: International Review of Philosophy of Knowledge, 71(4):669–671, December 2017. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1111/1746-8361.12209.

Anonymous:2018:IIa

[2126] Anonymous. Issue information. Dialectica: International Review of Philosophy of Knowledge, 72(1):1, March 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://onlinelibrary.

wiley.com/doi/10.1111/1746-8361.12184.

Hindriks:2018:CAM

[2127] Frank Hindriks. Collective agency: Moral and amoral. Dialectica: International Review of Philosophy of Knowledge, 72(1):3–23, March 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL

https://onlinelibrary.wiley.com/doi/10.1111/1746-8361.12215.

Vance:2018:CRP

[2128] Chad Vance. Causal relevance, permissible omissions, and famine relief. Dialectica: International Review of Philosophy of Knowledge, 72(1): 25–47, March 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361

(electronic). URL https://onlinelibrary.wiley.com/doi/10.1111/
1746-8361.12211.

Borgoni:2018:UB

[2129] Cristina Borgoni. Unendorsed beliefs. Dialectica: International Review of Philosophy of Knowledge, 72(1):49–68, March 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

onlinelibrary.wiley.com/doi/10.1111/1746-8361.12220.

Strandberg:2018:TET

[2130] Caj Sixten Strandberg. Towards an ecumenical theory of normative reasons. Dialectica: International Review of Philosophy of Knowledge, 72(1): 69–100, March 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361

(electronic). URL https://onlinelibrary.wiley.com/doi/10.1111/
1746-8361.12216.

Khani:2018:QIT

[2131] Ali Hossein Khani. Quine on the indeterminacy of translation: A dilemma for Davidson. Dialectica: International Review of Philosophy of Knowledge, 72(1):101–120, March 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://onlinelibrary.wiley.

com/doi/10.1111/1746-8361.12217.

Tooming:2018:TSA

[2132] Uku Tooming. There is something about the image: A defence of the two-component view of imagination. Dialectica: International Review of Philosophy of Knowledge, 72(1):121–139, March 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

onlinelibrary.wiley.com/doi/10.1111/1746-8361.12210.

Sorgiovanni:2018:RFP

[2133] Ben Sorgiovanni. Rule-following and primitive normativity. Dialectica: International Review of Philosophy of Knowledge, 72(1):141–150, March 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL

https://onlinelibrary.wiley.com/doi/10.1111/1746-8361.12218.

Lubrano:2018:BRP

[2134] Michele Lubrano. Book review: Pretense and Pathology: Philosophical Fictionalism and its Applications — By B. Armour-Garb & J. A. Woodbridge. Dialectica: International Review of Philosophy of Knowledge, 72(1):151–155, March 2018. CODEN ???? ISSN 0012-2017 (print),

1746-8361 (electronic). URL https://onlinelibrary.wiley.com/doi/
10.1111/1746-8361.12213.

Kaluzinski:2018:AEP

[2135] Bartosz Kaluziński. Advances in Experimental Philosophy of Language — By Jussi Haukioja. Dialectica: International Review of Philosophy of Knowledge, 72(1):155–159, March 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://onlinelibrary.wiley.

com/doi/10.1111/1746-8361.12212.

Rudnick:2018:BRB

[2136] Abraham Rudnick. Book review: The Bloomsbury Companion to Contemporary Philosophy of Medicine — By James A. Marcum. Dialectica: International Review of Philosophy of Knowledge, 72(1):159–162, March 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL

https://onlinelibrary.wiley.com/doi/10.1111/1746-8361.12214.

Anonymous:2018:IIb

[2137] Anonymous. Issue information. Dialectica: International Review of Philosophy of Knowledge, 72(2):163, June 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://onlinelibrary.

wiley.com/doi/10.1111/1746-8361.12185.

Hodgson:2018:MUW

[2138] Thomas Hodgson. Meaning underdetermines what is said, therefore utterances express many propositions. Dialectica: International Review of Philosophy of Knowledge, 72(2):165–189, June 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

onlinelibrary.wiley.com/doi/10.1111/1746-8361.12221.

Giubilini:2018:WWC

[2139] Alberto Giubilini and Neil Levy. What in the world is collective responsibility? Dialectica: International Review of Philosophy of Knowledge, 72 (2):191–217, June 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361

(electronic). URL https://onlinelibrary.wiley.com/doi/10.1111/
1746-8361.12228.

Muller:2018:SPM

[2140] Thomas Müller and Hans J. Briegel. A stochastic process model for free agency under indeterminism. Dialectica: International Review of Philosophy of Knowledge, 72(2):219–252, June 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://onlinelibrary.

wiley.com/doi/10.1111/1746-8361.12222.

Baumann:2018:WWB

[2141] Peter Baumann. What will be best for me? Big decisions and the problem of inter-world comparisons. Dialectica: International Review of Philosophy of Knowledge, 72(2):253–273, June 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://onlinelibrary.

wiley.com/doi/10.1111/1746-8361.12219.

Timmerman:2018:DNB

[2142] Travis Timmerman. Doomsday needn’t be so bad. Dialectica: International Review of Philosophy of Knowledge, 72(2):275–296, June 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL

https://onlinelibrary.wiley.com/doi/10.1111/1746-8361.12227.

Bacs:2018:MFE

[2143] Gábor Bács. Mental fictionalism and epiphenomenal qualia. Dialectica: International Review of Philosophy of Knowledge, 72(2):297–308, June 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL

https://onlinelibrary.wiley.com/doi/10.1111/1746-8361.12229.

Azzano:2018:BRM

[2144] Lorenzo Azzano. Book review: Massimiliano Carrara, Alexandra Arapinis, and Friederike Moltmann (eds.), Unity & Plurality: Logic, Philosophy, and Linguistics, Oxford: Oxford University Press, 2016, xv + 259 pp., £45 (hardback), ISBN 978-0-19-871632-7. Dialectica: International Review of Philosophy of Knowledge, 72(2):309–315, June 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

onlinelibrary.wiley.com/doi/10.1111/1746-8361.12230.

Jaquet:2018:BRM

[2145] François Jaquet. Book review: Marcus Arvan, Rightness as Fairness: A Moral and Political Theory, Basingstoke UK: Palgrave Macmillan, 2016, xi + 271 pp., £88.39 (hardback), ISBN 978-1-137-54180-2, eBook ISBN 978-1-137-54181-9. Dialectica: International Review of Philosophy of Knowledge, 72(2):315–320, June 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://onlinelibrary.wiley.

com/doi/10.1111/1746-8361.12231.

Pinzani:2018:BRS

[2146] Alessandro Pinzani. Book review: Stefan Müller–Doohm, Habermas: A Biography. Cambridge: Polity Press, 2016, 598 pp., £25 (hardback), ISBN 978-0-7456-8906-7. Dialectica: International Review of Philosophy of Knowledge, 72(2):320–324, June 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://onlinelibrary.wiley.

com/doi/10.1111/1746-8361.12232.

Michels:2018:BRD

[2147] Robert Michels. Book review: David Wiggins, Continuants: Their Activity, Their Being and Their Identity. Twelve Essays, Oxford/New York: Oxford University Press, 2016, 239 pp., $50.00 (£35.00) (hardback), ISBN: 978-0-19-871662-4. Dialectica: International Review of Philosophy of Knowledge, 72(2):325–328, June 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://onlinelibrary.wiley.

com/doi/10.1111/1746-8361.12233.

Anonymous:2018:IIc

[2148] Anonymous. Issue information. Dialectica: International Review of Philosophy of Knowledge, 72(3):329, September 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kalpokas:2018:NCM

[2149] Daniel E. Kalpokas. Non-conceptualism and the myth of the given. Dialectica: International Review of Philosophy of Knowledge, 72(3):331– 363, September 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zangwill:2018:BNN

[2150] Nick Zangwill. Brutalist non-naturalism and Hume’s principle. Dialectica: International Review of Philosophy of Knowledge, 72(3):365–383, September 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hommen:2018:FOP

[2151] David Hommen. Frames and the ontology of particular objects. Dialectica: International Review of Philosophy of Knowledge, 72(3):385–409, September 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Gillessen:2018:RVH

[2152] Jens Gillessen. Rationality, virtue and higher-order coherence. Dialectica: International Review of Philosophy of Knowledge, 72(3):411–436, September 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Clarke:2018:CCR

[2153] Henry Clarke. Concepts and communication: A reply to Onofri. Dialectica: International Review of Philosophy of Knowledge, 72(3):437–444, September 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Heyndels:2018:SRD

[2154] Sybren Heyndels and Benjamin De Mesel. On Shoemaker’s responsedependent theory of responsibility. Dialectica: International Review of Philosophy of Knowledge, 72(3):445–451, September 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marcoci:2018:DR

[2155] Alexandru Marcoci. On a dilemma of redistribution. Dialectica: International Review of Philosophy of Knowledge, 72(3):453–460, September 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Sachse:2018:BRL

[2156] Christian Sachse. Book review: L’esprit matériel. Réduction et émergence — by Max Kistler. Dialectica: International Review of Philosophy of Knowledge, 72(3):461–466, September 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Plate:2018:BRM

[2157] Jan Plate. Book review: Making Things Up — by Karen Bennett. Dialectica: International Review of Philosophy of Knowledge, 72(3):466– 473, September 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Plebani:2018:BRT

[2158] Matteo Plebani. Book review: Talking About Numbers: Easy Arguments for Mathematical Realism — by Katharina Felka. Dialectica: International Review of Philosophy of Knowledge, 72(3):473–479, September 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Siderits:2018:BRI

[2159] Mark Siderits. Book review: Indian Epistemology and Metaphysics — By Joerg Tuske. Dialectica: International Review of Philosophy of Knowledge, 72(3):479–484, September 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2018:II

[2160] Anonymous. Issue information. Dialectica: International Review of Philosophy of Knowledge, 72(4):485, December 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Verdejo:2018:TSC

[2161] Víctor M. Verdejo. Thought sharing, communication, and perspectives about the self. Dialectica: International Review of Philosophy of Knowledge, 72(4):487–507, December 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Boucher:2018:WRB

[2162] Sandy C. Boucher. What is the relation between a philosophical stance and its associated beliefs? Dialectica: International Review of Philosophy of Knowledge, 72(4):509–524, December 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fratantonio:2018:AAI

[2163] Giada Fratantonio. Armchair access and imagination. Dialectica: International Review of Philosophy of Knowledge, 72(4):525–547, December 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Butzer:2018:DPD

[2164] Tim Butzer. Dodging the perils of dogmatism: A response to Crispin Wright. Dialectica: International Review of Philosophy of Knowledge, 72(4):549–569, December 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Keinanen:2018:KTK

[2165] Markku Keinänen, Jani Hakkarainen, and Antti Keskinen. Kinds of tropes without kinds. Dialectica: International Review of Philosophy of Knowledge, 72(4):571–596, December 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Peterson:2018:ANW

[2166] Martin Peterson. The anti-nihilist wager. Dialectica: International Review of Philosophy of Knowledge, 72(4):597–602, December 2018. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Haraldsen:2018:YCS

[2167] Fredrik Haraldsen. You cannot steal something that doesn’t exist: Against fictionalism about fiction. Dialectica: International Review of Philosophy of Knowledge, 72(4):603–616, December 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Stern:2018:BRG

[2168] Johannes Stern. Book review: Gödel’s Basic Logic Course at Notre Dame — By Miloš Adžić and Kosta Došen. Dialectica: International Review of Philosophy of Knowledge, 72(4):617–622, December 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Georgi:2018:BRL

[2169] Geoff Georgi. Book review: Logical Form: Between Logic and Natural Language — By Andrea Iacona. Dialectica: International Review of Philosophy of Knowledge, 72(4):622–628, December 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Carrara:2018:BRM

[2170] Massimiliano Carrara and Filippo Mancini. Book review: Mereology: A Philosophical Introduction — By Giorgio Lando. Dialectica: International Review of Philosophy of Knowledge, 72(4):628–633, December 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Vaccarezza:2018:BRE

[2171] Maria Silvia Vaccarezza. Book review: Exemplarist Moral Theory — By Linda T. Zagzebski. Dialectica: International Review of Philosophy of Knowledge, 72(4):633–640, December 2018. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2019:IIa

[2172] Anonymous. Issue information. Dialectica: International Review of Philosophy of Knowledge, 73(1–2):1, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Roca-Royes:2019:CEE

[2173] Sonia Roca-Royes. Concepts and the epistemology of essence. Dialectica: International Review of Philosophy of Knowledge, 73(1–2):3–29, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kurbis:2019:AML

[2174] Nils Kürbis. An argument for minimal logic. Dialectica: International Review of Philosophy of Knowledge, 73(1–2):31–63, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Blumson:2019:NCC

[2175] Ben Blumson. Naturalness and convex class nominalism. Dialectica: International Review of Philosophy of Knowledge, 73(1–2):65–81, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Browning:2019:MCI

[2176] Jacob Browning. McDowell and the contents of intuition. Dialectica: International Review of Philosophy of Knowledge, 73(1–2):83–104, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Furst:2019:CCP

[2177] Martina Fürst. On the conceivability of a cognitive phenomenology zombie. Dialectica: International Review of Philosophy of Knowledge, 73 (1–2):105–127, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Marchesi:2019:BIC

[2178] Andrea Marchesi. Brentanian inner consciousness and the infinite regress problem. Dialectica: International Review of Philosophy of Knowledge, 73(1–2):129–147, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Eagle:2019:WL

[2179] Antony Eagle. Weak location. Dialectica: International Review of Philosophy of Knowledge, 73(1–2):149–181, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Riedener:2019:SBW

[2180] Stefan Riedener. The standing to blame, or why moral disapproval is what it is. Dialectica: International Review of Philosophy of Knowledge, 73(1–2):183–210, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Hennig:2019:SKK

[2181] Boris Hennig. Self-knowledge as knowledge of the good: Hugh of St. Victor on self-knowledge. Dialectica: International Review of Philosophy of Knowledge, 73(1–2):211–230, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Bowers:2019:TAS

[2182] Jason Bowers. A teleological answer to the special composition question. Dialectica: International Review of Philosophy of Knowledge, 73(1–2):

231–246, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Solomyak:2019:PSP

[2183] Olla Solomyak. Presentism and the specious present: From temporal experience to meta-metaphysics. Dialectica: International Review of Philosophy of Knowledge, 73(1–2):247–266, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Fano:2019:BRN

[2184] Vincenzo Fano and Gino Tarozzi. Book review: Niels Bohr and the Philosophy of Physics: Twenty-First Century Perspectives — By J. Faye and H. J. Folse. Dialectica: International Review of Philosophy of Knowledge, 73(1–2):267–273, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Zanetti:2019:BRF

[2185] Luca Zanetti. Book review: Free Will and Epistemology: A Defence of the Transcendental Argument for Freedom — By Robert Lockie. Dialectica: International Review of Philosophy of Knowledge, 73(1–2):273–279, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Logins:2019:BRB

[2186] Arturs Logins. Book review: Basic Knowledge and Conditions on Knowledge — By Mark McBride. Dialectica: International Review of Philosophy of Knowledge, 73(1–2):280–285, March 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2019:IIb

[2187] Anonymous. Issue information. Dialectica: International Review of Philosophy of Knowledge, 73(3):287, September 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rattan:2019:USC

[2188] Gurpreet Rattan. Understanding semantic coordination in cognition. Dialectica: International Review of Philosophy of Knowledge, 73(3):289– 313, September 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Baras:2019:HCI

[2189] Dan Baras. How close are impossible worlds? A critique of Brogaard and Salerno’s account of counterpossibles. Dialectica: International Review of Philosophy of Knowledge, 73(3):315–329, September 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Schulte:2019:CLR

[2190] Peter Schulte. Challenging liberal representationalism: a reply to Artiga. Dialectica: International Review of Philosophy of Knowledge, 73(3):331– 348, September 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). See [2067].

Thorn:2019:FSR

[2191] Paul D. Thorn. A formal solution to Reichenbach’s reference class problem. Dialectica: International Review of Philosophy of Knowledge, 73 (3):349–366, September 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Kim:2019:PFE

[2192] Joongol Kim. The problem of Fregean equivalents. Dialectica: International Review of Philosophy of Knowledge, 73(3):367–394, September 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

DeRizzo:2019:HAA

[2193] Julio De Rizzo. How (not) to argue against brute fundamentalism. Dialectica: International Review of Philosophy of Knowledge, 73(3):395– 410, September 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Tersman:2019:SAR

[2194] Folke Tersman. From scepticism to anti-realism. Dialectica: International Review of Philosophy of Knowledge, 73(3):411–427, September 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Choi:2019:RCI

[2195] Sungho Choi. Reverse-cycle intrinsic finks. Dialectica: International Review of Philosophy of Knowledge, 73(3):429–455, September 2019. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Anonymous:2019:IIc

[2196] Anonymous. Issue information. Dialectica: International Review of Philosophy of Knowledge, 73(4):457, December 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Jansen:2019:FMT

[2197] Charles M. Jansen. De-fining material things. Dialectica: International Review of Philosophy of Knowledge, 73(4):459–477, December 2019. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Moore:2019:CEP

[2198] Dwayne Moore. Causal exclusion and physical causal completeness. Dialectica: International Review of Philosophy of Knowledge, 73(4):479– 505, December 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Christias:2019:TRL

[2199] Dionysis Christias. Towards a reformed liberal and scientific naturalism. Dialectica: International Review of Philosophy of Knowledge, 73(4):507– 534, December 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Mitchell:2019:EEP

[2200] Jonathan Mitchell. Emotional experience and propositional content. Dialectica: International Review of Philosophy of Knowledge, 73(4):535– 561, December 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Rettler:2019:QOR

[2201] Bradley Rettler. Quantification in the ontology room. Dialectica: International Review of Philosophy of Knowledge, 73(4):563–585, December 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Panajoli:2019:BRT

[2202] Tommaso Panajoli. Book review: Thomas Kuhn’s Revolutions: An Historical and an Evolutionary Philosophy of Science? — By James A. Marcum. Dialectica: International Review of Philosophy of Knowledge, 73(4):587–590, December 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Richard:2019:BRE

[2203] Sébastien Richard. Book review: Essais et conférences I — By Franz Brentano. Dialectica: International Review of Philosophy of Knowledge, 73(4):591–597, December 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Humbert-Droz:2019:BRS

[2204] Steve Humbert-Droz. Book review: Supposition and the Imaginative Realm. A Philosophical Inquiry — By Margherita Arcangeli. Dialectica: International Review of Philosophy of Knowledge, 73(4):598–602, December 2019. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Blum:2020:EDG

[2205] Philipp Blum. Editorial: Dialectica goes open access. Dialectica: International Review of Philosophy of Knowledge, 74(1):1–2, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL

https://dialectica.philosophie.ch/dialectica/article/view/2.

Conitzer:2020:PTT

[2206] Vincent Conitzer. The personalized A-theory of time and perspective. Dialectica: International Review of Philosophy of Knowledge, 74(1):3–31, ???? 2020. CODEN ???? ISSN 0012-2017 (print),

1746-8361 (electronic). URL https://dialectica.philosophie.ch/
dialectica/article/view/3.

Elzein:2020:DOI

[2207] Nadine Elzein. Determinism, “Ought” implies “Can” and moral obligation. Dialectica: International Review of Philosophy of Knowledge, 74(1):33–60, ???? 2020. CODEN ???? ISSN 0012-2017 (print),

1746-8361 (electronic). URL https://dialectica.philosophie.ch/
dialectica/article/view/4.

Roelofs:2020:CRC

[2208] Luke Roelofs. Consciousness, revelation, and confusion: Are constitutive panpsychists hoist by their own petard? Dialectica: International Review of Philosophy of Knowledge, 74(1):61–93, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://dialectica.

philosophie.ch/dialectica/article/view/5.

OConaill:2020:DFG

[2209] Donnchadh O’Conaill. In defence of facts: Grounding, essential properties and the unity problem. Dialectica: International Review of Philosophy of Knowledge, 74(1):95–123, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://dialectica.

philosophie.ch/dialectica/article/view/6.

Goh:2020:SCL

[2210] Joshua Goh. Strevens’s counterexample to Lewis’s “Causation as influence”, and degrees of causation. Dialectica: International Review of Philosophy of Knowledge, 74(1):125–137, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://dialectica.

philosophie.ch/dialectica/article/view/7.

Daoust:2020:COA

[2211] Marc-Kevin Daoust. Consistency, obligations, and accuracy-dominance vindications. Dialectica: International Review of Philosophy of Knowledge, 74(1):139–156, ???? 2020. CODEN ???? ISSN 0012-2017 (print),

1746-8361 (electronic). URL https://dialectica.philosophie.ch/
dialectica/article/view/8.

MacBride:2020:RS

[2212] Fraser MacBride. Review of Soames (2018). Dialectica: International Review of Philosophy of Knowledge, 74(1):157–162, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

dialectica.philosophie.ch/dialectica/article/view/9.

Scharli:2020:RO

[2213] Mario Schärli. Review of Oppy (2018). Dialectica: International Review of Philosophy of Knowledge, 74(1):163–169, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

dialectica.philosophie.ch/dialectica/article/view/10.

Taieb:2020:RA

[2214] Hamid Taieb. Review of Antonelli (2018). Dialectica: International Review of Philosophy of Knowledge, 74(1):171–176, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

dialectica.philosophie.ch/dialectica/article/view/11.

Muller:2020:PAF

[2215] F. A. Muller. The paradox of the arche-fossil: an analysis of Meillassoux’s challenge to correlationism, idealism included. Dialectica: International Review of Philosophy of Knowledge, 74(3):429–462, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

dialectica.philosophie.ch/dialectica/article/view/21.

Vollet:2020:CA

[2216] Jacques-Henri Vollet. Certainty and assertion. Dialectica: International Review of Philosophy of Knowledge, 74(3):463–494, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

dialectica.philosophie.ch/dialectica/article/view/22.

Dixon:2020:RKR

[2217] Jonathan Dixon. Reliable knowledge: a reply to Turri. Dialectica: International Review of Philosophy of Knowledge, 74(3):495–508, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (elec-

tronic). URL https://dialectica.philosophie.ch/dialectica/
article/view/23.

Baratella:2020:TOC

[2218] Riccardo Baratella. Are there occurrent continuants?: a reply to Stout’s “The category of occurrent continuants”. Dialectica: International Review of Philosophy of Knowledge, 74(3):509–519, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

dialectica.philosophie.ch/dialectica/article/view/24.

Zimmerman:2020:RRA

[2219] Dean Zimmerman. Robinson’s regress argument from vagueness to dualism: Critical notice to Robinson (2016). Dialectica: International Review of Philosophy of Knowledge, 74(3):521–549, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

dialectica.philosophie.ch/dialectica/article/view/25.

Campdelacreu:2020:HTS

[2220] Marta Campdelacreu, Ramón García-Moya, Genoveva Martí, and Enrico Terrone. How to test the ship of Theseus. Dialectica: International Review of Philosophy of Knowledge, 74(3):551–559, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

dialectica.philosophie.ch/dialectica/article/view/26.

Morvarid:2020:MGC

[2221] Mahmoud Morvarid. Mereology is not a guide to (in)conceivability: a reply to Giberman. Dialectica: International Review of Philosophy of Knowledge, 74(3):561–574, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://dialectica.

philosophie.ch/dialectica/article/view/27.

Stephenson:2020:RW

[2222] Andrew Stephenson. Review of Willaschek (2018). Dialectica: International Review of Philosophy of Knowledge, 74(3):575–581, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (elec-

tronic). URL https://dialectica.philosophie.ch/dialectica/
article/view/28.

Wilson:2020:RED

[2223] Alastair Wilson. Review of Esfeld and Deckert (2018). Dialectica: International Review of Philosophy of Knowledge, 74(3):583–589, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (elec-

tronic). URL https://dialectica.philosophie.ch/dialectica/
article/view/29.

Daly:2020:RM

[2224] Chris Daly. Review of MacBride (2018). Dialectica: International Review of Philosophy of Knowledge, 74(3):591–595, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

dialectica.philosophie.ch/dialectica/article/view/30.

Stevens:2020:MMD

[2225] Graham Stevens. Metalinguistic monstrosity and displaced communications. Dialectica: International Review of Philosophy of Knowledge, 74(4):597–634, ???? 2020. CODEN ???? ISSN 0012-2017 (print),

1746-8361 (electronic). URL https://dialectica.philosophie.ch/
dialectica/article/view/31.

Cox:2020:CAI

[2226] Ryan Cox. Constitutivism about instrumental desire and introspective belief. Dialectica: International Review of Philosophy of Knowledge, 74(4):635–664, ???? 2020. CODEN ???? ISSN 0012-2017 (print),

1746-8361 (electronic). URL https://dialectica.philosophie.ch/
dialectica/article/view/32.

Liberman:2020:MSF

[2227] Alida Liberman. The mental states first theory of promising. Dialectica: International Review of Philosophy of Knowledge, 74(4): 665–692, ???? 2020. CODEN ???? ISSN 0012-2017 (print),

1746-8361 (electronic). URL https://dialectica.philosophie.ch/
dialectica/article/view/33.

Finocchiaro:2020:PAP

[2228] Peter Finocchiaro. A puzzle about parsimony. Dialectica: International Review of Philosophy of Knowledge, 74(4):693–708, ???? 2020. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

dialectica.philosophie.ch/dialectica/article/view/34.

Donaldson:2020:DAM

[2229] Thomas Donaldson. David Armstrong on the metaphysics of mathematics. Dialectica: International Review of Philosophy of Knowledge, 74(4):709–731, ???? 2020. CODEN ???? ISSN 0012-2017 (print),

1746-8361 (electronic). URL https://dialectica.philosophie.ch/
dialectica/article/view/35.

Nicolas:2021:MMT

[2230] David Nicolas. Mixtures and mass terms. Dialectica: International Review of Philosophy of Knowledge, 75(1):1–14, ???? 2021. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

dialectica.philosophie.ch/dialectica/article/view/40.

Schneider:2021:AFA

[2231] Leonhard Schneider. An analysis of Fink’s argument in favour of normative process-requirements. Dialectica: International Review of Philosophy of Knowledge, 75(1):15–23, ???? 2021. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://dialectica.

philosophie.ch/dialectica/article/view/19.

Burnston:2021:PLC

[2232] Daniel C. Burnston. Perceptual learning, categorical perception, and cognitive permeation. Dialectica: International Review of Philosophy of Knowledge, 75(1):25–58, ???? 2021. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://dialectica.

philosophie.ch/dialectica/article/view/36.

Weintraub:2021:DUH

[2233] Ruth Weintraub. The dis-unity of Humean space. Dialectica: International Review of Philosophy of Knowledge, 75(1):59–83, ???? 2021. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https:

//dialectica.philosophie.ch/dialectica/article/view/37.

Vanrie:2021:BBG

[2234] Wim Vanrie and Maarten Van Dyck. Boghossian, Bellarmine, and Galileo: Adjudication and epistemic relativism. Dialectica: International Review of Philosophy of Knowledge, 75(1):85–118, ???? 2021. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL https://

dialectica.philosophie.ch/dialectica/article/view/38.

Laasik:2021:PRP

[2235] Kristjan Laasik. Perspectivity and rationality of perception. Dialectica: International Review of Philosophy of Knowledge, 75(1): 119–136, ???? 2021. CODEN ???? ISSN 0012-2017 (print),

1746-8361 (electronic). URL https://dialectica.philosophie.ch/
dialectica/article/view/39.

Heisenberg:1974:NCT

[2236] Werner Heisenberg. The notion of a “closed theory” in modern science. In Across the Frontiers [2238], pages 39–46. ISBN 0-06-011824-5. LCCN

???? URL http://alsos.wlu.edu/information.aspx?id=941.

Brogaard:2013:RC

[2237] Berit Brogaard and Joe Salerno. Remarks on counterpossibles. Synthese, 190(4):639–660, March 2013. CODEN SYNTAE. ISSN 0039-7857 (print),

1573-0964 (electronic). URL http://link.springer.com/article/10.
1007/s11229-012-0196-6;   http://link.springer.com/content/
pdf/10.1007/s11229-012-0196-6.pdf. See critique [2189].

Heisenberg:1974:AF

[2238] Werner Heisenberg. Across the Frontiers, volume 48 of World perspectives. Harper & Row, New York, NY, USA, 1974. ISBN 0-06-011824-5.

xxii + 229 pp. LCCN ???? URL http://alsos.wlu.edu/information.
aspx?id=941.